### v1

# From Many Models to ONE THEORY From the Dark Ages to Enlightenment J. Iliopoulos

ABSTRACT
The year 2024 marks two anniversaries: The 70th anniversary of CERN and the 50th anniversary of the $J/\Psi$ discovery. At this occasion I have been asked to give review talks on the significance of these anniversaries. This article is an expanded version of these talks.
1 Introduction
By a worth noting coincidence, the year 2024 marks two anniversaries: The 70th anniversary of CERN[1] and the 50th anniversary of the $J/\Psi$ discovery[2].
CERN’s founding fathers were among Europe’s leading scientists. CERN became the most successful International Institution. Scientists built the Scientific European Union long before politicians thought of the Economic European Union. If the Scientific Institution appears to be more solid than the Economic one, it is because its foundations are made with ideas.
Regarding the second anniversary, I want to argue that $J/\Psi$ was not just a new resonance; we knew already a large number of them. It was not even only the first indirect evidence for the existence of a new quark flavor. It was all that, but it was also much more. It was the final proof which convinced the large majority of our community that we were witnessing a radical change of paradigm in our understanding of microscopic physics. From phenomenological models and specific theories, each one applied to a restricted set of experimental data, we had to think in terms of a fundamental theory of universal validity. From many models to the STANDARD THEORY of particle physics. For most of us this transition was a revelation, for some others it was a painful experience. It is appropriate to combine it with CERN’s anniversary because the experimental verification of the Standard Model started at CERN with the neutral currents and was completed also at CERN with the BEH scalar boson. It is this story that I will attempt to narrate in this note, although I do not consider myself to be an unbiased observer.
In order to avoid any misunderstanding, I want to start by emphasising what this article is not meant to be. It is not a review article. I will not cite all important papers which were published in the last one hundred years. Although it tells a story, it is not a scholarly article on the history of high energy physics. I do not feel capable of writing such a history. My purpose is to present a personal view of some ideas among those which contributed to the establishment of the Standard Theory. It is neither exhaustive nor objective.
2 Elementary Particles: The origins
[…]
Ιφ Ι ηαδ το αςςιγν α ψεαρ το τηε βιρτη οφ εξπεριμενταλ παρτιςλε πηψςιςς Ι ωουλδ ςηοοςε 1950, τηε διςςοερψ οφ τηε νευτραλ πιον βψ Η.Θ. Στεινβεργερ, Ω.Κ.Η. Πανοφςϰψ ανδ Θ.Σ. Στελλερ ατ τηε Βερϰελεψ ελεςτρον ςψνςηροτρον[4]. Ιτ ις τηε φιρςτ “ελεμενταρψ” παρτιςλε διςςοερεδ ωιτη αν αςςελερατορ (τηε ςηαργεδ πιον ωας διςςοερεδ ιν 1947 ιν ςοςμις ραψς [^2] The first cyclotron was built at Berkeley by E. O. Laurence and his student M. S. Livingstone in 1930. It was a toy accelerator with a diameter of only 4 inches, but in the following years Laurence built in his laboratory a series of larger and larger accelerators reaching a diameter of 184 inches. They were capable of producing pions abundantly but the experimentalists had not yet developed detection techniques suitable to the new environment. It is not enough to acquire a new toy, you must also learn how to play with it. It is also true that Laurence was more interested in building accelerators than in making experiments and did not allocate sufficient resources to detectors.). Τηε εξιςτενςε οφ $\pi^{0}$ ωας πρεδιςτεδ ιν 1938 βψ Ν. Κεμμερ[5] ωηο ωροτε τηε φιρςτ ιςοςπιν ιναριαντ τηεορψ φορ τηε νυςλεαρ φορςες [^3] Kemmer has not received the appropriate recognition for his groundbreaking work. His equations for the pion-nucleon interaction can be found in every textbook, but his name is rarely mentioned.. Σο, $\pi^{0}$ ις τηε φιρςτ παρτιςλε ωηοςε εξιςτενςε ωας πρεδιςτεδ βψ αν αργυμεντ βαςεδ ον αν ιντερναλ ςψμμετρψ ανδ αλςο τηε φιρςτ παρτιςλε το βε διςςοερεδ ιν αν αςςελερατορ. Σινςε τηατ τιμε αςςελερατορς βεςαμε τηε μαιν ενγινες οφ διςςοερψ ιν παρτιςλε πηψςιςς[6]. Τηε φιρςτ ςονςεχυενςε ωας τηε ςεπαρατιον οφ τηε τωο ςομμυνιτιες, παρτιςλε ανδ ςοςμις ραψ πηψςιςιςτς [^4] The last conference which was common to both was held in France, at Bagnères-de-Bigorre, in July 1953. It was organised by the French L. Leprince-Ringuet and the cosmic ray physicists complained claiming that he allocated too much time to the still scarce results coming from accelerators. C.F. Powell, the discoverer of the charged pion and a leading figure in cosmic ray research, said in his closing lecture: “Gentlemen, we have been invaded …The accelerators are here”. An immediate measure of this “invasion”: In the same conference the Australian R.H. Dalitz presented a simple method to study the properties of the newly discovered particles which exhibited three-body decays. It is the famous “Dalitz plot” in which every event is represented by a point in a graph. In 1953 Dalitz had 13 points in his plot, all coming from cosmic rays. At the 1955 Rochester Conference he had 53, with 42 still coming from cosmic rays. The following year he had more than 600, mostly coming from accelerators. Note however, that in recent years the two communities often joined forces again with the emergence of a new discipline, the “Astroparticles”.. Ωιτη τηε υςε οφ αςςελερατορς τηε φιελδ εξπανδεδ ερψ ραπιδλψ ανδ τηε νατυρε οφ εξπεριμενταλ ωορϰ ςηανγεδ [^5] We should also acknowledge the fact that this rapid expansion was to a certain extent due to the success of the Manhattan project during the war.. Δεδιςατεδ ρεςεαρςη ςεντερς ωερε βυιλτ – Βροοϰηαεν ιν τηε ϒΣΑ, ῝ΕΡΝ ιν Ευροπε – ωηιςη ωερε νοτ αςςοςιατεδ ωιτη α παρτιςυλαρ ϒνιερςιτψ. Τηε δριε το βυιλδ λαργερ ανδ μορε ςομπλεξ δετεςτορς, ανδ τηε ρατε ωιτη ωηιςη δατα ωερε ςολλεςτεδ, λεδ το τηε ςρεατιον οφ λαργε ιντερνατιοναλ ςολλαβορατιονς. Ωε εντερεδ τηε ερα οφ “Βιγ Σςιενςε”. Α μεαςυρε οφ τηις εξπονεντιαλ εξπανςιον ις γιεν βψ τηε νυμβερ οφ ελεμενταρψ παρτιςλες. Ιν 1940 ωε ϰνεω τηε προτον ανδ τηε νευτρον, τηε ελεςτρον ανδ τηε νευτρινο (νοτ ψετ δετεςτεδ), τηε πηοτον ανδ τηε παρτιςλε ωηιςη ωας βελιεεδ το βε τηε Ψυϰαωα μεςον (ιτ ωας ιν φαςτ τηε μυον). Ιν 1960 τηε Ταβλε οφ Ελεμενταρψ Παρτιςλες ηαδ ςεεραλ δοζεν εντριες ανδ τηατ οφ τοδαψ ςεεραλ ηυνδρεδ, αλτηουγη ωε ϰνοω τηατ ερψ φεω αμονγ τηεμ αρε “ελεμενταρψ”.
Ιφ εξπεριμενταλ παρτιςλε πηψςιςς φολλοωεδ α μονοτονιςαλλψ ριςινγ τραϑεςτορψ, τηατ οφ ιτς τηεορετιςαλ ςουντερπαρτ ωας μορε ςιρςυιτους. Μοδερν τηεορετιςαλ πηψςιςς ηας α πρεςιςε δατε οφ βιρτη: Θυνε 2 1947, τηε δατε οφ τηε Σηελτερ Ιςλανδ ῝ονφερενςε. Ιτ ις τηε φιρςτ ποςτωαρ μεετινγ ον τηε Φουνδατιονς οφ Χυαντυμ Μεςηανιςς, ςπονςορεδ βψ τηε ϒΣΑ Νατιοναλ Αςαδεμψ οφ Σςιενςες. Ιτ ωας ηελδ ατ Λονγ Ιςλανδ´ς Σηελτερ Ιςλανδ ατ Ραμ´ς Ηεαδ Ινν ανδ γατηερεδ 24 παρτιςιπαντς αρουνδ Θ.Ρ. Οππενηειμερ. Ωιτη τηε αυρα οφ τηε Μανηατταν Προϑεςτ, ηε ωας ςονςιδερεδ ας τηε φουνδινγ φατηερ οφ τηε Αμεριςαν Σςηοολ οφ τηεορετιςαλ πηψςιςς. Ρεγαρδινγ τηε ῝ονφερενςε, Οππενηειμερ λατερ δεςλαρεδ ιτ ωας τηε μοςτ ςυςςεςςφυλ ςςιεντιφις μεετινγ ηε ηαδ εερ αττενδεδ. Ινδεεδ, ιν ρετροςπεςτ Σηελτερ Ιςλανδ ις α λανδμαρϰ οφ τηεορετιςαλ πηψςιςς. Τηε μοςτ ιμπορταντ ςοντριβυτιονς ωηιςη ωερε πρεςεντεδ ιν τηατ ςονφερενςε ωερε νοτ τηεορετιςαλ βρεαϰτηρουγης βυτ τωο εξπεριμενταλ ρεςυλτς: τηε νον-ζερο αλυες οφ τηε Λαμβ ςηιφτ ανδ οφ τηε ελεςτρον ανομαλους μαγνετις μομεντ. Βοτη τηεςε ρεςυλτς ωερε ιμπορταντ βεςαυςε, φορ τηε φιρςτ τιμε, τηεψ ςηοωεδ, βεψονδ ανψ ποςςιβλε δουβτ, τηατ τηε πρεδιςτιονς βαςεδ ον τηε Διρας εχυατιον φορ τηε ελεςτρον αρε ονλψ αππροξιματελψ ςορρεςτ. Τηις ιν τυρν μοτιατεδ α ςεριους ςτυδψ οφ χυαντυμ φιελδ τηεορψ. Ας Σ. Ωεινβεργ ηας πυτ ιτ[7]: “Τηε γρεατ τηινγ αςςομπλιςηεδ βψ τηε διςςοερψ οφ τηε Λαμβ ςηιφτ ωας νοτ ςο μυςη τηατ ιτ φορςεδ υς το ςηανγε ουρ πηψςιςαλ τηεοριες, ας τηατ ιτ φορςεδ υς το ταϰε τηεμ ςεριουςλψ”. Ινδεεδ, τηε φορμαλιςμ οφ χυαντυμ φιελδ τηεορψ εξιςτεδ αλρεαδψ φορ βοτη βοςονς ανδ φερμιονς. Βυτ τηε ενορμους πρεςτιγε οφ τηε Διρας τηεορψ ον τηε ονε ηανδ, ανδ τηε αβςενςε οφ α ςλεαρ πηψςιςαλ μοτιατιον ον τηε οτηερ, διςςουραγεδ τηεοριςτς το φαςε ςεριουςλψ τηε προβλεμ οφ τηε διεργενςες οφ περτυρβατιον τηεορψ. Τηε φαςτ τηατ τηε τηεορετιςαλ ιδεας εξιςτεδ ιν α ςυβςονςςιους φορμ ις ωιτνεςςεδ βψ τηε φαςτ τηατ ιτ τοοϰ πραςτιςαλλψ νο τιμε το δεελοπ τηεμ. Τηε φιρςτ εςτιματιον οφ τηε Λαμβ ςηιφτ ιν α νον-ρελατιιςτις αππροξιματιον ωας δονε βψ Η.Α. Βετηε ιν τηε τραιν ωηιςη βρουγητ ηιμ βαςϰ φρομ Νεω Ψορϰ το Ιτηαςα. Ιν τηε φολλοωινγ μοντης Ρ.Π. Φεψνμαν ανδ Θ.Σ. Σςηωινγερ, ινδεπενδεντλψ, υςινγ αππαρεντλψ διφφερεντ φορμυλατιονς, ςετ υπ τηε προγραμ φορ τηε ρενορμαλιςεδ περτυρβατιον εξπανςιον οφ Χυαντυμ Ελεςτροδψναμιςς ανδ Σςηωινγερ γαε τηε φιρςτ ςαλςυλατιον οφ $g-2$. Ας ιτ τυρνεδ ουτ, ςιμιλαρ ρεςυλτς ωερε οβταινεδ ινδεπενδεντλψ ιν Θαπαν βψ Σιν-Ιτιρο Τομοναγα, ωηο οβταινεδ αλςο τηε φιρςτ ςομπλετε ςαλςυλατιον οφ τηε Λαμβ ςηιφτ. Τηε εχυιαλενςε οφ αλλ τηεςε αππροαςηες ωας φορμαλλψ ςηοων βψ Φ. Δψςον ιν 1948. Ραρε αρε τηε εξαμπλες ιν πηψςιςς ιν ωηιςη ςο μυςη προγρεςς ωας αςςομπλιςηεδ ιν ςυςη α ςηορτ τιμε.
Τηε Σηελτερ Ιςλανδ ςονφερενςε ωας μεαντ το βε τηε φιρςτ ιν α ςεριες ον τηε ςαμε ςυβϑεςτ. Ινδεεδ α φολλοω υπ ςονφερενςε ωας οργανιςεδ ιν Ποςονο (Πεννςψλανια) ιν 1948 ανδ α τηιρδ ονε ιν Ολδςτονε (Νεω Ψορϰ) ιν 1949. Αφτερ τηατ τηε προγραμμε οφ τηε Φουνδατιονς οφ Χυαντυμ Μεςηανιςς ωας δεςλαρεδ ςομπλετε. Τοδαψ ωε ϰνοω τηατ τηις ωας α ρατηερ οπτιμιςτις ιεω, βυτ ατ τηε τιμε ιτ ςεεμεδ το βε ϑυςτιφιεδ. Τηε ρεςυλτς οβταινεδ βψ αππλψινγ το χυαντυμ ελεςτροδψναμιςς τηε νεωλψ εςταβλιςηεδ πρινςιπλες οφ ρενορμαλιςεδ περτυρβατιον τηεορψ ωερε ιν ςπεςταςυλαρ αγρεεμεντ ωιτη εξπεριμεντ.
Τηε φιρςτ ηιντ τηατ εερψτηινγ ωας νοτ περφεςτ ςαμε φρομ α ςεεμινγλψ ματηεματιςαλ προβλεμ. Ωιτη τηε εξςεπτιον οφ ςομε ερψ ςιμπλε τοψ μοδελς, ωε ηαε νο εξαςτ ςολυτιονς ιν χυαντυμ φιελδ τηεοριες. Εερψ ρεςυλτ ις οβταινεδ ιν τηε φορμ οφ α φορμαλ ποωερ ςεριες ιν α παραμετερ, ϰνοων ας τηε ςουπλινγ ςονςταντ, ωηοςε αλυε ρεφλεςτς τηε ςτρενγτη οφ τηε ιντεραςτιον. Φορ ΧΕΔ ιτ ις τηε φινε ςτρυςτυρε ςονςταντ $\alpha\approx 1/137$. Τηε τηεορψ οφ ρενορμαλιςατιον γυαραντεες τηατ εερψ τερμ ιν τηις εξπανςιον ις ςαλςυλαβλε βυτ, ιν πραςτιςε, τηε αςτυαλ ςαλςυλατιονς αρε ςο ςομπλιςατεδ τηατ, εεν τοδαψ, ονλψ ερψ φεω τερμς ηαε βεεν ςομπυτεδ. Ηοωεερ, ςινςε $\alpha<<1$, τηε ηιγηερ ορδερ τερμς αρε βελιεεδ το γιε νεγλιγιβλε ςοντριβυτιονς. Νεερτηελεςς, φρομ τηε ματηεματιςαλ ποιντ οφ ιεω, τηερε ρεμαινς τηε οβιους χυεςτιον ςονςερνινγ τηε ςονεργενςε προπερτιες οφ τηε εξπανςιον. Ιν ηις φυνδαμενταλ παπερς, ωηιςη φορμυλατεδ τηε τηεορψ οφ χυαντυμ ελεςτροδψναμιςς, Δψςον ρεπεατεδλψ εμπηαςιςεδ τηατ τηε τηεορψ ωας ματηεματιςαλλψ ινςομπλετε βεςαυςε τηε ςονεργενςε προπερτιες οφ τηε ςεριες ωερε νοτ ϰνοων. Ηε αδδρεςςεδ ςπεςιφιςαλλψ τηις χυεςτιον ιν τηε ψεαρς αρουνδ 1950 ανδ, ιν τηε πρεαιλινγ ευπηορια, ηε ωας ςονινςεδ ηε ωουλδ προε τηατ τηε ςεριες ωας ςονεργεντ, ατ λεαςτ φορ τηε παρτ ωηιςη ςορρεςπονδς το ηιγη φρεχυενςιες [^6] It seems that W. Pauli had warned him against such expectation.. Ηοωεερ πρεττψ ςοον ηε φουνδ α ερψ ςιμπλε αργυμεντ ςηοωινγ τηατ τηις ωας νοτ τρυε· τηε περτυρβατιον ςεριες διεργες[8]. Τηε αργυμεντ ις βαςεδ ον τηε οβιους πηψςιςαλ ρεμαρϰ τηατ, φορ $\alpha<0$, α ςτατε ωιτη αν αρβιτραριλψ λαργε νυμβερ οφ ελεςτρον-ποςιτρον παιρς ωιλλ ηαε νεγατιε ενεργψ ωηιςη ις υνβουνδεδ φρομ βελοω. Τηις προες τηατ τηε ςεριες ιν ποωερς οφ $\alpha$ ςαννοτ βε αβςολυτελψ ςονεργεντ. Αλτηουγη τηις ρεςυλτ δοες νοτ εξςλυδε ωεαϰερ φορμς οφ ςονεργενςε, εεν τοδαψ ωε ηαε νο ριγορους προοφς φορ α πηψςιςαλ τηεορψ λιϰε ΧΕΔ. Το χυοτε Δψςον[9], “Σο μψ προγραμ φορ μαϰινγ χυαντυμ φιελδ τηεορψ ςονςιςτεντ, βψ πυρελψ ματηεματιςαλ μανιπυλατιονς ωιτηουτ ανψ νεω πηψςιςς, ςαμε το αν ιγνομινιους ενδ. Τηε ηαρδ-εαρνεδ λεςςον, …ις τηατ νεω πηψςιςς ις υναοιδαβλε.”
Γοινγ βεψονδ περτυρβατιον, Α.Σ. Ωιγητμαν φορμυλατεδ ιν 1955 α ριγορους ματηεματιςαλ φραμεωορϰ φορ α χυαντυμ φιελδ τηεορψ[10]. Ιτ ςονςιςτς οφ α ςετ οφ ςιξ αξιομς τηατ εερψ πηψςιςαλλψ αςςεπταβλε χυαντυμ φιελδ τηεορψ ις ςυπποςεδ το ςατιςφψ ατ τηε νον-περτυρβατιε λεελ. Αλτηουγη τηεψ αρε ερψ γενεραλ – Λορεντζ ιναριανςε, λοςαλιτψ, ποςιτιιτψ οφ τηε ενεργψ ετς – τηεψ ηαε τωο ιμπορταντ πηψςιςαλ ςονςεχυενςες: τηεψ ιμπλψ τηε ιναριανςε οφ ανψ φιελδ τηεορψ υνδερ τηε $CPT$ τρανςφορματιον [^7] The product of the three discrete transformations “charge conjugation” ($C$), “parity” ($P$) and “time reversal” ($T$)., ας ωελλ ας τηε ςοννεςτιον βετωεεν ςπιν ανδ ςτατιςτιςς. Βοτη αρε ιν εξςελλεντ αγρεεμεντ ωιτη οβςερατιονς, βυτ τηε προβλεμ ις τηατ ωε δο νοτ ϰνοω ωηετηερ ανψ αμονγ τηε φιελδ τηεοριες ωε υςε ιν παρτιςλε πηψςιςς αδμιτς νον-τριιαλ ςολυτιονς ςατιςφψινγ τηε Ωιγητμαν αξιομς. Τηις ωας ςερταινλψ α δραωβαςϰ φορ α ματηεματιςιαν λιϰε Δψςον, βυτ Ι δο νοτ τηινϰ τηατ τηε ηιγη ενεργψ πηψςιςς ςομμυνιτψ ωορριες τοο μυςη αβουτ ιτ. Σινςε τηε λοω ορδερ ρεςυλτς αγρεε ςο ωελλ ωιτη εξπεριμεντ, ωε ςαν λεαε τηε ςτυδψ οφ τηε ματηεματιςαλ προπερτιες φορ λατερ. Ι ωιλλ ςομε βαςϰ το τηις ρεμαρϰ ιν α λατερ ςεςτιον.
Τηε νατυραλ νεξτ ςτεπ ωας το αππλψ τηις αππροαςη το τηε οτηερ τωο ιντεραςτιονς, το ωιτ τηε ςτρονγ ανδ τηε ωεαϰ ονες. Τηε φορμερ ωερε ρεπρεςεντεδ βψ τηε ιςοςπιν ιναριαντ πιον-νυςλεον ιντεραςτιον ανδ ιτ ωας ςηοων τηατ τηε ρενορμαλιςατιον προγραμ αππλιεδ το ιτ. Ηοωεερ, τηε ρεςυλτς ωερε υςελεςς βεςαυςε τηε εφφεςτιε πιον-νυςλεον ςουπλινγ ςονςταντ τυρνεδ ουτ το βε ερψ λαργε, ον τηε ορδερ οφ 10, μαϰινγ τηε ποωερ ςεριες εξπανςιον μεανινγλεςς. Τηε δψναμιςς αππεαρς το βε δομινατεδ βψ πηενομενα ςυςη ας ρεςονανςε προδυςτιον, ωηιςη ςαννοτ βε δεςςριβεδ βψ ορδιναρψ περτυρβατιον τηεορψ. Τηερε ρεμαινεδ τηε ωεαϰ ιντεραςτιονς. Νατυραλλψ, τηεψ αρε μυςη ωεαϰερ, βυτ νοω α νεω προβλεμ αππεαρεδ: τηε ρενορμαλιςατιον προγραμ ις α χυιτε ςομπλεξ προςεςς ανδ αππλιες ονλψ το α ςμαλλ νυμβερ οφ χυαντυμ φιελδ τηεοριες. Τηε Φερμι τηεορψ, ωηιςη δεςςριβες ερψ ωελλ τηε λοω ενεργψ ωεαϰ ιντεραςτιονς, ις νοτ ονε οφ τηεμ.
Τηις δουβλε φαιλυρε ςοον ταρνιςηεδ τηε γλορψ οφ χυαντυμ φιελδ τηεορψ. Τηε διςαπποιντμεντ ωας ςυςη τηατ τηε ςυβϑεςτ ωας νοτ εεν ταυγητ ιν μανψ υνιερςιτιες. Βψ τηε λατε φιφτιες τηε τηεορετιςαλ ηιγη ενεργψ πηψςιςς λανδςςαπε ωας φραγμεντεδ ιν μανψ διςςοννεςτεδ δομαινς, ηαινγ νο ςομμον τρενδς ανδ οφτεν ιγνορινγ εαςη οτηερ. Φορ ςτρονγ ιντεραςτιον προςεςςες τηε μαιν αππροαςη ωας βαςεδ ον τηε αςςυμεδ αναλψτις προπερτιες οφ τηε $S$-ματριξ ελεμεντς [^8] This approach, which from exaggerated heights of almost religious faith has fallen into equally unjustified depths of oblivion, has produced many fundamental concepts in particle physics, such as Regge theory, bootstrap methods, dual models, the Veneziano amplitude and string theory., βυτ ωε ηαδ αλςο ςεεραλ ςιμπλε μοδελς, νονε ωιτη ανψ ςολιδ τηεορετιςαλ βαςις, εαςη ονε αππλιεδ το α παρτιςυλαρ ςορνερ οφ πηαςε ςπαςε. Φορ ωεαϰ ιντεραςτιονς τηε Φερμι τηεορψ προεδ το βε α ερψ γοοδ πηενομενολογιςαλ μοδελ, βυτ ιτ ηαδ νο λογιςαλ ϑυςτιφιςατιον ανδ νο οβιους ςοννεςτιον ωιτη ανψτηινγ ελςε. Τηε μοςτ ιμπορταντ προγρεςς ιν ουρ υνδερςτανδινγ οφ Νατυρε´ς φυνδαμενταλ λαως ςαμε μαινλψ φρομ τηε αππλιςατιον οφ ςψμμετρψ πρινςιπλες ανδ νοτ φρομ δψναμιςαλ ςαλςυλατιονς. Χυαντυμ φιελδ τηεορψ ωας νοτιςεαβλε εςςεντιαλλψ βψ ιτς αβςενςε. Α τοταλλψ μαργιναλ ςυβϑεςτ ςονφινεδ το ερψ φεω πρεςιςιον ςαλςυλατιονς ιν χυαντυμ ελεςτροδψναμιςς. Μανψ πηψςιςιςτς ηαδ ονλψ αγυε ανδ οφτεν ερρονεους ιδεας αβουτ ιτ ανδ, το α ςερταιν εξτεντ, τηις μιςυνδερςτανδινγ ηας ςυριεδ εεν τοδαψ.
[…]
Τηε τιτλε οφ τηις ςεςτιον ςουλδ ηαε βεεν: Χυαντυμ Φιελδ Τηεορψ ςτριϰες βαςϰ, βυτ ιτ ωουλδ ηαε βεεν μιςλεαδινγ. Τηε νεω τηεορψ διδ νοτ εμεργε ουτ οφ τηε βλυε. Ι ωαντ το ςηοω τηατ ιτ ις βυιλτ ον ρεςυλτς ςομινγ φρομ αλλ λινες οφ ρεςεαρςη ανδ εαςη ονε ςοντριβυτεδ ιτς φαιρ ςηαρε το ιτ. Τηε ςονςτρυςτιον οφ τηε Στανδαρδ Μοδελ, ωηιςη βεςαμε γραδυαλλψ τηε Στανδαρδ Τηεορψ οφ ελεμενταρψ παρτιςλε πηψςιςς, ις, προβαβλψ, τηε μοςτ ρεμαρϰαβλε αςηιεεμεντ οφ μοδερν τηεορετιςαλ πηψςιςς [^9] For a historical account of the Standard Model see, for example[11] and [12].. Ιτ ςταρτεδ ωιτη τηε ωεαϰ ιντεραςτιονς. Ιτ μαψ ςουνδ ςτρανγε τηατ α ρεολυτιον ιν τηε πηψςιςς οφ ελεμενταρψ παρτιςλες ωας ινιτιατεδ βψ τηε ςτυδψ οφ τηε ωεαϰεςτ αμονγ τηειρ ιντεραςτιονς (τηε εφφεςτς οφ τηε γραιτατιοναλ ιντεραςτιονς αρε νοτ μεαςυραβλε ιν ηιγη ενεργψ πηψςιςς), βυτ τηρουγη τηε ηιςτορψ οφ ουρ φιελδ, τηε ωεαϰ ιντεραςτιονς ηαε τριγγερεδ μανψ ςυςη ρεολυτιονς ανδ ωε οφτεν ηαδ τηε οςςαςιον το μεδιτατε ον τηε φυνδαμενταλ ςιγνιφιςανςε οφ “τινψ” εφφεςτς.
Ιτ ις υςυαλλψ ςαιδ τηατ προγρεςς ιν ςςιενςε οςςυρς ωηεν αν υνεξπεςτεδ εξπεριμενταλ ρεςυλτ ςοντραδιςτς τηε ςυρρεντ τηεορετιςαλ βελιεφς. Τηις φορςες ςςιεντιςτς το ςηανγε τηειρ ιδεας ανδ λεαδς το α νεω τηεορψ. Τηις ηας οφτεν βεεν τηε ςαςε ιν τηε παςτ, βυτ τηε ρεολυτιον ωε αρε γοινγ το δεςςριβε ηερε ηαδ α τηεορετιςαλ, ρατηερ αν αεςτηετις μοτιατιον. Ιτ ωας α τριυμπη οφ αβςτραςτ τηεορετιςαλ τηουγητ ωηιςη βρουγητ γεομετρψ ιντο πηψςιςς. Τηε ροαδ ηας βεεν λονγ ανδ ςιρςυιτους ανδ μανψ α τιμε ιτ γαε τηε ιμπρεςςιον οφ λεαδινγ το α δεαδ ενδ. Ιν φαςτ ιτ ωας α λονγ ςεριες οφ ιςολατεδ ανδ μοςτλψ ςονφιδεντιαλ ςοντριβυτιονς. Εςςεντιαλλψ αλλ τηε μιλεςτονες ωεντ υννοτιςεδ ωηεν τηεψ ωερε φιρςτ προποςεδ. Μανψ ιμπορταντ ιδεας ηαδ το βε ρεδιςςοερεδ αγαιν ανδ αγαιν. Τηις ςομπλιςατες τηε ταςϰ οφ τηε ναρρατορ βεςαυςε α ςηρονολογιςαλ ορδερ ωουλδ ηαε βεεν ηαρδ το φολλοω. Ι ςηοοςε ινςτεαδ το γρουπ τογετηερ αλλ ςοντριβυτιονς ιντο α φεω ςεντραλ τοπιςς: Ι ςταρτ ωιτη α ςεριες οφ ςεεμινγλψ υνςορρελατεδ διςςοεριες, μοςτ οφ ωηιςη ηαδ μοτιατιονς υνρελατεδ το τηειρ φιναλ αππλιςατιον. Τηεψ φορμ ωηατ Ι ςαλλ “τηε βαςϰγρουνδ”. Τηεν Ι μοε το τηε τωο μαιν ςομπονεντς οφ τηε Στανδαρδ Μοδελ, ναμελψ τηε ελεςτροωεαϰ τηεορψ, ωηιςη παεδ τηε ωαψ το τηε νεω ερα, φολλοωεδ βψ τηε ςτρονγ ιντεραςτιονς ανδ Χ῝Δ.
[…]
Ιν τηε ςεςρετ ροαδ τηερε ωας νο ωελλ δεφινεδ διρεςτιον. Σεεραλ μιλεςτονες διδ νοτ ςεεμ το ποιντ το α ςινγλε πατη. Τηε πιονεερς ωερε οφτεν υναωαρε οφ εαςη οτηερ´ς ωορϰ ανδ ιτ ις ονλψ νοω τηατ ωε ςαν ςεε α ςοηερεντ πιςτυρε. Ι ωιλλ μεντιον ονλψ ερψ φεω ςοντριβυτιονς, παρτιςυλαρλψ ςομε τηατ αρε νοτ ςο ωελλ-ϰνοων, βυτ α ςομπλετε ςτυδψ ςηουλδ ινςλυδε μανψ μορε [^10] The way we learn physics we often get the impression that it progresses towards a clear and well defined goal. This is almost unavoidable because we learn only those attempts which have succeeded. However, scientific research resembles more a monte-carlo algorithm. Many directions in the phase space of ideas are explored and we pursue only the most promising ones. Occasionally we find later that they lead to a dead end and we have to come back and start a new direction. Text books never present this real story, thus creating the illusion of a unique line of thought..
[…]
$\bullet$ ῝λαςςιςαλ ελεςτροδψναμιςς. Τηε φιρςτ ςονςεπτ οφ γαυγε ιναριανςε γοες βαςϰ το ςλαςςιςαλ ελεςτροδψναμιςς[13] βυτ ιτς ηιςτορψ ανδ εολυτιον ις χυιτε ςομπλιςατεδ. Ι αμ νοτ αβλε το πρεςεντ α φυλλ ςτορψ ανδ Ι ωιλλ μεντιον ονλψ α φεω ςτεπς ωηιςη, το μψ υνδερςτανδινγ, ηαε πλαψεδ αν ιμπορταντ ρολε ιν ςηαπινγ ουρ πρεςεντ ιδεας ον γαυγε ιναριανςε. Τηεψ αρε νοτ νεςεςςαριλψ αμονγ τηε μοςτ ιμπορταντ ςοντριβυτιονς το ςλαςςιςαλ ελεςτροδψναμιςς.
Βψ τηε ενδ οφ τηε 18τη ςεντυρψ τηε φιελδ οφ ελεςτροςτατιςς ωας χυιτε ωελλ υνδερςτοοδ. Ιν Φρανςε τηε μαιν αςτορς ωερε ῝. ῝ουλομβ, ανδ ηις φαμους εξπεριμεντς τηατ εςταβλιςηεδ τηε $1/r^{2}$ λαω, ανδ Σ. Δ. Ποιςςον ωηο ωροτε τηε διφφερεντιαλ εχυατιον φορ τηε ελεςτροςτατις ποτεντιαλ $\Delta V+4\pi\rho=0$, ιν αναλογψ ωιτη Νεωτονιαν δψναμιςς. Σιμιλαρ ρεςυλτς, αλτηουγη ςταρτινγ φρομ αππαρεντλψ διφφερεντ αςςυμπτιονς, ωερε οβταινεδ ιν Ενγλανδ βψ Β. Φρανϰλιν. Τηε χυεςτιον οφ αςτιον ατ α διςτανςε πυζζλεδ πεοπλε φορ μανψ ψεαρς, ας ιτ ηαδ πυζζλεδ Νεωτον [^11] In a letter addressed to Richard Bentley in 1687, Newton writes “That one body may act upon another at a distance through a vacuum, without the mediation of something else, by and through which their action and force may be conveyed from one to another, is to me so great an absurdity, that I believe no man, who has in philosophical matters a competent faculty of thinking, can ever fall into it…”. μορε τηαν ονε ςεντυρψ εαρλιερ. Ον τηε οτηερ ηανδ μαγνετις πηενομενα ωερε βελιεεδ το ηαε α διφφερεντ οριγιν ανδ ωερε νοτ ςονςιδερεδ ας εχυαλλψ ιμπορταντ. Τηις βελιεφ ωας ςηυττερεδ δυρινγ τηε εαρλψ ψεαρς οφ τηε 19τη ςεντυρψ φολλοωινγ αριους εξπεριμεντς περφορμεδ νοτ ονλψ ωιτη ςτατις ςηαργες, βυτ αλςο ωιτη ελεςτρις ςυρρεντς ανδ μαγνετς. Οφ παρτιςυλαρ ιντερεςτ φορ τηε ςονςεπτ οφ γαυγε ιναριανςε τυρνεδ ουτ το βε τηε εξπεριμεντς οφ Η.῝. Οερςτεδ ανδ Α.-Μ. Αμπèρε, ιν παρτιςυλαρ τηοςε μεαςυρινγ τηε ιντεραςτιονς αμονγ ςλοςεδ ελεςτρις ςιρςυιτς ςαρρψινγ ςτεαδψ ςυρρεντς.
Ι δο νοτ ϰνοω ωηο ωας τηε φιρςτ το ρεμαρϰ τηατ τηε δψναμιςαλ ςψςτεμ δεςςριβεδ βψ τηε ςομπονεντς οφ τηε ελεςτρις ανδ μαγνετις φιελδς ${\bm{E}}$ ανδ ${\bm{B}}$ – τηατ ις, ςιξ δεγρεες οφ φρεεδομ ιν ουρ ςουντινγ – ωας ιν φαςτ ρεδυνδαντ βεςαυςε ςομε οφ τηε εχυατιονς δο νοτ ινολε ανψ τιμε δεριατιες ανδ ςηουλδ βε ςονςιδερεδ ας ςονςτραιντς. Ιτ ςεεμς τηατ τηε φιρςτ περςον ωηο αττεμπτεδ το ρεδυςε τηε ρεδυνδανςψ ωας ῝.Φ. Γαυςς ωηο, ιν ςομε μανυςςριπτ νοτες ιν 1835, ιντροδυςεδ τηε ςονςεπτ οφ τηε “εςτορ ποτεντιαλ” ${\bm{A}}$. Ιτ ωας φυρτηερ δεελοπεδ βψ ςεεραλ αυτηορς ανδ ωας φυλλψ ωριττεν βψ Γ. Κιρςηοφφ ιν 1857, φολλοωινγ εαρλιερ ωορϰ, ιν παρτιςυλαρ βψ Φ. Νευμανν. Τηε ςομπονεντς οφ τηε ελεςτρις ανδ μαγνετις φιελδς ςουλδ βε εξπρεςςεδ ιν τερμς οφ τηε εςτορ ανδ ςςαλαρ ποτεντιαλς, τηυς ρεδυςινγ τηε νυμβερ οφ δεγρεες οφ φρεεδομ φρομ ςιξ το φουρ. Ιτ ωας ςοον νοτιςεδ τηατ ιτ ςτιλλ ςαρριεδ ρεδυνδαντ αριαβλες ανδ ςεεραλ ``γαυγε ςονδιτιονς´´ ωερε υςεδ. Τηε ςονδιτιον, ωηιςη ιν μοδερν νοτατιον ις ωριττεν ας $\partial_{\mu}A^{\mu}=0$, ωας προποςεδ βψ τηε Δανιςη ματηεματιςαλ πηψςιςιςτ Λ.῞. Λορενζ ιν 1867. Τηε φυννψ ρεμαρϰ Ι οφτεν μαϰε ις τηατ μοςτ πηψςιςς βοοϰς μιςςπελλ Λορενζ´ς ναμε ας Λορεντζ, αδδινγ α “τ”, τηυς ερρονεουςλψ αττριβυτινγ τηε ςονδιτιον το τηε φαμους Δυτςη Η.Α. Λορεντζ, οφ τηε Λορεντζ τρανςφορματιονς [^12] In French: On ne prête qu’aux riches.. Ιτ ςεεμς τηατ Μαξωελλ φαορεδ τηε ςονδιτιον ${\bm{\nabla}}\cdot{\bm{A}}(x)\!=\!0$ ωηιςη τοδαψ ωε ςαλλ “τηε ῝ουλομβ γαυγε” [^13] It is obvious that Coulomb never wrote such a condition. The name comes from the fact that, in this gauge, the scalar potential satisfies the Poisson equation $\Delta V(x)+\rho(x)=0$ which, for $\rho$ given by a static point source, reproduces the Coulomb potential.. ϒςινγ Μαξωελλ´ς εχυατιονς ωε ςαν ιμμεδιατελψ ςεε τηε ρεδυνδανςψ οφ τηε ςψςτεμ $(\Phi,{\bm{A}})$ βεςαυςε τηε εχυατιον φορ $\Phi$ δοες νοτ ινολε ανψ τιμε δεριατιε. Λορενζ αρριεδ το τηις ςονςλυςιον βεςαυςε ηε ηαδ α φορμυλατιον οφ ςλαςςιςαλ ελεςτροδψναμιςς εχυιαλεντ το Μαξωελλ´ς.
Ιν τηις ςοντεξτ, αν ιντερεςτινγ ςτορψ ις τηε φολλοωινγ[13]: Αρουνδ τηε ψεαρς 1840 Φ.Ε. Νευμανν ανδ, ινδεπενδεντλψ, Ω.Ε. Ωεβερ, ςτυδιεδ τηε ιντεραςτιον βετωεεν τωο ςλοςεδ ελεςτρις ςιρςυιτς ςαρρψινγ ςυρρεντς $I$ ανδ $I^{\prime}$, ρεςπεςτιελψ. Τηειρ μετηοδς ανδ πηψςιςαλ αςςυμπτιονς ωερε διφφερεντ ανδ τηεψ πρεςεντ ονλψ α ηιςτοριςαλ ιντερεςτ τοδαψ, ςο Ι γιε ονλψ τηειρ φιναλ εξπρεςςιονς φορ τηε μαγνετις ιντεραςτιον ενεργψ $W$ βετωεεν τηε τωο ςιρςυιτς, υςινγ τηε ςαμε μοδερν νοτατιον φορ βοτη.
${\rm d}W_{N}=\frac{II^{\prime}}{c^{2}}\frac{{\bm{n}}\cdot{\bm{n}^{\prime}}}{r}{\rm d}s{\rm d}s^{\prime}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\rm d}W_{W}=\frac{II^{\prime}}{c^{2}}\frac{({\bm{n}}\cdot{\hat{{\bm{r}}}})({\bm{n}^{\prime}}\cdot{\hat{{\bm{r}}}})}{r}{\rm d}s{\rm d}s^{\prime}$ (1)
ωηερε τηε ςυβςςριπτς $N$ ανδ $W$ ςτανδ φορ Νευμανν ανδ Ωεβερ, ρεςπεςτιελψ. Τηε νοτατιον ις τηε φολλοωινγ: ${\bm{x}}$ ανδ ${\bm{x}^{\prime}}$ δενοτε τωο ποιντς ον τηε ςιρςυιτς $C$ ανδ $C^{\prime}$, ${\bm{r}}={\bm{x}}-{\bm{x}^{\prime}}$, ${\hat{{\bm{r}}}}$ ις τηε υνιτ εςτορ ιν τηε διρεςτιον οφ ${\bm{r}}$, ανδ τηε λινε ελεμεντς ${\rm d}s$ ανδ ${\rm d}s^{\prime}$ αρε παραμετριςεδ ας ${\rm d}{\bm{s}}={\bm{n}}{\rm d}s$ ανδ ${\rm d}{\bm{s}^{\prime}}={\bm{n}^{\prime}}{\rm d}s^{\prime}$. Τηε τοταλ ιντεραςτιον ενεργψ $W$ ςαν βε οβταινεδ βψ ιντεγρατινγ τηε διφφερεντιαλ εξπρεςςιονς (1) αλονγ τηε τωο ςιρςυιτς. Νειτηερ Νευμανν νορ Ωεβερ ωροτε εξπλιςιτλψ τηε ςορρεςπονδινγ εςτορ ποτεντιαλς βυτ, ηαδ τηεψ δονε ςο, τηεψ ωουλδ ηαε αρριεδ ατ εξπρεςςιονς οφ τηε φορμ:
${\bm{A}}_{N}({\bm{x}},t)=\frac{1}{c}\int{\rm d}^{3}x^{\prime}\frac{1}{r}{\bm{J}}({\bm{x}^{\prime}},t)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\bm{A}}_{W}({\bm{x}},t)=\frac{1}{c}\int{\rm d}^{3}x^{\prime}\frac{1}{r}{\hat{{\bm{r}}}}({\hat{{\bm{r}}}}\cdot{\bm{J}}({\bm{x}^{\prime}},t))$ (2)
ωιτη ${\bm{J}}({\bm{x}^{\prime}},t)$ α γενεραλ ςυρρεντ δενςιτψ. Τηε ιντερεςτινγ παρτ οφ τηε ςτορψ ις τηατ, ιν τηε ψεαρς αφτερ 1870, Η.Λ.Φ. ον Ηελμηολτζ ςριτιςιςεδ ανδ ςομπαρεδ τηεςε εξπρεςςιονς. Ιν παρτιςυλαρ, ηε νοτιςεδ τηατ τηε τωο φορμυλαε φορ τηε ελεμενταρψ μαγνετις ιντεραςτιον ενεργψ διφφερ βψ α χυαντιτψ ωηιςη ςαν βε εξπρεςςεδ ας α μυλτιπλε οφ τηε περφεςτ διφφερεντιαλ
${\rm d}s{\rm d}s^{\prime}\frac{\partial^{2}r}{\partial s\partial s^{\prime}}={\rm d}s{\rm d}s^{\prime}\frac{({\bm{n}}\cdot{\hat{{\bm{r}}}})({\bm{n}^{\prime}}\cdot{\hat{{\bm{r}}}})-({\bm{n}}\cdot{\bm{n}^{\prime}})}{r}$ (3)
ςο, τηεψ γιε τηε ςαμε ρεςυλτ ωηεν ιντεγρατεδ οερ τηε ςλοςεδ ςιρςυιτς. Ιν ουρ πρεςεντ τερμινολογψ ηε ςηοωεδ τηατ τηε τωο εξπρεςςιονς αρε γαυγε εχυιαλεντ. Ηε εεν ωεντ α ςτεπ φυρτηερ: ηε γενεραλιςεδ Νευμανν´ς ανδ Ωεβερ´ς ρεςυλτς βψ εξηιβιτινγ α ονε-παραμετερ φαμιλψ οφ εξπρεςςιονς φορ τηε μαγνετις ενεργψ ωηιςη ιντερπολατε βετωεεν Νευμανν ανδ Ωεβερ:
${\rm d}W_{\alpha}=\frac{II^{\prime}}{2c^{2}r}[(1+\alpha)({\bm{n}}\cdot{\bm{n}^{\prime}})+(1-\alpha)({\bm{n}}\cdot{\hat{{\bm{r}}}})({\bm{n}^{\prime}}\cdot{\hat{{\bm{r}}}})]{\rm d}s{\rm d}s^{\prime}$ (4)
Τηις λινεαρ ςομβινατιον διφφερς φρομ ειτηερ ονε οφ τηε εξπρεςςιονς ιν εχ. (1) βψ α μυλτιπλε οφ τηε περφεςτ διφφερεντιαλ οφ εχ. (3) ανδ, τηερεφορε, ιτ ις ςονςιςτεντ ωιτη Αμπèρε’ς οβςερατιονς φορ ανψ αλυε οφ τηε παραμετερ $\alpha$. Τηε αναλογυε φορ τηε εςτορ ποτεντιαλς οφ εχ. (2), ις τηε εξπρεςςιον
${\bm{A}}_{\alpha}=\frac{1}{2}(1+\alpha){\bm{A}}_{N}+\frac{1}{2}(1-\alpha){\bm{A}}_{W}$ (5)
Ηελμηολτζ μαδε τηε γαυγε εχυιαλενςε ςλεαρ βψ ωριτινγ
${\bm{A}}_{\alpha}={\bm{A}}_{N}+\frac{1-\alpha}{2}{\bm{\nabla}}\Psi\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \Psi=-\frac{1}{c}\int{\hat{{\bm{r}}}}\cdot{\bm{J}}({\bm{x}^{\prime}},t){\rm d}^{3}x^{\prime}$ (6)
Ηε τηεν ωεντ ον το ςηοω τηατ, ιφ $\Phi$ δενοτες τηε ελεςτροςτατις ποτεντιαλ, $\Psi$, ${\bm{A}}_{\alpha}$ ανδ $\Phi$ ςατιςφψ τηε εχυατιονς
$\Delta\Psi=\frac{2}{c}\frac{\partial\Phi}{\partial t}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\bm{\nabla}}\cdot{\bm{A}}_{\alpha}=-\frac{\alpha}{c}\frac{\partial\Phi}{\partial t}$ (7)
Νοτιςε τηατ τηε ςεςονδ εχυατιον ις Λορενζ´ς γαυγε ςονδιτιον φορ τηε ποτεντιαλς υςεδ βψ Ηελμηολτζ. Ηε εεν ρεμαρϰεδ τηατ τηε ςηοιςε $\alpha=0$ γιες τηε ῝ουλομβ γαυγε υςεδ βψ Μαξωελλ. Ιτ ωας τηε φιρςτ εξαμπλε οφ α φαμιλψ οφ γαυγες.
Βψ τηε ενδ οφ τηε ςεντυρψ Η.Α. Λορεντζ πυβλιςηεδ α βοοϰ ανδ ςομε ενςψςλοπεδια αρτιςλες ωιτη τηε φυλλ ςλαςςιςαλ ελεςτρομαγνετις τηεορψ. Ιτ τοοϰ μορε τηαν α ςεντυρψ το οβταιν ιτ ανδ τηε ιναριανςε υνδερ γαυγε τρανςφορματιονς οφ τηε εςτορ ανδ ςςαλαρ ποτεντιαλς ις αν ιντεγραλ παρτ οφ ιτ.
$\bullet$ Γαυγε ιναριανςε ιν Γενεραλ Ρελατιιτψ ανδ τηε αττεμπτς το υνιφψ ελεςτρομαγνετιςμ ανδ γραιτατιον. Τηε δεελοπμεντ οφ τηε γενεραλ τηεορψ οφ ρελατιιτψ οφφερεδ α νεω παραδιγμ φορ α γαυγε τηεορψ. Ι βελιεε τηατ ιτς ματηεματιςαλ φορμυλατιον ας α τηεορψ ιναριαντ υνδερ λοςαλ τρανςλατιονς (μορε πρεςιςελψ διφφεομορπηιςμς) ωας φιρςτ υνδερςτοοδ βψ Δ. Ηιλβερτ[14]. Βυτ απαρτ φρομ τηις ποιντ, τηε γαυγε τηεοριες ωε υςε ιν τηε Στανδαρδ Μοδελ αρε λογιςαλλψ υνρελατεδ το γενεραλ ρελατιιτψ ανδ γραιτατιον βεςαυςε τηεψ δεαλ ωιτη ιντερναλ ςψμμετριες ανδ νοτ ωιτη τηοςε οφ ςπαςε ανδ τιμε. Τηερεφορε Ι ωιλλ νοτ πρεςεντ τηεςε δεελοπμεντς ιν ανψ δεταιλ, αλτηουγη φορ μανψ δεςαδες τηε τωο ωερε εντανγλεδ ανδ γενεραλ ρελατιιτψ ωας τηε ςταρτινγ ποιντ φορ μανψ ςτυδιες οφ τηεοριες ιναριαντ υνδερ λοςαλ τρανςφορματιονς.
Σινςε τηε τωο ςλαςςιςαλλψ ϰνοων ιντεραςτιονς, ελεςτρομαγνετιςμ ανδ γραιτατιον, ωερε φουνδ το οβεψ α γαυγε πρινςιπλε, ιτ ωας νορμαλ το ςεαρςη φορ α υνιφιεδ δεςςριπτιον. Τηε αττεμπτ ωηιςη ηας ςυριεδ τοδαψ ςαρριες τηε ναμες οφ Τ. Καλυζα ανδ Ο. Β. Κλειν[15]. Τηειρ αππροαςη ςονςιςτς ιν ωριτινγ γενεραλ ρελατιιτψ ιν α φιε διμενςιοναλ ςπαςε-τιμε. Τηεψ ςηοω τηατ, ατ λεαςτ ιν ςομε ςιμπλε ςαςες, τηε ςολυτιον φορ τηε μετρις τενςορ ςπλιτς ιντο α φουρ διμενςιοναλ ςπαςε, φορ εξαμπλε Μινϰοωςϰι, ανδ α ςομπαςτ φιφτη διμενςιον ρεπρεςεντινγ α ςιρςλε οφ ραδιυς $R$. Ατ τηε λιμιτ ωηεν $R$ γοες το ζερο, α φουρ-διμενςιοναλ οβςερερ ωιλλ ςεε τηε ςομπονεντς οφ τηε μετρις αλονγ τηε φιφτη διμενςιον ας α φουρ-διμενςιοναλ εςτορ ενδοωεδ ωιτη α γαυγε πρινςιπλε, τηυς μιμιςϰινγ α τηεορψ οφ γραιτατιον ανδ ελεςτρομαγνετιςμ. Ιτ ις τηε αππροαςη ωηιςη ις μοςτλψ υςεδ τοδαψ ιν ςυπεργραιτψ ανδ ςυπερςτρινγ τηεοριες. Ωηατ ις λεςς ϰνοων ις τηατ τηε ιδεα ωας ιντροδυςεδ εαρλιερ βψ τηε Φιννιςη Γυνναρ Νορδςτρο̈μ[16] ωηο ηαδ ςονςτρυςτεδ α ςςαλαρ τηεορψ οφ γραιτατιον. Ιν 1914 ηε ωροτε α φιε-διμενςιοναλ τηεορψ οφ ελεςτρομαγνετιςμ ανδ ςηοωεδ τηατ, ιφ ονε αςςυμες τηατ τηε φιελδς αρε ινδεπενδεντ οφ τηε φιφτη ςοορδινατε, τηε αςςυμπτιον μαδε λατερ βψ Καλυζα, τηε ελεςτρομαγνετις εςτορ ποτεντιαλ ςπλιτς ιντο α φουρ διμενςιοναλ ονε ανδ α ςςαλαρ φιελδ ιδεντιφιεδ το ηις ςςαλαρ γραιτον. Ιν ςομε ςενςε, Καλυζα ανδ Κλειν ωροτε τηε μιρρορ ιμαγε οφ Νορδςτρο̈μ´ς τηεορψ.
Τηε φιρςτ αττεμπτ το αν ε.μ. – γραιτατιον υνιφιςατιον ιν τηε φραμεωορϰ οφ γενεραλ ρελατιιτψ ις δυε το Η.Κ.Η. Ωεψλ ιν 1919[17]. Ηε προποςεδ το ενλαργε τηε ιναριανςε οφ γενεραλ ρελατιιτψ βψ ινςλυδινγ λοςαλ ςςαλε τρανςφορματιονς ιν ωηιςη τηε μετρις τρανςφορμς ας $g\rightarrow{\rm e}^{2\lambda(x)}g$, ωιτη $\lambda(x)$ αν αρβιτραρψ φυνςτιον οφ τηε ςπαςε-τιμε ποιντ $x$. Ωεψλ ωαντεδ το αςςοςιατε τηις νεω γαυγε τρανςφορματιον το ελεςτρομαγνετιςμ. Γεομετριςαλλψ, ιτ ςορρεςπονδς το α ςηανγε οφ τηε υνιτ οφ λενγτη φρομ ποιντ το ποιντ, τηερεφορε, ιτ ωας νατυραλ το ςαλλ τηε ρεςυλτινγ ιναριανςε ειςηιναριανζ ιν Γερμαν, ωηιςη ωας τρανςλατεδ ιν Ενγλιςη ας γαυγε ιναριανςε. Βεφορε λοοϰινγ ατ τηε πηψςιςς οφ τηις τηεορψ, λετ με ςτρεςς τηατ, φρομ τηε ματηεματιςαλ ποιντ οφ ιεω, ιτ ις ερψ ιντερεςτινγ. Ωεψλ ωας αν αςςομπλιςηεδ ματηεματιςιαν ανδ ιν τηις ωορϰ ηε δεελοπς ανδ ιντροδυςες μανψ ςονςεπτς φρομ διφφερεντιαλ γεομετρψ ωηιςη τοοϰ δεςαδες το ρεαςη τηε τηεορετιςαλ πηψςιςς ςομμυνιτψ. Ηε ωας προβαβλψ τηε φιρςτ περςον το υνδερςτανδ τηε υνδερλψινγ γεομετρψ οφ γαυγε ιναριανςε. Ιν τηις αρτιςλε ηε δεφενδς ηις ποιντ οφ ιεω βψ νοτινγ “…Ριεμαννιαν γεομετρψ … [ςονταινς] α λαςτ ελεμεντ οφ γεομετρψ ‘‘ατ α διςτανςε’’ (φερνγεομετριςςηες Ελεμεντ)—ωιτη νο γοοδ ρεαςον …Τηε μετρις αλλοως τηε τωο μαγνιτυδες οφ τωο εςτορς το βε ςομπαρεδ, νοτ ονλψ ατ τηε ςαμε ποιντ, βυτ ατ ανψ αρβιτραριλψ ςεπαρατεδ ποιντς…” Φορ Ωεψλ α τρυε ινφινιτεςιμαλ γεομετρψ ςηουλδ νοτ αλλοω το ςομπαρε τηε λενγτης οφ τωο εςτορς ατ α διςτανςε, τηε ςαμε ωαψ ιτ δοες νοτ αλλοω το ςομπαρε τηειρ διρεςτιονς. ῝ονςερνινγ τηε πηψςιςς, ηε ωας ερψ εντηυςιαςτις. Ιν α λεττερ το Εινςτειν ηε ωροτε: “…Ι ςυςςεεδεδ, ας Ι βελιεε, το δεριε ελεςτριςιτψ ανδ γραιτατιον φρομ α ςομμον ςουρςε…”. Τηε ρεαςτιον οφ τηε λεαδινγ πηψςιςιςτς οφ τηε τιμε ωας μορε ρεςερεδ. Αμονγ τηεμ Εινςτειν, ωηο λιϰεδ τηε ματηεματιςαλ παρτ οφ τηε τηεορψ, ςποττεδ ιμμεδιατελψ α πηψςιςαλ ινςονςιςτενςψ. Ιν α ποςτςαρδ ηε ςεντ το Ωεψλ ηε ποιντεδ ουτ τηατ, ιφ ωε αςςεπτ α νονιντεγραβλε λενγτη ςοννεςτιον, τηεν τηε βεηαιορ οφ ςλοςϰς ωουλδ δεπενδ ον τηειρ ηιςτορψ, ιν ςλεαρ ςοντραδιςτιον ωιτη τηε εξιςτενςε οφ ςταβλε ατομις ςπεςτρα. Παυλι, ιν ηις υςυαλ ερψ αγγρεςςιε ςτψλε, ωροτε το Ωεψλ “…ψου ηαε βεεν γιεν α ςηαιρ ιν ‘‘Πηψςιςς’’ ιν Αμεριςα. Ι αδμιρε ψουρ ςουραγε· ςινςε τηε ςονςλυςιον ις ινειταβλε τηατ ψου ωιςη το βε ϑυδγεδ, νοτ φορ ςυςςεςς ιν πυρε ματηεματιςς, βυτ φορ ψουρ τρυε βυτ υνηαππψ λοε φορ πηψςιςς.” Ιν ρετροςπεςτ, Ι τηινϰ τηατ, ιν ηις 1919 παπερ, Ωεψλ αςϰεδ τηε ριγητ ματηεματιςαλ χυεςτιον βυτ γαε τηε ωρονγ πηψςιςαλ ανςωερ [^14] I realise that I am making the mistake a real historian should not make: to interpret old papers in the light of today’s knowledge. But, first I am not a real historian and second, every time I make this mistake I will try to make it clear.. Ταϰε τηε εξαμπλε οφ γενεραλ ρελατιιτψ: ωε ςταρτ φρομ α τηεορψ ιναριαντ υνδερ α γρουπ οφ γλοβαλ τρανςφορματιονς, ιν τηις ςαςε τηε Ποινςαρé τρανςφορματιονς οφ ςπεςιαλ ρελατιιτψ. Προμοτινγ τηε ιναριανςε φρομ γλοβαλ το λοςαλ ιντροδυςες νεω γεομετριςαλ χυαντιτιες – ςαλλεδ “ςοννεςτιονς” – ωηιςη αρε ιντερπρετεδ πηψςιςαλλψ ας γιινγ ριςε το τηε φορςες οφ γραιτατιον. Ωεψλ υνδερςτοοδ τηατ ιν τηις ρεςπεςτ, τηε ςλαςςιςαλ ελεςτρομαγνετις τηεορψ ωας ματηεματιςαλλψ ινςομπλετε. Πεοπλε ηαδ διςςοερεδ βψ τριαλ ανδ ερρορ ιτς προπερτψ οφ γαυγε ιναριανςε, βυτ τηε υνδερλψινγ γλοβαλ ςψμμετρψ ωας μιςςινγ. Τηις ωας τηε ριγητ ματηεματιςαλ χυεςτιον. Τηε ωρονγ πηψςιςαλ ανςωερ ωας το ιδεντιφψ ιτ ωιτη ςςαλε τρανςφορματιονς. Τηε λαως οφ πηψςιςς αρε ιν νο ωαψ ιναριαντ υνδερ γλοβαλ ςςαλε τρανςφορματιονς. Ας ιτ τυρνεδ ουτ, τηε ςορρεςτ ανςωερ ωας φουνδ ιν τηε φραμεωορϰ οφ χυαντυμ μεςηανιςς το ωηιςη ωε τυρν νεξτ.
[…]
${\bm{A}}={\bm{A}}_{1}+{\bm{\nabla}}f\leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \Phi=\Phi_{1}-\frac{\partial f}{\partial t}\leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \theta=\theta_{1}-ef\leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \Psi=\Psi_{1}{\rm e}^{-{\rm i}ef}$ (8)
ωηερε ${\bm{A}}$ ανδ $\Phi$ αρε τηε εςτορ ανδ ςςαλαρ ποτεντιαλς, $e$ ις τηε ελεςτρις ςηαργε ανδ $f=f({\bm{x}},t)$ αν αρβιτραρψ φυνςτιον οφ ςπαςε ανδ τιμε. Ιν α ςεςονδ ςεςτιον οφ ηις αρτιςλε Φοςϰ εξτενδς τηε αναλψςις το ινςλυδε γενεραλ ρελατιιτψ. Ι δο νοτ ϰνοω ωηετηερ ηε ωας αωαρε οφ Ωεψλ´ς 1919 αρτιςλε, βυτ ηε δοες νοτ ρεφερ το ιτ [^15] He refers instead to a preliminary version of a later paper, presumably ref[19], which Weyl had communicated to him in manuscript form..
Τηε ςαμε ςονςλυςιον ωας ρεαςηεδ αλςο βψ Φριτζ Λονδον ιν 1927[20]. Ηε ςταρτεδ φρομ Ωεψλ´ς ςςαλε τρανςφορματιον $g\rightarrow{\rm e}^{2\lambda(x)}g$ ανδ νοτεδ α φορμαλ αναλογψ ωιτη τηε πηαςε τρανςφορματιον ιν χυαντυμ μεςηανιςς προιδεδ $\lambda$ ταϰες ςομπλεξ αλυες. Τηε τιτλες οφ ηις αρτιςλες αρε ρεεαλινγ: Διε Τηεοριε ον Ωεψλ υνδ διε Χυαντενμεςηανιϰ ανδ Χυαντενμεςηανιςςηε Δευτυνγ δερ Τηεοριε ον Ωεψλ. Ηε ςιτες βοτη Ωεψλ ανδ Φοςϰ. Αππαρεντλψ Ωεψλ ςηαρεδ Λονδον´ς ποιντ οφ ιεω. Ιν τηε 1928 εδιτιον οφ ηις βοοϰ ον γρουπ τηεορψ ανδ χυαντυμ μεςηανιςς[17] ηε ωριτες: “…Ι νοω βελιεε τηατ τηις γαυγε ιναριανςε δοες νοτ τιε τογετηερ ελεςτριςιτψ ανδ γραιτατιον, βυτ ρατηερ ελεςτριςιτψ ανδ ματτερ…”. Τηε φολλοωινγ ψεαρ ηε πυβλιςηεδ α ρεμαρϰαβλε αρτιςλε[19] ιν ωηιςη ηε εξτενδς τηε πρινςιπλε οφ γαυγε ιναριανςε το τηε Διρας ελεςτρον, ιν βοτη τηε ςπεςιαλ ανδ τηε γενεραλ τηεορψ οφ ρελατιιτψ. Ιτ ις ονε οφ τηε ςλαςςις παπερς οφ τηεορετιςαλ πηψςιςς ιντροδυςινγ μανψ νεω ςονςεπτς, ςυςη ας τηε Ωεψλ τωο-ςομπονεντ ςπινορς ανδ τηε ιερβειν ανδ ςπιν-ςοννεςτιον φορμαλιςμ [^16] It seems that after the publication of this article, Pauli withdrew his sarcastic comments regarding Weyl’s contributions to physics.. Αλτηουγη τηε τηεορψ ις νο μορε ςςαλε ιναριαντ, ηε ςτιλλ υςεδ τηε τερμ ειςηιναριανζ –γαυγε ιναριανςε – α τερμ ωηιςη ηας ςυριεδ εερ ςινςε.
$\bullet$ Νον-αβελιαν ιντερναλ ςψμμετριες. Τηε ςονςεπτ οφ νον-αβελιαν ιντερναλ ςψμμετριες ιν παρτιςλε πηψςιςς ις αττριβυτεδ το Ω. Ηειςενβεργ ωιτη ηις ιντροδυςτιον οφ τηε ιςοςπιν φορμαλιςμ ιν 1932[21], βυτ τηε ρεαλ ςτορψ ις μορε ςομπλιςατεδ. Ηειςενβεργ´ς 1932 παπερς αρε αν ινςρεδιβλε μιξτυρε οφ τηε ολδ ανδ τηε νεω. Τηε νευτρον ηαδ ϑυςτ βεεν διςςοερεδ, βυτ φορ μανψ πεοπλε ιτ ωας α νεω βουνδ ςτατε οφ α προτον ανδ αν ελεςτρον, λιϰε α “ςμαλλ᾽ ηψδρογεν ατομ. Ηειςενβεργ δοες νοτ ρεϑεςτ τηις ιδεα. Αλτηουγη φορ ηις ωορϰ ηε ςονςιδερς τηε νευτρον ας α ςπιν ονε-ηαλφ Διρας φερμιον, ςομετηινγ ινςομπατιβλε ωιτη α προτον-ελεςτρον βουνδ ςτατε, ηε νοτες τηατ “…υνδερ ςυιταβλε ςιρςυμςτανςες τηε νευτρον ωιλλ βρεαϰ υπ ιντο α προτον ανδ αν ελεςτρον ιν ωηιςη ςαςε τηε ςονςερατιον λαως οφ ενεργψ ανδ μομεντυμ προβαβλψ δο νοτ αππλψ.᾽ Ον τηε $\beta$-δεςαψ ςοντροερςψ ανδ τηε εξιςτενςε οφ τηε νευτρινο ηε δοες νοτ ταϰε ανψ ςλεαρ ςτανδ, βυτ ηε ςιδες μορε ωιτη ηις μαςτερ Βοηρ τηαν ωιτη ηις φριενδ Παυλι: “…Τηε αδμιττεδλψ ηψποτηετιςαλ αλιδιτψ οφ Φερμι ςτατιςτιςς φορ νευτρονς ας ωελλ ας τηε φαιλυρε οφ τηε ενεργψ λαω ιν $\beta$-δεςαψ προες τηε ιναππλιςαβιλιτψ οφ πρεςεντ Χυαντυμ Μεςηανιςς το τηε ςτρυςτυρε οφ τηε νευτρον.᾽ Ιν φαςτ, Ηειςενβεργ´ς φυνδαμενταλ ςοντριβυτιον ςηουλδ βε αππρεςιατεδ νοτ δεςπιτε τηεςε ςηορτςομινγς, βυτ πρεςιςελψ βεςαυςε οφ τηεμ. Ωε ςηουλδ ρεμεμβερ τηατ ιν 1932 εξπεριμενταλ δατα ον νυςλεαρ φορςες ωερε αλμοςτ εντιρελψ αβςεντ. Ηειςενβεργ ηαδ το γυεςς τηε αλυες οφ τηε νυςλεαρ αττραςτιε φορςες βετωεεν νυςλεον παιρς βψ υςινγ α ςτρανγε αναλογψ ωιτη μολεςυλαρ φορςες. Ηε ποςτυλατεδ α $p-n$ ανδ αν $n-n$ νυςλεαρ φορςε, βυτ νοτ α $p-p$ ονε, ςο ηις τηεορψ ωας νοτ ρεαλλψ ιςοςπιν ιναριαντ. Νεερτηελεςς, ηε μαδε τηε ςονςεπτυαλ ςτεπ το δεςςριβε τηε νυςλεον ωαε φυνςτιονς ιν τερμς οφ α νεω τωο-ςομπονεντ οβϑεςτ
$\Psi(x)=\begin{pmatrix}\Psi_{p}(x)\cr\Psi_{n}(x)\cr\end{pmatrix}$ (9)
ανδ τηε νυςλεαρ ποτεντιαλ ιν τηε φορμ οφ τωο-βψ-τωο Παυλι ματριςες. Τηε ςομπονεντς ιν εχυατιον (9) δενοτε τηε ωαε φυνςτιονς οφ α προτον ανδ α νευτρον. $\Psi(x)$ ηας τηε φορμ οφ α τωο-ςομπονεντ ςπινορ, βυτ, ηερε ις τηε νεω ελεμεντ, ιτ ις ΝΟΤ α ςπινορ ιν ουρ τηρεε διμενςιοναλ ςπαςε. Ωε ςηαλλ αςςυμε τηατ τηερε εξιςτς α ςεςονδ ςυςη ςπαςε, ιςομορπηις το, βυτ διςτινςτ φρομ τηε ονε ιν ωηιςη ωε λιε. Α ροτατιον ιν τηις νεω ςπαςε μιξες $\Psi_{p}$ ανδ $\Psi_{n}$ τηε ςαμε ωαψ τηατ α ροτατιον ιν ουρ ςπαςε μιξες τηε υπ ανδ δοων ςομπονεντς οφ α ςπινορ. Ιτ ωας τηε φιρςτ ιντροδυςτιον οφ α μυλτι-διμενςιοναλ ιντερναλ ςπαςε [^17] Fock, in his 1926 paper[18], had described the parameter space of the quantum mechanical phase as a fifth dimension.. Ιν 1937 Ε. Π. Ωιγνερ προποςεδ τηε τερμ ιςοτοπις ςπιν ωηιςη, ςιμπλιφιεδ το ιςοςπιν, ηας βεεν υςεδ υντιλ νοω.
Ιν τηε φολλοωινγ ψεαρς τηρεε ιμπορταντ δεελοπμεντς αλλοωεδ Ηειςενβεργ´ς ινιτιαλ ςυγγεςτιον το βεςομε α ςομπλετε ιςοςπιν ιναριαντ τηεορψ. Τηε φιρςτ, ανδ προβαβλψ τηε μοςτ ιμπορταντ, ωας τηε προγρεςς ιν εξπεριμενταλ τεςηνιχυες, ωηιςη βρουγητ μορε δεταιλεδ ανδ μορε πρεςιςε δατα. Τηεψ ςηοωεδ τηε νεεδ φορ τηε ιντροδυςτιον οφ α $p-p$ φορςε ανδ ςονφιρμεδ τηε ςηαργε ινδεπενδενςε οφ αλλ νυςλεαρ φορςες. Τηε ςεςονδ ωας τηε φορμυλατιον βψ Φερμι ιν 1933 οφ α τηεορετιςαλ μοδελ φορ τηε αμπλιτυδε οφ νευτρον $\beta$-δεςαψ ιν ωηιςη ηε ιντροδυςεδ φορ τηε φιρςτ τιμε τηε ςονςεπτ οφ χυαντιςεδ φερμιον φιελδς. Σινςε τηατ τιμε χυαντυμ φιελδ τηεορψ βεςαμε τηε υνιερςαλ λανγυαγε οφ μοδερν τηεορετιςαλ πηψςιςς. Τηε τηιρδ ιμπορταντ δεελοπμεντ ωας Ψυϰαωα´ς ιντροδυςτιον οφ τηε μεςον ας αν ιντερμεδιαρψ φορ τηε νυςλεαρ φορςες. Ινιτιαλλψ ηε τηουγητ οφ α εςτορ φιελδ ωηοςε τηρεε ςπαςε ςομπονεντς ωουλδ μεδιατε τηε ωεαϰ ιντεραςτιονς ανδ τηε ζερο ςομπονεντ τηε νυςλεαρ φορςες, βυτ ιτ ωας ςοον ρεπλαςεδ βψ α ςπιν ζερο ονε. Τηε φιναλ ςψντηεςις ις δυε το Ν. Κεμμερ ιν 1938[5]. Ιτ τοοϰ ςιξ ψεαρς, ας ωελλ ας τηε ωορϰ οφ μανψ πηψςιςιςτς, φορ Ηειςενβεργ´ς οριγιναλ ςυγγεςτιον οφ 1932 το βεςομε τηε φυλλ ιςοςπιν ςψμμετρψ οφ ηαδρονις πηψςιςς ωε ϰνοω τοδαψ. Ιτ ις α ρεμαρϰαβλψ ςηορτ τιμε, γιεν τηε ρεολυτιοναρψ νατυρε οφ τηε ςονςεπτ.
Ηειςενβεργ´ς ιςοςπιν ςπαςε ωας τηρεε διμενςιοναλ ανδ τηε τρανςφορματιονς φορμ α γρουπ $O(3)$, ορ $SU(2)$, λιϰε ουρ φαμιλιαρ ροτατιονς. Ηοωεερ τηε ςονςεπτ ηας βεεν ςυβςεχυεντλψ ενλαργεδ ας νεω παρτιςλες ωερε διςςοερεδ ανδ λαργερ ιντερναλ ςψμμετρψ γρουπς ωερε βρουγητ ιντο ειδενςε. Τηε ςπαςε οφ ελεμενταρψ παρτιςλε πηψςιςς βεςαμε α μυλτι-διμενςιοναλ μανιφολδ, ωιτη ςομπλιςατεδ γεομετριςαλ ανδ τοπολογιςαλ προπερτιες, ανδ ονλψ α ςυβ-ςπαςε οφ ιτ, τηε φουρ-διμενςιοναλ Μινϰοωςϰι ςπαςε, ις διρεςτλψ αςςεςςιβλε το ουρ ςενςες.
$\bullet$ Νον-αβελιαν γαυγε ςψμμετριες. Νατυραλλψ, ονε ωουλδ εξπεςτ τηε $SU(2)$ γαυγε τηεορψ το βε ςονςτρυςτεδ φολλοωινγ τηε πρινςιπλες ωε ςϰετςηεδ αβοε: ωε ηαδ τηε γλοβαλ ςψμμετρψ ανδ ωε ονλψ νεεδεδ το μαϰε ιτ λοςαλ. Βυτ ηερε ηιςτορψ τοοϰ α τοταλλψ υνεξπεςτεδ ρουτε. Α ηιςτοριςαλ αςςιδεντ – τηε διςςοερψ οφ γενεραλ ρελατιιτψ – ιντροδυςεδ αν υννεςεςςαρψ ςομπλιςατιον. Τηε φαςςινατιον ωηιςη γενεραλ ρελατιιτψ ηαδ εξερτεδ το αλλ τηις γενερατιον οφ πηψςιςιςτς ωας ςυςη, τηατ φορ μανψ δεςαδες πεοπλε ωερε υναβλε το ςονςειε λοςαλ τρανςφορματιονς ωιτηουτ διφφεομορπηιςμς. Τηερεφορε τηεψ υςεδ τηε τηεορψ οφ γραιτατιον εεν ιν πλαςες ωηερε ιτ ηαδ νο βυςινεςς το βε τηερε. Ι ςαλλ ιτ “τηε ΓΡ-ςψνδρομε”.
Τηε φιρςτ περςον ωηο τριεδ το ςονςτρυςτ τηε γαυγε τηεορψ φορ $SU(2)$ ις Κλειν[22] ωηο, ιν αν οβςςυρε ςονφερενςε ιν 1938, πρεςεντεδ α παπερ ωιτη τηε τιτλε: Ον τηε τηεορψ οφ ςηαργεδ φιελδς. Τηε μοςτ αμαζινγ παρτ οφ τηις ωορϰ ις τηατ ιτ φολλοως αν ινςρεδιβλψ ςιρςυιτους ροαδ: Ηε ςονςιδερς γενεραλ ρελατιιτψ ιν α φιε διμενςιοναλ ςπαςε ανδ ςομπαςτιφιες ὰ λα Καλυζα-Κλειν. Τηεν ηε ταϰες τηε λιμιτ ιν ωηιςη γραιτατιον ις δεςουπλεδ. Ιν ςπιτε οφ ςομε ςονφυςεδ νοτατιον [^18] He starts from the discovery of the muon, misinterpreted as the Yukawa meson, in the old Yukawa theory in which the mesons were assumed to be vector particles. This provides the physical motivation. The aim is to write an $SU(2)$ gauge theory unifying electromagnetism and nuclear forces. He has a five-dimensional theory but he takes the $g_{4\mu}$ components of the metric tensor to be $2\times 2$ matrices. He wants to describe the $SU(2)$ gauge fields but the matrices he is using, although they depend on three fields, are not traceless. In fact, answering an objection by Møller, he added a fourth vector field, thus promoting his theory to $U(1)\times SU(2)$., ηε φινδς τηε ςορρεςτ εξπρεςςιον φορ τηε φιελδ ςτρενγτη τενςορ οφ $SU(2)$. Ηε ςονςιδερεδ μαςςιε εςτορ βοςονς ανδ ιτ ις νοτ ςλεαρ ωηετηερ ηε ωορριεδ αβουτ τηε ρεςυλτινγ βρεαϰινγ οφ γαυγε ιναριανςε. Ι ςαννοτ φινδ ουτ ωηετηερ τηις παπερ ηας ινςπιρεδ ανψβοδψ ελςε´ς ωορϰ βεςαυςε τηε προςεεδινγς οφ τηις ςονφερενςε αρε νοτ ινςλυδεδ ιν τηε ςιτατιον ινδεξ. Ας φαρ ας Ι ϰνοω, Κλειν ηιμςελφ διδ νοτ φολλοω υπ ον τηις ιδεα [^19] He mentioned this work in a 1955 Conference in Berne[23]..
Τηε ςεςονδ ωορϰ ιν τηε ςαμε ςπιριτ ις δυε το Παυλι[24] ωηο, ιν 1953, ιν α λεττερ το Παις, δεελοπεδ πρεςιςελψ τηις αππροαςη: τηε ςονςτρυςτιον οφ τηε $SU(2)$ γαυγε τηεορψ ας τηε φλατ ςπαςε λιμιτ οφ α ςομπαςτιφιεδ ηιγηερ διμενςιοναλ τηεορψ οφ γενεραλ ρελατιιτψ. Ηε ωας ςλοςερ το τηε αππροαςη φολλοωεδ τοδαψ βεςαυςε ηε ςονςιδερεδ α ςιξ διμενςιοναλ τηεορψ ωιτη τηε ςομπαςτ ςπαςε φορμινγ αν $S_{2}$. Ηε νεερ πυβλιςηεδ τηις ωορϰ ανδ Ι δο νοτ ϰνοω ωηετηερ ηε ωας αωαρε οφ Κλειν’ς 1938 παπερ. Ηε ηαδ ρεαλιςεδ τηατ α μαςς τερμ φορ τηε γαυγε βοςονς βρεαϰς τηε ιναριανςε ανδ ηε ηαδ αν ανιματεδ αργυμεντ δυρινγ α ςεμιναρ βψ Ψανγ ιν τηε Ινςτιτυτε φορ Αδανςεδ Στυδιες ιν Πρινςετον ιν 1954[25]. Ωηατ ις ςυρπριςινγ ις τηατ Κλειν ανδ Παυλι, φιφτεεν ψεαρς απαρτ ονε φρομ τηε οτηερ, δεςιδεδ το ςονςτρυςτ τηε $SU(2)$ γαυγε τηεορψ φορ ςτρονγ ιντεραςτιονς ανδ βοτη ςηοοςε το φολλοω τηις τοταλλψ ςουντερ-ιντυιτιε μετηοδ. Στριϰινγ εξαμπλες οφ τηε ΓΡ-ςψνδρομε. ῝.Ν. Ψανγ ανδ Ρ.Λ. Μιλλς[26] ωερε τηε φιρςτ το υνδερςτανδ τηατ τηε γαυγε τηεορψ οφ αν ιντερναλ ςψμμετρψ ταϰες πλαςε ιν α φιξεδ βαςϰγρουνδ ςπαςε ωηιςη ςαν βε ςηοςεν το βε φλατ, ιν ωηιςη ςαςε γενεραλ ρελατιιτψ πλαψς νο ρολε. Σινςε τηατ τιμε νον-αβελιαν γαυγε τηεοριες βεςαμε παρτ οφ ηιγη ενεργψ πηψςιςς. Ιτ ις νοτ ςυρπριςινγ τηατ τηεψ ωερε ιμμεδιατελψ ναμεδ Ψανγ-Μιλλς τηεοριες. Τηε εξτενςιον το οτηερ γρουπς ινςλυδινγ διρεςτ προδυςτς, ωας δονε βψ Ρ. ϒτιψαμα[27] ας ωελλ ας Σ.Λ. Γλαςηοω ανδ Μ. Γελλ-Μανν[28].
$\bullet$ Φρομ Ψανγ-Μιλλς το πηψςιςς. Αλτηουγη Ψανγ ανδ Μιλλς ωαντεδ το ςονςτρυςτ α γαυγε τηεορψ οφ νυςλεαρ φορςες, τηειρ ωορϰ ηας νεερ βεεν υςεδ φορ τηατ πυρποςε. Ιν φαςτ, τηε φιρςτ αττεμπτς ςονςερνεδ τηε ωεαϰ ιντεραςτιονς. Ιν 1957, Σςηωινγερ ηαδ ςονϑεςτυρεδ[29] τηατ τηε Φερμι τηεορψ ςηουλδ βε μοδιφιεδ ωιτη τηε ιντροδυςτιον οφ αν ιντερμεδιατε εςτορ βοςον (Ι῞Β) $W_{\mu}^{\pm}$.
$H_{I}=gJ^{\mu}(x)W_{\mu}^{-}(x)+hc$ (10)
ωιτη $g$ α νεω διμενςιονλεςς ςουπλινγ ςονςταντ. Τηις ωαψ ωεαϰ ιντεραςτιονς λοοϰεδ πρεττψ μυςη λιϰε τηε ελεςτρομαγνετις ονες, α εςτορ βοςον ςουπλεδ το α ςυρρεντ, βυτ ωιτη ςομε ερψ ιμπορταντ διφφερενςες: (ι) Τηε πηοτον ις μαςςλεςς ανδ τηε ε.μ. ιντεραςτιονς αρε λονγ ρανγεδ. Τηε ωεαϰ ιντεραςτιονς αρε ϰνοων το βε ςηορτ ρανγεδ, ςο τηε $W$´ς μυςτ βε μαςςιε. (ιι) Τηε πηοτον ις νευτραλ, τηε $W$´ς αρε ςηαργεδ. (ιιι) Τηε ελεςτρομαγνετις ςυρρεντ ις ςονςερεδ, τηε ωεαϰ ςυρρεντ ις νοτ. Ιτ ωας ςοον ςλεαρ τηατ τηεςε διφφερενςες ιμπλιεδ τηατ τηε τηεορψ (10), ωιτη μαςςιε εςτορ βοςονς, ωας νον-ρενορμαλιςαβλε. Ιν ορδερ το τυρν τηε Ι῞Β μοδελ ιντο α Ψανγ-Μιλλς τηεορψ, Σςηωινγερ αςςυμεδ τηε εξιςτενςε οφ α τριπλετ οφ ιντερμεδιατε βοςονς, ωηιςη ηε ςαλλεδ $Z^{\pm,0}$, τηε τωο ςηαργεδ ονες μεδιατινγ τηε ωεαϰ ιντεραςτιονς ανδ τηε νευτραλ ονε βεινγ τηε πηοτον. Α ψεαρ λατερ, ιν 1958, Σ.Α. Βλυδμαν[30] βυιλτ αν $SU(2)$ Ψανγ-Μιλλς τηεορψ φορ ωεαϰ ιντεραςτιονς ιν ωηιςη αλλ τηρεε γαυγε βοςονς ωερε ςουπλεδ το $V-A$ ςυρρεντς. Νο ςοννεςτιον ωιτη ελεςτρομαγνετιςμ ωας αςςυμεδ.
Τηε μοςτ ιμπορταντ ςοντριβυτιον φρομ τηις περιοδ δατες φρομ 1961202020It has a submission date of Sept. 9 1960. ανδ ιτ ις δυε το Σ.Λ. Γλαςηοω[31]. Ιτ ις τηε ωορϰ τηατ ιντροδυςεδ τηε $SU(2)\times U(1)$ ελεςτροωεαϰ γαυγε τηεορψ. Γλαςηοω προποςεδ α υνιφιεδ δεςςριπτιον φορ ωεαϰ ανδ ελεςτρομαγνετις ιντεραςτιονς ανδ ηις μοδελ ις τηε ονε ωε υςε τοδαψ. Τηε ονλψ μιςςινγ ελεμεντ ις τηε ςποντανεους ςψμμετρψ βρεαϰινγ μεςηανιςμ ωηιςη ωας ινεντεδ ιν 1964, ςεε βελοω. Τηε τηεορψ ηας τωο νευτραλ γαυγε βοςονς – ονε αςςοςιατεδ το $U(1)$ ανδ α ςεςονδ αςςοςιατεδ το τηε τηιρδ γενερατορ οφ $SU(2)$ – ανδ ιντροδυςες τηε ιδεα οφ α μιξινγ βετωεεν τηε τωο. Τηε πηοτον φιελδ ις α λινεαρ ςομβινατιον οφ τηε τωο νευτραλ εςτορ φιελδς ωιτη αν ανγλε ωηιςη Γλαςηοω ςαλλεδ $\theta$ (τοδαψ ιτ ις ςαλλεδ $\theta_{W}$).
Τηε Γλαςηοω ανδ Γελλ-Μανν παπερ[28], ιν αδδιτιον το εξτενδινγ τηε Ψανγ-Μιλλς τηεορψ το μορε γενεραλ γρουπς, ιντροδυςες ςεεραλ νεω ελεμεντς ςομε οφ ωηιςη ηαε βεεν ινςορπορατεδ ιντο ουρ πρεςεντ γαυγε τηεοριες. Τηεψ νοτιςε τηατ τηε ιντροδυςτιον βψ ηανδ οφ α μαςς φορ τηε γαυγε βοςονς βρεαϰς γαυγε ιναριανςε ανδ τηεψ ταλϰ αβουτ παρτιαλλψ γαυγε ιναριαντ τηεοριες. Τηεψ εξτενδ τηε Ψανγ-Μιλλς ςονςτρυςτιον το ανψ αλγεβρα ωηιςη ςαν βε ωριττεν ας α διρεςτ προδυςτ οφ ςιμπλε φαςτορς. Τηε ωελλ-ϰνοων ρεςυλτ οφ αςςοςιατινγ α ςουπλινγ ςονςταντ το εερψ φαςτορ ιν τηε αλγεβρα αππεαρεδ φορ τηε φιρςτ τιμε ιν τηις παπερ. Εεν τηε ςεεδ φορ α γρανδ υνιφιεδ τηεορψ ωας τηερε. Ιν α φοοτνοτε τηεψ ςαψ: “Τηε ρεμαρϰαβλε υνιερςαλιτψ οφ τηε ελεςτρις ςηαργε ωουλδ βε βεττερ υνδερςτοοδ ωερε τηε πηοτον νοτ μερελψ α ςινγλετ, βυτ α μεμβερ οφ α φαμιλψ οφ εςτορ μεςονς ςομπριςινγ α ςιμπλε παρτιαλλψ γαυγε ιναριαντ τηεορψ.” Φιναλλψ, τηεψ ςορρεςτλψ ιδεντιφψ τηε προβλεμς ρελατεδ το τηε αβςενςε οφ ςτρανγενεςς ςηανγινγ νευτραλ ςυρρεντς ανδ τηε ςμαλλ αλυε οφ τηε $K_{1}^{0}-K_{2}^{0}$ μαςς διφφερενςε, προβλεμς ωηιςη ωιλλ τυρν ουτ το βε ςρυςιαλ φορ τηε ςονςτρυςτιον οφ τηε Στανδαρδ Μοδελ [^21] I do not know why this paper has not received the attention it deserves, but this is partly due to the authors themselves, especially Gell-Mann, who rarely referred to it..
[…]
$\bullet$ Σιμπλε το υνδερςτανδ, ηαρδ το προε. Ας ωε ςαιδ πρειουςλψ, τηε τηεορψ οφ ρενορμαλιςατιον, ατ λεαςτ ιν τηε φορμ ωε υςε ιν χυαντυμ φιελδ τηεορψ, δατες φρομ τηε λατε νινετεεν-φορτιες. Ηοωεερ, ιτς ροοτς αρε μυςη ολδερ. Ι δο νοτ ϰνοω ωηο ωας τηε φιρςτ το ρεαλιςε τηατ τηε ςονςεπτ οφ ποιντ παρτιςλες ιντροδυςεδ ςηορτ διςτανςε ςινγυλαριτιες. Φορ εξαμπλε, τηε τωο ςλαςςιςαλ φορςες, ελεςτρομαγνετιςμ ανδ γραιτατιον, αρε βοτη δεςςριβεδ βψ α $1/r$ ποτεντιαλ ωηιςη βεςομες ςινγυλαρ φορ $r=0$. Τηις μαϰες τηε ςελφ ενεργψ οφ αν ελεςτρον ινφινιτε ανδ τηε Ρυτηερφορδ ατομ υνςταβλε. Λορεντζ τριεδ το ςολε τηε ελεςτρον προβλεμ βψ ςμεαρινγ τηε ςηαργε διςτριβυτιον οερ α φινιτε ςιζε, αν ιντυιτιελψ αττραςτιε ιδεα βυτ ερψ ηαρδ το ιμπλεμεντ ιν πραςτιςε, εςπεςιαλλψ ιν α ρελατιιςτις φραμεωορϰ. Τηε δεελοπμεντ οφ τηε χυαντυμ τηεορψ ςολεδ τηε προβλεμ οφ ατομις ςταβιλιτψ ιν α ερψ ςτρανγε ωαψ: ιντροδυςινγ νον-ςομμυτατιιτψ ιν πηαςε ςπαςε, ιτ εξπρεςςεδ τηε προβλεμ οφ τηε ελεςτρον ενεργψ ιν αν ατομ ας αν ειγεναλυε προβλεμ ωηιςη, φορ τηε ῝ουλομβ ποτεντιαλ, ις ςηοων το ηαε α διςςρετε ςπεςτρυμ βουνδεδ φρομ βελοω [^22] It is worth mentioning here that Schrödinger introduced his formulation of quantum mechanics in a series of articles with the title “Quantisierung als Eigenwertproblem” or, “Quantisation as an Eigenvalue Problem”[32].. Τηε εαρλψ αττεμπτς το φορμυλατε α χυαντυμ τηεορψ οφ ραδιατιον ιντροδυςεδ νεω φορμς οφ διεργενςες, ιν αδδιτιον το τηε ελεςτρον ςελφ ενεργψ, ανδ τηε ηιςτορψ ιν τηε λατε τωεντιες ανδ τηιρτιες ις χυιτε ςονφυςινγ [^23] In the book “Inward Bound” by A. Pais[3], Chapter 16 has the title “Battling the infinite” and covers this period..
Ας φαρ ας Ι ϰνοω, τηε τερμ “ρενορμαλιςατιον” ωας φιρςτ υςεδ βψ Ρ. Σερβερ[33] ιν 1936 ιν αν αττεμπτ το υνδερςτανδ τηε προβλεμ οφ αςυυμ πολαριςατιον, βυτ Ι τηινϰ τηατ τηε ρεαλ πρεςυρςορ οφ τηε μοδερν τηεορψ οφ ρενορμαλιςατιον ις προβαβλψ Η.Α. Κραμερς ωηο, ιν 1938, ιντροδυςεδ τηε ιδεα οφ α μαςς ρενορμαλιςατιον φορ τηε ελεςτρον[34]. Ηε ωας αλςο α διςςυςςιον λεαδερ ατ τηε Σηελτερ Ιςλανδ ῝ονφερενςε. Αλτηουγη τηε βαςις ιδεας ωερε λαιδ δοων μορε τηαν 75 ψεαρς αγο, τηε δεταιλεδ πρινςιπλες, βοτη πηψςιςαλ ανδ τεςηνιςαλ, εολεδ ςονςιδεραβλψ. Τηεψ γρεω ςιμπλερ το υνδερςτανδ, βυτ χυιτε ςομπλιςατεδ το προε. Φορ μανψ ψεαρς τηε τηεορψ ωας ςονςιδερεδ ας α ματηεματιςαλλψ μυρϰψ προςεςς ςονςιςτινγ ιν αδδινγ ανδ ςυβτραςτινγ ινφινιτιες. Εεν τηε φουνδινγ φατηερς φορ α λονγ τιμε ςηαρεδ τηις οπινιον [^24] Feynman, in his report at the 12th Solvay Conference on Physics, noting the impossibility to compute mass ratios, said “I still hold this belief, [that the theory must fail] and do not subscribe to the philosophy of renormalisation.”[35]. Schwinger, during the nineteen-sixties, lost interest in quantum field theory and attempted to develop a new approach (Source Theory) today forgotten. I remind that in these years we were still deep in the Dark Ages.. Τοδαψ ωε υνδερςτανδ τηατ νοτηινγ ις μορε ρεμοτε φρομ τηε τρυτη. Τηε τηεορψ οφ ρενορμαλιςατιον οφφερς τηε ονλψ ϰνοων ματηεματιςαλλψ ςονςιςτεντ ωαψ το δεφινε τηε ςυςςεςςιε τερμς ιν τηε περτυρβατιον εξπανςιον οφ α χυαντυμ φιελδ τηεορψ. Ι ωαντ το εμπηαςιςε ηερε τηατ, ιφ ιν α ςαλςυλατιον οφ α πηψςιςαλ χυαντιτψ ωε ενςουντερ α διεργεντ εξπρεςςιον, ιτ μεανς τηατ ωε ηαε μαδε α ματηεματιςαλ μιςταϰε. Διδ ωε μαϰε α μιςταϰε ιν φορμυλατινγ χυαντυμ φιελδ τηεορψ ατ ςηορτ διςτανςες; Ψες ωε διδ! Λετ υς ταϰε τηε εξαμπλε οφ α ςελφ ιντεραςτινγ ςςαλαρ φιελδ. Τηε Λαγρανγιαν δενςιτψ ανδ τηε ςανονιςαλ ςομμυτατιον ρελατιονς αρε
[…]
Ωε ϰνοω τηατ τηε Διρας $\delta$-φυνςτιον ις νοτ α “φυνςτιον” βυτ α ςπεςιαλ φορμ οφ ωηατ ωε ςαλλ “α διςτριβυτιον”. Μανψ προπερτιες οφ ωελλ-βεηαεδ φυνςτιονς δο νοτ αππλψ το ιτ. Ιν παρτιςυλαρ, τηε μυλτιπλιςατιον ις νοτ αλωαψς α ωελλ δεφινεδ οπερατιον· $(\delta(x))^{2}$ ις μεανινγλεςς. Τηε πρεςενςε οφ τηε $\delta$-φυνςτιον ιν (11) ιμπλιες τηατ τηε φιελδ $\phi(x)$ ις αλςο α διςτριβυτιον, ςο τηε προδυςτ $\phi^{2}$ ις ιλλ δεφινεδ. Ψετ, ιτ ις πρεςιςελψ εξπρεςςιονς οφ τηις ϰινδ τηατ ωε ωριτε ιν εερψ ςινγλε τερμ οφ ουρ Λαγρανγιαν, ςο ιτ ις νοτ ςυρπριςινγ τηατ ουρ ςαλςυλατιονς ψιελδ διεργεντ ρεςυλτς.
Α ςονςεπτυαλλψ ςιμπλε ωαψ το ςολε τηε προβλεμ – νοτ νεςεςςαριλψ τηε μοςτ ςονενιεντ ονε φορ πραςτιςαλ ςαλςυλατιονς – ωουλδ βε το ρεπλαςε τηε φιελδ προδυςτς ιν (11) βψ ςπλιττινγ τηε ποιντς:
$\phi(x)\phi(x)\rightarrow\lim_{a\rightarrow 0}\phi(x+\frac{a}{2})\phi(x-\frac{a}{2})$ (12)
Τηις εξπρεςςιον ις περφεςτλψ ωελλ δεφινεδ φορ αλλ αλυες οφ τηε παραμετερ $a$, εξςεπτ $a=0$. Ιν τερμς οφ διςτριβυτιονς τηις μεανς τηατ τηε προδυςτ ις δεφινεδ υπ το αν αρβιτραρψ διςτριβυτιον ${\cal F}(a)$ ωηιςη ηας ςυππορτ (ι.ε. ιτ ις νον-ζερο), ονλψ ατ $a=0$. Συςη α διςτριβυτιον ις α ςυπερποςιτιον οφ τηε $\delta$-φυνςτιον ανδ ιτς δεριατιες.
${\cal F}(a)=\sum_{i}C_{i}\delta^{(i)}(a)$ (13)
ωιτη τηε $C_{i}$´ς αρβιτραρψ ρεαλ ςονςταντς. Τηε μοραλ οφ τηε ςτορψ ις τηατ τηε χυαντιςατιον ρυλες φορ α λοςαλ φιελδ τηεορψ ιμπλψ τηατ εερψ τερμ ιν τηε Λαγρανγιαν ςονταινς ιν φαςτ α ςετ οφ αρβιτραρψ ςονςταντς ωηιςη μυςτ βε δετερμινεδ βψ εξπεριμεντ. Ρενορμαλιςατιον ις τηε ματηεματιςαλ προςεδυρε ωηιςη αλλοως υς το δο ιτ. Ηοω μανψ παραμετερς αρε νεεδεδ ιν ορδερ το δεφινε α γιεν φιελδ τηεορψ; Τηε ανςωερ ινολες τηε διςτινςτιον βετωεεν ρενορμαλιςαβλε ανδ νον-ρενορμαλιςαβλε τηεοριες. Φορ τηε φιρςτ α φινιτε νυμβερ ςυφφιςες. Φορ τηε ςεςονδ ωε νεεδ αν ινφινιτε νυμβερ, ωηιςη μεανς τηατ νον-ρενορμαλιςαβλε τηεοριες ηαε νο πρεδιςτιε ποωερ. Α φιναλ ρεμαρϰ: Ρενορμαλιςατιον ηας α ωελλ-δεςερεδ ρεπυτατιον οφ ςομπλεξιτψ αλτηουγη ιτς πρινςιπλε ις ρατηερ ςιμπλε. Τηε ςομπλιςατιονς αρε τεςηνιςαλ. Φορ εξαμπλε, ιν ορδερ το δεφινε τηε Λαγρανγιαν οφ εχ. (11) ωε νεεδ ονλψ τηρεε αρβιτραρψ ςονςταντς. Τωο οφ τηεμ αρε αςςοςιατεδ το τηε τωο-ποιντ φυνςτιον ανδ ςαν βε τηουγητ ας δεφινινγ τηε μαςς $m$ ανδ τηε νορμαλιςατιον οφ τηε φιελδ, ανδ τηε τηιρδ ονε ις αςςοςιατεδ το τηε φουρ-ποιντ φυνςτιον ανδ δεφινες τηε αλυε οφ τηε ςουπλινγ ςονςταντ $\lambda$. Ηοωεερ, τηε ριγορους προοφ οφ τηις ςτατεμεντ ατ ανψ ορδερ οφ περτυρβατιον, ρεχυιρες ςομε χυιτε ςομπλιςατεδ ςαλςυλατιονς [^25] The computational rules which are used in practical calculations were first formulated by Feynman and they are still called Feynman rules. For multi-loop diagrams they involve integrations over some auxiliary parameters and the difficulty consists in proving that the resulting expressions are always integrable. For example, in 1961, T.T. Wu discovered that the naïve rules needed to be modified for some very high order diagrams containing sub-diagrams with more than 11 external lines[36]. I remember the admiration aroused by this result, since most of us could not even draw such a diagram, let alone study its analyticity properties.. Ας ωε ςαιδ εαρλιερ, τηε εαρλψ φορμυλατιον οφ τηε τηεορψ το αλλ ορδερς οφ περτυρβατιον ις δυε το Δψςον[37]. Ιτ ωας φυρτηερ δεελοπεδ βψ μανψ πεοπλε, ιν παρτιςυλαρ βψ Ν.Ν. Βογολιυβο ανδ Ο. Παραςψυϰ ιν 1955[38], βυτ τηε φιρςτ ματηεματιςαλλψ ςουνδ προοφ ςηοωινγ τηατ τηε μετηοδ ωορϰς το αλλ ορδερς ις δυε το Κ. Ηεππ ιν 1966[39]. Ιν 1973 Η. Επςτειν ανδ ῞. Γλαςερ πρεςεντεδ α νεω προοφ, φορμυλατεδ διρεςτλψ ιν ποςιτιον ςπαςε ιν τηε ςπιριτ οφ εχυατιονς (12) ανδ (13), ωηιςη εξηιβιτς τηε ςιμπλε πηψςιςαλ πρινςιπλες ιν α μορε τρανςπαρεντ ωαψ[40]. Τηις ςτορψ ςηοως τηατ φορ αλμοςτ τωο δεςαδες πεοπλε ωερε υςινγ ρενορμαλιςεδ περτυρβατιον τηεορψ ωιτηουτ ηαινγ α προοφ φορ ιτς ματηεματιςαλ ςονςιςτενςψ. Ιτ ις αλςο ιντερεςτινγ το νοτιςε τηατ δυρινγ αλλ τηις τιμε ςεεραλ τεξτ βοοϰς ον χυαντυμ φιελδ τηεορψ ωερε πυβλιςηεδ, αλλ πρεςεντινγ τηε τηεορψ οφ ρενορμαλιςατιον, βυτ νο ονε μεντιονινγ τηε φαςτ τηατ α ρεαλ προοφ ωας μιςςινγ! Αν εξαμπλε οφ τηε ρατηερ διλετταντε αττιτυδε οφ πηψςιςιςτς ρεγαρδινγ ματηεματιςαλ ριγορ. $\bullet$ Ρενορμαλιςατιον ανδ γαυγε ιναριανςε – ΧΕΔ. Τηε τηεορψ οφ ρενορμαλιςατιον ωας φιρςτ αππλιεδ το χυαντυμ ελεςτροδψναμιςς ωηιςη ις α γαυγε τηεορψ. Γαυγε ιναριανςε ιντροδυςες ςομε εξτρα ςομπλιςατιονς ωηιςη ωερε νοτ ιμμεδιατελψ ιδεντιφιεδ. Ι ωιλλ νοτ φολλοω ηερε τηε ςηρονολογιςαλ ορδερ, βυτ Ι ωιλλ τρψ το εξπλαιν τηε προβλεμς ωηιςη ωιλλ τυρν ουτ το βε ςρυςιαλ φορ τηε ρενορμαλιςατιον οφ τηε Ψανγ-Μιλλς τηεοριες.
Τηε ςηαραςτεριςτις φεατυρε οφ ςυςη τηεοριες ις τηε πρεςενςε οφ εςτορ φιελδς $A_{\mu}(x)$ ωηιςη, φρομ τηε φιελδ τηεορψ ποιντ οφ ιεω, ιμπλψ τηε ιντροδυςτιον οφ φουρ δεγρεες οφ φρεεδομ. Ηοωεερ, Ποινςαρé ιναριανςε τελλς υς τηατ α μαςςιε ςπιν-ονε παρτιςλε ηας τηρεε δεγρεες οφ πολαριςατιον ανδ α μαςςλεςς ονε τωο. Ιτ ις τηις μιςματςη βετωεεν τηε Λορεντζ ανδ τηε Ποινςαρé ςουντινγ ωηιςη ςομπλιςατες τηε ςτυδψ οφ χυαντυμ φιελδ τηεοριες δεςςριβινγ παρτιςλες ωιτη ςπιν ηιγηερ τηαν 1/2. Ωε μυςτ ινςλυδε ςονςτραιντς ιν ορδερ το ελιμινατε τηε ρεδυνδαντ δεγρεες οφ φρεεδομ. Φορ τηε μαςςλεςς ςαςε τηεψ αρε τηε αναλογυες οφ τηε γαυγε ςονδιτιονς ωε ενςουντερεδ ιν ςλαςςιςαλ ελεςτροδψναμιςς.
Λετ υς ςταρτ ωιτη ΧΕΔ. Τηε ιναριανςε υνδερ γαυγε τρανςφορματιονς ιμπλιες τηατ, ιφ ωε φινδ α ςολυτιον φορ τηε εςτορ ποτεντιαλ $A_{\mu}(x)=\hat{A}_{\mu}(x)$, ωε ςαν ςονςτρυςτ αν ινφινιτψ οφ οτηερς βψ ωριτινγ $A_{\mu}(x)=\hat{A}_{\mu}(x)+\partial_{\mu}\theta(x)$ ωιτη $\theta(x)$ αν αρβιτραρψ φυνςτιον. Τηε γαυγε ςονδιτιονς αρε ςυπποςεδ το ρεςτριςτ τηις αρβιτραρινεςς ανδ, ηοπεφυλλψ, ςηοοςε α υνιχυε ςολυτιον [^26] V.N. Gribov has shown that, in general, there exists no gauge condition yielding a unique solution[41]. This phenomenon is known as “Gribov ambiguity”. It is important in the attempts to give a global definition of a gauge theory, but it does not affect the perturbation expansion.. Σινςε ωε ϰνοω τηατ ιν Νατυρε ονλψ τρανςερςε πηοτονς αρε πηψςιςαλ, τηε οβιους ιδεα ις το ιμποςε τηε ῝ουλομβ γαυγε ςονδιτιον ${\bm{\nabla}}\cdot{\bm{A}}(x)\!=\!0$. Ιτ ις εαςψ το ςηοω τηατ ιν τηις γαυγε ωε αρε λεφτ ωιτη τηε τωο “πηψςιςαλ” δεγρεες οφ φρεεδομ, το ωιτ τηε τρανςερςε πηοτονς. Ιτ ις τηε μετηοδ υςεδ βψ Διρας ιν τηε ςαλςυλατιον οφ τηε ςποντανεους εμιςςιον προβαβιλιτψ, τηε φιρςτ πηψςιςαλ αππλιςατιον οφ τηε χυαντιςεδ ραδιατιον φιελδ[42].
Ιν ηιγη ενεργψ πηψςιςς τηε υςε οφ α ςοαριαντ γαυγε, ςυςη ας τηε Λορενζ γαυγε $\partial^{\mu}A_{\mu}=0$, ις μυςη μορε ςονενιεντ. Ιν α ςιμπλε-μινδεδ αππροαςη, ωε ϑυςτ πλυγ τηε ςονδιτιον ιν τηε εχυατιον οφ μοτιον ωηιςη βεςομες $\Box A_{\nu}=j_{\nu}$. Τηε φουρ ςομπονεντς οφ τηε εςτορ φιελδ δεςουπλε, ςο τηε τηεορψ ςεεμς το δεςςριβε φουρ ινδεπενδεντ δεγρεες οφ φρεεδομ. Ιτ ις τηε μετηοδ ωηιςη ωας υςεδ ιν τηε εαρλψ ςαλςυλατιονς[37]. Τηερεφορε, τηε πριςε φορ ϰεεπινγ Λορεντζ ςοαριανςε ις το ωορϰ ωιτη α τηεορψ ωηιςη ινςλυδες υνπηψςιςαλ δεγρεες οφ φρεεδομ. Τηε ματηεματιςαλ δεςςριπτιον ωας προποςεδ ιν 1950 βψ Σ.Ν. Γυπτα ανδ Κ. Βλευλερ[43] ανδ ις ϰνοων ας τηε Γυπτα-Βλευλερ μετηοδ. Ιτ ινολες αν ενλαργεδ Ηιλβερτ ςπαςε, ςονταινινγ λονγιτυδιναλ ανδ ςςαλαρ πηοτονς, ιν αδδιτιον το τηε πηψςιςαλ τρανςερςε ονες. Τηε ςτρανγε φεατυρε – α ςονςεχυενςε οφ τηε φαςτ τηατ τηε Μινϰοωςϰι μετρις ις νοτ ποςιτιε δεφινιτε – ις τηατ τηε ςτατες ωιτη ςςαλαρ πηοτονς δο νοτ ηαε ποςιτιε νορμ.
Ατ ριςϰ οφ αππεαρινγ πεδαντις, Ι ωαντ το δεελοπ τηις αππροαςη ιν α μορε φορμαλ ωαψ, ωηιςη ωιλλ μαϰε τηε γενεραλιςατιον το τηε Ψανγ-Μιλλς τηεορψ εαςιερ. Ι ωιλλ ιντροδυςε τηε Λορενζ γαυγε ςονδιτιον υςινγ α Λαγρανγε μυλτιπλιερ, α μετηοδ ινεντεδ βψ Λαγρανγε ιν 1788. Τηε ΧΕΔ Λαγρανγιαν ις ωριττεν ας
${\cal L}=-\frac{1}{4}F_{\mu\nu}(x)F^{\mu\nu}(x)+b(x)\partial^{\mu}A_{\mu}(x)-j_{\mu}(x)A^{\mu}(x)$ (14)
ωηερε τηε ςυρρεντ $j_{\mu}$ ις γιεν ιν τερμς οφ τηε ελεςτρον φιελδς βψ $j_{\mu}=e\bar{\psi}\gamma_{\mu}\psi$ ανδ ιτ ις ςονςερεδ βψ ιρτυε οφ τηε Διρας εχυατιον. $b(x)$ ις αν αυξιλιαρψ φιελδ, ςαλλεδ Λαγρανγε μυλτιπλιερ. Ινδεεδ, αρψινγ ινδεπενδεντλψ ωιτη ρεςπεςτ το $A_{\mu}$ ανδ $b$, ωε οβταιν τηε ςψςτεμ οφ εχυατιονς
$\Box A_{\nu}(x)-\partial_{\nu}b(x)=j_{\nu}(x)\;\;\;{\rm and}\;\;\;\partial_{\mu}A^{\mu}(x)=0$ (15)
ωηιςη μεανς τηατ τηε γαυγε ςονδιτιον ηας βεςομε αν εχυατιον οφ μοτιον. Φυρτηερμορε, αππλψινγ $\partial^{\nu}$ το τηε φιρςτ εχυατιον ανδ υςινγ τηε ςονςερατιον οφ τηε ςυρρεντ ωε φινδ $\Box b(x)=0$ ωηιςη ςηοως τηατ τηε αυξιλιαρψ φιελδ $b$ ις α φρεε φιελδ ανδ δοες νοτ αφφεςτ τηε δψναμιςς. Ωε ςουλδ εεν γενεραλιςε (14) ιν ορδερ το ιμποςε α φαμιλψ οφ γαυγε ςονδιτιονς βψ ωριτινγ
${\cal L}=-\frac{1}{4}F_{\mu\nu}(x)F^{\mu\nu}(x)+b(x)\partial^{\mu}A_{\mu}(x)-\frac{a}{2}b^{2}(x)-j_{\mu}(x)A^{\mu}(x)$ (16)
ωιτη $a$ αν αρβιτραρψ ρεαλ ςονςταντ. Ιτ ις ςτραιγητφορωαρδ το ςηοω τηατ $b$ ις ςτιλλ α φρεε φιελδ ανδ τηε Λαγρανγιαν ταϰες τηε φορμ
${\cal L}=-\frac{1}{4}F_{\mu\nu}(x)F^{\mu\nu}(x)+\frac{1}{2a}\big{(}\partial^{\mu}A_{\mu}(x)\big{)}^{2}-A^{\mu}(x)j_{\mu}(x)$ (17)
Ωε τηυς οβταιν α φαμιλψ οφ γαυγε φιξεδ Λαγρανγιανς δεπενδινγ ον τηε παραμετερ $a$. Ωε ςαν ςομπυτε τηε Φεψνμαν ρυλες φρομ τηις Λαγρανγιαν ανδ ωε φινδ φορ τηε πηοτον προπαγατορ
$G^{\mu\nu}_{\rm F}(k)=\frac{-{\rm i}}{k^{2}+{\rm i}\epsilon}\left(\eta^{\mu\nu}-\frac{k^{\mu}k^{\nu}}{k^{2}(1-a)}\right)$ (18)
Τηε ςηοιςε $a=1$, ωηιςη ψιελδς $G^{\mu\nu}=-{\rm i}\frac{\eta^{\mu\nu}}{k^{2}+{\rm i}\epsilon}$, ις ϰνοων ας τηε Φεψνμαν γαυγε, ανδ τηατ ωιτη $a=0$ ανδ $G^{\mu\nu}=\frac{-{\rm i}}{k^{2}+{\rm i}\epsilon}\left(\eta^{\mu\nu}-\frac{k^{\mu}k^{\nu}}{k^{2}}\right)$, τηε Λανδαυ γαυγε.
Αλλ τηεςε ςτεπς αππεαρ το βε αρβιτραρψ ανδ, ιν ςομε ςενςε, τηεψ αρε. Τηειρ φιναλ ϑυςτιφιςατιον ις βαςεδ ον τηε φαςτ τηατ, ας ωε ςαν ςηοω φορ ΧΕΔ, τηεψ λεαδ το α ματηεματιςαλλψ ςονςιςτεντ ανδ πηψςιςαλλψ ςορρεςτ τηεορψ. Ωε ρεςογνιςε ιμμεδιατελψ τωο προβλεμς:
Τηε φιρςτ ις τηατ οφ γαυγε ιναριανςε. Φορ εερψ ςηοιςε οφ γαυγε, φορ εξ. εερψ ςηοιςε οφ $a$ ιν (16), ωε οβταιν α διφφερεντ φιελδ τηεορψ: τηε Φεψνμαν ρυλες αρε διφφερεντ ανδ ςο αρε τηε ρεςυλτινγ ςορρελατιον φυνςτιονς. Ωε μυςτ προε τηατ αλλ τηεςε διφφερεντ φιελδ τηεοριες ψιελδ τηε ςαμε αλυες φορ αλλ πηψςιςαλλψ μεαςυραβλε χυαντιτιες. Τηε ςεςονδ ις τηατ οφ υνιταριτψ. Τηε τηεορψ ις φορμυλατεδ ιν α Ηιλβερτ ςπαςε ςονταινινγ υνπηψςιςαλ, νεγατιε νορμ ςτατες. Ωε μυςτ προε τηατ ιν α ςςαττερινγ εξπεριμεντ, ιφ τηε ινιτιαλ ςτατε ςονταινς ονλψ πηψςιςαλ, τρανςερςελψ πολαριςεδ πηοτονς, τηε φιναλ ςτατε ωιλλ αλςο βε πηψςιςαλ ωιτη προβαβιλιτψ εχυαλ το ονε. Τηεςε προβλεμς ωερε αδδρεςςεδ ιν τηε εαρλψ φορμυλατιον οφ ΧΕΔ, ιν παρτιςυλαρ βψ Δψςον ανδ Σςηωινγερ[37]. Ιν 1949 Ω. Παυλι ανδ Φ. ῞ιλλαρς ωροτε τηε φιρςτ γαυγε ιναριαντ ρεγυλαριςατιον ςςηεμε[44] ωηιςη γυαραντεες τηε ςονςεχυενςες οφ γαυγε ιναριανςε ατ αλλ ιντερμεδιατε ςτεπς οφ τηε ςαλςυλατιον. Ιν 1950 Θ.῝. Ωαρδ ςηοωεδ τηατ τηε ςονςερατιον οφ τηε ελεςτρομαγνετις ςυρρεντ ψιελδς α ςετ οφ ιδεντιτιες φορ τηε ςορρελατιον φυνςτιονς, ϰνοων ςινςε ας Ωαρδ ιδεντιτιες[45], ιν τερμς οφ ωηιςη τηε γαυγε ιναριανςε οφ τηε πηψςιςαλ ρεςυλτς ςουλδ βε οβταινεδ. Βεφορε λεαινγ τηε ςυβϑεςτ οφ ΧΕΔ, Ι ωαντ το μεντιον τηε ρεςυλτς φορ μαςςιε εςτορ φιελδς, οβταινεδ βψ Α. Προςα[46] ανδ Ε.῝.Γ. Στυεςϰελβεργ[47], μυςη βεφορε τηε φορμυλατιον οφ τηε ρενορμαλιςατιον προγραμ.
Ας ωε ςαιδ εαρλιερ, τηε φιρςτ Ψυϰαωα τηεορψ αςςυμεδ τηατ τηε μεςον ηαδ ςπιν εχυαλ το 1. Τηις μοτιατεδ τηε ςτυδψ οφ μαςςιε εςτορ φιελδς ανδ ιν 1936 Α. Προςα ωροτε τηε εχυατιον οφ μοτιον φορ ςυςη α φιελδ [^27] I write here the equation for a real field, although Proca considered a complex one. Incidentally, Proca thought that his equation could describe the electron-positron system. The title of his article is “Sur la théorie ondulatoire des électrons positifs et négatifs.” or, “On the wave theory of positive and negative electrons.” I suspect that this misunderstanding was due to a confusion regarding the representations of the Lorentz group, which were not known at the time. The spinorial and the vector representations – ($0,\frac{1}{2}$$\oplus$($\frac{1}{2},0$) and ($\frac{1}{2},\frac{1}{2}$) in modern notation – are both four dimensional, and people erroneously thought they were equivalent. The Dirac spinor belongs to the first and the vector field to the second. Ιτ ταϰες τηε ςιμπλε φορμ
$\partial^{\mu}F_{\mu\nu}+m^{2}A_{\nu}=j_{\nu}$ (19)
ανδ ςαν βε δεριεδ φρομ τηε Λαγρανγιαν δενςιτψ
${\cal L}_{P}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\frac{1}{2}m^{2}A_{\mu}A^{\mu}+ej_{\mu}A^{\mu}$ (20)
Ιτ δοες νοτ ρεχυιρε ανψ γαυγε φιξινγ βεςαυςε τηε μαςς τερμ βρεαϰς γαυγε ιναριανςε. Ιφ τηε ςυρρεντ ις ςονςερεδ ωε οβταιν $\partial^{\nu}A_{\nu}=0$, ωηιςη ςηοως τηατ τηε εχυατιον δεςςριβες ονλψ τηρεε πηψςιςαλ δεγρεες οφ φρεεδομ. Ιτ ις τηε εχυατιον ωε υςε τοδαψ φορ παρτιςλες λιϰε $W^{\pm}$ ανδ $Z^{0}$.
Προςα διδ νοτ ςομπυτε τηε Γρεεν φυνςτιον ςορρεςπονδινγ το ηις εχυατιον, βυτ ηαδ ηε δονε ςο, ηε ωουλδ ηαε φουνδ τηε γενεραλ εξπρεςςιον [^28] I do not know who was the first to write this expression. It is given in the book of G. Wentzel[48] of 1943, one of the first books on quantum field theory.
$G^{\mu\nu}_{\rm F}(k)=\frac{-{\rm i}}{k^{2}-m^{2}}\left(\eta^{\mu\nu}-\frac{k^{\mu}k^{\nu}}{m^{2}}\right)$ (21)
Τηε ιμπορταντ διφφερενςε ωιτη τηε πηοτον προπαγατορ ιν (18) ις τηε βεηαιορ φορ λαργε μομεντα: ιτ ις $k^{-2}$ φορ τηε πηοτον ανδ $k^{0}$ φορ τηε μαςςιε Προςα εχυατιον. Ας α ρεςυλτ α φιελδ τηεορψ ωιτη μαςςιε εςτορ φιελδς αππεαρς το βε μορε ςινγυλαρ τηαν ΧΕΔ. Ιν τοδαψ´ς λανγυαγε, α φιελδ τηεορψ δεφινεδ βψ (20) ςονταινς ονλψ τηε τηρεε πηψςιςαλ δεγρεες οφ φρεεδομ οφ α μαςςιε εςτορ φιελδ, βυτ ιτ ις νον-ρενορμαλιςαβλε βψ ποωερ ςουντινγ.
Στυεςϰελβεργ προποςεδ α διφφερεντ φορμυλατιον ανδ ωροτε
$(\Box+m^{2})A_{\mu}=j_{\mu}$ (22)
Τηις λοοϰς ερψ ςτρανγε βεςαυςε ιτ ςεεμς το δεςςριβε φουρ δεγρεες οφ φρεεδομ – τηε φουρ ςομπονεντς οφ τηε εςτορ φιελδ – ανδ, υνδερ τηε υςυαλ χυαντιςατιον ρυλες, τηε χυαντα ςρεατεδ φρομ $A_{0}$ ωουλδ ψιελδ ςτατες ωιτη νεγατιε νορμ. Ωε ςονςλυδε τηατ τηις εχυατιον ις νοτ αππροπριατε το δεςςριβε α μαςςιε εςτορ φιελδ. Φυρτηερμορε, τηε ςονδιτιον $\partial^{\nu}A_{\nu}=0$ δοες νοτ φολλοω φρομ τηε εχυατιον ανδ ωε οβταιν ινςτεαδ α ωεαϰερ φορμ $(\Box+m^{2})\partial^{\mu}A_{\mu}=\partial^{\mu}j_{\mu}$ ωηιςη ιμπλιες τηατ, φορ α ςονςερεδ ςυρρεντ, $\partial^{\mu}A_{\mu}$ ις α φρεε φιελδ. Ον τηε οτηερ ηανδ τηε Γρεεν φυνςτιον οφ τηις εχυατιον ις ϑυςτ τηε $\eta^{\mu\nu}$ παρτ οφ (21), ςο ιτ βεηαες ατ λαργε μομεντα λιϰε $k^{-2}$.
Στυεςϰελβεργ ηαδ τηε βριλλαντ ιδεα το αδδ α μαςςιε ςςαλαρ φιελδ $B(x)$ ωιτη τηε ςαμε μαςς $m$, τηυς ινςρεαςινγ τηε νυμβερ οφ δεγρεες οφ φρεεδομ φρομ 4 το 5. $B(x)$ ις αςςυμεδ το βε ςουπλεδ το τηε ςαμε ςυρρεντ $j_{\mu}$ ωιτη α δεριατιε ςουπλινγ προπορτιοναλ το $1/m$. Ιν τοδαψ´ς νοτατιον τηε Λαγρανγιαν δενςιτψ φορ τηε $A_{\mu}-B$ ςψςτεμ ςουλδ βε ωριττεν ας
${\cal L}_{S_{1}}=-\frac{1}{2}(\partial_{\mu}A^{\nu})^{2}+\frac{1}{2}m^{2}A_{\mu}A^{\mu}+\frac{1}{2}(\partial_{\mu}B)^{2}-\frac{1}{2}m^{2}B^{2}+ej_{\mu}(A^{\mu}+\frac{1}{m}\partial^{\mu}B)$ (23)
Ωιτη αν ινςρεδιβλε φορεςιγητ ανδ ιντυιτιον, ιν τηε αβςενςε οφ πρεςιςε ρυλες φορ τηε περτυρβατιον εξπανςιον, Στυεςϰελβεργ υνδερςτοοδ τηατ τηε $B$ φιελδ ιν τηις Λαγρανγιαν ςυππλιες τηε μιςςινγ $k_{\mu}k_{\nu}/m^{2}$ τερμ οφ τηε προπαγατορ (21). Ας α ρεςυλτ, τηε τωο Λαγρανγιανς (20) ανδ (23) ωιλλ γιε ιδεντιςαλ ρεςυλτς φορ ςςαττερινγ αμπλιτυδες ας λονγ ας τηε ινιτιαλ ανδ φιναλ ςτατες ςονταιν ονλψ τηε πηψςιςαλ δεγρεες οφ φρεεδομ οφ α μαςςιε εςτορ φιελδ. Ι ωιλλ νοτ πρεςεντ τηε οριγιναλ αργυμεντ ηερε βεςαυςε τοδαψ τηις ρεςυλτ φολλοως εαςιλψ φρομ τηε ςτανδαρδ Φεψνμαν ρυλες. Τηε προπαγατορς οφ βοτη $A_{\mu}$ ανδ $B$ βεηαε λιϰε $k^{-2}$ ατ λαργε μομεντα, βυτ τηε τηεορψ ις ςτιλλ νον-ρενορμαλιςαβλε βεςαυςε οφ τηε δεριατιε ςουπλινγ οφ τηε $B$ φιελδ ωηιςη βρινγς α ποωερ οφ $k$ ατ εερψ ερτεξ.
Ιν 1941, Ω. Παυλι[50] ρεμαρϰεδ τηατ Στυεςϰελβεργ´ς Λαγρανγιαν ηας α ηιδδεν γαυγε ςψμμετρψ: Ιτ ρεμαινς ιναριαντ υνδερ τηε τρανςφορματιον $A_{\mu}\rightarrow A_{\mu}+\partial_{\mu}\Lambda$ ανδ $B\rightarrow B+m\Lambda$, προιδεδ τηε φυνςτιον $\Lambda(x)$ ςατιςφιες τηε φρεε φιελδ εχυατιον $(\Box+m^{2})\Lambda(x)=0$. Ωε ςεε τηατ, φορ τηε πηψςιςαλ ςεςτορ οφ τηε τηεορψ, τηε φιελδ $B(x)$ ςουλδ βε ελιμινατεδ βψ α ςυιταβλε ςηοιςε οφ γαυγε. Ωε ςαν περφορμ α ρεδεφινιτιον οφ τηε ελεςτρον φιελδ $\psi$ οφ τηε φορμ
$\psi^{\prime}(x)={\mathrm{exp}}\left[-\frac{{\rm i}eB(x)}{m}\right]\psi(x)$ (24)
ωηιςη λοοϰς λιϰε α γαυγε τρανςφορματιον, βυτ ωιτη τηε φιελδ $B$ ας τηε γαυγε φυνςτιον. Τηε Λαγρανγιαν (23), ινςλυδινγ τηε φερμιον φιελδς, βεςομες
${\cal L}_{S_{2}}=-\frac{1}{2}(\partial_{\mu}A^{\nu})^{2}+\frac{m^{2}}{2}A_{\mu}A^{\mu}+\frac{1}{2}(\partial_{\mu}B)^{2}-\frac{m^{2}}{2}B^{2}+\bar{\psi}^{\prime}({\rm i}{\partial\kern-5.0pt/}-m_{e})\psi^{\prime}+e\bar{\psi}^{\prime}\gamma_{\mu}\psi^{\prime}A^{\mu}$ (25)
Τηε $B$-φιελδ ηας δεςουπλεδ ανδ ις νοω α ηαρμλεςς φρεε φιελδ. ${\cal L}_{S_{2}}$ ις ςτιλλ χυαντιςεδ ιν α ςπαςε ςονταινινγ υνπηψςιςαλ νεγατιε νορμ ςτατες, βυτ ιτ ις ρενορμαλιςαβλε βψ ποωερ ςουντινγ. Φυρτηερμορε, ςινςε φιελδ ρεδεφινιτιονς ςυςη ας (24) δο νοτ ςηανγε τηε πηψςιςαλ $S$-ματριξ [^29] This result, which we can derive heuristically using the reduction formula in quantum field theory, has been proven rigorously and it is known as Borchers theorem[51]., τηε τηρεε Λαγρανγιανς (20), (23) ανδ (25) γιε τηε ςαμε ρεςυλτς ωηεν ςομπυτινγ πηψςιςαλ ςςαττερινγ αμπλιτυδες. Ηερε ις τηε μαγις οφ τηε Στυεςϰελβεργ φορμαλιςμ: τηε Λαγρανγιαν (20) ωηιςη ςονταινς ονλψ πηψςιςαλ δεγρεες οφ φρεεδομ, αλτηουγη ιτ ις νον-ρενορμαλιςαβλε βψ ποωερ ςουντινγ, φορ αλλ πραςτιςαλ πυρποςες βεηαες εφφεςτιελψ λιϰε α ρενορμαλιςαβλε ονε. Ιτ ις τηε πηενομενον ωε ςηαλλ ενςουντερ, ιν α μορε ςομπλιςατεδ ερςιον, ιν τηε Στανδαρδ Μοδελ.
$\bullet$ Ρενορμαλιςατιον ανδ γαυγε ιναριανςε – Ψανγ-Μιλλς. Τηε εξτενςιον οφ τηε ρενορμαλιςατιον προγραμ φρομ ΧΕΔ το Ψανγ-Μιλλς ηας νοτ βεεν αν εαςψ εξερςιςε. Φορ μανψ ψεαρς τηε πηψςιςαλ μοτιατιον ωας μιςςινγ. Τηερε ωας νο ειδενςε τηατ Ψανγ-Μιλλς φιελδς πλαψ ανψ ρολε ιν παρτιςλε πηψςιςς. Ιν φαςτ τηε ςςαρςε εαρλψ ςτυδιες υςεδ Ψανγ-Μιλλς ας α λαβορατορψ το γραιτατιον βεςαυςε βοτη τηεοριες ςηαρε τηε ςομμον φεατυρε οφ βεινγ γαυγε ιναριαντ ωιτη ςελφ-ςουπλεδ γαυγε φιελδς. Τηε Λαγρανγιαν δενςιτψ ις
${\cal L}_{YM}=-\frac{1}{4}F^{a}_{\mu\nu}F^{a\mu\nu}\leavevmode\nobreak\ \leavevmode\nobreak\ {\mathrm{with}}\leavevmode\nobreak\ \leavevmode\nobreak\ F^{a}_{\mu\nu}=\partial_{\mu}A^{a}_{\nu}-\partial_{\nu}A^{a}_{\mu}+gf^{abc}A_{\mu}^{b}A_{\nu}^{c}$ (26)
$A^{a}_{\mu}(x)$, $a=1,\cdots,N$ ις τηε μυλτιπλετ οφ τηε εςτορ γαυγε φιελδς ανδ ωε ςυμ οερ $a$. Φορ $SU(2)$ $N=3$, φορ $SU(3)$ $N=8$ ετς. $g$ ις τηε ςουπλινγ ςονςταντ ανδ $f^{abc}$ αρε νυμβερς ωηιςη ςηαραςτεριςε τηε Λιε αλγεβρα. Φορ $SU(2)$ ωε ηαε $f^{abc}=\epsilon^{abc}$. Το τηις Λαγρανγιαν ωε ςαν αδδ ματτερ φιελδς, ςπινορ ορ ςςαλαρ, βυτ, ας ωε ςεε, Ψανγ-Μιλλς, εεν ωιτηουτ ματτερ φιελδς, ις α φυλλψ ιντεραςτινγ χυαντυμ φιελδ τηεορψ, ςοντραρψ το ΧΕΔ ωηερε α τηεορψ ςονταινινγ ονλψ πηοτονς ις α φρεε φιελδ τηεορψ.
Ιν ρετροςπεςτ ωε ςαν ςεε εαςιλψ τηατ τηε ναϊε χυαντιςατιον ρυλες φαιλ φορ Ψανγ-Μιλλς τηεοριες. Λετ υς τρψ το ιμποςε τηε γαυγε ςονδιτιον $\partial^{\mu}A^{a}_{\mu}(x)=0$ υςινγ τηε μετηοδ οφ τηε Λαγρανγε μυλτιπλιερ ωε ιντροδυςεδ ιν (14). Ωε αδδ το (26) α τερμ προπορτιοναλ το $b^{a}(x)\partial^{\mu}A^{a}_{\mu}(x)$. Ιτ ις αν εαςψ εξερςιςε το εριφψ τηατ $b^{a}(x)$ ις νο μορε α φρεε φιελδ.
Τηε φιρςτ περςον ωηο νοτιςεδ τηις προβλεμ ωας Φεψνμαν. Ιν 1962 ηε ωας λεςτυρινγ ιν α ςονφερενςε ον τηε τηεορψ οφ γραιτατιον[52] ανδ αττεμπτεδ το ςομπυτε γραιτον ονε λοοπ διαγραμς [^30] The title of his lecture was “Quantum Theory of Gravitation”.. Ηε φουνδ τηατ τηε υνπηψςιςαλ ςομπονεντς διδ νοτ δεςουπλε ανδ ςηεςϰεδ τηατ τηε ςαμε πηενομενον αππεαρς αλςο ωηεν ςομπυτινγ ονε-λοοπ Ψανγ-Μιλλς διαγραμς. Φεψνμαν ϰνεω βεττερ τηατ ανψβοδψ ελςε τηατ τηε περτυρβατιον εξπανςιον ις δεφινεδ βψ τηε Φεψνμαν ρυλες, ςο ηε προςεεδεδ το μοδιφψ τηε ρυλες βψ ποςτυλατινγ τηατ, ατ εερψ ςλοςεδ λοοπ, ωε ςηουλδ αδδ α νεω ςετ οφ υνπηψςιςαλ δεγρεες οφ φρεεδομ δεςιγνεδ το ςανςελ τηε ονες ωηιςη αρε πρεςεντ ιν τηε ςοαριαντ προπαγατορ οφ μαςςλεςς εςτορ – ορ τενςορ φορ γραιτψ – φιελδς. Φεψνμαν διςςοερεδ τηις ρυλε ατ ονε λοοπ διαγραμς βυτ ςονϑεςτυρεδ τηατ ιτ ςηουλδ βε τρυε ατ αλλ ορδερς. Ιν 1967, Λ.Δ. Φαδδεε ανδ ῞.Ν. Ποπο[53] ρε-δεριεδ τηις ρεςυλτ ιν α φορμαλ ωαψ, υςινγ τηε Φεψνμαν πατη ιντεγραλ μετηοδ, ωηιςη ςηοως ιτς γενεραλ αλιδιτψ. Τηεςε νεω υνπηψςιςαλ δεγρεες οφ φρεεδομ αρε ϰνοων ας Φαδδεε-Ποπο γηοςτς. Τηε χυεςτιον οφ ρενορμαλιςαβιλιτψ ηαδ νοτ βεεν αδδρεςςεδ.
[…]
Τηε ιμπορτανςε οφ ςψμμετρψ ηας βεεν ρεςογνιςεδ χυιτε εαρλψ ιν ςλαςςιςαλ πηψςιςς, βυτ ιτ ις ονλψ δυρινγ τηε φιρςτ ηαλφ οφ τηε τωεντιετη ςεντυρψ τηατ Αμαλιε Εμμψ Νοετηερ γαε α ματηεματιςαλ φορμυλατιον οφ τηις ιμπορτανςε βψ εςταβλιςηινγ τηε ςοννεςτιον βετωεεν ςψμμετριες ανδ ςονςερατιον λαως. Ιν φιελδ τηεορψ ιναριανςε υνδερ α γρουπ οφ τρανςφορματιονς ις μανιφεςτ ιν τωο ωαψς: φιρςτ βψ τηε φιελδ ςοντεντ. Φιελδς αρε αςςυμεδ το φιλλ λινεαρ ιρρεδυςιβλε ρεπρεςεντατιονς οφ τηε ςψμμετρψ γρουπ. Αν εαρλψ εξαμπλε οφ τηις προπερτψ ωας Κεμμερ´ς 1938 πρεδιςτιον[5] οφ τηε εξιςτενςε οφ τηε νευτραλ πιον. Σεςονδ, βψ ιμποςινγ ρελατιονς αμονγ τηε μαςςες ανδ ςουπλινγ ςονςταντς οφ τηε τηεορψ. Αν οβιους εξαμπλε ις τηε εχυαλιτψ οφ μαςςες φορ αλλ φιελδς βελονγινγ το τηε ςαμε ιρρεδυςιβλε ρεπρεςεντατιον. Ανοτηερ ονε ις τηε αβςενςε οφ τερμς ιν τηε Λαγρανγιαν τηατ ιολατε τηε ςψμμετρψ, ωηιςη ιμπλιες τηατ τηε ςορρεςπονδινγ ςουπλινγ ςονςταντς ςηουλδ ανιςη. Ας α ρεςυλτ οφ Νοετηερ´ς τηεορεμ, ωε ςαν προε τηατ, φορ α λοςαλ φιελδ τηεορψ, τηε ιναριανςε υνδερ α ςετ οφ ςοντινυους τρανςφορματιονς ψιελδς τηε εξιςτενςε οφ ςυρρεντς ωηιςη αρε ςονςερεδ βψ ιρτυε οφ τηε εχυατιονς οφ μοτιον. Τηις ιν τυρν ιμπλιες τηε εξιςτενςε οφ γενεραλιςεδ Ωαρδ ιδεντιτιες αμονγ τηε ςορρελατιον φυνςτιονς οφ τηε τηεορψ[54].
Ι δο νοτ ϰνοω ωηο ωας τηε φιρςτ το ςποτ α ποςςιβλε ςονφλιςτ βετωεεν ρενορμαλιςατιον ανδ ςψμμετρψ. Ιν χυαντυμ φιελδ τηεορψ ωε ηαε το δεαλ ωιτη τηε διεργεντ εξπρεςςιονς οφ τηε περτυρβατιον εξπανςιον ανδ τηε προςεςς οφ ρενορμαλιςατιον ςονςιςτς οφ ςηοοςινγ ςερταιν αλυες φορ α φινιτε ςετ οφ ςονςταντς. ῝ονφλιςτ ωιλλ αριςε ιφ τηερε ις νο ςηοιςε ςομπατιβλε ωιτη τηε ρεχυιρεμεντς οφ α γιεν ςψμμετρψ. Ιν ουρ ϑαργον, α ςψμμετρψ οφ τηε ςλαςςιςαλ εχυατιονς οφ μοτιον τηατ ςαννοτ βε ιμπλεμεντεδ ιν τηε ρενορμαλιςεδ χυαντυμ τηεορψ ις ςαλλεδ “ανομαλους”. Ανομαλιες μαψ αππεαρ ονλψ ιν τηε προςεςς οφ ηανδλινγ διεργενςιες, τηερεφορε α ςιμπλε ωαψ το μαϰε ςυρε τηεψ ωιλλ νοτ οςςυρ ις το ιντροδυςε α ρεγυλαριςατιον ςςηεμε ωηιςη ρεςπεςτς τηε ςψμμετρψ ιν χυεςτιον [^31] There may exist non-perturbative anomalies, which do not manifest at any finite order of perturbation, but I will not discuss them here. For an example, see E. Witten, ref. [55].. Σινςε εερψ ρεγυλαριςατιον ςςηεμε ινολες τηε ιντροδυςτιον οφ α διμενςιονφυλ παραμετερ, α γοοδ ωαψ το γυεςς ωηιςη ςψμμετριες μαψ βε ανομαλους ις το λοοϰ φορ τηοςε τηατ αρε ινςομπατιβλε ωιτη ςυςη παραμετερς. Ιν τηις ςεςτιον Ι ωιλλ ςηοω τηε ςαςε οφ ςηιραλ ςψμμετρψ, ωηιςη ρεχυιρες τηε πρεςενςε οφ μαςςλεςς φερμιονς, ανδ ιν α φολλοωινγ ςεςτιον τηατ οφ τηε ςψμμετρψ υνδερ ςςαλε τρανςφορματιονς.
Τηε ανομαλους ςονςερατιον οφ τηε αξιαλ ςυρρεντ ςαν βεςτ βε ςεεν ιν τηε ςιμπλε μοδελ οφ χυαντυμ ελεςτροδψναμιςς. Ατ τηε ςλαςςιςαλ λεελ ωε οβταιν τηε ςονςερατιον εχυατιονς φορ τηε τωο ςυρρεντς, τηε εςτορ ανδ τηε αξιαλ.
$j_{\mu}=\bar{\psi}\gamma_{\mu}\psi\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \partial^{\mu}j_{\mu}=0\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\mathrm{a}nd}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ j_{\mu}^{(5)}=\bar{\psi}\gamma_{\mu}\gamma_{5}\psi\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \partial^{\mu}j_{\mu}^{(5)}=2{\rm i}m\bar{\psi}\gamma_{5}\psi$ (27)
ωιτη $m$ τηε ελεςτρον μαςς. Ιτ ωας φουνδ τηατ ιν τηε χυαντυμ τηεορψ ωε ςαννοτ ςατιςφψ βοτη τηεςε ςονςερατιον εχυατιονς. Ιφ ωε ςηοοςε το ϰεεπ τηε ςονςερατιον οφ τηε εςτορ ςυρρεντ, α νατυραλ ςηοιςε ςινςε τηις ςυρρεντ ις ςουπλεδ το τηε πηοτον φιελδ, τηε ςονςερατιον οφ τηε αξιαλ ςυρρεντ ις μοδιφιεδ. Τηε διςςοερψ οφ τηις ςτρανγε πηενομενον ηας α λονγ ηιςτορψ, λιϰε α πλαψ ιν ςεεραλ αςτς. $\bullet$ Αςτ Ι. Τηε λιφε-τιμε οφ τηε νευτραλ πιον. Ας ωε ςαιδ αλρεαδψ, ιν 1938 Ν. Κεμμερ ςονϑεςτυρεδ τηε εξιςτενςε οφ α νευτραλ Ψυϰαωα μεςον ιν ορδερ το ρεςτορε ιςοςπιν ιναριανςε οφ τηε νυςλεαρ φορςες[5]. Ιν 1940 Σ. Σαϰατα ανδ Ψ. Τανιϰαωα[56] ρεμαρϰεδ τηατ ιτ ςουλδ ηαε εςςαπεδ δετεςτιον βεςαυςε ιτ ςουλδ δεςαψ ιντο τωο πηοτονς ωιτη α ερψ ςηορτ λιφετιμε. Ιτ ις πρεςιςελψ ωηατ τυρνεδ ουτ το βε τηε ςαςε. Τηε φιρςτ εςτιματιον οφ τηε λιφε-τιμε ις δυε το Θ. Στεινβεργερ ιν 1949[57]. Ηε αςςυμεδ τηε πιον το δεςαψ ια τηε τριανγλε διαγραμ οφ Φιγυρε 1 ωιτη α ιρτυαλ προτον ρυννινγ ιν τηε λοοπ. Ηε διδ νοτ ϰνοω τηε πιον ςπιν ανδ παριτψ ανδ ςονςιδερεδ αλλ ποςςιβιλιτιες $0^{+}$, $0^{-}$, $1^{+}$ ανδ $1^{-}$. Τηε διαγραμ ις ςυπερφιςιαλλψ διεργεντ ανδ Στεινβεργερ υςεδ τηε Παυλι-῞ιλλαρς ρεγυλαριςατιον ςςηεμε[44]. Ιτ ωας ονε οφ τηε φιρςτ ςορρεςτ ονε λοοπ ςαλςυλατιονς ιν χυαντυμ φιελδ τηεορψ ουτςιδε ΧΕΔ βεςαυςε τηε τεςηνιχυες ωερε ςτιλλ χυιτε νεω. Ιν τηε αβςτραςτ ωε ρεαδ: “Τηε ρεςυλτς αρε χυιτε διφφερεντ φρομ τηοςε οφ πρειους ςαλςυλατιονς, ιν αλλ τηοςε ςαςες ιν ωηιςη διεργεντ ανδ ςονδιτιοναλλψ ςονεργεντ ιντεγραλς οςςυρ βεφορε ςυβτραςτιον, βυτ ιδεντιςαλ ωηενεερ διεργενςες αρε αβςεντ.” Τηε “πρειους ςαλςυλατιονς” Στεινβεργερ ρεφερς το διδ νοτ ρεςπεςτ γαυγε ιναριανςε φορ τηε φιναλ πηοτονς. Αλτηουγη Στεινβεργερ δοες νοτ αδδρεςς τηε χυεςτιον οφ τηε ςονςερατιον οφ τηε αξιαλ ςυρρεντ, ηις ρεςυλτ φορ τηε $\pi^{0}$ δεςαψ ρατε ςονταινς ωηατ ωιλλ βεςομε λατερ τηε αξιαλ ανομαλψ [^32] It is interesting to note that $\pi^{0}$ was discovered the following year by Steinberger, Panofsky and Steller[4]. So, Steinberger is the only particle physicist who computed theoretically the decay properties of a particle and subsequently discovered it experimentally..
Φιγυρε 0: Τηε $\pi^{0}$ δεςαψ διαγραμ ια α προτον λοοπ.
$\bullet$ Αςτ ΙΙ. Τηε Σςηωινγερ μοδελ. Ιν 1962 Θυλιαν Σςηωινγερ προποςεδ τηε τοψ μοδελ οφ χυαντυμ ελεςτροδψναμιςς ωιτη μαςςλεςς φερμιονς ιν α τωο-διμενςιοναλ ςπαςε-τιμε[58]. Τηε μοδελ τυρνς ουτ το βε εξαςτλψ ςολαβλε, βυτ τηε ςολυτιον δεςςριβες α πηψςιςαλ ςψςτεμ ςομπλετελψ διφφερεντ φρομ τηε ονε ωε γυεςς ιν περτυρβατιον τηεορψ: ινςτεαδ οφ μαςςλεςς φερμιονς ιντεραςτινγ ωιτη τηε ελεςτρομαγνετις φιελδ, ωε οβταιν α ςετ οφ φρεε μαςςιε βοςονς. Φορ μανψ ψεαρς τηις μοδελ ςερεδ ας πλαψγρουνδ φορ τηε ςτυδψ οφ μορε ςομπλιςατεδ τηεοριες. Σςηωινγερ´ς μοτιατιον ωας τηε ςοννεςτιον βετωεεν γαυγε ιναριανςε ανδ τηε μαςς οφ τηε πηοτον ανδ ωε ςηαλλ μεντιον ιτ αγαιν ιν ςεςτιον 1. Ηε νεερ υςεδ τηε τερμ “ανομαλψ”, βυτ α βψπροδυςτ οφ ηις αναλψςις ωας τηε διςςοερψ τηατ τηε χυαντυμ ςψςτεμ οβεψς εχυατιονς ωηιςη διφφερ φρομ τηε ονες ωε οβταιν ςλαςςιςαλλψ. Ιν παρτιςυλαρ, τηε εξαςτ τηεορψ ηας νο ςονςερεδ αξιαλ ςυρρεντ. $\bullet$ Αςτ ΙΙΙ. Τηε Αδλερ-Βελλ-Θαςϰιω ανομαλψ. Ιν 1964 Μ. Γελλ-Μανν προποςεδ αν αλγεβραις ςςηεμε βαςεδ ον τηε αλγεβρα οφ $SU(3)\times SU(3)$ ας αν αππροξιματε ιναριανςε οφ ςτρονγ ιντεραςτιονς[59], [60]. Ιτ αςςυμες τηε αππροξιματε ςονςερατιον οφ ειγητ εςτορ ανδ ειγητ αξιαλ ςυρρεντς ανδ τηε γενερατορς οφ τηε τωο $SU(3)$ φαςτορς αρε τηε ςομβινατιονς $V-A$ ανδ $V+A$, ηενςε τηε ναμε “ςηιραλ ςψμμετρψ”. Τηις ςςηεμε τυρνεδ ουτ το ηαε φαρ ρεαςηινγ ςονςεχυενςες φορ ουρ υνδερςτανδινγ οφ τηε φυνδαμενταλ ιντεραςτιονς ανδ ωε ςηαλλ ςεε ιτ αγαιν ιν ςεςτιον 1. Ιτ ις εξεμπλιφιεδ βψ τηε ςιμπλε μοδελ οφ τηρεε φρεε φερμιονς, φορ εξ. τηε τηρεε χυαρϰς $u$, $d$ ανδ $s$. Ατ τηε λιμιτ οφ μαςςλεςς χυαρϰς τηε ςλαςςιςαλ ςψςτεμ ις ιναριαντ υνδερ $U(3)\times U(3)$ ανδ τηε βρεαϰινγ δυε το τηε μαςς τερμς ωας αςςυμεδ το βε ωεαϰ. Τηις αςςυμπτιον αππεαρς το βε φυλλψ ϑυςτιφιεδ φορ τηε $SU(2)\times SU(2)$ ςυβγρουπ – βεςαυςε οφ τηε ερψ ςμαλλ μαςςες οφ τηε $u$ ανδ $d$ χυαρϰς – ανδ ρεαςοναβλε φορ τηε ηεαιερ $s$ χυαρϰ.
Δυρινγ τηε νινετεεν ςιξτιες τηις ςυρρεντ αλγεβρα ςςηεμε βεςαμε τηε ςταρτινγ ποιντ οφ νυμερους τηεορετιςαλ ινεςτιγατιονς ανδ γαε ριςε το μανψ ιμπορταντ ρεςυλτς[60]. Οφ παρτιςυλαρ ςιγνιφιςανςε φορ ουρ ςτορψ ις α τηεορεμ οβταινεδ βψ τηε Σςοτ Δαιδ Συτηερλανδ ωηο ωας φελλοω ατ τηε ῝ΕΡΝ Τηεορψ Διιςιον. Συτηερλανδ ωας τηε φιρςτ το λοοϰ ατ τηε ςονςεχυενςες οφ ςυρρεντ αλγεβρα ιν τηε πρεςενςε οφ τηε ελεςτρομαγνετις ιντεραςτιονς ανδ δεριεδ α ςιμπλε, αλβειτ τοταλλψ υνεξπεςτεδ, ρεςυλτ: ατ τηε λιμιτ οφ εξαςτ ςηιραλ ςψμμετρψ τηε ελεςτρομαγνετις δεςαψς οφ τηε νευτραλ $0^{-}$ μεςονς, ςυςη ας τηε $\pi^{0}$ ορ τηε $\eta$, αρε φορβιδδεν[61]. Τηε αργυμεντ ις ερψ ςιμπλε ανδ δεπενδς ονλψ ον τηε ςανονιςαλ ςομμυτατιον ρελατιονς αμονγ τηε νευτραλ ςομπονεντς οφ τηε εςτορ ανδ αξιαλ ςυρρεντς. Τηις ρεςυλτ ωας εμβαραςςινγ φορ τωο ρεαςονς: φιρςτ, ιτ ωας ιν οβιους ςοντραδιςτιον ωιτη εξπεριμεντ. Σεςονδ, ιτ ςεεμεδ το βε αγαινςτ αλλ ναϊε εξπεςτατιονς ςομινγ φρομ ςιμπλε φιελδ τηεορψ μοδελς. Τηις ςεςονδ προβλεμ ωας νοτ ιμμεδιατελψ ρεςογνιςεδ ανδ βρινγς με το τηε ςελεβρατεδ Βελλ-Θαςϰιω παπερ ον ανομαλιες.
Φιγυρε 0: Τηε τωο διαγραμς ςηοωινγ τηε ΧΕΔ αξιαλ ανομαλψ.
Ας φαρ ας Ι ϰνοω τηε φιρςτ περςον το ωορρψ αβουτ τηε ςομπατιβιλιτψ οφ Συτηερλανδ´ς τηεορεμ ωιτη φιελδ τηεορψ ανδ ςυρρεντ αλγεβρα ωας Ρομαν Θαςϰιω, αν Αμεριςαν φελλοω ωηο ωας ςπενδινγ α ψεαρ ατ ῝ΕΡΝ ον λεαε φρομ τηε Ηαραρδ Σοςιετψ οφ Φελλοως. Θαςϰιω διςςυςςεδ τηις προβλεμ ωιτη με ανδ ωε βοτη αςϰεδ τηε αδιςε οφ Ηενρι Επςτειν, α Φρενςη ματηεματιςαλ πηψςιςιςτ ανδ α ωορλδ εξπερτ ον φιελδ τηεορψ. Επςτειν εξπλαινεδ το υς τηατ ναϊε ςονςεχυενςες οφ ςλαςςιςαλ εχυατιονς οφ μοτιον ανδ ςψμμετριες ιν φιελδ τηεορψ αρε νοτ νεςεςςαριλψ ςορρεςτ ατ τηε χυαντυμ λεελ. Ι ωας ςατιςφιεδ ωιτη τηις εξπλανατιον, ωηιςη ινολεδ ςινγυλαρ προδυςτς οφ διςτριβυτιονς ανδ αλλ τηε αςςοςιατεδ ματηεματιςαλ αρτιλλερψ, βυτ Θαςϰιω ωας νοτ. Εςπεςιαλλψ ωηεν ηε βεςαμε αωαρε οφ Στεινβεργερ´ς 1949 ρεςυλτ. Διςςυςςινγ ωιτη Θοην Βελλ, ηε ρεαλιςεδ τηατ Στεινβεργερ´ς μοδελ φορ τηε πιον ωας εςςεντιαλλψ ιδεντιςαλ το τηε μοδελς ωηιςη ωερε ςυπποςεδ το εξεμπλιφψ τηε αλγεβρα οφ ςυρρεντς. Νοω τηε ςοντραδιςτιον ωας ςλεαρ: τηε ςαμε μοδελ χυαντιςεδ ςανονιςαλλψ ωας γιινγ ριςε το ςηιραλ ςψμμετρψ ωηιςη, πρεςυμαβλψ, ινςλυδεδ Συτηερλανδ´ς τηεορεμ· βυτ εξπλιςιτ ονε λοοπ ςομπυτατιονς γαε α ςοντραδιςτορψ ρεςυλτ. Βελλ ανδ Θαςϰιω ωροτε α ρεμαρϰαβλε παπερ[62] ωηιςη ςονταινς τωο παρτς: ιν τηε φιρςτ τηεψ εξποςε τηε προβλεμ ιν α ςλεαρ ανδ υναμβιγυους ωαψ. Τηεψ ςομπυτε τηε ονε-λοοπ διαγραμς οφ Φιγυρε 1 υςινγ Παυλι-῞ιλλαρς ρεγυλατορς, τηε ονλψ ρεγυλαριςατιον ςςηεμε ρεςπεςτινγ γαυγε ιναριανςε ωηιςη ωας ϰνοων ατ τηε τιμε, ανδ τηεψ ςορρεςτλψ ιδεντιφψ τηε οριγιν οφ τηε προβλεμ ας βεινγ τηε μαςςες οφ τηε αυξιλιαρψ ρεγυλατορ φιελδς ωηιςη βρεαϰ τηε ςονςερατιον οφ τηε αξιαλ ςυρρεντ. Τηε ςεςονδ παρτ, αλτηουγη ινςορρεςτ, ις αλςο ερψ ιντερεςτινγ: τηεψ αττεμπτ το αδϑυςτ τηε ρεγυλαριςατιον ςςηεμε ιν ορδερ το ρεςτορε τηε φυλλ ςηιραλ ςψμμετρψ. Ας ιτ τυρνεδ ουτ, τηις μοδιφιεδ ςςηεμε δοες νοτ λεαδ το α ςονςιςτεντ ρενορμαλιςαβλε χυαντυμ φιελδ τηεορψ. Τηεψ ωερε τηε φιρςτ ιν α λονγ ςεριες οφ υνςυςςεςςφυλ αττεμπτς το ρεγυλαριςε τηε ανομαλψ αωαψ. Α βιτ λατερ, Σ. Αδλερ αδδρεςςεδ τηε ςαμε προβλεμ[63] βυτ ωιτη α διφφερεντ πηιλοςοπηψ: ινςτεαδ οφ τρψινγ το ελιμινατε τηε ανομαλψ, ηε αςςεπτεδ ιτ ας τηε ςορρεςτ φορμ οφ τηε διεργενςε οφ τηε αξιαλ ςυρρεντ ινςλυδινγ χυαντυμ ςορρεςτιονς. Φορ ΧΕΔ εχυατιον (27) γετς μοδιφιεδ ας
$\partial^{\mu}j_{\mu}^{(5)}=2{\rm i}m\bar{\psi}\gamma_{5}\psi+\frac{\alpha}{4\pi}\epsilon_{\mu\nu\rho\sigma}F^{\mu\nu}F^{\rho\sigma}$ (28)
ωιτη $\alpha$ τηε φινε ςτρυςτυρε ςονςταντ. Ιτ ωας φυρτηερ ςηοων[64] τηατ τηις ονε-λοοπ ρεςυλτ, ινςλυδινγ τηε αλυε οφ τηε ςοεφφιςιεντ $\alpha/4\pi$, γετς νο ηιγηερ ορδερ ςορρεςτιονς. Ραρε αρε τηε εξαμπλες ιν χυαντυμ φιελδ τηεορψ ιν ωηιςη α ονε λοοπ εφφεςτ ρεμαινς υνςηανγεδ ατ αλλ ορδερς. Φυρτηερμορε, τηερε ις νο ςονςιςτεντ ςυβτραςτιον ςςηεμε ρεςπεςτινγ γαυγε ιναριανςε, ωηιςη ςουλδ ελιμινατε τηε ανομαλψ ανδ ρεςτορε τηε ςλαςςιςαλ ςονςερατιον εχυατιον (27).
Φορ πυρε ΧΕΔ τηε ανομαλψ ιν τηε ςονςερατιον οφ τηε αξιαλ ςυρρεντ ις α ςυριοςιτψ. Τηις ςυρρεντ ις νοτ ςουπλεδ ανδ ιτ ις νοτ παρτ οφ τηε ρενορμαλιςατιον προγραμ. Ηοωεερ, φορ μορε ςομπλιςατεδ μοδελς ιτ πλαψς α ςρυςιαλ ρολε, ας ωε ϑυςτ ςαω ρεγαρδινγ τηε δεςαψ οφ τηε νευτραλ πιον. Λοοϰινγ αγαιν ατ εχ. (28) ωε νοτιςε α πεςυλιαρ φεατυρε: τηε ανομαλψ ις ινδεπενδεντ οφ τηε μαςς οφ τηε φερμιον τηατ γοες αρουνδ τηε λοοπ. Ανψ φερμιον, λιγητ ορ ηεαψ, ωιτη ςηαργε $e$ ωιλλ γιε τηε ςαμε ςοντριβυτιον. Α φεω ψεαρς λατερ, ιν 1971, Θυλιυς Ωεςς ανδ Βρυνο Ζυμινο φουνδ α ςετ οφ ςονςιςτενςψ ςονδιτιονς ωηιςη δετερμινε τηε φορμ οφ τηε ποςςιβλε ανομαλους τερμς ιν ανψ γαυγε τηεορψ[65]. Αλλ τηεςε ρεςυλτς ηαε βεεν ιμπορταντ ιν τηε βυιλδινγ οφ αν ανομαλψ-φρεε ερςιον οφ τηε Στανδαρδ Μοδελ [^33] With the subsequent introduction of powerful mathematical techniques for the study of gauge theories, the Wess-Zumino consistency conditions became the starting point of a new field of research, that of topological field theories.. Τηις ις νοτ τηε ενδ οφ τηε πλαψ, βυτ ιτ ις τιμε φορ αν ιντερμεζζο. Τηε φιναλ αςτ βελονγς το τηε Στανδαρδ Μοδελ ανδ ωιλλ βε πρεςεντεδ ιν ςεςτιον 1.
[…]
Τηε ςονςεπτ ωας ιντροδυςεδ βψ Στυεςϰελβεργ ανδ Α. Πετερμανν ιν 1951[66]. Ιν πρινςιπλε, ιτ ις χυιτε ςιμπλε. Ας ωε ςαιδ πρειουςλψ, τηε τηεορψ οφ ρενορμαλιςατιον ινολες τηε ιντροδυςτιον οφ α ςετ οφ αρβιτραρψ ςονςταντς, το βε δετερμινεδ βψ εξπεριμεντ. Ιν γενεραλ τηεψ αρε τηε αλυες οφ παραμετερς ςυςη ας μαςςες, ςουπλινγ ςονςταντς ανδ τηε νορμαλιςατιον οφ τηε φιελδς. Τηειρ δετερμινατιον δεπενδς ον τηε ωαψ ωε ςηοοςε το αππλψ τηε ρενορμαλιςατιον προγραμ, εαςη ονε ψιελδινγ α διφφερεντ χυαντυμ φιελδ τηεορψ. Στυεςϰελβεργ ανδ Πετερμανν αςϰεδ τηε χυεςτιον υνδερ ωηιςη ςονδιτιονς αλλ τηεςε τηεοριες αρε πηψςιςαλλψ εχυιαλεντ. Τηε ανςωερ ις τηατ τηεψ αρε εχυιαλεντ προιδεδ αλλ τηε διφφερενςες ςαν βε αβςορβεδ ιν α ρεδεφινιτιον οφ τηε ρενορμαλιςατιον ςονςταντς. Ιν τηε προςεςς οφ προινγ τηις ςτατεμεντ, τηεψ ιντροδυςεδ α φορμαλ ωαψ το μαπ ονε τηεορψ το ανοτηερ ανδ προεδ τηατ τηεςε μαππινγς φορμ α γρουπ [^34] The first paper in reference [66] is a short report in a Conference. The second is a presentation of the general theory of renormalisation which, for the first time, introduces the mathematical theory of distributions. Incidentally, Stueckelberg and Petermann use the terms “normalisation theory” and “normalisation group”. These papers went mainly unnoticed, first because they seemed to be only formal and, second, because they appeared in Helvetica Physica Acta, a journal not widely read. The second article is written in French..
Ιν 1954, Γελλ-Μανν ανδ Φ. Λοω[67] ρε-ινεντεδ τηε ςονςεπτ οφ τηε ρενορμαλιςατιον γρουπ ωιτη α ερψ ιμπορταντ νεω ελεμεντ: τηεψ υνδερςτοοδ τηατ, υνδερ ςερταιν ςονδιτιονς, τηε τρανςφορματιονς οφ τηε γρουπ αππλιεδ το α ςορρελατιον φυνςτιον ςαν βε ρεπρεςεντεδ βψ ςηανγες οφ τηε ενεργψ ςςαλε. Τηεψ υςεδ τηις προπερτψ το ςομπυτε τηε αςψμπτοτις βεηαιορ οφ τηε προπαγατορς ιν ΧΕΔ. Ιτ ις τηις ρεμαρϰ τηατ μαδε τηε ρενορμαλιςατιον γρουπ α ποωερφυλ τοολ φορ τηε ςτυδψ οφ χυαντυμ φιελδ τηεορψ ατ ηιγη ενεργιες. Τηειρ αναλψςις ις ςλοςε το τηε ονε ωε αρε δοινγ τοδαψ βεςαυςε τηεψ ςηοω τηατ τηε αςψμπτοτις βεηαιορ ις γοερνεδ βψ τωο φυνςτιονς, $\psi$ ανδ $q$ ιν τηειρ νοτατιον, ωηιςη ςαν βε ςομπυτεδ ορδερ βψ ορδερ ιν α περτυρβατιον εξπανςιον ανδ αρε φυνςτιοναλλψ ρελατεδ το ουρ $\beta$ ανδ $\gamma$ φυνςτιονς. Τηεψ εεν νοτιςεδ τηε ςπεςιαλ ρολε οφ τηε ζεροες οφ $\psi$, ωηιςη αρε τηε ςαμε ωιτη τηε ζεροες οφ ουρ $\beta$-φυνςτιον, ανδ ςονϑεςτυρεδ τηατ τηε πηψςιςαλ αλυε οφ $\alpha$ ιν ΧΕΔ ςουλδ βε ςομπυτεδ τηις ωαψ. Ιν τηειρ ωορϰ τηερε ις νο ρεφερενςε το Στυεςϰελβεργ ανδ Πετερμανν.
Ιν λινε ωιτη τηε γενεραλ διςιντερεςτ ιν χυαντυμ φιελδ τηεορψ, ρενορμαλιςατιον γρουπ ιδεας ωερε λαργελψ ιγνορεδ βψ τηε ηιγη ενεργψ πηψςιςς ςομμυνιτψ υντιλ τηε ενδ οφ τηε 1960ς [^35] Some important results were obtained by the Landau school, see below. There exists also a dedicated chapter in the book by N. N. Bogoliubov and D. V. Shirkov[68], first published in Russian in 1957.. Τηεψ βεςαμε φαςηιοναβλε βψ τηε ενδ οφ τηε δεςαδε, φολλοωινγ τηε ρεβιρτη οφ χυαντυμ φιελδ τηεορψ. Σομε ιμπορταντ παπερς πυβλιςηεδ ατ τηατ τιμε ςοντριβυτεδ το τηις εφφεςτ.
Προβαβλψ τηε μοςτ ςιγνιφιςαντ φορ παρτιςλε πηψςιςς ωερε τηε ονες βψ Κ. Σψμανζιϰ ανδ ῝.Γ. ῝αλλαν[69] ωιτη τηε ςελεβρατεδ ῝αλλαν-Σψμανζιϰ εχυατιον. Στριςτλψ ςπεαϰινγ, τηεψ βοτη αδδρεςς α χυεςτιον ωηιςη ις νοτ διρεςτλψ ρελατεδ το τηε ρενορμαλιςατιον γρουπ, αλτηουγη ιτ λεαδς το τηε ςαμε πηψςιςαλ ρεςυλτς. Ι ωιλλ πρεςεντ τηεμ ςηορτλψ. Ηερε Ι ωαντ το μεντιον αν ιμπορταντ ωορϰ βψ Κ.Γ. Ωιλςον ωηο φιρςτ αππλιεδ τηε ρενορμαλιςατιον γρουπ ιδεας το ςτρονγ ιντεραςτιονς[70]. Αλτηουγη τηε ςονςεπτ οφ αςψμπτοτις φρεεδομ ωας νοτ ϰνοων, Ωιλςον πρεςεντεδ α ςλεαρ αναλψςις οφ τηε ηιγη ενεργψ βεηαιορ οφ ςουπλινγ ςονςταντς ωιτη τηε ποςςιβιλιτψ οφ ρυννινγ ιντο φιξεδ ποιντς ορ εεν λιμιτινγ ςψςλες. Ιν α ςεριες οφ ςυβςεχυεντ πυβλιςατιονς[71] ηε αππλιεδ τηε ρενορμαλιςατιον γρουπ ιδεας το τηε πηψςιςς οφ πηαςε τρανςιτιονς ανδ ςριτιςαλ πηενομενα.
ϒςινγ μοδερν νοτατιον, ωε ςαν ωριτε τηε ρενορμαλιςατιον γρουπ εχυατιον φορ τηε ςιμπλεςτ ςςαλαρ φιελδ τηεορψ οφ ρεφερενςε (11) ας φολλοως: Τηε πηψςιςαλ τηεορψ, ι.ε. τηε ονε ωηιςη ςαν βε υςεδ το ςομπαρε ωιτη τηε εξπεριμενταλ δατα, ις τηε ονε ιν ωηιςη τηε αλυε οφ τηε μαςς εχυαλς τηε πηψςιςαλλψ μεαςυρεδ αλυε $m_{\mathrm{ph}}$, τηε φιελδ ις νορμαλιςεδ το ονε ον τηε μαςς-ςηελλ, ανδ τηε ςουπλινγ ςονςταντ ις τηε αλυε οφ τηε φουρ-ποιντ φυνςτιον ατ ςομε πηψςιςαλλψ μεαςυραβλε ποιντ, φορ εξαμπλε τηε ςςαττερινγ αμπλιτυδε ατ τηρεςηολδ. Τηε ςορρεςπονδινγ τηρεε ρενορμαλιςατιον ςονδιτιονς ταϰε τηε φορμ [^36] The renormalisation program is usually formulated in terms of the “amputated, one-particle-irreducible $n$-particle correlation functions” $\Gamma^{(n)}$. They are defined as the sum of all connected diagrams with $n$ external lines which cannot be separated into two disconnected parts by cutting a single internal line (the meaning of “one-particle-irreducible”) and from which we have dropped the propagators of the external lines (the meaning of “amputated”).
$\Gamma^{(2)}(p^{2}=m_{\rm{ph}}^{2})=0\leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \frac{\partial\Gamma^{(2)}(p^{2})}{\partial p^{2}}\Bigg{|}_{p^{2}=m_{\rm{ph}}^{2}}={\rm i}\leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \Gamma^{(4)}(p_{1},\ldots,p_{4})\Big{|}_{m_{\rm{ph}}^{2}}={\rm i}\lambda$ (29)
Τηε ςονςεπτ οφ τηε ρενορμαλιςατιον γρουπ ςτεμς φρομ τηε ρεαλιςατιον τηατ, φρομ τηε πυρελψ τεςηνιςαλ ποιντ οφ ιεω, νοτηινγ φορςες υς το ςηοοςε τηεςε πηψςιςαλ ρενορμαλιςατιον ςονδιτιονς· ανψ οτηερ ςηοιςε ωουλδ βε αςςεπταβλε, προιδεδ ιτ ψιελδς ωελλ-δεφινεδ αλυες φορ τηε τηρεε αρβιτραρψ ςονςταντς. Τηε εξαμπλε ωηιςη ις μοςτ ςομμονλψ υςεδ ις το ιντροδυςε α μαςς παραμετερ $M$ ανδ ρεωριτε τηε ςεςονδ ανδ τηιρδ οφ τηε εχυατιονς (29 ας [^37] We could use even more general conditions, for example not keeping the physical mass in the first of the equations (29), but no substantially new insight is gained. Examples of such more general renormalisation group equations are given in reference [72].
$\frac{\partial\Gamma^{(2)}(p^{2})}{\partial p^{2}}\Bigg{|}_{p^{2}=M^{2}}={\rm i}\leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \Gamma^{(4)}(p_{1},\ldots,p_{4})\Big{|}_{\mathrm{point}\,M}=-{\rm i}\lambda^{(M)}$ (30)
Ιν τηε λαςτ ςονδιτιον βψ “ποιντ $M$” ωε μεαν ςομε ποιντ ιν τηε ςπαςε οφ τηε φουρ μομεντα $p_{a}^{2}=M^{2}$, $a=1,\ldots,4$, προιδεδ ιτ ις α ποιντ ατ ωηιςη $\Gamma^{(4)}$ ις ρεγυλαρ. Σιμιλαρλψ, $\Gamma^{(2)}(p^{2})$ ανδ ιτς δεριατιε ςηουλδ βε ωελλ δεφινεδ ατ $p^{2}=M^{2}$. Φορ ηιςτοριςαλ ρεαςονς τηε ποιντ $M$ ιν μομεντυμ ςπαςε ις ςαλλεδ τηε ςυβτραςτιον ποιντ.
Σινςε $M$ ηας νο πηψςιςαλ μεανινγ, ωε ςαν ςηανγε ιτ βψ α τρανςφορματιον $M_{1}\longrightarrow M_{2}=\eta M_{1}$ ωιτη $\eta$ α νυμεριςαλ φαςτορ, ωηιςη ςαν ταϰε ανψ αλυε ιν α δομαιν ινςλυδινγ τηε αλυε $\eta=1$ ανδ ςυςη τηατ αλλ ρεςυλτινγ ποιντς $M_{2}$ βελονγ το τηε δομαιν οφ αναλψτιςιτψ οφ $\Gamma^{(2)}$ ανδ $\Gamma^{(4)}$. Τηεςε τρανςφορματιονς φορμ α γρουπ, ςαλλεδ τηε ρενορμαλιςατιον γρουπ. Φορ εερψ αλυε οφ $M$ ωε οβταιν α διφφερεντ ςετ οφ ςορρελατιον φυνςτιονς ανδ τηε ςονδιτιον τηατ αλλ τηεςε τηεοριες δεςςριβε τηε ςαμε πηψςιςς ις τηε ωελλ-ϰνοων εχυατιον οφ τηε ρενορμαλιςατιον γρουπ.
$\left[M\frac{\partial}{M}+\beta\left(\lambda,\frac{m}{M}\right)\frac{\partial}{\partial\lambda}-n\gamma\left(\lambda,\frac{m}{M}\right)\right]\Gamma^{(n)}(p_{1},...,p_{n};m,\lambda;M)=0$ (31)
$\beta$ ανδ $\gamma$ ςαν βε ςομπυτεδ ορδερ-βψ-ορδερ ιν τηε περτυρβατιον εξπανςιον ανδ, ςινςε τηεψ αρε διμενςιονλεςς, τηεψ δεπενδ ονλψ ον $\lambda$ ανδ τηε ρατιο $m/M$ [^38] Here $m$ is the physical mass. In order to simplify the notation we have dropped the subscript ${\mathrm{ph}}$..
Τηε ρενορμαλιςατιον γρουπ εχυατιον (31) δεςςριβες τηε ρεςπονςε οφ τηε ςψςτεμ υνδερ α ρεςςαλινγ οφ τηε ςυβτραςτιον ποιντ. Ονε ςουλδ αργυε τηατ ιτ ις οφ λιμιτεδ ιντερεςτ βεςαυςε ιτ ονλψ ςολες α προβλεμ ωε ηαε ςρεατεδ φορ ουρςελες βψ νοτ μαϰινγ τηε πηψςιςαλ ςηοιςε $M=m_{\rm{ph}}$. Τηις ις υνδουβτεδλψ ςορρεςτ ανδ ωε νεεδ ςομε μορε ινπυτ ιν ορδερ το εξτραςτ πηψςιςαλλψ ιντερεςτινγ ινφορματιον ουτ οφ τηις εχυατιον. Ιτ ωας τηε ςοντριβυτιον οφ Γελλ-Μανν ανδ Λοω ανδ Ι ωιλλ ςομε βαςϰ το ιτ πρεςεντλψ.
Λετ με ςομε νοω το τηε ψεαρ 1970 ανδ τηε ωορϰ οφ Κ. Σψμανζιϰ ανδ ῝.Γ. ῝αλλαν[69]. Τηε βαςις ρεμαρϰ ις τηατ ιν α φουρ-διμενςιοναλ ρενορμαλιςαβλε φιελδ τηεορψ ιν τηε ςλαςςιςαλ αππροξιματιον, τηε ονλψ τερμς ιν τηε Λαγρανγιαν ωηιςη βρεαϰ ςςαλε ιναριανςε αρε τηοςε ωιτη διμενςιονφυλ παραμετερς, ςυςη ας μαςς τερμς. Ιφ τηεψ αρε αβςεντ, φορ εξαμπλε ιφ ωε ςονςιδερ τηε τηεορψ (11) ωιτη $m=0$, ωε εξπεςτ τηε δεπενδανςε οφ τηε ςορρελατιον φυνςτιονς ον τηε εξτερναλ μομεντα το βε δετερμινεδ βψ διμενςιοναλ αναλψςις. Τηε 2-ποιντ φυνςτιον ςηουλδ βε γιεν βψ $\Gamma^{(2)}(p,\lambda)=f(\lambda)p^{2}$ ωιτη $f$ ςομε φυνςτιον οφ τηε ςουπλινγ ςονςταντ. Τηις ις ινδεεδ ςορρεςτ ατ τηε ςλαςςιςαλ λεελ ωιτη $f={\rm i}$. Ωε ϰνοω ηοωεερ τηατ τηις βεηαιουρ ις ωρονγ φορ τηε χυαντυμ τηεορψ βεςαυςε τηε διεργεντ διαγραμς ιν τηε περτυρβατιον εξπανςιον ωιλλ φορςε υς το ιντροδυςε α διμενςιονφυλ ςυτ-οφφ $\Lambda$ ανδ ωε φινδ τερμς προπορτιοναλ το ποωερς οφ $\log(p^{2}/\Lambda^{2})$. Ιν ουρ πρειους τερμινολογψ, ωε φινδ τηατ τηε ςψμμετρψ υνδερ ςςαλε τρανςφορματιονς ις ανομαλους. Τηε ῝αλλαν-Σψμανζιϰ εχυατιονς αρε τηε ςορρεςτ Ωαρδ ιδεντιτιες οφ βροϰεν ςςαλε ιναριανςε [^39] A very good pedagogical presentation can be found in S. Coleman, reference [73].. Τηεψ ταϰε τηε φορμ
$\left[m\frac{\partial}{m}+\beta(\lambda)\frac{\partial}{\partial\lambda}-n\gamma(\lambda)\right]\Gamma^{(n)}(p_{1},...,p_{n};m,\lambda)=m^{2}\delta(\lambda)\Gamma^{(n)}_{\phi^{2}}(p_{1},...,p_{n};m,\lambda)$ (32)
ωηερε $m$ δενοτες τηε πηψςιςαλ μαςς $m_{\rm{ph}}$. $\Gamma^{(n)}_{\phi^{2}}$ ις ςαλλεδ “τηε $n$-ποιντ φυνςτιον ωιτη α ζερο μομεντυμ $\phi^{2}$ ινςερτιον”. Ιτς μεανινγ ις εξπλαινεδ ιν ρεφερενςες [69] ανδ [73] βυτ ωε ωον´τ νεεδ ιτ ηερε. Τηε φυνςτιονς $\beta$, $\gamma$ ανδ $\delta$ ςαν βε ςομπυτεδ ιν περτυρβατιον τηεορψ. Τηεψ ςοδιφψ τηε ανομαλιες οφ ςςαλε ιναριανςε. ῝οντραρψ το ωηατ ωε φουνδ ιν τηε αξιαλ ανομαλψ, τηεψ ρεςειε νον-ζερο ςοντριβυτιονς ατ αλλ ορδερς οφ τηε περτυρβατιον εξπανςιον. Τηε $\gamma$ φυνςτιον ις οφτεν ςαλλεδ “τηε ανομαλους διμενςιον οφ τηε φιελδ” [^40] This terminology, although it is commonly used, it is misleading because it sounds as if the quantum theory violates dimensional analysis, which obviously cannot happen. For a discussion see reference [73]..
Τηε τωο εχυατιονς, (31) ανδ (32), αλτηουγη τηεψ λοοϰ αγυελψ ςιμιλαρ, τηεψ δεςςριβε τοταλλψ διφφερεντ προπερτιες οφ τηε τηεορψ. ῞αρψινγ τηε ςυβτραςτιον ποιντ $M$ ωε οβταιν α ςετ οφ χυαντυμ φιελδ τηεοριες, βυτ ονλψ ονε οφ τηεμ, οβταινεδ φορ $M=m_{\rm{ph}}$, ις τηε πηψςιςαλ ονε. Ιτ φολλοως τηατ τηε ρενορμαλιςατιον γρουπ τρανςφορματιονς μαπ ονε υνπηψςιςαλ τηεορψ το ανοτηερ. Σςαλε τρανςφορματιονς ον τηε οτηερ ηανδ ςηανγε τηε αλυε οφ τηε πηψςιςαλ μαςς, ςο τηεψ μαπ ονε πηψςιςαλ τηεορψ το ανοτηερ πηψςιςαλ τηεορψ.
Νοω τηατ ωε εξπλαινεδ ιν ωηατ τηεςε εχυατιονς αρε διφφερεντ, ωε ςαν τυρν ιντο ωηατ τηεψ ηαε ιν ςομμον. Τηεψ βοτη δεςςριβε τηε ρεςπονςε οφ τηε τηεορψ υνδερ α ςηανγε οφ α διμενςιονφυλ παραμετερ. Ωερε ιτ τηε ονλψ ςυςη παραμετερ, ωε ςουλδ αππεαλ το ορδιναρψ διμενςιοναλ αναλψςις ανδ τραδε τηις δεπενδενςε ωιτη τηατ ον τηε εξτερναλ μομεντα. Ωε εξπεςτ τηατ τηις ςαν βε αςηιεεδ ατ ερψ ηιγη ενεργιες ωηεν μαςςες ςουλδ βε νεγλεςτεδ. Ιτ ωας τηε ρεμαρϰ οφ Γελλ-Μανν ανδ Λοω. Τηις ναϊε εξπεςτατιον τυρνς ουτ το βε ςορρεςτ ορδερ βψ ορδερ ιν τηε περτυρβατιον εξπανςιον, προιδεδ ωε λοοϰ ατ χυαντιτιες ωηιςη αρε φρεε φρομ ινφραρεδ ςινγυλαριτιες[74]. Λετ υς ρεςςαλε αλλ εξτερναλ μομεντα $p_{i}\rightarrow\rho\leavevmode\nobreak\ p_{i}$ ωιτη α ρεαλ παραμετερ $\rho$ ανδ ταϰε $\rho>>1$. Σταρτινγ φρομ ειτηερ ονε οφ τηε τωο εχυατιονς (31) ορ (32), ωε οβταιν
$\left[-\rho\frac{\partial}{\partial\rho}+\beta(\lambda)\frac{\partial}{\partial\lambda}-n\gamma(\lambda)\right]\Gamma_{\mathrm{as}}^{(n)}(\rho\leavevmode\nobreak\ p_{i},\lambda)=0$ (33)
ωιτη $\Gamma_{\mathrm{as}}^{(n)}$ τηε αςψμπτοτις φορμ οφ τηε Γρεεν φυνςτιον ατ λαργε μομεντα ιν α ϰινεματιςαλ ρεγιον φρεε φρομ ινφραρεδ ςινγυλαριτιες.
Τηε διφφερεντιαλ εχυατιον (33) ωας ςολεδ αρουνδ τηε ενδ οφ τηε 18τη ςεντυρψ βψ Γαςπαρδ Μονγε, ωηο ωας ςτυδψινγ προβλεμς οφ λαμιναρ φλοω ιν ηψδροδψναμιςς. Ωε μαϰε α ςηανγε οφ αριαβλες φολλοωινγ τηε ηψδροδψναμις φλοω: $(\rho,\lambda)\rightarrow(\rho,\bar{\lambda})$ ωιτη $\bar{\lambda}(\lambda,\rho)$ ςολυτιον οφ τηε ςηαραςτεριςτις εχυατιον
$\left[-\rho\frac{\partial}{\partial\rho}+\beta(\lambda)\frac{\partial}{\partial\lambda}\right]\bar{\lambda}(\lambda,\rho)=0,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \bar{\lambda}(\lambda,\rho=1)=\lambda$ (34)
ωηιςη ις εχυιαλεντ το
$\rho\frac{\partial\bar{\lambda}}{\partial\rho}=\beta(\bar{\lambda})$ (35)
Τηε γενεραλ ςολυτιον οφ (34) ις νοω γιεν βψ
$\Gamma_{\rm as}^{(n)}(\rho p_{1},...,\rho p_{n};\lambda)=\Gamma_{\rm as}^{(n)}(p_{1},...,p_{n};\bar{\lambda})\exp\left\{-n\int_{1}^{\rho}{\rm d}\rho^{\prime}\gamma[\bar{\lambda}(\lambda,\rho^{\prime})]\right\}$ (36)
Τηε πηψςιςαλ μεανινγ οφ (36) ις ςλεαρ. Σςαλινγ αλλ μομεντα οφ α Γρεεν φυνςτιον βψ α ςομμον φαςτορ $\rho$ ανδ ταϰινγ $\rho$ λαργε, ηας τηε φολλοωινγ εφφεςτς: (ι) Ιτ μυλτιπλιες εερψ εξτερναλ λινε βψ τηε ςςαλε δεπενδεντ φαςτορ $\exp(-\int_{1}^{\rho}{\rm d}\rho^{\prime}\gamma(\bar{\lambda}(\lambda,\rho^{\prime})))$414141 We see that the Green functions get rescaled as if they had dimensions related to $\gamma(\lambda)$. This is the origin of the term “anomalous dimension”. ανδ (ιι) ιτ ρεπλαςες τηε πηψςιςαλ ςουπλινγ ςονςταντ $\lambda$ βψ αν εφφεςτιε ονε, $\bar{\lambda}=\bar{\lambda}(\lambda,\rho)$, ωηιςη ις τηε ςολυτιον οφ (35). Φορ τηις ρεαςον $\bar{\lambda}$ ις οφτεν ςαλλεδ ρυννινγ ςουπλινγ ςονςταντ.
Γελλ-Μανν ανδ Λοω, ιν τηειρ 1954 παπερ, ρεςογνιςεδ τηε ςιγνιφιςανςε οφ τηε ποςςιβλε ζερος οφ τηε $\beta$ φυνςτιον. Εχυατιον (35) ςηοως τηατ, ιφ $\beta(\lambda)=0$ φορ ςομε $\lambda=\lambda^{*}$, τηε εφφεςτιε ςουπλινγ ςονςταντ $\bar{\lambda}(\lambda^{*},\rho)$ βεςομες ινδεπενδεντ οφ $\rho$. Τηεψ ςονϑεςτυρεδ τηατ τηε φινε ςτρυςτυρε ςονςταντ $\alpha\approx 1/137$ ςουλδ βε τηε ςολυτιον οφ ςυςη αν ειγεναλυε εχυατιον $\beta_{\rm{QED}}(\alpha)=0$. Οβιουςλψ, ας λονγ ας περτυρβατιον τηεορψ ις ουρ ονλψ γυιδε, ωε ςαννοτ ςαψ ανψτηινγ αβουτ τηε προπερτιες οφ $\beta(\lambda)$ φορ αρβιτραρψ $\lambda$ ανδ ωε δο νοτ ϰνοω ωηετηερ ιτ ηας ανψ νον-τριιαλ ζερος. Τηερε ηαε βεεν ςεεραλ αττεμπτς το γυεςς τηε προπερτιες οφ $\beta$ βεψονδ περτυρβατιον – φορ εξ. υςινγ τηε μετηοδ οφ Παδé αππροξιμαντς – ωιτη ςομε γοοδ ρεςυλτς ιν προβλεμς οφ ςτατιςτιςαλ μεςηανιςς, βυτ νο ςυςςεςς ιν ηιγη ενεργψ πηψςιςς. Ωε ϰνοω τηατ $\beta(0)=0$, βεςαυςε $\lambda=0$ ςορρεςπονδς το α φρεε φιελδ τηεορψ. Ωε αλςο ϰνοω τηατ τηε βεηαιορ οφ τηε εφφεςτιε ςουπλινγ ςονςταντ ατ ηιγη ενεργιες δεπενδς ον τηε ςιγν οφ τηε φιρςτ νον-ζερο τερμ ιν τηε εξπανςιον οφ $\beta$.
$\beta(\lambda)=b_{0}\lambda^{2}+b_{1}\lambda^{3}+\dots\leavevmode\nobreak\ \leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ b_{0}=\frac{3}{16\pi^{2}}$ (37)
Τηε φιρςτ τερμ $b_{0}$ ις ποςιτιε, ςο $\bar{\lambda}$ γροως ωιτη $\rho$. Τηις βεηαιορ ις ςομμον το αλλ ρενορμαλιςαβλε χυαντυμ φιελδ τηεοριες ωηιςη ωερε ϰνοων ιν τηε φιφτιες ανδ εαρλψ ςιξτιες ανδ ιτ ωας οφτεν αςςυμεδ το βε υνιερςαλ. Λ.Δ. Λανδαυ ςονςλυδεδ[75] τηατ αλλ τηεςε τηεοριες ωερε λογιςαλλψ ινςονςιςτεντ, υνλεςς τηε ςουπλινγ ςονςταντ ωας ζερο, ιν ωηιςη ςαςε τηεψ ωερε τριιαλ. Ιν ηις αρτιςλε ωιτη Ι. Πομερανςηυϰ τηεψ νοτε “Ωε ρεαςη τηε ςονςλυςιον τηατ ωιτηιν τηε λιμιτς οφ φορμαλ ελεςτροδψναμιςς α ποιντ ιντεραςτιον ις εχυιαλεντ, φορ ανψ ιντενςιτψ ωηατεερ, το νο ιντεραςτιον ατ αλλ.” [^42] This was the famous zero charge problem. At high energies the effective charge $\bar{\alpha}(\alpha,\rho)$ grows without limits unless the physical charge $\alpha$ vanishes. Ιτ ις οβιους τηατ ςτατεμεντς οφ τηις ϰινδ, ςομινγ φρομ αυτηοριτιες ωιτη τηε ςτατυς οφ Λανδαυ, ςοντριβυτεδ ιν τηροωινγ α τοταλ διςςρεδιτ ον χυαντυμ φιελδ τηεορψ. Λετ με ρεμαρϰ ηοωεερ, τηατ Λανδαυ ωας ριγητ το ςαψ τηατ “…ωεαϰ ςουπλινγ ελεςτροδψναμιςς ις α τηεορψ ωηιςη ις, φυνδαμενταλλψ, λογιςαλλψ ινςομπλετε…” [^43] Today we expect the problem to be solved in some kind of ultraviolet completion of the theory, for example in the framework of a grand unified theory.. Τηε μιςταϰε, ηαρδ το αοιδ ωιτη τηε ϰνοωλεδγε οφ ηις τιμε, ωας το γενεραλιςε ιτ το αλλ χυαντυμ φιελδ τηεοριες ανδ αςςυμε τηατ νο οτηερ βεηαιορ ωας ποςςιβλε. “…Ωε αρε δριεν το τηε ςονςλυςιον τηατ τηε Ηαμιλτονιαν μετηοδ φορ ςτρονγ ιντεραςτιονς ις δεαδ ανδ μυςτ βε βυριεδ, αλτηουγη οφ ςουρςε ωιτη δεςερεδ ηονορ.”[76] Τηε πιςτυρε ςηανγεδ ωιτη τηε διςςοερψ οφ αςψμπτοτιςαλλψ φρεε χυαντυμ φιελδ τηεοριες, βυτ τηεψ βελονγ το τηε νεω ερα ανδ ωιλλ βε πρεςεντεδ ιν ςεςτιον 1.
[…]
Ιν ςεςτιον 1 ωε πρεςεντεδ τηε ςονςτρυςτιον οφ τηε ιςοςπιν ςψμμετρψ οφ νυςλεαρ φορςες βψ Ηειςενβεργ ανδ Κεμμερ. Ιτ ις βαςεδ ον τηε γρουπ $SU(2)$ ανδ τηε παιρ προτον-νευτρον βελονγς το τηε τωο-διμενςιοναλ φυνδαμενταλ ρεπρεςεντατιον. Τηε διςςοερψ οφ ςτρανγε παρτιςλες βρουγητ τηε νεεδ το ενλαργε τηις ςψμμετρψ ανδ τηε “οβιους” εξτενςιον ωας $SU(2)\rightarrow SU(3)$ βψ ενλαργινγ τηε δουβλετ το α τριπλετ. Ιτ ωας τηε μοδελ προποςεδ βψ Σ. Σαϰατα ιν 1956[77] ιν ωηιςη ηε αδδεδ τηε $\Lambda$-ηψπερον το τηε τωο νυςλεονς το φορμ α φυνδαμενταλ τριπλετ οφ $SU(3)$. Τηις ςςηεμε διδ νοτ φιτ τηε οβςερεδ ηαδρονς ανδ ιν τηε φολλοωινγ ψεαρς μανψ γρουπς ωερε ςονςιδερεδ ας ςανδιδατες φορ τηις λαργερ ςψμμετρψ. Φιναλλψ ιτ τυρνεδ ουτ τηατ Νατυρε ωας μορε ςυβτλε: $SU(3)$ ις τηε ριγητ ανςωερ, βυτ τηε φυνδαμενταλ τριπλετ ις νοτ φορμεδ βψ παρτιςλες ωε ςεε. Ιν 1961 Μ. Γελλ-Μανν ανδ Ψ. Νεέμαν φορμυλατεδ τηε μοδελ ωηιςη βεςαμε ϰνοων ας “Τηε ειγητφολδ ωαψ”[78] ιν ωηιςη τηε λοω-λψινγ ηαδρονς φορμ οςτετς οφ $SU(3)$. Τηε πεςυλιαρ προπερτψ οφ τηις ςςηεμε ις πρεςιςελψ τηε αβςενςε οφ πηψςιςαλ παρτιςλες βελονγινγ το τηε φυνδαμενταλ τριπλετ ρεπρεςεντατιον. Τηις λεδ Γελλ-Μανν ανδ, ινδεπενδεντλψ Γ. Ζωειγ, το προποςε α ςομποςιτε μοδελ ιν ωηιςη ηαδρονς αρε βουνδ ςτατες οφ ςονςτιτυεντς – ναμεδ “χυαρϰς” βψ Γελλ-Μανν ανδ “αςες” βψ Ζωειγ – ωηιςη φιλλ τηε μιςςινγ τριπλετ[79]. Ωηατ ις λεςς ϰνοων ις τηατ ςιμιλαρ ιδεας ηαδ βεεν εξπρεςςεδ εαρλιερ, Ι βελιεε φιρςτ βψ Λ.Β. Οϰυν[80]. Ηε ςταρτεδ φρομ τηε Σαϰατα μοδελ βυτ ηε ιντερπρετεδ τηε μεμβερς οφ τηε φυνδαμενταλ τριπλετ νοτ ας τηε πηψςιςαλ $p$, $n$ ανδ $\Lambda$, βυτ ας φιςτιτιους παρτιςλες [^44] He used the term “sakatons”.. Ιν φαςτ, ουρ ιδεας ον χυαρϰς φολλοωεδ α ςιμιλαρ τραϑεςτορψ. Ατ φιρςτ τηεψ ωερε ςονςιδερεδ ας ϑυςτ ανοτηερ λαψερ οφ ςομποςιτενεςς ανδ εξπεριμεντς ωερε δεςιγνεδ το διςςοερ τηεμ [^45] There have been even several unsubstantiated claims of discovery.. Ονε οφ τηε μοτιατιονς το βυιλτ ΙΣΡ, τηε φιρςτ ηαδρονις ςολλιδερ, ωας το διςςοερ χυαρϰς. Ηοωεερ, τηε πιςτυρε βεςαμε ςοον μορε ςομπλιςατεδ. Τηε νεεδ το ςατιςφψ τηε ρεχυιρεμεντς οφ Φερμι ςτατιςτιςς, λεδ το τηε ιντροδυςτιον οφ ςολορ[81] ανδ τηε ηψποτηεςις οφ “ςολορ βλινδνεςς” λεδ το τηε νοτιον οφ ςονφινεμεντ. Εεν τοδαψ ωε ηαε νο ρεαλ υνδερςτανδινγ οφ ιτ ανδ ωε ςαννοτ ριγορουςλψ δεριε ιτ φρομ φιρςτ δψναμιςαλ πρινςιπλες [^46] It is one of the unsolved problems endowed with a prize of one million dollars from the Clay Mathematics Institute. For a review of the various ideas around confinement see reference [82].. Ας α ρεςυλτ, ιν ςπιτε οφ ςομε πηενομενολογιςαλ ςυςςεςςες οφ τηε ςιμπλε νον-ρελατιιςτις χυαρϰ μοδελ, μοςτ πεοπλε ςηιφτεδ τηειρ ιεως το ςονςιδερινγ χυαρϰς ας μερε ματηεματιςαλ δειςες το ϰεεπ τραςϰ οφ ςψμμετρψ προπερτιες. Νατυραλλψ, Μ. Γελλ-Μανν, ωηο ςτρονγλψ αδοςατεδ τηις οπινιον, πλαψεδ α μαϑορ ρολε το ιτς γενεραλ ενδορςεμεντ. Τηε ατμοςπηερε ςηανγεδ ωιτη τηε ΣΛΑ῝ ρεςυλτς ον δεεπ ινελαςτις ςςαττερινγ, βυτ Ι ωιλλ διςςυςς τηεμ ιν ςεςτιον 1.
[…]
Ι δο νοτ ϰνοω ωηο ινεντεδ τηις τερμ, ωηιςη Ι φινδ ρατηερ μιςλεαδινγ [^47] Sidney Coleman, in one of his Erice lectures, wanted to replace it with “secret symmetry”[83], but it was not adopted., βυτ τηε πηενομενον ις ωελλ ϰνοων ιν μανψ προβλεμς οφ ςλαςςιςαλ ανδ χυαντυμ πηψςιςς, εςπεςιαλλψ τηοςε ρελατεδ το πηαςε τρανςιτιονς. Ιτς μαιν φεατυρες ςαν βε ςυμμαριςεδ ας φολλοως:
Ιφ τηε δψναμιςς οφ α πηψςιςαλ ςψςτεμ ις ιναριαντ υνδερ α γρουπ οφ τρανςφορματιονς, ωε εξπεςτ τηε εχυατιονς οφ μοτιον το αδμιτ ςψμμετρις ςολυτιονς. Το ςιτε Π. ῝υριε[84] “ Ωηεν ςερταιν ςαυςες προδυςε ςερταιν εφφεςτς, τηε ελεμεντς οφ ςψμμετρψ οφ τηε ςαυςες μυςτ βε φουνδ ιν τηε εφφεςτς προδυςεδ.” Εξπρεςςεδ τηις ωαψ, τηε ςτατεμεντ ςουνδς αλμοςτ τριιαλ, ηοωεερ ιτ μαψ ηαππεν τηατ, ιν α ςερταιν ρεγιον οφ τηε παραμετερ ςπαςε, τηε ςψμμετρις ςολυτιον βεςομες υνςταβλε. Ιτ ις οφτεν τηε ςαςε ιν ινφινιτε ςψςτεμς εξηιβιτινγ τηε πηενομενον οφ πηαςε τρανςιτιονς. Ιν ςυςη ςαςες τηε μινιμυμ ενεργψ ςολυτιονς, ωηιςη ρεπρεςεντ τηε ςταβλε ςτατες, αρε δεγενερατε ανδ ηαε λεςς ςψμμετρψ τηαν τηε εχυατιονς οφ μοτιον. Τηε δεγρεε οφ δεγενεραςψ ςορρεςπονδς το τηε ςψμμετρψ δεφιςιτ. Τηε ςψμμετρψ ις νοτ “βροϰεν”, βυτ ιτ ις “ηιδδεν” βεςαυςε ιτ ις νοτ μανιφεςτ ιν τηε ςπεςτρυμ οφ ςτατες. Φορ α χυαντυμ ςψςτεμ τηε δεγενεραςψ οφ τηε γρουνδ ςτατε ις τρανςλατεδ ιντο τηε αππεαρανςε οφ ζερο ενεργψ εξςιτατιονς. Ιν τηε φολλοωινγ Ι ωιλλ πρεςεντ ονλψ τηε αππλιςατιονς οφ τηις πηενομενον το παρτιςλε πηψςιςς.
$\bullet$ Σποντανεους βρεαϰινγ οφ γλοβαλ ιντερναλ ςψμμετριες. Τηε πηενομενον ωας φιρςτ ςτυδιεδ βψ Ψοιςηιρο Ναμβυ[85] ανδ Θεφφρεψ Γολδςτονε[86]. Ναμβυ ωας μοτιατεδ βψ α ςονςρετε πηψςιςαλ προβλεμ, τηατ οφ δετερμινινγ τηε ηαδρονις αξιαλ εςτορ ςυρρεντ οφ ωεαϰ ιντεραςτιονς. Φολλοωινγ τηε εξπεριμενταλ διςςοερψ οφ παριτψ ιολατιον ανδ τηε μεαςυρεμεντ οφ τηε νευτρινο ηελιςιτψ ιν $\beta$-δεςαψ, τηε Φερμι τηεορψ οφ τηε ωεαϰ ιντεραςτιονς ωας δετερμινεδ το βε οφ τηε ςυρρεντ$\times$ςυρρεντ φορμ. Τηε ςυρρεντ $J_{\mu}$ ωας τηε ςυμ οφ α ηαδρονις ανδ α λεπτονις πιεςε, εαςη ονε βεινγ α ςυπερποςιτιον οφ α εςτορ ανδ αν αξιαλ παρτ.
${\cal H}_{W}=\frac{G_{F}}{\sqrt{2}}J^{\mu}J^{\dagger}_{\mu}\leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ J_{\mu}=J_{\mu}^{(h)}+J_{\mu}^{(l)}\leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ J_{\mu}=V_{\mu}-A_{\mu}$ (38)
Ιν ωριτινγ (38) ωε μαδε τηε αςςυμπτιον οφ υνιερςαλιτψ ωηιςη, ιν α ρατηερ αγυε ςενςε, ιμπλιεδ τηατ τηε Φερμι ςουπλινγ ςονςταντ $G_{F}$ ωας τηε ςαμε φορ αλλ ςυρρεντς. Ιτ ις ςιμπλε το ωριτε τηε λεπτονις παρτ ιν τερμς οφ τηε φιελδς οφ ϰνοων λεπτονς [^48] Today we must add also the $\tau$ piece.
$J_{\mu}^{(l)}=\bar{\psi}_{e}\gamma_{\mu}(1+\gamma_{5})\psi_{\nu_{e}}+(e\rightarrow\mu,\nu_{e}\rightarrow\nu_{\mu})$ (39)
βυτ τηερε ις νοτηινγ αναλογους φορ τηε ηαδρονις παρτ φορ ωηιςη αν εξπρεςςιον ιν τερμς οφ φρεε φιελδς ις μεανινγλεςς. Τηερεφορε ωε ηαε το λοοϰ ατ παρτιςυλαρ ματριξ ελεμεντς ανδ τηε φιρςτ ονε ις νευτρον $\beta$-δεςαψ. Ιγνορινγ τηε ςμαλλ μομεντυμ τρανςφερ βετωεεν τηε νευτρον ανδ τηε προτον, ωε ςαν ωριτε τηε ματριξ ελεμεντ ας
$\bra{p}J_{\mu}^{(h)}\ket{n}=\bar{u}_{p}\gamma_{\mu}\big{[}g_{V}(0)+g_{A}(0)\gamma_{5}\big{]}u_{n}$ (40)
Ιν γενεραλ, $g_{V}$ ανδ $g_{A}$ αρε τωο φορμ φαςτορς ωηιςη δεπενδ ον τηε ςχυαρε οφ τηε μομεντυμ τρανςφερ [^49] If we relax the zero momentum transfer approximation, we obtain the contribution of additional form factors., βυτ ιν ουρ αππροξιματιον τηεψ αρε τωο ρεαλ νυμβερς το βε δετερμινεδ βψ εξπεριμεντ. Τηεψ αρε τηε εςτορ ανδ τηε αξιαλ ωεαϰ ςηαργες οφ τηε νυςλεον. Ιφ ωε φαςτορ ουτ τηε Φερμι ςουπλινγ ςονςταντ, ωηιςη ις δετερμινεδ βψ τηε ρατε οφ μυον δεςαψ, τηε εξπεριμενταλ ρεςυλτς γιε υς τηε αλυες
$g_{V}(0)\simeq 1\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ g_{A}(0)\simeq 1.25$ (41)
Τηε ςυρπριςινγ ρεςυλτ ις τηατ οφ $g_{V}$. Ι ρεμινδ τηατ, ιν τηε αβςενςε οφ ςτρονγ ιντεραςτιονς, ωε ωουλδ ηαε οβταινεδ $g_{V}=g_{A}=1$, τηερεφορε τηε αλυε ιν (41) μεανς τηατ, φορ τηε εςτορ παρτ οφ τηε ςυρρεντ, αλλ ςτρονγ ιντεραςτιον ςορρεςτιονς μαναγε το ςυμ υπ το ζερο! Ιτ ις τηε ρεςυλτ ωε οβταιν φορ τηε ελεςτρομαγνετις ςυρρεντ, ωηερε τηε ελεςτρις ςηαργες οφ τηε προτον ανδ τηε ποςιτρον αρε εξαςτλψ εχυαλ – ιν ςπιτε οφ τηε προτον ηαινγ ςτρονγ ιντεραςτιονς – ας α ςονςεχυενςε οφ τηε ςονςερατιον οφ τηε ελεςτρομαγνετις ςυρρεντ. Τηις ςτρονγλψ ςυγγεςτς τηατ τηε εςτορ ηαδρονις ςυρρεντ οφ τηε ωεαϰ ιντεραςτιονς μυςτ βε α ςονςερεδ ςυρρεντ οφ τηε ςτρονγ ιντεραςτιονς, ωηιςη ωε ςηουλδ ιδεντιφψ ωιτη τηε Νοετηερ ςυρρεντ οφ α ςψμμετρψ. Σ.Σ. Γερςητειν ανδ Ψ.Β. Ζελδοιςη ανδ, ινδεπενδεντλψ, Ρ.Π. Φεψνμαν ανδ Μ. Γελλ- Μανν[87] φορμυλατεδ τηε ῝ονςερεδ ῞εςτορ ῝υρρεντ (῝῞῝) ηψποτηεςις, αςςορδινγ το ωηιςη τηε εςτορ παρτ οφ $J_{\mu}^{(h)}$ βελονγς το τηε τριπλετ οφ τηε ςονςερεδ ιςοςπιν ςυρρεντς οφ ςτρονγ ιντεραςτιονς. Ις ιτ ποςςιβλε το μαϰε α ςιμιλαρ ηψποτηεςις φορ τηε αξιαλ ςυρρεντ; Ιν οτηερ ωορδς, ςαν ωε αςςυμε τηατ $\partial^{\mu}A^{(h)}_{\mu}=0$, τηυς προμοτινγ τηε ιςοςπιν ςψμμετρψ οφ τηε ςτρονγ ιντεραςτιονς φρομ $SU(2)$ το $SU(2)\times SU(2)$; Ατ φιρςτ ςιγητ τηε ανςωερ ις νο ανδ τηερε ις πλεντψ οφ ειδενςε αγαινςτ ιτ. Φιρςτ, α ςονςερεδ αξιαλ ςυρρεντ ωουλδ ιμπλψ $g_{A}(0)\simeq 1$ ιν ςοντραδιςτιον ωιτη (41). Σεςονδ, ςυςη α ηψποτηεςις αππεαρεδ το βε ινςομπατιβλε ωιτη τηε δατα ον λεπτονις πιον δεςαψ ωηιςη ηαδ βεεν ςονφιρμεδ βψ τηε οβςερατιον ατ ῝ΕΡΝ οφ τηε ελεςτρονις μοδε ωιτη τηε ςορρεςτ ρατε. Ινδεεδ, ςινςε τηε πιον ις α πςευδοςςαλαρ, ιτς λεπτονις δεςαψ αμπλιτυδε ις γιεν βψ τηε ματριξ ελεμεντ οφ τηε αξιαλ ςυρρεντ
$\bra{0}A^{(h)}_{\mu}(0)\ket{\pi(q)}=f_{\pi}\,q_{\mu}\leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \bra{0}\partial^{\mu}A^{(h)}_{\mu}(0)\ket{\pi(q)}={\rm i}f_{\pi}m_{\pi}^{2}$ (42)
ωιτη τηε πιον δεςαψ ςονςταντ $f_{\pi}\simeq 130\,{\rm MeV}\simeq 0.93\,m_{\pi}$. Τηιρδ, ανδ μοςτ ιμπορταντ, τηερε ις νο τραςε οφ ςυςη α ςηιραλ ςψμμετρψ ιν τηε ςπεςτρυμ οφ ηαδρονς. Τηε προτον ανδ τηε νευτρον φορμ α δουβλετ οφ ιςοςπιν βυτ τηεψ ηαε νο παρτνερς ωιτη οπποςιτε παριτψ. Φορ αλλ τηεςε ρεαςονς, τηε ςονςερατιον οφ τηε αξιαλ ςυρρεντ ςεεμεδ το βε εξςλυδεδ.
Ναμβυ ωας τηε φιρςτ το οβςερε τηατ αλλ τηις “ειδενςε” ωας ιν φαςτ μιςλεαδινγ. Λετ υς ςταρτ ωιτη τηε αλυε οφ $g_{A}(0)$ ανδ τηε πιον δεςαψ αμπλιτυδε. Τηεψ ςηαρε α ςτριϰινγ φεατυρε: τηεψ αρε βοτη νυμεριςαλλψ ωεαϰ. Ινδεεδ, ιν τηε ςςαλε οφ ςτρονγ ιντεραςτιονς τηε ςηανγε φρομ 1 το 1.25 ις α ρατηερ ςμαλλ ρενορμαλιςατιον, ανδ $m_{\pi}^{2}$ ις ερψ ςμαλλ ςομπαρεδ το τηε ςχυαρε οφ α τψπιςαλ ηαδρον μαςς, ωηιςη ις οφ ορδερ 1 Γε῞2. Σο, τηε χυεςτιον ις: ςουλδ ωε αςςυμε τηατ τηε αξιαλ ςυρρεντ ις αππροξιματελψ ςονςερεδ; Βυτ τηεν, ωηατ αβουτ τηε τηιρδ πιεςε οφ ειδενςε, τηε αβςενςε οφ ςηιραλ ςψμμετρψ ιν τηε ςπεςτρυμ οφ ηαδρονς; Ηερε τηε λεςςονς ωε λεαρνεδ φρομ τηε πηενομενον οφ ςποντανεους ςψμμετρψ βρεαϰινγ ιν ςτατιςτιςαλ μεςηανιςς ςομε ηανδψ. Τρανςλατεδ ιντο τηε λανγυαγε οφ παρτιςλε πηψςιςς, τηεψ ιμπλψ τηατ τηε ιναριανςε οφ τηε εχυατιονς οφ μοτιον υνδερ α γρουπ οφ τρανςφορματιονς μαψ βε ρεαλιςεδ ιν ατ λεαςτ τωο διφφερεντ ωαψς. Ιν τηε φιρςτ, ωηιςη ωε ςαλλ Ωιγνερ ρεαλιςατιον, τηε ειγενςτατες οφ τηε Ηαμιλτονιαν φορμ ιρρεδυςιβλε ρεπρεςεντατιονς οφ τηε ςψμμετρψ γρουπ ωιτη δεγενερατε ειγεναλυες. Τηις ις τηε ςαςε οφ τηε ιςοςπιν ςψμμετρψ. Ηοωεερ, ωε ςαω τηατ τηερε εξιςτς αν αλτερνατιε ωαψ, ωηιςη ιν ουρ ϑαργον ις ςαλλεδ τηε Ναμβυ-Γολδςτονε ρεαλιςατιον. Τηε ςψμμετρψ ις ςποντανεουςλψ βροϰεν ανδ ιτ ις νοτ μανιφεςτ ιν τηε ςπεςτρυμ οφ ςτατες. Ωε ςαω τηατ ιτ ςονταινς α μαςςλεςς εξςιτατιον ωηιςη, ιν τηε λανγυαγε οφ παρτιςλε πηψςιςς, ιμπλιες τηε πρεςενςε οφ α ζερο μαςς ςτατε, τηε Γολδςτονε παρτιςλε. Φολλοωινγ Ναμβυ, ωε ποςτυλατε τηατ τηις πηενομενον οςςυρς ιν τηε ςτρονγ ιντεραςτιονς. Το βε πρεςιςε, ωε αςςυμε τηατ τηε ςτρονγ ιντεραςτιον Ηαμιλτονιαν ςαν βε ωριττεν ιν τηε φορμ
$H_{\rm strong}=H_{\rm symmetric}+H_{\rm breaking}$ (43)
ωηερε $H_{\rm symmetric}$ ις ιναριαντ υνδερ ςηιραλ $SU(2)\times SU(2)$ τρανςφορματιονς [^50] The generators of the two $SU(2)$ factors are the combinations $V+A$ and $V-A$ respectively, hence the term “chiral symmetry”. ανδ $H_{\rm breaking}$ βρεαϰς τηις ιναριανςε, βυτ ιτς εφφεςτς αρε ρελατιελψ ςμαλλ. Φυρτηερμορε, ωε αςςυμε τηατ $SU(2)\times SU(2)$ ις ςποντανεουςλψ βροϰεν το τηε διαγοναλ ςυβγρουπ $SU(2)\times SU(2)\rightarrow SU(2)_{V}$, ωιτη $SU(2)_{V}$ ινολινγ ονλψ εςτορ ςυρρεντς, ωηιςη ωε ιδεντιφψ ωιτη ιςοςπιν. Ιν τηε αβςενςε οφ $H_{\rm breaking}$ τηις ςποντανεους βρεαϰινγ ωουλδ λεαδ το τηρεε μαςςλεςς πςευδο-ςςαλαρ Γολδςτονε βοςονς. Σωιτςηινγ ον $H_{\rm breaking}$ τυρνς τηεςε βοςονς ιντο ``πςευδο-Γολδςτονε βοςονς᾽ ωιτη ςμαλλ, νον-ζερο μαςςες. Ωε αςςυμε τηατ τηεψ ςαν βε ιδεντιφιεδ ωιτη τηε πιονς. Τηις ςςηεμε ις ϰνοων ας Π῝Α῝ φορ Παρτιαλ ῝ονςερατιον οφ τηε Αξιαλ ῝υρρεντ.
Α φεω μοντης αφτερ Ναμβυ, Γολδςτονε πυβλιςηεδ α παπερ ωιτη τηε τιτλε Φιελδ Τηεοριες ωιτη “Συπερςονδυςτορ” Σολυτιονς[86]. Ι ωιλλ εξπλαιν τηε μεανινγ οφ τηε ωορδ “Συπερςονδυςτορ” ιν τηε νεξτ ςεςτιον, βυτ τηε παπερ ςονταινς τηε φιρςτ ςονςρετε φιελδ τηεορψ μοδελς τηατ εξηιβιτ τηε πηενομενον οφ ΣΣΒ. Ιν παρτιςυλαρ, ιτ ςονταινς τηε μοδελ ωε υςε τοδαψ ιν ωηιςη τηε βρεαϰινγ ις τριγγερεδ βψ α ςςαλαρ φιελδ ωηοςε ςχυαρε μαςς βεςομες νεγατιε. Ιτ ις τηε “Στανδαρδ Μοδελ” οφ ςποντανεους ςψμμετρψ βρεαϰινγ. Γολδςτονε ςηοως τηατ αλλ ςυςη μοδελς ςονταιν α μαςςλεςς ςςαλαρ, ηενςε τηε οριγιν οφ τηε τερμ Γολδςτονε βοςον [^51] The connection between SSB and massless particles has been proven later under very general assumptions, independent of particular field theory models. See, for example reference [88].
Βεφορε ςλοςινγ τηις ςεςτιον Ι ωαντ το μεντιον α παπερ ωηιςη πλαψεδ α ερψ ιμπορταντ ρολε ιν αλλ ςυβςεχυεντ δεελοπμεντς. Ιτ ις τηε $\sigma$-μοδελ οφ Γελλ-Μανν ανδ Μαυριςε Λéψ[89]. Τηε τιτλε ις “Τηε αξιαλ εςτορ ςυρρεντ ιν βετα δεςαψ” ανδ, αλτηουγη ιτ δοες νοτ μεντιον τηε πηενομενον οφ ΣΣΒ – ιτ ωας ςυβμιττεδ α φεω δαψς βεφορε τηε παπερ βψ Ναμβυ – ιτ οφφερς α ςονςρετε φιελδ τηεορετις φραμεωορϰ φορ ιμπλεμεντινγ Ναμβυ´ς ιδεα. Βοτη παπερς, βψ Ναμβυ ανδ Γελλ-Μανν–Λéψ, ςταρτ βψ νοτιςινγ αν ιντριγυινγ ρελατιον, διςςοερεδ εαρλιερ βψ Μ.Λ. Γολδβεργερ ανδ Σ.Β. Τρειμαν[90], βετωεεν τωο αππαρεντλψ υνρελατεδ χυαντιτιες, το ωιτ $g_{A}$ οφ $\beta$-δεςαψ ανδ τηε πιον-νυςλεον ςουπλινγ ςονςταντ. Γελλ-Μανν ανδ Λéψ´ς μοδελ ςονταινς α δουβλετ οφ νυςλεον φιελδς $\Psi(x)$, ωηοςε ριγητ- ανδ λεφτ-ηανδεδ ςομπονεντς φορμ ρεπρεςεντατιονς (1/2, 0) ανδ (0, 1/2) οφ ςηιραλ $SU(2)\times SU(2)$, ρεςπεςτιελψ, ανδ α χυαρτετ οφ ςςαλαρ φιελδς τρανςφορμινγ λιϰε α (1/2, 1/2) ρεπρεςεντατιον οφ τηε ςηιραλ γρουπ. Σινςε $SU(2)\times SU(2)$ ις λοςαλλψ ιςομορπηις το $O(4)$, τηε $2\times 2$ ματριξ οφ τηε ςςαλαρς φιελδς ςαν βε ωριττεν ας $\Phi(x)=[\sigma(x){\rm 1\kern-3.99994pt1}+{\rm i}{\bm{\tau}}\cdot{\bm{\pi}}(x)]/\sqrt{2}$ ωιτη ${\bm{\tau}}$ τηε τηρεε Παυλι ματριςες. Τηε Λαγρανγιαν δενςιτψ ις γιεν βψ
${\cal L}=$ $\bar{\Psi}_{L}i{\partial\kern-5.0pt/}\Psi_{L}+\bar{\Psi}_{R}i{\partial\kern-5.0pt/}\Psi_{R}+g\bar{\Psi}_{L}\Phi\Psi_{R}+h.c.$ (44)
$+\frac{1}{2}Tr[(\partial^{\mu}\Phi)(\partial_{\mu}\Phi^{{\dagger}})]-\frac{1}{2}M^{2}Tr[\Phi\Phi^{{\dagger}}]-\frac{\lambda}{4!}Tr[\Phi\Phi^{{\dagger}}]^{2}$
Νοτε τηατ ςηιραλ ςψμμετρψ φορβιδς α μαςς τερμ φορ τηε φερμιονς. Γελλ-Μανν ανδ Λéψ βρεαϰ τηε ςψμμετρψ εξπλιςιτλψ βψ αδδινγ α τερμ λινεαρ ιν τηε $\sigma$ φιελδ, βυτ ωε ςαν φολλοω Γολδςτονε ανδ ςηοοςε $M^{2}<0$ ωιτη τηε ρεςυλτ τηατ τηε φερμιονς αςχυιρε α μαςς ανδ τηε τηρεε πιον φιελδς βεςομε μαςςλεςς. Τηεψ αρε τηε Γολδςτονε βοςονς. Ιτ ις τηε μεςηανιςμ ωε ςηαλλ αππλψ το τηε Στανδαρδ Μοδελ. Τηε χυαντιτψ, ωηιςη ςορρεςπονδς το τηε ορδερ παραμετερ οφ τηε ςποντανεους ςψμμετρψ βρεαϰινγ, ις τηε αςυυμ εξπεςτατιον αλυε οφ τηε $\sigma$ φιελδ, ωηιςη ις νον-ζερο ιν τηε βροϰεν πηαςε. Ιν αδδιτιον το τηε τηρεε νεαρλψ μαςςλεςς πιονς, τηε μοδελ πρεδιςτς τηε εξιςτενςε οφ α $0^{+}$ παρτιςλε, τηε $\sigma$, οφ υνϰνοων μαςς [^52] Notice the analogy with the Standard Model.. Ηοωεερ, ας ιτ ωας αλρεαδψ ποιντεδ ουτ βψ Γελλ-Μανν ανδ Λéψ, ιν τηε ςπεςτρυμ οφ $0^{+}$ παρτιςλες, τηερε ις νο οβιους ςανδιδατε φορ $\sigma$. Τηυς ωε αςςυμε τηατ τηις ρολε ις πλαψεδ βψ α ςομποςιτε οπερατορ ανδ τηε ορδερ παραμετερ ις τηε αςυυμ εξπεςτατιον αλυε οφ α φερμιον–αντι-φερμιον παιρ $<\bar{\Psi}\Psi>_{0}$. Τηις πηενομενον ις ϰνοων ας “δψναμιςαλ ςψμμετρψ βρεαϰινγ” ανδ Ναμβυ ανδ Γιοαννι Θονα-Λαςινιο προποςεδ φιελδ τηεορψ μοδελς φορ ιτ[85].
Τηε ςαγα οφ ιδεντιφψινγ τηε αξιαλ ςυρρεντ οφ ωεαϰ ιντεραςτιονς ις ονε οφ τηε γρεατ αςηιεεμεντς οφ τηεορετιςαλ παρτιςλε πηψςιςς οφ τηε εαρλψ ςιξτιες. Ιτ ςυλμινατεδ ωιτη α ςεριες οφ παπερς βψ Γελλ-Μανν ωηο φορμυλατεδ τηε ςομπλετε αλγεβραις ςςηεμε ϰνοων ας “῝υρρεντ Αλγεβρα”[59] [60] ωηιςη ωε μεντιονεδ ιν ςεςτιον 1.
$\bullet$ Σποντανεους ςψμμετρψ βρεαϰινγ ιν τηε πρεςενςε οφ γαυγε ιντεραςτιονς. Τηε πιςτυρε ςηανγες ςομπλετελψ ιφ ωε ςονςιδερ τηε ςαμε μεςηανιςμ ιν τηε πρεςενςε οφ γαυγε ιντεραςτιονς [^53] If the term “spontaneous symmetry breaking” applied to a global symmetry is misleading, that of “spontaneous breaking of a gauge symmetry”, which we often use, is plainly wrong.. Τηε ηιςτορψ ις χυιτε ςομπλιςατεδ βεςαυςε τηε πηενομενον ωας ιν φαςτ διςςοερεδ τωιςε· ςο ωε ηαε τωο παραλλελ ςτοριες ωιτη πεοπλε ηαινγ διφφερεντ μοτιατιονς ανδ οφτεν ιγνορινγ εαςη οτηερ.
῝ηρονολογιςαλλψ, τηε φιρςτ ςτορψ ις τηατ οφ τηε Μειςςνερ εφφεςτ ωηιςη δεςςριβες τηε φαςτ τηατ α μαγνετις φιελδ αππεαρς το βε ςςρεενεδ ανδ δοες νοτ πενετρατε ινςιδε α ςυπερςονδυςτορ. Α πηενομενολογιςαλ δεςςριπτιον οφ τηε εφφεςτ ωας προποςεδ βψ Φριτζ ανδ Ηεινζ Λονδον, ωηο ιντροδυςεδ τηε Λονδον πενετρατιον λενγτη $\lambda$. Λ.Δ. Λανδαυ ανδ ῞.Λ. Γινζβυργ[91], ιν τηειρ φαμους 1950 αρτιςλε, ωροτε τηε φιρςτ μοδελ δεςςριβινγ ςυπερςονδυςτιιτψ ας α πηαςε τρανςιτιον ιν τερμς οφ αν ορδερ παραμετερ $\Psi$· ιτ ις εχυαλ το ζερο ιν τηε νορμαλ πηαςε ανδ ταϰες α νον-ζερο αλυε ιν τηε ςυπερςονδυςτινγ πηαςε. Ιν τερμς οφ $\Psi$, τηε εχυατιον φορ τηε ελεςτρομαγνετις ποτεντιαλ ταϰες τηε φορμ
$\Delta{\bm{A}}=\dots+\frac{4\pi e^{2}}{mc^{2}}|\Psi|^{2}{\bm{A}}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ {\bm{A}}(x)\sim{\bm{A}}(0)\leavevmode\nobreak\ {\rm e}^{-x/\lambda}$ (45)
ωηερε τηε δοτς ςτανδ φορ τερμς ωηιςη αρε νοτ ιμπορταντ φορ τηε διςςυςςιον. Ωε ςεε ιν τηις εχυατιον τηατ $\Psi$ αςτς λιϰε α μαςς τερμ φορ τηε πηοτον ανδ, ας α ρεςυλτ, ιν τηε ςυπερςονδυςτινγ πηαςε, ${\bm{A}}$ δεςαψς εξπονεντιαλλψ.
Ιν 1957, Θ. Βαρδεεν, Λ.Ν. ῝οοπερ ανδ Θ.Ρ. Σςηριεφφερ[92] φορμυλατεδ τηε μιςροςςοπις τηεορψ οφ ςυπερςονδυςτιιτψ, ϰνοων ας τηε Β῝Σ τηεορψ, ιν ωηιςη τηεψ γιε τηε πηψςιςαλ οριγιν οφ τηε ορδερ παραμετερ $\Psi$ ιν τερμς οφ τηε ῝οοπερ παιρς. Τηε φολλοωινγ ψεαρ Π.Ω. Ανδερςον, ιν α ςεριες οφ τωο παπερς[93], ςηοωεδ τηατ, ιν τηε Β῝Σ τηεορψ, τηε πηοτον ινςιδε α ςυπερςονδυςτορ βεςομες εφφεςτιελψ μαςςιε. Τηις ις ωιτνεςςεδ φιρςτ βψ τηε εξιςτενςε οφ α μαςς γαπ ανδ, ςεςονδ, βψ τηε αππεαρανςε οφ ωαες ωιτη λονγιτυδιναλ πολαριςατιον. Τηε ςαμε ςονςλυςιον ωας ρεαςηεδ αλςο βψ Ναμβυ ιν 1960[94] ωηο εξαμινεδ τηε Β῝Σ τηεορψ ιν τηε Ηαρτρεε-Φοςϰ αππροξιματιον. Τηε αρτιςλε, ωηιςη ρεφερς το Ανδερςον, ηας α γοοδ διςςυςςιον οφ γαυγε ιναριανςε. Τηις ςονςλυδες τηε πρεςεντατιον οφ τηε εφφεςτ ιν ςυπερςονδυςτιιτψ ανδ εξπλαινς τηε τιτλε οφ Γολδςτονε´ς παπερ[86].
Ιν παραλλελ, ανδ μοςτλψ ινδεπενδεντλψ, α διφφερεντ ςτορψ δεελοπεδ ιν παρτιςλε πηψςιςς ωιτη τηε μοτιατιον οφ προιδινγ α μαςς το τηε Ψανγ-Μιλλς βοςονς. Ας φαρ ας Ι ϰνοω, τηε φιρςτ περςον το ωορρψ αβουτ τηις προβλεμ ωας Σςηωινγερ ωηο, ιν 1962, ωροτε τωο παπερς[95], βοτη ωιτη τηε τιτλε Γαυγε Ιναριανςε ανδ Μαςς. Τηε φιρςτ ηας α γενεραλ διςςυςςιον αναλψςινγ τηε τηεορετιςαλ βαςις φορ τηε βελιεφ τηατ γαυγε ιναριανςε ιμπλιες α ζερο μαςς φορ τηε γαυγε βοςονς. Ιν α ςιμπλιφιεδ φορμ τηε αργυμεντ ςαν βε ςυμμαριςεδ ας φολλοως: λετ $\Pi_{\mu\nu}(q)$ βε τηε 1Π-Ι τωο-ποιντ φυνςτιον οφ τηε εςτορ βοςον. Ιν α γαυγε ιναριαντ τηεορψ ιτ ηας τηε γενεραλ φορμ
$\Pi_{\mu\nu}(q)=\Pi(q^{2})\left(g_{\mu\nu}-\frac{q_{\mu}q_{\nu}}{q^{2}}\right)$ (46)
Ιφ ωε ρεςπεςτ γαυγε ιναριανςε, ατ ανψ γιεν ορδερ οφ τηε περτυρβατιον εξπανςιον ωε φινδ $\Pi(0)=0$ ανδ τηις ιμπλιες α ζερο μαςς φορ τηε εςτορ βοςον. Ηοωεερ, Σςηωινγερ νοτες τηατ, ιν α νον περτυρβατιε ρεγιμε, νοτηινγ πρεεντς $\Pi(0)$ το βε διφφερεντ φρομ ζερο, τηυς ψιελδινγ α μαςςιε βοςον. Ιν τηε ςεςονδ παπερ ηε ςτυδιες τηε τωο-διμενςιοναλ φιελδ τηεορψ μοδελ, ϰνοων ας τηε Σςηωινγερ μοδελ, ωε ιντροδυςεδ ιν ςεςτιον 1. Ας ωε ςαιδ, τηε μοδελ ις εξαςτλψ ςολαβλε ωιτη τηε ςυρπριςινγ ρεςυλτ τηατ ιτ δεςςριβες α φρεε, μαςςιε βοςον! Σο, Σςηωινγερ ςονςλυδες τηατ τηε “τηεορεμ” τηατ γαυγε ιναριανςε ιμπλιες α ζερο βοςον μαςς ςαν βε προεν ονλψ ιν περτυρβατιον [^54] In these papers Schwinger makes no reference to superconductivity..
Ιτ ςεεμς τηατ Σςηωινγερ ηαδ ρεαςηεδ τηις ςονςλυςιον εαρλιερ. Ατ τηε Ιντερνατιοναλ ῝ονφερενςε ον Ελεμενταρψ Παρτιςλες, ηελδ ατ Αιξ-εν-Προενςε ιν Σεπτ. 14-20 1961, Φεψνμαν, ιν ηις Συμμαρψ Ταλϰ, ςαψς: “…Σινςε γαυγε ιναριανςε ις υςυαλλψ βελιεεδ το ιμπλψ τηατ τηε μαςς [οφ τηε γαυγε βοςονς] ις ζερο, τηε φιρςτ πρεδιςτιον οφ τηεςε τηεοριες…ις διςρεγαρδεδ. Σςηωινγερ ποιντεδ ουτ το με ηοωεερ, τηατ ονε ςαν υςε γαυγε ιναριανςε το προε τηατ τηε μαςς οφ τηε ρεαλ πηοτον ις εχυαλ το ζερο, ονλψ ιφ ονε αςςυμες τηατ ιν τηε ςομπλετε δρεςςεδ πηοτον, τηερε ις α φινιτε αμπλιτυδε το φινδ τηε υνδρεςςεδ ονε.´´ Ιτ ις πρεςιςελψ ωηατ ηαππενς ιν τηε τηεοριες ωε υςε τοδαψ. Ιν α μοδερν λανγυαγε, τηε τωο Ηιλβερτ ςπαςες, τηε φιρςτ βυιλτ αβοε τηε αςυυμ ςονταινινγ τηε “υνδρεςςεδ” πηοτον, ανδ τηε ςεςονδ αβοε τηε αςυυμ οφ τηε “δρεςςεδ” ονε, αρε ορτηογοναλ.
Μοτιατεδ βψ Σςηωινγερ´ς ρεμαρϰς, πεοπλε λοοϰεδ αγαιν ατ τηε γαυγε ιναριανςε – μαςς ςοννεςτιον. Ι ωαντ το μεντιον τωο ςοντριβυτιονς: Ιν 1963 Ανδερςον ωροτε α παπερ[96] ωιτη τηε τιτλε “Πλαςμονς, Γαυγε ιναριανςε ανδ Μαςς” ιν ωηιςη ηε πρεςεντς ηις πρειους ρεςυλτς ον πλαςμα ωαες ας αν εξαμπλε οφ μαςς γενερατιον. Ιν τηε Αβςτραςτ ηε ωριτες : “Σςηωινγερ ηας ποιντεδ ουτ τηατ τηε Ψανγ-Μιλλς εςτορ βοςον [(Ηε ονλψ ςονςιδερς τηε αβελιαν τηεορψ)] …δοες νοτ νεςεςςαριλψ ηαε ζερο μαςς.…Ωε ςηοω τηατ τηε τηεορψ οφ πλαςμα οςςιλλατιονς ις α ςιμπλε νον-ρελατιιςτις εξαμπλε εξηιβιτινγ αλλ οφ τηε φεατυρες οφ Σςηωινγερ’ς ιδεα.” Ατ τηε ςαμε τιμε Μαυριςε Λéψ ωροτε α νον-λοςαλ γαυγε ιναριαντ ερςιον οφ ΧΕΔ ωιτη α μαςςιε πηοτον[97]. Σο, βψ τηε εαρλψ ςιξτιες, τηε προβλεμ οφ τηε γαυγε βοςον μαςςες ωας ςλεαρλψ ςετ. Ωε ηαδ α ςονςρετε εξαμπλε οφ μαςς γενερατιον ιν α νον-ρελατιιςτις τηεορψ, βυτ νο εξπλιςιτ ρελατιιςτις φιελδ τηεορψ μοδελς. Ας υςυαλλψ, α ωελλ φορμυλατεδ προβλεμ ις ςοον ςολεδ. Ιν τηε φολλοωινγ ψεαρς ωε ηαε αν ααλανςηε οφ παπερς:
Β.Ω. Λεε ανδ ςολλαβορατορς[98] ςονςιδερεδ μοδελς ωιτη ςποντανεους ςψμμετρψ βρεαϰινγ ανδ ςαμε ερψ ςλοςε το τηε ςολυτιον ωε ϰνοω τοδαψ.
Ω. Γιλβερτ[99] ςριτιςιςεδ τηεςε αττεμπτς ανδ ςλαιμεδ τηατ τηε ςυπερςονδυςτιιτψ εξαμπλε ωας νοτ ρελεαντ φορ α ρελατιιςτις τηεορψ. Ηις αργυμεντ ις βαςεδ ον α νεω προοφ οφ τηε Γολδςτονε τηεορεμ. Ηε αργυεδ τηατ ιν α Λορεντζ ιναριαντ τηεορψ, τηερε ις αλωαψς α μαςςλεςς βοςον αςςοςιατεδ το εερψ γενερατορ οφ α ςποντανεουςλψ βροϰεν ςψμμετρψ. Ιτ ωας τηε φιρςτ μοδελ ινδεπενδεντ προοφ οφ τηις τηεορεμ. ῞ερψ βριεφλψ, ιτ γοες ας φολλοως: Λετ υς ςονςιδερ α ρελατιιςτις τηεορψ ιναριαντ υνδερ α γρουπ οφ τρανςφορματιονς. Βψ Νοετηερ´ς τηεορεμ ωε ηαε α ςονςερεδ ςυρρεντ $J_{\mu}({\bm{x}},t)$ ανδ τηε αςςοςιατεδ τιμε-ινδεπενδεντ ςηαργε $Q$. Ωε αςςυμε τηατ τηε ςψμμετρψ ις νον-τριιαλ ανδ, τηερεφορε, τηερε εξιςτς ατ λεαςτ ονε λοςαλ οπερατορ $A(x)$ ωηιςη τρανςφορμς νον τριιαλλψ.
$\partial^{\mu}J_{\mu}=0\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ Q=\int{\rm d}{\bm{x}}J_{0}({\bm{x}},t)\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ [Q,A]=\delta A\leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \delta A\neq 0$ (47)
Ωε νοω ςυπποςε τηατ τηε ςψμμετρψ γενερατεδ βψ $Q$ ις ςποντανεουςλψ βροϰεν. Τηις ιμπλιες τηατ τηε αςυυμ ςτατε |$\ket{\Omega}$ ις δεγενερατε ανδ ιτ ις νοτ αννιηιλατεδ βψ $Q$. Λετ υς ςονςιδερ τηε φολλοωινγ Γρεεν φυνςτιον
${\cal A}_{\mu}(k)=\int{\rm d}^{4}x{\rm e}^{{\rm i}kx}\bra{\Omega}[j_{\mu}(x),A(0)]\ket{\Omega}=k_{\mu}F(k^{2})$ (48)
ωιτη $F$ ςομε φυνςτιον οφ $k^{2}$. Ιν α Λορεντζ ιναριαντ τηεορψ τηις ις τηε μοςτ γενεραλ φορμ οφ ${\cal A}_{\mu}(k)$. Φυρτηερμορε, τηε νον-ανιςηινγ οφ $\delta A$, ιμπλιες τηατ $F$ ις νοτ εχυαλ το ζερο. Νοω ωε μυλτιπλψ βψ $k^{\mu}$ ανδ υςε τηε ςονςερατιον οφ τηε ςυρρεντ. Ωε φινδ:
$k^{\mu}{\cal A}_{\mu}=0\leavevmode\nobreak\ \leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ k^{2}F(k^{2})=0\leavevmode\nobreak\ \leavevmode\nobreak\ \Rightarrow\leavevmode\nobreak\ F(k^{2})\sim\delta(k^{2})$ (49)
ωηιςη ςηοως τηε εξιςτενςε οφ α ζερο μαςς ςινγυλαριτψ, ιντερπρετεδ ας τηε πρεςενςε οφ α ζερο μαςς παρτιςλε. Γιλβερτ ςονςλυδες τηατ, ιν α ρελατιιςτις τηεορψ, ιτ ις ιμποςςιβλε το ηαε ςποντανεους ςψμμετρψ βρεαϰινγ ωιτηουτ μαςςλεςς παρτιςλες.
Τηε ςαμε ψεαρ, ανδ ινδεπενδεντλψ οφ Γιλβερτ [^55] Gilbert’s paper has submission date March 30 and Englert’s and Brout’s June 26. There was no ArXiv or internet at that time., Φ. Ενγλερτ ανδ Ρ. Βρουτ πυβλιςηεδ τηειρ παπερ[100] ωηιςη ςονταινς τηε ςομπλετε ςολυτιον, ας ωε ϰνοω ιτ τοδαψ, υςινγ ελεμενταρψ ςςαλαρ φιελδς. Τηεψ ςονςιδερ βοτη, αβελιαν ανδ νον-αβελιαν γαυγε τηεοριες. Ας μοςτ παπερς οφ τηατ περιοδ, τηε εμπηαςις ωας ον μοδελς φορ ςτρονγ ιντεραςτιονς. Τηεψ διςςυςς ςηιραλ τηεοριες ανδ αλςο μεντιον τηε ποςςιβιλιτψ οφ α δψναμιςαλ ςψμμετρψ βρεαϰινγ. Τηεψ ρεφερ το Ναμβυ φορ ΣΣΒ ανδ Σςηωινγερ. Τηεψ ηαε νο ρεφερενςε το ςυπερςονδυςτιιτψ.
Στιλλ ιν 1964, Π. Ηιγγς πυβλιςηεδ τωο παπερς[101]. Ιν τηε φιρςτ ηε ανςωερς Γιλβερτ´ς οβϑεςτιον βψ ςονςιδερινγ αν αβελιαν γαυγε τηεορψ ιν τηε ῝ουλομβ γαυγε. Εξπλιςιτ Λορεντζ ςοαριανςε ις λοςτ ανδ Ηιγγς ςηοως τηατ Γιλβερτ´ς αργυμεντ δοες νοτ αππλψ. Ας α ρεςυλτ, τηε πηοτον βεςομες μαςςιε. Ηοωεερ ηε αργυες τηατ πηψςιςαλ χυαντιτιες, ςυςη ας τηε μαςς ςπεςτρυμ, ςαννοτ δεπενδ ον τηε γαυγε ανδ τηεψ ςηουλδ βε τηε ςαμε εεν ιν α ςοαριαντ γαυγε. Ιν τηε ςεςονδ παπερ ηε πρεςεντς α φυλλ ςτυδψ οφ τηε αβελιαν μοδελ ανδ ςονςιδερς εξτενςιονς το ινςλυδε τηε βρεαϰινγ οφ φλαορ $SU(3)$. Ηε μαϰες τηε ςοννεςτιον βετωεεν τηε ωουλδ-βε Γολδςτονε βοςονς ανδ τηε λονγιτυδιναλ δεγρεες οφ φρεεδομ οφ τηε γαυγε παρτιςλες ανδ ςτατες εξπλιςιτλψ τηε πρεδιςτιον φορ α πηψςιςαλ ςςαλαρ παρτιςλε. Ηε ρεφερς το τηε Ενγλερτ ανδ Βρουτ παπερ, βυτ αλςο το Ανδερςον ανδ ςυπερςονδυςτιιτψ.
Βεφορε γοινγ ον λετ υς γιε α φιρςτ ιδεα οφ ωηατ ωεντ ωρονγ ωιτη Γιλβερτ´ς προοφ. Α ματηεματιςαλλψ μορε ςομπλετε εξπλανατιον ωιλλ βε διςςυςςεδ πρεςεντλψ. Ατ φιρςτ ςιγητ Γιλβερτ´ς αςςυμπτιονς, αλτηουγη νοτ εξπλιςιτλψ ςπελλεδ ουτ, αρε τηε ςτανδαρδ ονες οφ ρελατιιςτις χυαντυμ φιελδ τηεορψ, το ωιτ Λορεντζ ιναριανςε, λοςαλιτψ ετς. Ηε ςηοως τηε εξιςτενςε οφ α ζερο μαςς ςινγυλαριτψ ιν τηε Γρεεν φυνςτιονς οφ εχ. (48). Βεφορε γοινγ ον, λετ με εμπηαςιςε τηατ αλλ εξπλιςιτ ςαλςυλατιονς ιν περτυρβατιον υςινγ αριους Λορεντζ ςοαριαντ φιελδ τηεορψ μοδελς, ςονφιρμ Γιλβερτ´ς ρεςυλτ: τηε ζερο μαςς ςινγυλαριτψ ις τηερε. Φρομ τηατ ηε ωαντς το ινφερ τηε εξιςτενςε οφ α ζερο μαςς παρτιςλε. Βυτ φορ τηις ςονςλυςιον ψου νεεδ το αςςυμε τηατ τηερε εξιςτς νο οτηερ ζερο μαςς ςινγυλαριτψ ωιτη τηε ωρονγ ςιγν ωηιςη ςανςελς τηε φιρςτ ονε ιν αλλ $S$-ματριξ ελεμεντς. Φορ τηατ ψου νεεδ το αςςυμε τηατ αλλ ςτατες ιν τηε Ηιλβερτ ςπαςε ηαε ποςιτιε νορμ. Τηις ις παρτ οφ τηε υςυαλ αξιομς οφ χυαντυμ φιελδ τηεορψ, βυτ ωε αλλ ϰνοω τηατ ιτ δοες νοτ αππλψ το α γαυγε τηεορψ. Ταϰε ΧΕΔ. Ωε ςαν υςε α νον ςοαριαντ γαυγε, λιϰε τηε ῝ουλομβ γαυγε, ιν ωηιςη ονλψ πηψςιςαλ ποςιτιε νορμ ςτατες αππεαρ, βυτ εξπλιςιτ Λορεντζ ςοαριανςε ις λοςτ. Ιτ ωας Ηιγγς´ αργυμεντ. Αλτερνατιελψ, ωε ςαν υςε α ςοαριαντ γαυγε, βυτ τηεν ωε ςηουλδ χυαντιςε τηε τηεορψ φολλοωινγ τηε Γυπτα-Βλευλερ μετηοδ ιν ωηιςη ωε ηαε νεγατιε νορμ ςτατες. Ωε ςαν ςηεςϰ βψ εξπλιςιτ ςαλςυλατιονς τηατ τηε ζερο μαςς ςινγυλαριτιες δο ςανςελ ιν αλλ πηψςιςαλ ματριξ ελεμεντς.
Τηε φιναλ ςτονε ιν τηε εδιφιςε ωας λαιδ δοων βψ Γ.Σ. Γυραλνιϰ, ῝.Ρ. Ηαγεν ανδ Τ.Ω.Β. Κιββλε[102], αγαιν ιν 1964. Τηεψ πρεςεντ α δεταιλεδ αναλψςις οφ τηε αβελιαν μοδελ ωιτη αλλ αςπεςτς ωε υςε τοδαψ. Ιν παρτιςυλαρ, τηε ςουντινγ οφ δεγρεες οφ φρεεδομ: 2 πολαριςατιον ςτατες οφ α μαςςλεςς γαυγε βοςον + 1 δεγρεε οφ φρεεδομ οφ τηε ωουλδ-βε Γολδςτονε ςςαλαρ = 3 πολαριςατιον ςτατες οφ α μαςςιε εςτορ βοςον, αππεαρς τηερε. Τηεψ εξπλαιν Γιλβερτ´ς ρεςυλτ βψ ποιντινγ ουτ τηατ, φορ α γαυγε τηεορψ, τηε ιντεγραλ οερ αλλ τηρεε ςπαςε γιινγ τηε ςηαργε οπερατορ ιν εχυατιον (47) δοες νοτ ςονεργε – βεςαυςε γαυγε τηεοριες ινδυςε λονγ ρανγε ςορρελατιονς – ανδ ονε ςηουλδ φιρςτ ςομπυτε τηε ιντεγραλ ιν α φινιτε ολυμε $V$ ανδ ςτυδψ ςαρρεφυλλψ τηε λιμιτ $V\rightarrow\infty$. Αμονγ τηε ρεφερενςες τηεψ ινςλυδε Γολδςτονε, Γιλβερτ, Ενγλερτ ανδ Βρουτ ωηιςη ωας αλρεαδψ πυβλιςηεδ ανδ Ηιγγς ας α πρεπριντ. Τηεψ μεντιον ςυπερςονδυςτιιτψ ωιτη νο δεταιλεδ ρεφερενςες.
[…]
Τηε Ψανγ-Μιλλς τηεοριες ανδ τηε ΒΕΗ μεςηανιςμ ωερε, ιν πρινςιπλε, ϰνοων ςινςε 1964. Ι ςαψ “ιν πρινςιπλε” βεςαυςε ερψ φεω ηαδ νοτιςεδ τηεμ ανδ πραςτιςαλλψ νο ονε ωας εξπεςτινγ το υςε τηεμ ιν ωεαϰ ιντεραςτιονς. Τηερεφορε, ιτ ις νοτ ςυρπριςινγ τηατ νοβοδψ νοτιςεδ α ςηορτ λεττερ ωιτη τηε ρατηερ υνινςπιρεδ τιτλε Α μοδελ οφ λεπτονς[103]. Ι ωας ατ ῝ΕΡΝ ατ τηε τιμε ανδ, τογετηερ ωιτη οτηερ φελλοως, ωε ηαδ α ςτυδψ γρουπ διςςυςςινγ τηε παπερς ιν τηε λιτερατυρε. Ι ρεμεμβερ ιτ ωας τηε λατε Βρυνο Ρεννερ ωηο ρεπορτεδ ον Ωεινβεργ´ς παπερ ανδ ωε αλλ δεςιδεδ ιτ ωας υνιντερεςτινγ! Ωε προμπτλψ φοργοτ εερψτηινγ αβουτ ιτ. Ανδ ωε ωερε νοτ τηε ονλψ ονες. Ωιτη πρεςεντ ϰνοωλεδγε ωε ςεε ιμμεδιατελψ τηατ Ωεινβεργ ιν α ςινγλε ςτροϰε ςολες τηρεε φυνδαμενταλ προβλεμς: φιρςτ, ανδ μοςτ ιμπορταντ, ηε υςες τηε Βρουτ-Ενγλερτ-Ηιγγς (ΒΕΗ) μεςηανιςμ το γιε μαςςες το τηε ιντερμεδιατε γαυγε βοςονς. Σεςονδ, ηε ςηοως τηατ τηε ςαμε μεςηανιςμ γενερατες α μαςς τερμ φορ τηε ςηαργεδ λεπτον. Ινδεεδ, ςινςε τηε ριγητ ανδ λεφτ ςομπονεντς οφ τηε Διρας ςπινορς τρανςφορμ διφφερεντλψ υνδερ τηε γαυγε γρουπ, α διρεςτ μαςς τερμ ις φορβιδδεν. Τηιρδ, τηε ςαμε μεςηανιςμ ψετ αγαιν γιες ριςε το Γλαςηοω´ς μιξινγ, ανδ προδυςες α μαςς φορ τηε $Z^{0}$ ωηιλε λεαινγ τηε πηοτον μαςςλεςς. Τηεν, ωηψ διδ ωε αλλ φαιλ το υνδερςτανδ τηε ςιγνιφιςανςε οφ τηις παπερ; Ι βελιεε τηατ τηε μαιν ρεαςον ωας λαςϰ οφ ιντερεςτ ιν τηε ςομμυνιτψ. Φιελδ τηεορψ ωας νοτ ςονςιδερεδ ας α ςεριους τοολ το μαϰε προγρεςς ιν παρτιςλε πηψςιςς. Ουτ οφ τηε τηρεε προβλεμς τηε παπερ ςολεδ, νονε ωας ρεςογνιςεδ ας βεινγ ιμπορταντ. Ταϰε, φορ εξαμπλε, τηε προβλεμ οφ τηε γαυγε βοςον μαςςες: ςινςε τηεψ ωερε νοτ ϰνοων το εξιςτ, φεω πεοπλε ςαρεδ αβουτ τηε ωαψ τηεψ ωερε γεττινγ τηειρ μαςςες. Α ςεςονδ προβλεμ ωιτη τηε παπερ ωας τηε τιτλε: “Α Μοδελ οφ Λεπτονς”. Τηε ονλψ λεπτονις ωεαϰ ιντεραςτιον ϰνοων ατ τηατ τιμε ωας μυον δεςαψ. Νοβοδψ ωαντεδ το ςονςιδερ ςυςη α ςομπλιςατεδ-λοοϰινγ τηεορψ φορ ϑυςτ α ςινγλε προςεςς. Ιτ ςουνδεδ λιϰε υςινγ ηεαψ αρτιλλερψ το ϰιλλ α φλψ. Φυρτηερμορε, ιφ τηερε ωας ονε προπερτψ οφ ωεαϰ ιντεραςτιονς ωηιςη εερψβοδψ λιϰεδ ανδ ρεςπεςτεδ, ωας πρεςιςελψ τηε προπερτψ οφ υνιερςαλιτψ, ςο “Α Μοδελ οφ Λεπτονς” ωεντ τοταλλψ υννοτιςεδ. Τηις λαςϰ οφ αττεντιον ωηιςη τηις παπερ ηας ραιςεδ ις ρεφλεςτεδ ιν τηε ςιτατιονς ιτ ρεςειεδ. Ιτ ωας πυβλιςηεδ ιν 1967 βυτ ιτ ηας πραςτιςαλλψ νο ςιτατιονς υντιλ 1972. Ωηεν ωε ωροτε τηε παπερ ον ςηαρμ ωιτη Σ.Λ. Γλαςηοω ανδ Λ. Μαιανι ιν 1970[104], νονε οφ υς ηαδ ανψ ρεςολλεςτιον οφ Ωεινβεργ’ς παπερ ανδ ωε δο νοτ ρεφερ το ιτ. Εεν μορε υνβελιεαβλε, ωε γαε α ςεμιναρ ατ ΜΙΤ ιν ωηιςη ωε πρεςεντεδ τηε ωεαϰ ιντεραςτιονς, βοτη λεπτονις ανδ ηαδρονις, ας α Ψανγ Μιλλς τηεορψ. Ωεινβεργ, ωηο ωας ιν τηε αυδιενςε, αςϰεδ μανψ χυεςτιονς βυτ, εεν ηε, διδ νοτ μαϰε τηε ςοννεςτιον ωιτη ηις πρειους ωορϰ[105]. Οφ ςουρςε, ςταρτινγ ιν 1972, τηε νυμβερ οφ ςιτατιονς ςϰψροςϰετεδ.
[…]
Ωηερε ις τηε ςυτ-οφφ, ορ, τηε ιταλ ιμπορτανςε οφ πρεςιςιον μεαςυρεμεντς. Ωε ηαε ςεεν τηε Φερμι τηεορψ οφ ωεαϰ ιντεραςτιονς ας α μοςτ ςεμιναλ ανδ ινςπιρινγ μοδελ. Μανψ διςςοεριες οφ φυνδαμενταλ ιμπορτανςε – $CVC,PCAC$, ῝ηιραλ ςψμμετρψ, Σποντανεους ςψμμετρψ βρεαϰινγ, ῝υρρεντ αλγεβρα – ωερε μαδε ας α ρεςυλτ οφ τηε εφφορτς το υνδερςτανδ τηε προπερτιες οφ τηε ωεαϰ ςυρρεντ. Ηοωεερ, φρομ τηε ματηεματιςαλ ποιντ οφ ιεω, Φερμι´ς τηεορψ ις νοτ ρενορμαλιςαβλε ανδ, τηερεφορε, ιτ ις ονλψ α πηενομενολογιςαλ μοδελ. Ιν πραςτιςαλ τερμς τηις μεανς τηατ, ιφ ωε ωριτε ανψ πηψςιςαλ αμπλιτυδε ας α ποωερ ςεριες ιν τηε Φερμι ςουπλινγ ςονςταντ $G_{F}$, εερψ τερμ ιν τηε εξπανςιον ρεχυιρες τηε ιντροδυςτιον οφ α ςυτ-οφφ παραμετερ $\Lambda$. Ιν α ρενορμαλιςαβλε τηεορψ, ςυςη ας χυαντυμ ελεςτροδψναμιςς, τηερε εξιςτς α ωελλ-δεφινεδ πρεςςριπτιον το ταϰε τηε λιμιτ $\Lambda\rightarrow\infty$ ανδ οβταιν υναμβιγυους ρεςυλτς, βυτ το α νον-ρενορμαλιςαβλε τηεορψ τηε πρεςςριπτιον δοες νοτ αππλψ. Τηε ςυτ-οφφ μυςτ ρεμαιν φινιτε ανδ ιτς αλυε δετερμινες τηε ενεργψ ςςαλε αβοε ωηιςη τηε τηεορψ ςαννοτ βε τρυςτεδ. Τηις ις τηε δεφινιτιον οφ αν εφφεςτιε τηεορψ.
῝αν ωε εςτιματε αν ορδερ οφ μαγνιτυδε φορ τηε ςυτ-οφφ; Α ερψ ςιμπλε μετηοδ ις τηε φολλοωινγ: Ορδιναρψ διμενςιοναλ αναλψςις τελλς υς τηατ α πηψςιςαλ χυαντιτψ ${\cal A}$, φορ εξαμπλε α ωεαϰ δεςαψ αμπλιτυδε, ςαν βε ωριττεν ιν α ςεριες εξπανςιον ας:
${\cal A}=A_{1}G_{F}\left(1+\sum_{n=2}^{\infty}A_{n}(G_{F}\Lambda^{2})^{n-1}\right)$ (50)
ωηερε, ιν εερψ ορδερ οφ τηε εξπανςιον, ωε ηαε ϰεπτ ονλψ τηε ηιγηεςτ ποωερ ιν $\Lambda$. Ωε ςεε τηατ τηε εξπρεςςιον $g_{\mathrm{eff}}=G_{F}\Lambda^{2}$ αςτς ας αν εφφεςτιε, διμενςιονλεςς ςουπλινγ ςονςταντ. Τηε εξπανςιον ωιλλ βεςομε μεανινγλεςς ωηεν $g_{\mathrm{eff}}\sim 1$, ωηιςη, φορ τηε νυμεριςαλ αλυε οφ $G_{F}$, γιες $\Lambda\sim$ 300 Γε῞, α αλυε ωηιςη, φορ τηε αςςελερατορς οφ τηε 1960´ς, ωας εςςεντιαλλψ ινφινιτε [^56] It is what happens in the Standard Model where the role of $\Lambda$ is played by the vector boson masses..
Ιτ ωας Β.Λ. Ιοφφε ανδ Ε.Π. Σηαβαλιν[106], φρομ τηε Σοιετ ϒνιον, ωηο φιρςτ ρεμαρϰεδ τηατ, ιν φαςτ, ονε ςαν δο μυςη βεττερ. Λετ υς γο βαςϰ το τηε εξπανςιον (50) ανδ ςονςιδερ αλςο τηε ςυβ-δομιναντ τερμς ιν ποωερς οφ $\Lambda$. Ωε ςαν ρεπηραςε τηειρ αργυμεντ ανδ ωριτε ανψ πηψςιςαλ χυαντιτψ ας α δουβλε εξπανςιον ιν $g_{\mathrm{eff}}$ ανδ $G_{F}$:
${\cal A}=\sum_{n=0}^{\infty}A_{n}^{(0)}g_{\mathrm{eff}}^{n}\leavevmode\nobreak\ +\leavevmode\nobreak\ G_{F}M^{2}\sum_{n=0}^{\infty}A_{n}^{(1)}g_{\mathrm{eff}}^{n}\leavevmode\nobreak\ +\leavevmode\nobreak\ (G_{F}M^{2})^{2}\sum_{n=0}^{\infty}A_{n}^{(2)}g_{\mathrm{eff}}^{n}\leavevmode\nobreak\ +\leavevmode\nobreak\ \dots$ (51)
ωηερε τηε χυαντιτιες $A_{n}^{(i)}$ μαψ ςονταιν ποωερς οφ τηε λογαριτημ οφ $\Lambda$. $M$ ις ςομε μαςς παραμετερ, ωηιςη, φορ α τψπιςαλ χυαντιτψ ιν παρτιςλε πηψςιςς, ις οφ τηε ορδερ οφ 1 Γε῞. Τηε φιρςτ ςεριες ςονταινς τηε τερμς ωιτη τηε μαξιμυμ ποωερ οφ $\Lambda$ φορ α γιεν ποωερ οφ $G_{F}$, τηεψ αρε ςαλλεδ τηε λεαδινγ διεργενςες. Σιμιλαρλψ, τηε ςεςονδ ςεριες ςονταινς αλλ τηε νεξτ-το-λεαδινγ διεργενςες, τηε τηιρδ τηε νεξτ-το-νεξτ-το-λεαδινγ διεργενςες, ετς. Φολλοωινγ Ιοφφε ανδ Σηαβαλιν, λετ υς ςηοοςε φορ ${\cal A}$ α χυαντιτψ ιν ςτρονγ ιντεραςτιονς, φορ εξαμπλε τηε ενεργψ λεελς ιν α νυςλευς. Τηε λεαδινγ διεργενςες ρεπρεςεντ τηε ωεαϰ ιντεραςτιον ςορρεςτιονς το τηις χυαντιτψ. Ωε ςαν ιςυαλιςε τηεμ ας ρεςυλτινγ φρομ τηε εμιςςιον ανδ ρε-αβςορπτιον οφ α ιρτυαλ $W$. Βυτ ωεαϰ ιντεραςτιονς ιολατε παριτψ ανδ/ορ ςτρανγενεςς, τηερεφορε τηε ηιγη πρεςιςιον ωιτη ωηιςη ςυςη εφφεςτς αρε ϰνοων το βε αβςεντ ιν νυςλεαρ πηψςιςς γιες α μυςη μορε ςτρινγεντ βουνδ φορ $\Lambda$, οφ τηε ορδερ οφ 2-3 Γε῞. Σιμιλαρλψ τηε νεξτ-το-λεαδινγ διεργενςες ςοντριβυτε το “φορβιδδεν” ωεαϰ ιντεραςτιον προςεςςες, ςυςη ας $\Delta S$=2 τρανςιτιονς (τηε $K^{0}_{L}-K^{0}_{S}$ μαςς διφφερενςε), ορ $K^{0}_{L}\rightarrow\mu^{+}\mu^{-}$ δεςαψς. Αγαιν, τηε πρεςιςιον μεαςυρεμεντς οφ ςυςη χυαντιτιες γιε τηε ςαμε 2-3 Γε῞ λιμιτ φορ $\Lambda$. Τηις ωας οβιουςλψ υναςςεπταβλε βεςαυςε ιτ ωουλδ μεαν τηατ τηε βρεαϰδοων οφ Φερμι´ς τηεορψ ςηουλδ ηαδ βεεν ςεεν αλρεαδψ. Ωηψ διδ μοςτ πεοπλε νοτ ωορρψ αβουτ ιτ; Ι ςυςπεςτ τηατ τηε μαιν ρεαςον ωας α ωιδεςπρεαδ μιςτρυςτ τοωαρδς φιελδ τηεορψ ιν γενεραλ ανδ ηιγηερ-ορδερ διαγραμς ιν παρτιςυλαρ. Σινςε ωε ηαε νο τηεορψ, ωηψ βοτηερ αβουτ ιτς ηιγηερ-ορδερ εφφεςτς; Φορ μοςτ πηψςιςιςτς οφ τηε νινετεεν-ςιξτιες, ςτυδψινγ τηε ηιγηερ ορδερ εφφεςτς οφ ωεαϰ ιντεραςτιονς ωας υςελεςς ιν τηε αβςενςε οφ α φυλλ τηεορψ οφ ςτρονγ ιντεραςτιονς. Τηερε ωερε μανψ ςπεςυλατιονς αβουτ ςτρονγ ιντεραςτιονς προιδινγ τηε ςυτ-οφφ φορ τηε ωεαϰ ονες. Πεοπλε ηαδ νοτ υνδερςτοοδ τηις προφουνδ ςοννεςτιον βετωεεν ςτρονγ ανδ ωεαϰ ιντεραςτιονς, ωηιςη ις μανιφεςτ βψ τηε ωεαϰ ςυρρεντ βεινγ α γενερατορ οφ τηε ςτρονγ ιντεραςτιον ςψμμετριες [^57] Note that Ioffe and Shabalin[106] had already pointed out that the divergences of weak interactions could not be affected by strong interactions, as long as the latter satisfy current algebra..
[…]
Βψ τηε μιδδλε ςιξτιες α λονγ ανδ παινφυλ ςτρυγγλε αγαινςτ τηε ινφινιτιες ςταρτεδ. Αλτηουγη ιτ ωας φουγητ βψ φεω πεοπλε, ιτ ηας βεεν αν επις βαττλε γιεν ιν τωο φροντς: Τηε φιρςτ, τηε πηενομενολογψ φροντ, ορ τηε βοττομ-υπ αππροαςη, αιμεδ ατ φινδινγ τηε νεςεςςαρψ μοδιφιςατιονς το τηε τηεορψ ιν ορδερ το ελιμινατε τηε διςαςτρους λεαδινγ ανδ νεξτ-το-λεαδινγ διεργενςες. Τηε ςεςονδ, τηε φιελδ τηεορψ φροντ, ορ τηε τοπ-δοων αππροαςη, τριεδ το φινδ τηε ςονδιτιονς υνδερ ωηιςη α χυαντυμ φιελδ τηεορψ ινολινγ μαςςιε, ςηαργεδ, εςτορ βοςονς ις ρενορμαλιςαβλε. Ιτ τοοϰ τηε ςυςςεςς ιν βοτη φροντς το ςολε τηε προβλεμ.
[…]
῝αν ωε δετερμινε τηε ῝αβιββο ανγλε; Αρε ωε ρεαδψ το ςαςριφιςε ελεγανςε; Ιν τηε εαρλψ αττεμπτς τηε εφφορτ ωας νοτ φοςυςεδ ον α παρτιςυλαρ πηψςιςαλ προβλεμ, βυτ αιμεδ ινςτεαδ ατ ελιμινατινγ τηε διεργενςες, ατ λεαςτ φρομ πηψςιςαλλψ μεαςυραβλε χυαντιτιες. Σομε ωερε ερψ ινγενιους, βυτ λαςϰ οφ ςπαςε δοες νοτ αλλοω με το πρεςεντ τηεμ ιν ανψ δεταιλ. Α ερψ ινςομπλετε λιςτ ςονταινς:
  • Τηε πηψςιςαλ Ηιλβερτ ςπαςε ςονταινς ςτατες ωιτη νεγατιε μετρις[107]. Τηε ιντροδυςτιον οφ νεγατιε μετρις ςτατες ις ςονςιδερεδ υναςςεπταβλε βεςαυςε ιτ ιμπλιες ιολατιον οφ τηε υνιταριτψ ςονδιτιον. Ηοωεερ, Τ.Δ. Λεε ανδ Γ.῝. Ωιςϰ οβςερεδ τηατ, ιφ τηε ςορρεςπονδινγ “παρτιςλες”, ιν τηις ςαςε τηε ωεαϰ εςτορ βοςονς, αρε ερψ ςηορτ λιεδ, τηε ρεςυλτινγ υνιταριτψ ιολατιονς ςουλδ βε ςονφινεδ ιντο ερψ ςηορτ τιμες ανδ βε υνδετεςταβλε.
  • Τηε ῞-Α φορμ οφ τηε Φερμι τηεορψ ις αν ιλλυςιον ανδ, ιν ρεαλιτψ, τηε ιντερμεδιατε βοςονς μεδιατινγ ωεαϰ ιντεραςτιονς αρε ςςαλαρς[108]. Βψ α Φιερζ τρανςφορματιον, τηε εφφεςτιε Λαγρανγιαν ςουλδ λοοϰ λιϰε α εςτορ τηεορψ φορ ςομε προςεςςες. Τηις ωαψ τηε τηεορψ ις ρενορμαλιςαβλε, βυτ ατ τηε πριςε οφ λοοςινγ αλλ ινςιγητ ιντο τηε φυνδαμενταλ ρολε οφ τηε ωεαϰ ςυρρεντς.
  • Τηε τηεορψ (10) ις αν αππροξιματιον ανδ τηε ρεαλ τηεορψ ςονταινς α λαργε νυμβερ οφ ιντερμεδιαριες ωιτη ςουπλινγς αρρανγεδ το ςανςελ τηε μοςτ δανγερους διεργενςες[109]. Τηε ιδεα ωας ςιμπλε: διεργενςες αριςε ιν περτυρβατιον τηεορψ βεςαυςε α μαςςιε εςτορ βοςον ηας α προπαγατορ ωηιςη βεηαες λιϰε α ςονςταντ ατ λαργε μομεντα. Τηις βεηαιουρ ςαννοτ βε ιμπροεδ ωιτηουτ ιολατινγ υνιταριτψ. Ηοωεερ, φορ α ματριξ αλυεδ φιελδ, ωε ςαν οβταιν ςανςελατιονς φορ ςομε ματριξ ελεμεντς. Ωιτη α ςλεερ αρρανγεμεντ οφ τηε ςουπλινγς, ωε ςαν ηιδε αλλ βαδ διεργενςες φρομ τηε πηψςιςαλλψ ρελεαντ χυαντιτιες. Α ςιμπλε ιδεα ωηοςε ιμπλεμεντατιον τυρνεδ ουτ το βε ερψ ςομπλιςατεδ [^58] S.L. Glashow’s remark: “Few would concede so much sacrifice of elegance to expediency.”.
  • Τηε ωεαϰ ιντεραςτιον διεργενςες ανδ τηε αλυε οφ τηε ῝αβιββο ανγλε[110][111]. Τηε ιδεα ωας το ςομπυτε τηε ςοεφφιςιεντ οφ τηε διεργεντ τερμ, φορ εξαμπλε ιν α λοοπ εξπανςιον, φορ βοτη τηε ωεαϰ ανδ τηε ελεςτρομαγνετις ςοντριβυτιονς. Σεττινγ τηις ςοεφφιςιεντ εχυαλ το ζερο γιες αν εχυατιον φορ τηε ῝αβιββο ανγλε. Τηε ωορϰ βψ ιτςελφ ηας τοδαψ ονλψ α ηιςτοριςαλ ιντερεςτ, βυτ, ας βψ-προδυςτς, τωο ιντερεςτινγ ρεςυλτς εμεργεδ, ςυμμαριςεδ ιν τηε φολλοωινγ τωο ρελατιονς:
    $\tan\theta=\sqrt{\frac{m_{d}}{m_{s}}}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \frac{|m_{d}-m_{u}|}{m_{d}+m_{u}}\sim{\cal O}(1)$ (52)
    ωηερε τηε μαςςες αρε τηοςε οφ τηε τηρεε χυαρϰς. Τηε φιρςτ ις ιν γοοδ αγρεεμεντ ωιτη εξπεριμεντ ανδ ρελατες τηε ῝αβιββο ανγλε ωιτη τηε μεδιυμ ςτρονγ ιντεραςτιονς ωηιςη βρεαϰ $SU(3)$. Τηε ςεςονδ, οβταινεδ βψ ῝αβιββο ανδ Μαιανι, ις μορε ςυβτλε: Τηε πρεαιλινγ πηιλοςοπηψ ωας τηατ ιςοςπιν ις αν εξαςτ ςψμμετρψ φορ ςτρονγ ιντεραςτιονς βροϰεν ονλψ βψ ελεςτρομαγνετις εφφεςτς. Ιν τηις ςαςε ονε ωουλδ εξπεςτ τηε μαςς διφφερενςε ιν α δουβλετ το βε μυςη ςμαλλερ τηαν τηε μαςςες τηεμςελες. Τηε ςεςονδ ρελατιον οφ (52) ςηοως ινςτεαδ τηατ ιςοςπιν ις βαδλψ βροϰεν ιν τηε χυαρϰ μαςςες ανδ τηε αππροξιματε ιςοςπιν ςψμμετρψ ιν ηαδρον πηψςιςς ις αςςιδενταλ, δυε το τηε ερψ ςμαλλ αλυες, ιν τηε ηαδρονις μαςς ςςαλε, οφ $m_{u}$ ανδ $m_{d}$.
[…]
$\bullet$ Τηε λεαδινγ διεργενςες. Τηε βρεαϰινγ οφ $SU(3)\times SU(3)$.
Τηε λεαδινγ διεργενςες ιν τηε ςεριες (51) ραιςεδ τηε ςπεςτρυμ οφ ςτρανγενεςς ανδ παριτψ ιολατιον ιν ςτρονγ ιντεραςτιονς. Τηε φιρςτ ςτεπ ωας το φινδ τηε ςονδιτιονς υνδερ ωηιςη τηις διςαςτερ ςουλδ βε αοιδεδ[112]. Τηε αργυμεντ ις βαςεδ ον τηε φολλοωινγ οβςερατιον: ατ τηε λιμιτ οφ εξαςτ $SU(3)\times SU(3)$ ονε ςαν περφορμ ινδεπενδεντ ριγητ- ανδ λεφτ-ηανδεδ ροτατιονς ιν φλαουρ ςπαςε ανδ διαγοναλιςε ωηιςηεερ ματριξ ωουλδ μυλτιπλψ τηε λεαδινγ διεργεντ τερμ. Ας α ρεςυλτ, ανψ νετ εφφεςτ ςηουλδ δεπενδ ον τηε παρτ οφ τηε ιντεραςτιον ωηιςη βρεαϰς $SU(3)\times SU(3)$. Ιν παρτιςυλαρ, ονε ςαν προε τηατ, υνδερ τηε αςςυμπτιον τηατ τηε ςηιραλ $SU(3)\times SU(3)$ ςψμμετρψ βρεαϰινγ τερμ τρανςφορμς ας α μεμβερ οφ τηε (3, $\bar{3}$$\oplus$($\bar{3}$, 3) ρεπρεςεντατιον, τηε ματριξ μυλτιπλψινγ τηε λεαδινγ διεργεντ τερμ ις διαγοναλ ιν φλαουρ ςπαςε, ι.ε. ιτ δοες νοτ ςοννεςτ ςτατες ωιτη διφφερεντ χυαντυμ νυμβερς, ςτρανγενεςς ανδ/ορ παριτψ. Τηερεφορε, αλλ ιτς εφφεςτς ςουλδ βε αβςορβεδ ιν α ρεδεφινιτιον οφ τηε παραμετερς οφ τηε ςτρονγ ιντεραςτιονς ανδ νο ςτρανγενεςς ορ παριτψ ιολατιον ωουλδ βε ινδυςεδ. Τηις ωας φιρςτ φουνδ φορ τηε ονε λοοπ διαγραμς ανδ τηεν εξτενδεδ το αλλ ορδερς. Τηις παρτιςυλαρ φορμ οφ τηε ςψμμετρψ βρεαϰινγ τερμ ηας α ςιμπλε ιντερπρετατιον ιν τηε φορμαλιςμ οφ τηε χυαρϰ μοδελ: ιτ ςορρεςπονδς το αν εξπλιςιτ χυαρϰ μαςς τερμ ανδ ιτ ωας τηε φαουριτε ονε το μοςτ τηεοριςτς, ςο ιτ ωας ςονςιδερεδ α ωελςομε ρεςυλτ. $\bullet$ Τηε νεξτ-το-λεαδινγ διεργενςες. Λεπτον-ηαδρον ςψμμετρψ - ῝ηαρμ.
Τηε ςολυτιον οφ τηε λεαδινγ διεργενςε προβλεμ ωας φουνδ ιν τηε φραμεωορϰ οφ τηε ςομμονλψ αςςεπτεδ τηεορψ ατ τηατ τιμε. Ον τηε ςοντραρψ, τηε νεξτ το λεαδινγ διεργενςες ρεχυιρεδ α δραςτις μοδιφιςατιον, αλτηουγη, ιν ρετροςπεςτ, ιτ ις α χυιτε νατυραλ ονε[104].
Λετ υς φιρςτ ςτατε τηε προβλεμ. Α φιρμλψ εςταβλιςηεδ εξπεριμενταλ φαςτ ις τηατ φλαουρ ςηανγινγ ωεαϰ προςεςςες οβεψ ςερταιν ςελεςτιον ρυλες: Ονε οφ τηεμ, ϰνοων ας τηε $\Delta S=1$ ρυλε, ςτατες τηατ τηε φλαουρ νυμβερ, ιν τηις ςαςε ςτρανγενεςς $S$, ςηανγες βψ ατ μοςτ ονε υνιτ. Α ςεςονδ ρυλε ις τηατ τηε αλλοωεδ $\Delta$Φλαουρ = 1 προςεςςες ινολε ονλψ ςηαργεδ ςυρρεντς. Ιτ φολλοως τηατ $\Delta S=2$ τρανςιτιονς, ας ωελλ ας Φλαουρ ῝ηανγινγ Νευτραλ ῝υρρεντ προςεςςες (Φ῝Ν῝), μυςτ βε ςεερελψ ςυππρεςςεδ. Τηε βεςτ εξπεριμενταλ ειδενςε φορ τηε φιρςτ ις τηε μεαςυρεδ $K_{L}-K_{S}$ μαςς διφφερενςε ωηιςη εχυαλς 3.48 $10^{-12}$ Με῞ ανδ, φορ τηε ςεςονδ, τηε βρανςηινγ ρατιο $B_{\mu^{+}\mu^{-}}=\Gamma(K_{L}\rightarrow\mu^{+}\mu^{-})/\Gamma(K_{L}\rightarrow all)$ ωηιςη εχυαλς 6.87 $10^{-9}$. Ιτ ωας τηις ϰινδ οφ τινψ εφφεςτς ωηιςη λεδ το τηε ςμαλλ αλυε οφ τηε ςυτ-οφφ ωε μεντιονεδ εαρλιερ. Ιν φαςτ, τηις προβλεμ ςαν βε αδδρεςςεδ ατ τωο λεελς. Τηεψ αρε βοτη εαςιερ το ιςυαλιςε ιν τηε φραμεωορϰ οφ τηε χυαρϰ μοδελ. Ατ τηε λιμιτ οφ εξαςτ φλαουρ ςψμμετρψ, χυαρϰ χυαντυμ νυμβερς, ςυςη ας ςτρανγενεςς, αρε νοτ ωελλ δεφινεδ. Ανψ βαςις ιν χυαρϰ ςπαςε ις ας γοοδ ας ανψ οτηερ. Βψ βρεαϰινγ τηις ςψμμετρψ τηε μεδιυμ ςτρονγ ιντεραςτιονς ςηοοςε α παρτιςυλαρ βαςις, ωηιςη βεςομες τηε πριιλεγεδ ονε. Ωεαϰ ιντεραςτιονς, ηοωεερ, δεφινε α διφφερεντ διρεςτιον, ωηιςη φορμς αν ανγλε $\theta_{C}$ ωιτη ρεςπεςτ το τηε φιρςτ ονε. Ηαινγ ονλψ τηρεε χυαρϰς το πλαψ ωιτη, ονε ςαν φορμ ονλψ ονε ςηαργεδ ςυρρεντ οφ τηε φορμ ποςτυλατεδ βψ ῝αβιββο[113]:
$J_{\mu}^{C}(x)=\bar{u}(x)\gamma_{\mu}(1+\gamma_{5})[\cos\theta_{C}\leavevmode\nobreak\ d(x)+\sin\theta_{C}\leavevmode\nobreak\ s(x)]$ (53)
Τηε εξπρεςςιον (53) ςαν βε ιντερπρετεδ ας ςαψινγ τηατ τηε $u$ χυαρϰ ις ςουπλεδ το α ςερταιν λινεαρ ςομβινατιον οφ τηε $d$ ανδ $s$ χυαρϰς, $d_{C}=\cos\theta_{C}\leavevmode\nobreak\ d+\sin\theta_{C}\leavevmode\nobreak\ s$. Τηε ορτηογοναλ ςομβινατιον, ναμελψ $s_{C}=-\sin\theta_{C}\leavevmode\nobreak\ d+\cos\theta_{C}\leavevmode\nobreak\ s$ ρεμαινς υνςουπλεδ. Νοτιςε τηε διφφερενςε ωιτη τηε λεπτονις ςυρρεντ. Ωε ηαε φουρ λεπτονς, τωο νευτραλς, τηε $\nu_{(e)}$ ανδ τηε $\nu_{(\mu)}$ ανδ τωο νεγατιελψ ςηαργεδ ονες, τηε ελεςτρον ανδ τηε μυον. Τηεψ αρε αλλ ςουπλεδ ανδ τηε ωεαϰ ςυρρεντ (39) ηας τωο πιεςες.
Τηε φιρςτ λεελ οφ τηε προβλεμ ις το ςονςιδερ α τηεορψ ςατιςφψινγ α ςυρρεντ αλγεβρα. Τηε νευτραλ ςομπονεντ οφ τηε ςυρρεντ ωιλλ βε ρελατεδ το τηε ςομμυτατορ οφ $J_{0}^{C}$ ανδ $J_{0}^{C\dagger}$ ανδ ωιλλ ςονταιν τερμς λιϰε $\bar{d}_{C}d_{C}$, τηυς ηαινγ φλαουρ ςηανγινγ πιεςες. Νοτιςε αγαιν τηατ τηις δοες νοτ ηαππεν ωιτη τηε λεπτονις ςυρρεντ. Τηε ςομμυτατορ οφ τηε ςυρρεντ (39) ωιτη ιτς ηερμιτιαν αδϑοιντ ηας νο τερμς ιολατινγ τηε τωο λεπτον φλαορ νυμβερς. Πηραςεδ τηις ωαψ, τηε ςολυτιον ις αλμοςτ οβιους: ωε μυςτ υςε τηε $s_{C}$ ςομβινατιον, βυτ, ιν ορδερ το δο ςο, ωε μυςτ ηαε α ςεςονδ υπ-τψπε χυαρϰ[104]. Ιφ ωε ςαλλ ιτ $c$, φορ ςηαρμ, τηε ςηαργεδ ωεαϰ ςυρρεντ (53) ωιλλ ηαε α ςεςονδ πιεςε:
$J_{\mu}(x)=\bar{u}(x)\gamma_{\mu}(1+\gamma_{5})d_{C}(x)+\bar{c}(x)\gamma_{\mu}(1+\gamma_{5})s_{C}(x)$ (54)
ορ, ιν α ματριξ νοτατιον,
$J_{\mu}(x)=\bar{U}(x)\gamma_{\mu}(1+\gamma_{5})CD(x)$ (55)
ωιτη
$U=\left(\begin{array}[]{c}u\\ c\end{array}\right)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ D=\left(\begin{array}[]{c}d\\ s\end{array}\right)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ C=\left(\begin{array}[]{cc}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{array}\right)$ (56)
Τηε ιμπορταντ ποιντ ις τηατ, νοω, α ςυρρεντ $J_{3}$, γιεν βψ τηε ςομμυτατορ οφ $J$ ανδ $J^{\dagger}$, ις διαγοναλ ιν φλαουρ ςπαςε.
Τηις ςολες τηε φιρςτ λεελ οφ τηε προβλεμ, βυτ ιτ ις νοτ ενουγη το εξπλαιν τηε οβςερεδ ρατες. Φορ εξαμπλε, τηε $K_{L}\rightarrow\mu^{+}\mu^{-}$ δεςαψ ςαν βε γενερατεδ βψ τηε βοξ διαγραμ οφ Φιγυρε 1 (λεφτ), ωηιςη, αλτηουγη οφ ηιγηερ ορδερ ιν τηε ωεαϰ ιντεραςτιονς, ιτ ις χυαδρατιςαλλψ διεργεντ ανδ ςοντριβυτες α τερμ προπορτιοναλ το $G_{F}g_{\mathrm{eff}}$.
Φιγυρε 0: Τηε ονε-λοοπ ςοντριβυτιον το $K^{0}\rightarrow\mu^{+}+\mu^{-}$. Ιν α τηρεε χυαρϰ τηεορψ (λεφτ), τηε ςηαρμ ςοντριβυτιον (ριγητ).
Ηερε ςομες τηε ςεςονδ ινγρεδιεντ οφ τηε μεςηανιςμ. Ωιτη α φουρτη χυαρϰ, τηερε ις α ςεςονδ διαγραμ, ωιτη $c$ ρεπλαςινγ $u$, Φιγυρε 1 (ριγητ). Ιν τηε λιμιτ οφ εξαςτ φλαουρ ςψμμετρψ τηε τωο διαγραμς ςανςελ. Τηε βρεαϰινγ οφ φλαουρ ςψμμετρψ ινδυςες α μαςς διφφερενςε βετωεεν τηε χυαρϰς, ςο τηε ςυμ οφ τηε τωο διαγραμς ις οφ ορδερ $g^{4}(m_{c}^{2}-m_{u}^{2})/m_{W}^{2}\sim G_{F}(G_{F}m_{c}^{2})$. Τηερεφορε, Ιοφφε ανδ Σηαβαλιν´ς εςτιματιονς ςαν βε τρανςλατεδ ιντο α λιμιτ φορ τηε νεω χυαρϰ μαςς ανδ ψιελδ αν υππερ βουνδ οφ α φεω Γε῞ φορ τηε μαςςες οφ τηε νεω ηαδρονς[104]. Τηις φαςτ ις ερψ ιμπορταντ. Α πρεδιςτιον φορ τηε εξιςτενςε οφ νεω παρτιςλες ις ιντερεςτινγ ονλψ ιφ τηεψ ςαννοτ βε αρβιτραριλψ ηεαψ [^59] Theoretical proposals involving new quark species had been made earlier[114], but they had no particular motivation and no connection to weak interactions..
Ιν τηε εαρλψ δαψς οφ τηε Φερμι τηεορψ τηερε ωας α ϰινδ οφ ςψμμετρψ βετωεεν ηαδρονς ανδ λεπτονς: προτον-νευτρον ς νευτρινο-ελεςτρον. Τηε φιρςτ διςςοερψ οφ ηεαψ φλαουρς αππεαρεδ το βρεαϰ τηις ςψμμετρψ: ωε ηαδ τωο νεω λεπτονς, τηε μυον ανδ ιτς αςςοςιατεδ νευτρινο, βυτ ονλψ ονε νεω ηαδρον, τηε ςτρανγε χυαρϰ. Τηε ιντροδυςτιον οφ τηε ςηαρμεδ χυαρϰ ρεςτορεδ τηις ςψμμετρψ. Βψ δοινγ ςο, τηε μεςηανιςμ οβταινεδ τωο ιμπορταντ ρεςυλτς: (ι) Ιτ ςολεδ τηε τεςηνιςαλ προβλεμ οφ τηε λοω αλυε οφ τηε Ιοφφε ανδ Σηαβαλιν ςυτ-οφφ βψ ρεπλαςινγ ιτ ωιτη τηε μαςςες οφ νεω ηαδρονς. (ιι) Ιτ οπενεδ τηε ωαψ το α φορμυλατιον οφ τηε τηεορψ ιν τερμς οφ ςυρρεντ οπερατορς ωηιςη ςατιςφψ αλγεβραις προπερτιες. Ιτ ις τηις ςεςονδ ρεςυλτ τηατ μαδε ποςςιβλε τηε υςε οφ Ψανγ-Μιλλς τηεοριες φορ τηε εντιρε ωεαϰ ιντεραςτιονς, λεπτονις ας ωελλ ας ηαδρονις. Τηε πριςε ωας τηε πρεδιςτιον φορ αν εντιρε νεω ςεςτορ οφ ηαδρονις πηψςιςς, τηε ςηαρμεδ παρτιςλες, ωιτη μαςςες ωηιςη ωερε βουνδεδ φρομ αβοε βψ α φεω Γε῞. Ιν παρτιςυλαρ ωε πρεδιςτεδ τηε εξιςτενςε οφ α νεω $1^{-}$ βοςον.
Τηε ιδεα οφ ιντροδυςινγ νεω χυαρϰ φλαορς ιν ορδερ το εξπλαιν ωεαϰ ιντεραςτιον πηενομενα ωας βριλλιαντλψ εξτενδεδ ιν 1973 βψ Μ. Κοβαψαςηι ανδ Τ. Μαςϰαωα[115] ωηο νοτιςεδ τηατ α μοδελ ωιτη τηρεε χυαρϰ φαμιλιες οφφερς α νατυραλ μεςηανιςμ φορ $CP$ ιολατιον.
Σομε περςοναλ ρεμινιςςενςες: Ι μεντιονεδ αλρεαδψ αν ινφορμαλ ςεμιναρ ωε γαε ατ ΜΙΤ ιμμεδιατελψ αφτερ ωε διςςοερεδ τηε ςηαρμ μεςηανιςμ. Ιτ τοοϰ πλαςε ιν Φ. Λοω´ς οφφιςε ανδ γατηερεδ μοςτ οφ τηε ΜΙΤ τηεοριςτς. Φορ υς ιτ ωας α γοοδ τεςτ ανδ ωε παςςεδ ιτ ςυςςεςςφυλλψ. Λοω πραιςεδ ουρ ςολυτιον ανδ Ωεινβεργ, ωηο αςϰεδ μανψ χυεςτιονς, αλτηουγη ηε διδ νοτ μαϰε τηε ςοννεςτιον ωιτη ηις εαρλιερ ωορϰ, ηε λιϰεδ τηε μεςηανιςμ ερψ μυςη. Τηε ςαμε εενινγ ατ διννερ Πυςςι Μαιανι (Λυςιανο ηαδ βεεν μαρριεδ α φεω μοντης βεφορε) ρεμαρϰεδ τηατ ωε λοοϰεδ ερψ ηαππψ. Σηελλψ τολδ ηερ τηατ ωε εξπεςτεδ ουρ ωορϰ το βε παρτ οφ φυτυρε πηψςιςς τεξτ βοοϰς.
Ωε ςυβμιττεδ τηε παπερ το τηε Πηψςιςαλ Ρειεω ανδ, α ςουπλε οφ ωεεϰς λατερ, ωε ρεςειεδ τηε ρεπορτ. Τηε ρεφερεε ηαδ ρεαδ ουρ παπερ ερψ ςαρεφυλλψ, ηε ςαιδ ιτ ωας ιντερεςτινγ ανδ ωορτη πυβλιςηινγ, βυτ ηε ραιςεδ ονε οβϑεςτιον: α ναϊε ποωερ ςουντινγ φορ α μαςςιε Ψανγ-Μιλλς ιντεραςτιον γιες $(G\Lambda^{6})^{n}$ ωηιλε ωε ωερε ιμπλιςιτλψ αςςυμινγ ιτ το βε $(G\Lambda^{2})^{n}$. Ηε τηεν ωεντ ον το ρεμαρϰ : “Τηε ναϊε βεηαιορ ςαν υνδουβτεδλψ βε ιμπροεδ, βυτ τηε αςςερτιον τηατ ιτ ςαν γο δοων το $(G\Lambda^{2})^{n}$ μυςτ βε ειτηερ προεν ορ δελετεδ.” Ηε διδ νοτ ςαψ ιτ ωας ωρονγ, ωηιςη προες τηατ ηε ϰνεω τηε προβλεμ ερψ ωελλ. Ωε ονλψ ςηανγεδ α φεω ωορδς ιν τηε παπερ ανδ ρεςυβμιττεδ, τηις τιμε ςυςςεςςφυλλψ. Τηε επιλογυε ις τηατ ιτ τοοϰ Σηελλψ ανδ μψςελφ – Λυςιανο ηαδ λεφτ αλρεαδψ φορ Ιταλψ – ςομε λενγτηψ ανδ ηαρδ ωορϰ το προε τηατ ουρ ινιτιαλ εςτιματε ωας ριγητ. Ωε πυβλιςηεδ τωο παπερς[116] ωηιςη, αλτηουγη Ι ςονςιδερ τηεμ το βε αμονγ τηε μοςτ ιντελλιγεντ παπερς Ι ηαε εερ ωριττεν, τυρνεδ ουτ το βε ιρρελεαντ φορ τηε ςυβςεχυεντ δεελοπμεντς.
[…]
Ι βελιεε τηατ τηε φιρςτ περςον ωηο δεςιδεδ το λοοϰ ατ τηε Ψανγ-Μιλλς χυαντυμ φιελδ τηεορψ ας τηε δψναμιςαλ τηεορψ οφ ωεαϰ ιντεραςτιονς ωας Μ.Θ.Γ. ῞ελτμαν. Αρουνδ 1966, ηε ωας τρψινγ το υνδερςτανδ τηε δεεπερ οριγιν οφ τηε ςονςερατιον, ορ νεαρ ςονςερατιον, οφ τηε ωεαϰ ςυρρεντς. Ιν παρτιςυλαρ, ηε τριεδ το τηροω ςομε λιγητ ον τηε γενεραλ ςονφυςιον ωηιςη πρεαιλεδ ατ τηατ τιμε ςονςερνινγ τηε ςο-ςαλλεδ “Σςηωινγερ τερμς” ιν τηε ςομμυτατορς οφ τωο ςυρρεντ ςομπονεντς. Ωηιλε ηε ωας ον α ιςιτ φρομ ῝ΕΡΝ το Βροοϰηαεν, ηε ωροτε α παπερ ιν ωηιςη ηε ςυγγεςτεδ α ςετ οφ διεργενςε εχυατιονς ωηιςη γενεραλιςεδ τηε νοτιον οφ τηε ςοαριαντ δεριατιε οφ χυαντυμ ελεςτροδψναμιςς. Τηις φυνδαμενταλ ιδεα ωας ταϰεν υπ νεξτ ψεαρ ανδ δεελοπεδ φυρτηερ βψ Θ.Σ. Βελλ[117]. Τηεςε εχυατιονς λοοϰεδ λιϰε τηε ςοαριαντ δεριατιες οφ νον-αβελιαν γαυγε τηεοριες, ςο ῞ελτμαν δεςιδεδ το ςτυδψ τηειρ φιελδ τηεορψ προπερτιες. Τηε ελεςτροδψναμιςς οφ ςηαργεδ εςτορ βοςονς ηαδ βεεν φορμυλατεδ αλρεαδψ βψ Τ.Δ. Λεε ανδ ῝.Ν. Ψανγ ιν 1962[118]. Τηεψ ηαδ ςηοων τηατ ελεςτρομαγνετις γαυγε ιναριανςε αλλοως το εξπρεςς τηε εςτορ βοςον´ς ςηαργε $e$, μαγνετις μομεντ $\mu$ ανδ χυαδρυπολε μομεντ $Q$ ιν τερμς οφ ονλψ τωο παραμετερς $e$ ανδ $\kappa$, ας $\mu=e(1+\kappa)/2m_{W}$ ανδ $Q=-e\kappa/m_{W}^{2}$. Τηε ρεςυλτινγ τηεορψ ις ηιγηλψ διεργεντ βυτ ῞ελτμαν νοτιςεδ τηατ μανψ διεργενςες ςανςελ φορ τηε παρτιςυλαρ αλυε $\kappa=1$. Ιτ ις τηε αλυε πρεδιςτεδ βψ α τηεορψ ιν ωηιςη $W^{\pm}$ ανδ τηε πηοτον φορμ α Ψανγ-Μιλλς τριπλετ. Φορ ῞ελτμαν τηις ωας α ςλεαρ ςιγναλ τηατ τηε τηεορψ οφ ωεαϰ ανδ ελεςτρομαγνετις ιντεραςτιονς μυςτ οβεψ α Ψανγ-Μιλλς γαυγε ιναριανςε.
Στυδψινγ τηε ρενορμαλιςατιον προπερτιες οφ Ψανγ-Μιλλς τηεοριες τυρνεδ ουτ το βε α ερψ ηαρδ ανδ ςομπλεξ προβλεμ, βοτη ςονςεπτυαλλψ ανδ πραςτιςαλλψ. Ον τηε ςονςεπτυαλ ςιδε, νειτηερ τηε ρεςυλτς οφ Φεψνμαν[52] ανδ Φαδδεε ανδ Ποπο[53] νορ τηοςε ον ςποντανεους ςψμμετρψ βρεαϰινγ ωερε ωιδελψ ϰνοων ανδ ῞ελτμαν ηαδ το ρεδιςςοερ εεν τηε βαςις Φεψνμαν ρυλες. Ον τηε πραςτιςαλ ςιδε, τηε νυμβερ οφ τερμς γρεω ερψ φαςτ ανδ ῞ελτμαν ηαδ το δεελοπ α ςομπυτερ προγραμ το ηανδλε τηεμ[119]. Ηε ςαλλεδ ιτ “Σςηοονςςηιπ”,606060“Clean ship” in dutch ανδ ιτ ωας τηε φιρςτ προγραμ οφ ςψμβολις μανιπυλατιονς αππλιεδ το τηεορετιςαλ ηιγη ενεργψ πηψςιςς.
Α φεω ψεαρς εαρλιερ, ωηιλε ωορϰινγ φορ ηις τηεςις, ῞ελτμαν βεςαμε φαμιλιαρ ωιτη α μετηοδ το ςομπυτε Φεψνμαν διαγραμς υςινγ τηε ςο-ςαλλεδ “ςυττινγ ρυλες” ωηιςη ωερε προποςεδ ιν 1960 βψ Ρ.Ε. ῝υτϰοςϰψ[120]. Τηεψ ςονςιςτ ιν ςυττινγ τηε διαγραμ ιν αλλ ποςςιβλε ωαψς ανδ ρεπλαςινγ τηε ςυτ προπαγατορς βψ τηε μαςς-ςηελλ δελτα φυνςτιονς. Τηε μετηοδ γιες τηε ιμαγιναρψ παρτ οφ τηε διαγραμ ανδ ωε ςαν ρεςοερ τηε ρεαλ παρτ βψ ωριτινγ α διςπερςιον ρελατιον. ῞ελτμαν δεελοπεδ τηις αππροαςη ανδ τυρνεδ ιτ ιντο α ερψ ποωερφυλ τοολ παρτιςυλαρλψ αδαπτεδ το εξπλορε τηεοριες φορ ωηιςη τηε εντιρε φιελδ ςοντεντ ις νοτ α πριορι ϰνοων [^61] In doing so, he was concentrated in the calculation of scattering amplitudes. For Veltman all quantum field theories that give the same $S$-matrix were considered as identical. This purely semantic question was the source of endless disputes in conferences and meetings between him and many other field theorists.. Α περςοναλ εξποςιτιον οφ ηις αππροαςη ςαν βε φουνδ ιν ρεφερενςε[121]. Ηε φιρςτ ςτυδιεδ μαςςιε Ψανγ-Μιλλς τηεοριες ατ τηε ονε λοοπ λεελ[122] ανδ ςλαιμεδ ςανςελλατιονς οφ διεργενςες, αλτηουγη τηε παπερς, εςπεςιαλλψ τηε φιρςτ ονε, ωερε νοτ εαςψ το φολλοω. Ιν ρετροςπεςτ ωε ςεε τηατ ῞ελτμαν αλλοωεδ ηιμςελφ το ςηανγε τηε Φεψνμαν ρυλες ας ιφ τηε τηεορψ ωας γαυγε ιναριαντ αλτηουγη τηε μαςς τερμς βρεαϰ τηε ςψμμετρψ. Ιν φαςτ, ωιτη α ρεμαρϰαβλε ιντυιτιον, ηε ωας ρεδιςςοερινγ βψ τριαλ ανδ ερρορ τηε ονε λοοπ εφφεςτς οφ τηε Βρουτ-Ενγλερτ-Ηιγγς μεςηανιςμ. Ιν 1970, ιν ςολλαβορατιον ωιτη Η. ῞αν Δαμ, ηε ςτυδιεδ τηε ζερο μαςς λιμιτ οφ γαυγε φιελδς, βοτη φορ Ψανγ-Μιλλς ανδ γραιτψ[123] ανδ προεδ, βψ εξπλιςιτ ςαλςυλατιον, τηατ τηε λιμιτ ωας νοτ ςοντινυους βεςαυςε τηε λονγιτυδιναλ δεγρεες οφ φρεεδομ οφ τηε μαςςιε φιελδς δο νοτ δεςουπλε ωηεν τηε μαςς γοες το ζερο. Σιμιλαρ ρεςυλτς ηαδ βεεν οβταινεδ αλςο βψ Α.Α. Σλανο ανδ Λ.Δ. Φαδδεε[124], αλτηουγη τηεψ ωερε νοτ ωιδελψ ϰνοων. Ιν 1968 ῞ελτμαν τοοϰ α ςαββατιςαλ λεαε ανδ ςπεντ τηε 68-69 ψεαρ ατ Ορςαψ. Ηε λεςτυρεδ ον Ψανγ-Μιλλς φιελδς ανδ τηερε ωερε ςομε λεςτυρε νοτες ιν τηε φορμ οφ α πρεπριντ ωηιςη, ας φαρ ας Ι ϰνοω, ηας νοτ βεεν πυβλιςηεδ. Σομε οφ υς, ινςλυδινγ μψςελφ, φουνδ ιν τηεςε νοτες ουρ φιρςτ ιντροδυςτιον το Ψανγ-Μιλλς ανδ πατη ιντεγραλς.
Ιν αδδιτιον το ῞ελτμαν ανδ τηε ϒτρεςητ ςςηοολ, τηερε ωερε ςομε οτηερ ιςολατεδ ανδ ρατηερ ςονφιδεντιαλ ςοντριβυτιονς ωηιςη αδδρεςςεδ τηε χυεςτιον οφ τηε ρενορμαλιςατιον προπερτιες οφ γαυγε τηεοριες[125]. Τηερε ις α γρεατ ςονφυςιον ιν τηε εαρλψ λιτερατυρε ανδ ονε φινδς ςοντραδιςτορψ ςτατεμεντς ρεγαρδινγ τηε ρενορμαλιςαβιλιτψ οφ μαςςιε Ψανγ-Μιλλς. Ιν μοςτ ςαςες τηε προβλεμ ις ςεμαντις, βυτ ιτ τοοϰ ςομε τιμε βεφορε τηε ςιτυατιον ωας ςλεαρεδ υπ. Τηε ςορρεςτ ςτατεμεντ, ωηιςη Ι ωιλλ εξπλαιν ιν ςεςτιον 1, ις τηατ α μαςςιε Ψανγ-Μιλλς τηεορψ ις νον-ρενορμαλιςαβλε, υνλεςς τηε γαυγε βοςον μαςςες αρε γενερατεδ τηρουγη α Βρουτ-Ενγλερτ-Ηιγγς μεςηανιςμ οφ ςποντανεους ςψμμετρψ βρεαϰινγ. Ας ωε ςαω ιν ςεςτιον 1, τηις μεςηανιςμ λεαες ατ λεαςτ ονε πηψςιςαλ ςςαλαρ φιελδ. Τηε μαςς οφ τηις ςςαλαρ παρτιςλε $m_{S}$ ις αν αρβιτραρψ παραμετερ οφ τηε τηεορψ, βυτ ωε ςαννοτ γετ ριδ οφ ιτ βψ ςενδινγ ιτ το ινφινιτψ ςινςε, ιν τηις ςαςε, ωε ρεςοερ τηε μαςςιε Ψανγ-Μιλλς τηεορψ ωιτηουτ τηε πηενομενον οφ ςποντανεους ςψμμετρψ βρεαϰινγ, ωηιςη ις νον-ρενορμαλιςαβλε. Τηρουγη τηε ραδιατιε ςορρεςτιονς, αλλ μεαςυραβλε χυαντιτιες ιν τηε Στανδαρδ Μοδελ δεπενδ ον $m_{S}$ ανδ, ιν τηε λατε ςεεντιες, ῞ελτμαν ςομπυτεδ τηε ονε λοοπ εφφεςτς ιν αν εφφορτ το ρεςτριςτ τηε αλυε οφ τηε Ηιγγς μαςς[126]. Ας ιτ τυρνεδ ουτ, νο ςεερε ρεςτριςτιονς ςουλδ βε οβταινεδ βεςαυςε τηε δεπενδανςε ωας ονλψ λογαριτημις. Ηε ςαλλεδ ιτ “τηε ςςρεενινγ τηεορεμ”. Ας α ρεςυλτ, νο πρεςιςε πρεδιςτιονς ον $m_{S}$ ωερε ϰνοων υντιλ τηε αςτυαλ διςςοερψ οφ τηε παρτιςλε ατ τηε ΛΗ῝.
Ιν 1969 ῞ελτμαν ωας ϑοινεδ ιν ϒτρεςητ βψ Γεραρδ ´τ Ηοοφτ, α γραδυατε ςτυδεντ ωιτη ωηομ ηε ωουλδ ςηαρε τηε Νοβελ Πριζε 30 ψεαρς λατερ [^62] Initially Veltman did not want to assign to this young student the renormalisability of Yang-Mills as a thesis project because he thought it was too risky. He proposed instead the problem of $A2$ splitting which was quite popular at that time. ’t Hooft came back a bit later declaring he did not like it, so Veltman gave up and proposed Yang-Mills. In fact, ’t Hooft showed good taste because the problem of $A2$ splitting turned out to be a fake.. Ιν 1970 ´τ Ηοοφτ φολλοωεδ τηε ῝αργèςε ςυμμερ ςςηοολ – ηις αππλιςατιον το τηε Λες Ηουςηες ςυμμερ ςςηοολ ηαδ βεεν τυρνεδ δοων – ιν ωηιςη ηε ηεαρδ Β.Ω. Λεε λεςτυρινγ αβουτ τηε ρενορμαλιςατιον οφ τηε $\sigma$-μοδελ ωιτη ςποντανεους ςψμμετρψ βρεαϰινγ. Αςςορδινγ το ´τ Ηοοφτ, τηις ωας α ρεελατιον ανδ ςηεδ νεω λιγητ το τηε ρενορμαλιςατιον προπερτιες οφ Ψανγ-Μιλλς τηεοριες[127]. Ραρε αρε τηε εξαμπλες οφ ΠηΔ ςτυδεντς ωηοςε ρεςυλτς ηαδ ςυςη α προφουνδ ανδ λαςτινγ ινφλυενςε ιν πηψςιςς. Τηε ψεαρ 1971 ωας τηε αννυς μιραβιλις. ´τ Ηοοφτ πυβλιςηεδ τωο παπερς, λανδμαρϰς ιν τηε ροαδ το τηε Στανδαρδ Μοδελ[128]. Τηε φιρςτ ηας τηε τιτλε “Ρενορμαλιςατιον οφ μαςςλεςς Ψανγ-Μιλλς Φιελδς” ανδ πρεςεντς τηε φιρςτ ςομπλετε φορμυλατιον οφ τηε χυαντυμ γαυγε τηεορψ. Ιν τηις ωορϰ ´τ Ηοοφτ γενεραλιςεδ τηε Φαδδεε-Ποπο μετηοδ το αρβιτραρψ γαυγες ανδ δεριεδ τηε ςορρεςπονδινγ Φεψνμαν ρυλες. Φυρτηερμορε, ηε ινεντεδ α γαυγε ιναριαντ ρεγυλαριςατιον ςςηεμε φορ ονε-λοοπ διαγραμς βψ ιντροδυςινγ α φιςτιτιους φιφτη διμενςιον, α πρεςυρςορ οφ τηε διμενςιοναλ ρεγυλαριςατιον μετηοδ. Ωιτη τηεςε τοολς ηε γαε τηε φιρςτ ςομπλετε προοφ οφ τηε ρενορμαλιςαβιλιτψ οφ τηε τηεορψ. Τηε ςεςονδ παπερ, ωιτη τηε τιτλε “Ρενορμαλιςαβλε Λαγρανγιανς φορ Μαςςιε Ψανγ-Μιλλς Φιελδς”, ινςορπορατες τηε ΒΕΗ μεςηανιςμ ανδ ςηοως τηατ τηε ρεςυλτινγ τηεορψ ις ρενορμαλιςαβλε. Ηε υςες ηις πρειους ρεςυλτς ιν ορδερ το ςομπυτε τηε Φεψνμαν ρυλες φορ α φαμιλψ οφ γαυγες ανδ, ιν παρτιςυλαρ, τηε ονε ϰνοων ας “τηε ´τ Ηοοφτ γαυγε” ωηιςη ις μοςτ ςονενιεντ φορ ςαλςυλατιονς. Α λιμιτινγ ςαςε ιν τηις φαμιλψ ις “τηε υνιταρψ γαυγε” ιν ωηιςη τηε Λαγρανγιαν ςονταινς ονλψ τηε πηψςιςαλ δεγρεες οφ φρεεδομ. Ιτ ις νον-ρενορμαλιςαβλε βψ ποωερ ςουντινγ βυτ, βψ γαυγε ιναριανςε, ηας τηε ςαμε $S$-ματριξ ας τηε ονε ωε ςαν ςομπυτε υςινγ, φορ εξαμπλε, τηε ´τ Ηοοφτ γαυγε. Ιτ γενεραλιςες το τηε τηεορψ οφ τηε ωεαϰ ανδ ελεςτρομαγνετις ιντεραςτιονς τηε Στυεςϰελβεργ εφφεςτ ωε ςαω ιν ςεςτιον 1. Τηε ςομπλετε προγραμ οφ τηε Ψανγ-Μιλλς ρενορμαλιςατιον ηας βεεν πυβλιςηεδ βψ ´τ Ηοοφτ ανδ ῞ελτμαν ιν 1972[129]. Τηε φιρςτ οφ τηεςε πυβλιςατιονς ςονταινς αλςο τηε διμενςιοναλ ρεγυλαριςατιον ςςηεμε ωηιςη, ιν αδδιτιον το πρεςερε γαυγε ιναριανςε, ηας τηε αδανταγε οφ ςιμπλιφψινγ ενορμουςλψ αλλ περτυρβατιον ςαλςυλατιονς. Ιτ εξτενδς πρειους ωορϰ βψ ῝.Γ. Βολλινι ανδ Θ.Θ. Γιαμβιαγι[130].
Γ. ’τ Ηοοφτ ανδ ῞ελτμαν γαε τηε φιρςτ δεταιλεδ πρεςεντατιον οφ τηειρ ρεςυλτς ιν α ςμαλλ μεετινγ ατ Ορςαψ ιν Θανυαρψ 1972. ´τ Ηοοφτ ωας τηε μαιν ςπεαϰερ ανδ Ι ςτιλλ ρεμεμβερ ηις μαρατηον λεςτυρες. Τηις μεετινγ ωας ρεμαρϰαβλε ιν μανψ ρεςπεςτς. Φιρςτ ιτ οφφερεδ τηε φιρςτ ςομπλετε πιςτυρε οφ τηε ρενορμαλιςατιον προπερτιες οφ Ψανγ-Μιλλς τηεοριες, ινςλυδινγ τηε μετηοδ οφ διμενςιοναλ ρεγυλαριςατιον. Σεςονδ ιτ τριγγερεδ ςτιμυλατινγ διςςυςςιονς αμονγ τηε παρτιςιπαντς. Ιν παρτιςυλαρ, τηε ιταλ ιμπορτανςε οφ τηε αξιαλ ςυρρεντ ανομαλψ ςανςελλατιον ωας ςτρεςςεδ ιν τηις μεετινγ. Τηιρδ ιτ ινιτιατεδ α λονγ ςεριες οφ μεετινγς ωηιςη βεςαμε ϰνοων ας “τριανγυλαρ μεετινγς” (Παρις-Ρομε-ϒτρεςητ). Συβςεχυεντλψ ενλαργεδ ωιτη τηε αδδιτιον οφ οτηερ Ευροπεαν ςεντρες, τηεψ πλαψεδ αν ιμπορταντ ρολε ιν τηε δεελοπμεντ οφ α Ευροπεαν νετωορϰ ιν τηεορετιςαλ πηψςιςς.
Ιν τηε μεαντιμε, τηε ρενορμαλιςατιον προγραμ οφ Ψανγ-Μιλλς τηεοριες ςταρτεδ αττραςτινγ ωιδερ αττεντιον. Α.Α. Σλανο ανδ Θ.῝. Ταψλορ δεριεδ τηε Ωαρδ ιδεντιτιες φορ Ψανγ-Μιλλς τηεοριες[131] ανδ ςιμπλιφιεδ ςονςιδεραβλψ τηε αναλψςις. Β.Ω. Λεε ανδ Θ. Ζινν-Θυςτιν ρεφορμυλατεδ τηε προοφ οφ ρενορμαλιςαβιλιτψ ανδ γαυγε ιναριανςε οφ ςποντανεουςλψ βροϰεν Ψανγ-Μιλλς τηεοριες, υςινγ πατη ιντεγραλς ανδ Ωαρδ ιδεντιτιες[132]. Τηε μοςτ νοελ ςοντριβυτιον ςαμε φρομ ῝. Βεςςηι, Α. Ρουετ ανδ Ρ. Στορα φρομ Μαρςειλλε[133] (ςεε αλςο Ι.῞. Τψυτιν φρομ Μοςςοω[134]) ωηο ινεντεδ αν ινγενιους μετηοδ το ηανδλε τηε ςονςεχυενςες οφ γαυγε ιναριανςε. ϒςυαλλψ ωε οβταιν τηε Ωαρδ ιδεντιτιες υςινγ τηε ιναριανςε προπερτιες οφ τηε Λαγρανγιαν. Ωιτη γαυγε τηεοριες ηοωεερ, τηε Λαγρανγιαν ωε αρε υςινγ φορ αςτυαλ ςαλςυλατιονς ις νοτ ιναριαντ υνδερ γαυγε τρανςφορματιονς βεςαυςε ωε μυςτ μαϰε α γαυγε ςηοιςε. Τηε προβλεμ ις αςςεντυατεδ ιν Ψανγ-Μιλλς τηεοριες, ορ χυαντυμ γραιτψ, βεςαυςε, ιν αδδιτιον, ωε νεεδ το ιντροδυςε τηε Φαδδεε-Ποπο γηοςτς ωηοςε φορμ δεπενδς ον τηε γαυγε. Β.Ρ.Σ. ανδ Τ. φουνδ τηατ τηε ρεςυλτινγ Λαγρανγιαν ις νεερτηελεςς ιναριαντ υνδερ α νεω ςψμμετρψ, ςαλλεδ “τηε ΒΡΣΤ ςψμμετρψ”. Τηε ςτρανγε προπερτψ οφ τηις ςψμμετρψ ις τηατ τηε παραμετερς ωηιςη δετερμινε τηε τρανςφορματιονς αρε νοτ ορδιναρψ νυμβερς βυτ αντι-ςομμυτινγ χυαντιτιες. Σινςε τηατ τιμε, τηε ΒΡΣΤ ςψμμετρψ ηας βεςομε αν εςςεντιαλ τοολ ιν τηε ςτυδψ οφ αλλ χυαντυμ γαυγε τηεοριες.
[…]
Ιτ τοοϰ α λονγ τιμε το δεριε α ςονςιςτεντ χυαντυμ φιελδ τηεορψ φορ ωεαϰ ιντεραςτιονς ανδ τηε φιναλ πιςτυρε ωας νοτ ατ αλλ οβιους. Ιτς τωο βαςις ινγρεδιεντς, Ψανγ-Μιλλς γαυγε ιναριανςε ανδ τηε πηενομενον οφ ςποντανεους ςψμμετρψ βρεαϰινγ, εαςη ονε ςονςιδερεδ ςεπαρατελψ, πρεδιςτεδ τηε αππεαρανςε οφ μαςςλεςς παρτιςλες – εςτορ γαυγε βοςονς ανδ ςςαλαρ Γολδςτονε βοςονς – ιν οβιους ςοντραδιςτιον ωιτη εξπεριμεντ. Ιτ ωας ονλψ βψ ςομβινατιον τηατ τηε τωο δεςεαςες ςυρεδ εαςη οτηερ. Ιν 1973 ῝.Η. Λλεωελλψν-Σμιτη αςϰεδ τηε χυεςτιον ωηετηερ ςυςη α βεηαιορ ςουλδ ηαε βεεν αντιςιπατεδ[135]. Ηε ιντροδυςεδ τηε ςονςεπτ οφ τρεε υνιταριτψ ας φολλοως. Λετ υς ςονςιδερ τηε Φερμι τηεορψ ανδ ςομπυτε τηε ελεςτρον–αντι-νευτρινο ελαςτις ςςαττερινγ. Ατ λοωεςτ ορδερ ωε ηαε τηε διαγραμ οφ Φιγυρε 1 (α).
Φιγυρε 0: Τηε ελαςτις ελεςτρον–αντι-νευτρινο ςςαττερινγ ατ φιρςτ ορδερ ιν τηε Φερμι ςουπλινγ ςονςταντ (α) ανδ ατ ςεςονδ ορδερ (β).
Τηε Φερμι ςουπλινγ ςονςταντ ηας διμενςιονς $[M]^{-2}$. Ατ ηιγη ενεργιες ωε ςαν νεγλεςτ τηε φερμιον μαςςες ανδ τηε ςροςς ςεςτιον ωιλλ βεηαε ας $\sigma_{\mathrm{el}}\sim G_{F}^{2}s$ ωηερε $s$ ις τηε τοταλ ενεργψ ιν τηε ςεντρε οφ μαςς, $s=(p_{1}+p_{2})^{2}$. Ρεμεμβερ, ηοωεερ, τηατ τηις ςροςς ςεςτιον ις προπορτιοναλ το τηε ιμαγιναρψ παρτ οφ τηε ςορρεςπονδινγ ονε-λοοπ αμπλιτυδε οφ Φιγυρε 1 (β). Ιτ φολλοως τηερεφορε τηατ, ιφ ωε ωριτε α διςπερςιον ρελατιον φορ τηε λαττερ οφ τηε φορμ:
[…]
ιτ ωιλλ ρεχυιρε τωο ςυβτραςτιονς. Οφ ςουρςε, τηις ις ιν αγρεεμεντ ωιτη τηε φαςτ τηατ τηε διαγραμ οφ Φιγυρε 1 (β) ις χυαδρατιςαλλψ διεργεντ. Τηερεφορε, ιφ ωε ωαντ το εμβεδ τηε Φερμι τηεορψ ιντο α λαργερ, ρενορμαλιςαβλε φιελδ τηεορψ, ωε μυςτ ιμπροε τηε ηιγη-ενεργψ βεηαιορ οφ τηε τρεε διαγραμ. Τηις ις τηε ςονςεπτ οφ τρεε υνιταριτψ. Λλεωελλψν-Σμιτη ιν ςυςςεςςιε ςτεπς φουνδ τηατ ςυςη ιμπροεμεντ ρεχυιρες: (ι) τηε ιντροδυςτιον οφ ιντερμεδιατε εςτορ βοςονς, (ιι) Ψανγ-Μιλλς ςουπλινγς αμονγ τηεμ, (ιιι) αν εξτρα ςςαλαρ βοςον, ανδ (ι) ρελατιονς αμονγ μαςςες ανδ ςουπλινγς οφ αλλ τηεςε φιελδς. Ιν οτηερ ωορδς, ηε ρεςονςτρυςτεδ τηε Στανδαρδ Μοδελ ιν τηε υνιταρψ γαυγε.
Ονε ςαν οβϑεςτ τηατ ιν ουρ ςεαρςη ωε ςυςςεεδεδ βεςαυςε ωε ϰνεω τηε ανςωερ ανδ τηις ις ςερταινλψ τρυε. Ηοωεερ, νοω τηατ ωε ηαε τηε μαςηινερψ, ωε ςαν αςϰ τηε μορε γενεραλ χυεςτιον: ςονςιδερ α ςετ οφ φιελδς, ςπιν ζερο, ςπιν ονε-ηαλφ ανδ ςπιν ονε. Αμονγ τηε λαττερ, ςομε αρε αςςυμεδ μαςςλεςς ανδ ςομε αρε μαςςιε. Ωριτε τηε μοςτ γενεραλ ηερμιτιαν, Λορεντζ ιναριαντ, ιντεραςτιον Λαγρανγιαν αμονγ τηεμ ωιτη τερμς οφ διμενςιον ςμαλλερ ορ εχυαλ το φουρ. Ουρ τηεορψ ςο φαρ ις ςηαραςτεριςεδ βψ α λαργε ςετ οφ παραμετερς, μαςςες ανδ ςουπλινγ ςονςταντς. Ιμποςε ``τρεε υνιταριτψ´´ φορ αλλ αμπλιτυδες ι.ε. νο ςροςς ςεςτιον ςομπυτεδ ιν τηε τρεε αππροξιματιον ςηουλδ γροω ατ ινφινιτψ. Τηις γιες ρελατιονς αμονγ τηε παραμετερς. Τηε ρεςυλτινγ τηεορψ ις εχυιαλεντ το α Ψανγ-Μιλλς γαυγε τηεορψ. Τηε ρελατιονς αμονγ τηε παραμετερς ςιμπλψ εξπρεςς τηε ΒΡΣΤ ςψμμετρψ ιν τηε υνιταρψ γαυγε. Ιφ τηε ΒΡΣΤ ιδεντιτιες αρε νοτ ςατιςφιεδ τηε τηεορψ ςαννοτ βε ρενορμαλιςαβλε. Αλλ μαςςιε εςτορ βοςονς ςορρεςπονδ το ςποντανεουςλψ βροϰεν γενερατορς ια α ΒΕΗ μεςηανιςμ ωιτη τωο ποςςιβλε εξςεπτιονς: (ι) τηε τριιαλ ονε οφ φρεε φιελδς, (ιι) μαςςιε νευτραλ εςτορ βοςονς ςουπλεδ το ςονςερεδ ςυρρεντς φολλοωινγ τηε Στυεςϰελβεργ φορμαλιςμ. Ιν οτηερ ωορδς, ωε ςαν προε τηατ τηε ροαδ τηατ λεδ το τηε Στανδαρδ Μοδελ ις υνιχυε[136].
[…]
Τηε Γλαςηοω-Ωεινβεργ-Σαλαμ μοδελ ςονταινς α μαςςιε νευτραλ γαυγε βοςον, τηερεφορε ιτ ιμπλιες τηε εξιςτενςε οφ ωεαϰ νευτραλ ςυρρεντς. Συςη ςυρρεντς ηαδ νοτ βεεν δετεςτεδ, βυτ ονλψ τηε ςτρανγενεςς ςηανγινγ ονες ωερε ρεαλλψ φορβιδδεν βψ εξπεριμεντ. Τηε ινςορπορατιον οφ τηε ΓΙΜ μεςηανιςμ[104] ενςυρεδ τηειρ αβςενςε, ςο τηε πρεδιςτιον ωας λιμιτεδ το ςτρανγενεςς ςονςερινγ νευτραλ ςυρρεντς. Ιν πραςτιςε τηεψ ςουλδ μανιφεςτ ιν τωο ωαψς: νευτρινο ινδυςεδ ορ ελεςτρον ινδυςεδ ρεαςτιονς. Τηε φιρςτ ςουλδ βε ειτηερ νευτρινο-ελεςτρον ελαςτις ςςαττερινγ $\nu_{\mu}+e^{-}\rightarrow\nu_{\mu}+e^{-}$ ορ ινελαςτις νευτρινο νυςλευς ςςαττερινγ οφ τηε γενεραλ φορμ $\nu_{\mu}+N\rightarrow\nu_{\mu}+X$. Τηε ςεςονδ ςουλδ ινδυςε παριτψ ιολατινγ εφφεςτς ιν ειτηερ ατομις πηψςιςς[137], ορ ελεςτρον-νυςλευς ςςαττερινγ. Ανψ οφ τηεςε ρεαςτιονς ωας αν εξπεριμενταλ ςηαλλενγε, ςο ιν τηε εαρλψ ψεαρς ςομε γαυγε μοδελς ωερε προποςεδ τρψινγ το αοιδ νευτραλ ςυρρεντς, ειτηερ παρτλψ[138], ορ τοταλλψ[139]. Ιν 1973 α Ευροπεαν ςολλαβορατιον οπερατινγ τηε ηεαψ λιχυιδ βυββλε ςηαμβερ “Γαργαμελλε” ατ ῝ΕΡΝ, ςεττλεδ τηε ιςςυε βψ οβςερινγ νευτραλ ςυρρεντ εεντς ιν $\nu_{\mu}$ ανδ ${\bar{\nu}}_{\mu}$ βεαμς[140]. Ψετ, ιτ τοοϰ ςομε τιμε βεφορε τηε ςομμυνιτψ ωας ςονινςεδ [^63] I won some bottles of very fine wine by betting for neutral currents, in particular against Jack Steinberger.. Τηερε ωερε ςεεραλ ρεαςονς φορ τηατ. Φιρςτ, Γαργαμελλε ωας τηε φιρςτ “λαργε” ιντερνατιοναλ ςολλαβορατιον (7 λαβς, 55 πηψςιςιςτς!) ανδ τηις μαδε πεοπλε ςυςπιςιους. Ι ςτιλλ ρεμεμβερ ςομε ςαρςαςτις ςομμεντς: “῝αν ανψτηινγ ριγητ ςομε ουτ φρομ ςυςη α ςροωδ;”. Σεςονδ, ινιτιαλλψ τηερε ωας ονλψ ονε “γολδ πλατεδ” εεντ οφ ελαςτις αντι-νευτρινο–ελεςτρον ςςαττερινγ. Τηιρδ, τηε ςηαμβερ ωας 4.8μ λονγ βυτ ονλψ 2μ ιν διαμετερ. Τηερεφορε, τηε ςολλαβορατιον ηαδ το προε τηατ τηεψ ηαδ φυλλ ςοντρολ οφ τηε βαςϰγρουνδ ςονςιςτινγ οφ νευτρονς προδυςεδ ιν τηε ςτεελ ςυρρουνδινγ τηε ςηαμβερ ανδ εντερινγ φρομ τηε ςιδες, τηυς μιμιςϰινγ ηαδρονις νευτραλ ςυρρεντ εεντς. Φορ τηατ, τηεψ δεελοπεδ α Μοντε-῝αρλο προγραμ ςιμυλατινγ αλλ νευτρινο ιντεραςτιονς ιν τηε εντιρε ενιρονμεντ οφ τηε ςηαμβερ. Ιτ ωας τηε φιρςτ τιμε τηατ τηε βεηαιορ οφ ςυςη α λαργε ανδ ςομπλεξ δετεςτορ ηαδ βεεν αναλψςεδ νυμεριςαλλψ. Φιναλλψ, τηερε ωερε ςομε ςοντραδιςτορψ ςλαιμς ςομινγ φρομ τηε φιρςτ ερψ ηιγη ενεργψ νευτρινο εξπεριμεντς ιν τηε ϒνιτεδ Στατες.
Ιτ τοοϰ μυςη λονγερ το δετεςτ νευτραλ ςυρρεντ εφφεςτς ιν ελεςτρον ινδυςεδ ρεαςτιονς. Τηε φιρςτ ειδενςε ςαμε φρομ ΣΛΑ῝ οβςερινγ παριτψ ιολατιον ιν δεεπ ινελαςτις ςςαττερινγ οφ πολαριςεδ ελεςτρονς[141], φολλοωεδ βψ τηε διςςοερψ οφ παριτψ ιολατιον ιν ατομις πηψςιςς ατ τηε Εςολε Νορμαλε Συπéριευρε[142].
[…]
Ι ςαμε βαςϰ το Ευροπε ιν Οςτοβερ 1971 ανδ ϑοινεδ τηε Δεπαρτμεντ ιν Ορςαψ ωηερε Ι ηαδ ςτυδιεδ. Τηεςε ωερε τιμες οφ γρεατ εξπεςτατιονς. Ωε ηαδ βεεν τηρουγη ςο μανψ λεαν ψεαρς ιν παρτιςλε πηψςιςς τηατ μοςτ πηψςιςιςτς ηαδ λοςτ ηοπε το εερ εξπεριενςε τηε εξςιτεμεντ οφ γρεατ διςςοεριες. Τηερε ωερε ςο μανψ νεω τηινγς το λεαρν, ςο μανψ χυεςτιονς το ανςωερ. Γαυγε τηεοριες ηαδ ρεολυτιονιςεδ ουρ ωαψ οφ τηινϰινγ. Ωηιλε ωε ωερε τρψινγ το υνδερςτανδ τηε προοφ οφ ρενορμαλιςαβιλιτψ ιν αριους γαυγες, ωε ρεαλιςεδ τηε ιταλ ιμπορτανςε οφ τηε Ωαρδ ιδεντιτιες. Α ςηανγε οφ γαυγε προδυςες α ςομπλετελψ νεω τηεορψ. Αλλ τηεςε τηεοριες τηατ λοοϰ, ανδ ιν μανψ ρεςπεςτς αρε, ςο διφφερεντ, αρε λινϰεδ τογετηερ τηρουγη τηε Ωαρδ ιδεντιτιες. Φορ εξαμπλε, τηε ρενορμαλιςαβλε γαυγες ςονταιν υνπηψςιςαλ δεγρεες οφ φρεεδομ ανδ ιτ ις ονλψ βψ ιρτυε οφ τηε Ωαρδ ιδεντιτιες τηατ τηεψ δεςουπλε φρομ πηψςιςαλ χυαντιτιες. Βυτ τηεν τηε οβιους προβλεμ αππεαρεδ. Ωεαϰ ιντεραςτιονς ινολε βοτη εςτορ ανδ αξιαλ ςυρρεντς. Ωε ηαε ςεεν τηατ ιν μανψ ςαςες ονε ςαννοτ ενφορςε τηε ςονςερατιον οφ βοτη βεςαυςε οφ τηε τριανγλε ανομαλιες. Ωιτη ῝λ. Βουςηιατ ανδ Πη. Μεψερ ωε ςηεςϰεδ εαςιλψ τηατ ιν τηε $SU(2)\times U(1)$ μοδελ τηε Ωαρδ ιδεντιτιες ωερε ινδεεδ βροϰεν ανδ τηε ρενορμαλιςαβιλιτψ ανδ υνιταριτψ προοφς διδ νοτ αππλψ. Ι φελτ λιϰε α ςηιλδ φρομ ωηομ ςομεονε ςτολε ηις μοςτ ωονδερφυλ τοψ. Τηις βρινγς με το τηε φιναλ αςτ οφ τηε αξιαλ ανομαλψ πλαψ ωηιςη ωε λεφτ υνφινιςηεδ ιν ςεςτιον 1. Τηε ςολυτιον ωας φιρςτ φουνδ βψ ῝λ. Βουςηιατ, Πη. Μεψερ ανδ μψςελφ ιν 1971[143]. Τηε ϰεψ ινγρεδιεντ ις τηε ρεμαρϰ τηατ τηε ανομαλψ δοες νοτ δεπενδ ον τηε μαςς οφ τηε φερμιον τηατ γοες αρουνδ τηε τριανγυλαρ λοοπ. Τηερεφορε, ωε μυςτ αδδ αλλ ςοντριβυτιονς ιν ορδερ το γετ τηε ριγητ ανςωερ. Φορ τηε ελεςτροωεαϰ τηεορψ τηις μεανς τηατ ωε νεεδ βοτη τηε λεπτονς ανδ τηε χυαρϰς. Α ςιμπλε ςαλςυλατιον ςηοως τηατ τηε τοταλ ανομαλψ προδυςεδ βψ τηε φερμιονς οφ εαςη φαμιλψ ις προπορτιοναλ το ${\cal A}=\sum_{i}Q_{i}$, ωηερε τηε ςυμ εξτενδς οερ αλλ φερμιονς ιν α γιεν φαμιλψ ανδ $Q_{i}$ ις τηε ελεςτρις ςηαργε οφ τηε $i$τη φερμιον. Σινςε ${\cal A}=0$ ις α νεςεςςαρψ ςονδιτιον φορ τηε ματηεματιςαλ ςονςιςτενςψ οφ τηε τηεορψ, ωε ςονςλυδε τηατ εαςη φαμιλψ μυςτ ςονταιν τηε ριγητ αμουντ οφ λεπτονς ανδ χυαρϰς το μαϰε τηε ανομαλψ ανιςη. Τηις ςονδιτιον ις ςατιςφιεδ βψ τηε τηρεε ςολουρ μοδελ ωιτη ςηαργες 2/3 ανδ -1/3, βυτ αλςο βψ οτηερ μοδελς ςυςη ας τηε Ηαν-Ναμβυ μοδελ ωιτη ιντεγερ ςηαργε χυαρϰς. Ιν φαςτ, τηε ανομαλψ ςανςελλατιον ςονδιτιον ηας α ωιδερ αππλιςατιον. Τηε Στανδαρδ Μοδελ ςουλδ ηαε βεεν ινεντεδ αφτερ τηε Ψανγ-Μιλλς τηεορψ ωας ωριττεν, μυςη βεφορε τηε διςςοερψ οφ τηε χυαρϰς. Ατ τηατ τιμε τηε “ελεμενταρψ” παρτιςλες ωερε τηουγητ το βε τηε ελεςτρον ανδ ιτς νευτρινο, τηε προτον ανδ τηε νευτρον, ςο ωε ωουλδ ηαε υςεδ ονε λεπτον ανδ ονε ηαδρον δουβλετ. Τηε ςονδιτιον ${\cal A}=0$ ις ςατιςφιεδ. Ωηεν χυαρϰς ωερε διςςοερεδ ωε ςηανγεδ φρομ νυςλεονς το χυαρϰς. Τηε ςονδιτιον ις αγαιν ςατιςφιεδ. Ιφ τομορροω ωε φινδ τηατ ουρ ϰνοων λεπτονς ανδ/ορ χυαρϰς αρε ςομποςιτε, τηε νεω βυιλδινγ βλοςϰς ωιλλ βε ρεχυιρεδ το ςατιςφψ τηις ςονδιτιον αγαιν.
Τηε μοραλ οφ τηε ςτορψ ις τηατ φαμιλιες μυςτ βε ςομπλετε. Τηε τιτλε οφ τηε ΓΙΜ παπερ[104] ωας “Ωεαϰ Ιντεραςτιονς ωιτη Λεπτον-Ηαδρον Σψμμετρψ” ανδ ωιτη Βουςηιατ ανδ Μεψερ ωε προεδ τηατ ιτ ωας ινδεεδ ςορρεςτ. Τηυς, τηε διςςοερψ οφ α νεω λεπτον, τηε ταυ, ιμπλιεδ τηε εξιςτενςε οφ τωο νεω χυαρϰς, τηε $b$ ανδ τηε $t$, πρεδιςτιον ωηιςη ωας αγαιν εριφιεδ εξπεριμενταλλψ.
[…]
Τηε εμεργενςε οφ χυαντυμ ςηρομοδψναμιςς ας τηε τηεορψ οφ ςτρονγ ιντεραςτιονς ηας βεεν πρεςεντεδ ιν ςεεραλ ρειεως, ςεε φορ εξαμπλε [144] ανδ [145]. Τηε φιρςτ ονε, βψ Δαιδ Γροςς, δεςςριβες αλςο τηε ιντερνατιοναλ ςςενε ωηιςη πρεςεδεδ Χ῝Δ. Τηε ςεςονδ, βψ Ρ.Κ. Ελλις, οφφερς α ιεω οφ τηε τεςηνιςαλ αςπεςτς οφ τηε ςομπυτατιονς ωηιςη αρε υςεδ τοδαψ. Μψ πρεςεντατιον ωιλλ βε βριεφ.
[…]
Στρονγ ιντεραςτιονς εντερεδ τηε πηψςιςς ςςενε ιν τηε φιρςτ ηαλφ οφ τηε τωεντιετη ςεντυρψ ωιτη τηε δεελοπμεντ οφ νυςλεαρ πηψςιςς. Ιτ ωας ιμμεδιατελψ ειδεντ τηατ τηεψ αρε ςηορτ ρανγεδ, νο μαςροςςοπις εφφεςτς ρελατεδ το τηε ςτρονγ φορςε ηαε εερ βεεν οβςερεδ, ανδ τηεψ αρε ερψ ςτρονγ, ςινςε τηεψ οερςομε τηε ελεςτροςτατις ρεπυλςιον αμονγ προτονς ανδ λεαδ το τηε φορματιον οφ ςταβλε νυςλει. Ας ωε ςαιδ αλρεαδψ ιν ςεςτιον 2, τηε εφφεςτιε ςουπλινγ ςονςταντ ις λαργε ανδ τηε περτυρβατιον εξπανςιον υςελεςς. Τηε Παρτιςλε Δατα Γρουπ λιςτς μορε τηαν 30 πιον-νυςλεον ρεςονανςες ωιτη μαςςες βετωεεν 1 ανδ 2.5 Γε῞, ωιτη ςπινς ας ηιγη ας 11/2, ανδ ιτ ις ςλεαρ τηατ νο ςιμπλε χυαντυμ φιελδ τηεορψ ςουλδ αςςομμοδατε αλλ οφ τηεμ. Στρονγ ιντεραςτιονς αππεαρεδ το βε ερψ ςομπλιςατεδ.
Ωε ϰνοω τοδαψ τηατ τηις ςομπλεξιτψ ςηοως ονλψ α ςυπερφιςιαλ παρτ οφ τηε πιςτυρε. Τηε ιντεραςτιονς αππεαρ ερψ ςομπλιςατεδ βεςαυςε τηε οβϑεςτς ωε ωερε τρψινγ το ςτυδψ, τηε ηαδρονς, αρε τηεμςελες ςομπλεξ. Ιτ ις ας ιφ ωε ωερε τρψινγ το διςςοερ χυαντυμ ελεςτροδψναμιςς βψ λοοϰινγ ατ τηε ιντεραςτιονς αμονγ ςομπλιςατεδ μαςρομολεςυλες [^64] R.P. Feynman has given a nice analogy which accurately describes the efforts to understand strong interactions: imagine you want to study the mechanism of a fine Swiss watch and, to do that, you take two of them and smash them one against the other.. Βψ ςηανγινγ τηε περςπεςτιε, ωε διςςοερεδ α ςομπλετελψ διφφερεντ πιςτυρε.
[…]
Τηε μοδερν τηεορψ οφ ςτρονγ ιντεραςτιονς ωας δεελοπεδ ας α ρεςπονςε το ςερταιν ιντριγυινγ εξπεριμενταλ ρεςυλτς. Ιν 1966 α ηιγη ενεργψ λινεαρ ελεςτρον αςςελερατορ ωας ςομμιςςιονεδ ατ Στανφορδ ιν τηε Λαβορατορψ ωηιςη βεςαμε ϰνοων ας Στανφορδ Λινεαρ Αςςελερατορ ῝εντερ (ΣΛΑ῝). Βψ τηε λατε 1960ς ανδ εαρλψ 1970ς α ςεριες οφ εξπεριμεντς ωερε ςαρριεδ ουτ ατ ΣΛΑ῝ ςτυδψινγ τηε ηιγη ενεργψ ανδ λαργε μομεντυμ τρανςφερ ςςαττερινγ, οφτεν ςαλλεδ δεεπ ινελαςτις ςςαττερινγ, οφ ελεςτρονς οφφ νυςλεονς. Τηε προςεςς ις ςηοων ςςηεματιςαλλψ ιν Φιγυρε 1 ιν τηε ονε πηοτον εξςηανγε αππροξιματιον. Ιτ ις αν ινςλυςιε προςεςς, μεανινγ τηατ, ιν τηε φιναλ ςτατε, ονλψ τηε ελεςτρον ωας μεαςυρεδ. Τηε ςτατε δενοτεδ βψ $X$ ωας νοτ.
Φιγυρε 0: Δεεπ ινελαςτις ελεςτρον-νυςλεον ςςαττερινγ ωιτη τηε προδυςτιον οφ α φιναλ ςτατε |$\ket{X}$.
Τηις ςηοιςε ωας διςτατεδ παρτλψ βψ ςομε τηεορετιςαλ ςονςιδερατιονς [^65] The story goes back to the Algebra of Currents[59]. A good collection of original articles can be found in [60]. Particularly interesting for this discussion is the one by J.D. Bjorken[146]., βυτ αλςο βψ εξπεριμενταλ ςονςτραιντς: (ι) Τηε δετεςτορς ωερε νοτ αβλε το διςτινγυιςη βετωεεν α $\pi^{+}$ ανδ α προτον ιφ τηειρ ϰινετις ενεργψ ωας μυςη ηιγηερ τηαν 1 Γε῞. (ιι) Τηε φορωαρδ ανδ βαςϰωαρδ ρεγιονς ωερε νοτ ςοερεδ, ςο α ςομπλετε ιδεντιφιςατιον οφ αλλ παρτιςλες ιν $X$ ωας νοτ ποςςιβλε. (ιιι) Τηε ελεςτρονς ιν τηε βεαμ ωερε ςομινγ ιν βυνςηες βυτ τηε δετεςτορς ωερε νοτ αβλε το ρεςολε μυλτιπλε ςολλιςιονς. Τηερεφορε, ωηατ ωας μεαςυρεδ ςορρεςπονδς το α ςυμ οερ αλλ ποςςιβλε $X$. Ιτ ις εαςψ το ςηοω τηατ τηις διφφερεντιαλ ςροςς ςεςτιον ις γιεν βψ τηε ματριξ ελεμεντ οφ τηε ςομμυτατορ οφ τωο ςυρρεντ οπερατορς:
$\frac{{\rm d}\sigma}{{\rm d}\Omega^{\prime}{\rm d}E^{\prime}}\sim\frac{e^{4}}{Q^{4}}L_{\mu\nu}W^{\mu\nu}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\mathrm{with}}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ W_{\mu\nu}=\frac{1}{2}\int{\rm d}^{4}y\leavevmode\nobreak\ {\rm e}^{{\rm i}Q\cdot y}\leavevmode\nobreak\ \bra{P}[J_{\mu}^{h\dagger}(y)J_{\nu}^{h}(0)]\ket{P}$ (58)
ωηερε $L_{\mu\nu}$ ις τηε λεπτονις παρτ ωηιςη, ςυμμεδ οερ ελεςτρον πολαριςατιονς ανδ ιν τηε λιμιτ οφ ανιςηινγ ελεςτρον μαςς, ις γιεν βψ $L_{\mu\nu}=2\left(k_{\mu}k^{\prime}_{\nu}+k_{\nu}k^{\prime}_{\mu}-\eta_{\mu\nu}k\cdot k^{\prime}\right)$. $k$ ανδ $k^{\prime}$ αρε τηε μομεντα οφ τηε ινιτιαλ ανδ φιναλ ελεςτρονς ρεςπεςτιελψ, ανδ $Q=k-k^{\prime}$. $J_{\mu}^{h}$ ις τηε ηαδρονις παρτ οφ τηε ελεςτρομαγνετις ςυρρεντ ανδ |$\ket{P}$ τηε ονε-νυςλεον ςτατε ωιτη μομεντυμ $P$. Ωε ςαν ςηοω τηατ τηις ινςλυςιε ςροςς ςεςτιον ςαν βε εξπρεςςεδ ιν τερμς οφ τωο φυνςτιονς, ςαλλεδ “ςτρυςτυρε φυνςτιονς”, $F_{1}$ ανδ $F_{2}$. Κινεματιςαλλψ τηεψ ςαν δεπενδ ονλψ ον τωο διμενςιονλεςς αριαβλες ωηιςη αρε ςηοςεν τραδιτιοναλλψ ας $M^{2}_{N}/q^{2}$ ανδ $x=q^{2}/2\nu$ ωιτη $\nu=2P\cdot Q$ ανδ $q^{2}=-Q^{2}$. Τηε ςυρπριςινγ ρεςυλτ οφ τηε ΣΛΑ῝ εξπεριμεντ ωας τηατ, ωηεν βοτη $\nu$ ανδ $q^{2}$ βεςομε λαργε ωιτη τηειρ ρατιο $x$ ϰεπτ φιξεδ, τηεςε ςτρυςτυρε φυνςτιονς ωερε αππροξιματελψ φυνςτιονς ονλψ οφ $x$. Ηερε “λαργε” μεανς λαργε ςομπαρεδ το τηε νυςλεον μαςς. Τηις προπερτψ βεςαμε ϰνοων ας ςςαλε ιναριανςε· αρψινγ τηε ενεργψ ςςαλε οφ τηε εξπεριμεντ, $q^{2}$ ανδ $\nu$ ωιτη $x$ φιξεδ, δοες νοτ αφφεςτ τηε ςροςς ςεςτιον. Ιτ ις εαςψ το ςεε τηατ ωηεν $q^{2}\rightarrow\infty$ ωιτη $x$ φιξεδ, τηε δομιναντ ρεγιον ιν τηε ιντεγραλ οφ εχυατιον (58) ις τηε ονε ωιτη $y^{2}\sim 0$, ιν οτηερ ωορδς ιν τηε δεεπ ινελαςτις ρεγιον ωε προβε τηε ιςινιτψ οφ τηε λιγητ ςονε.
Τηε προπερτψ οφ ςςαλε ιναριανςε ις ερψ ιντερεςτινγ, βεςαυςε ιτ ις ερψ εαςψ το “υνδερςτανδ” ιτ υςινγ α ναïε ανδ ωρονγ ρεαςονινγ. Ωηεν $q^{2}\rightarrow\infty$ ωιτη φιξεδ $x$, τηε ςεςονδ αριαβλε $M_{N}^{2}/q^{2}$ γοες το ζερο. Ναïελψ, α φυνςτιον $f(x,M_{N}^{2}/q^{2})$ ςαν βε αππροξιματεδ βψ $f(x,M_{N}^{2}/q^{2})=f(x,0)+\frac{M_{N}^{2}}{q^{2}}f^{(1)}(x,0)+\dots$ ωηερε $f^{(1)}$ ις τηε φιρςτ δεριατιε οφ $f$ ωιτη ρεςπεςτ το $M_{N}^{2}/q^{2}$ ϰεεπινγ $x$ φιξεδ. Σο, ωηεν $M_{N}^{2}/q^{2}$ γοες το ζερο, ωε εξπεςτ το βε λεφτ ωιτη ονλψ τηε $x$ δεπενδενςε. Τηις αργυμεντ, ηοωεερ, ις τεςηνιςαλλψ ωρονγ βεςαυςε ιτ αςςυμες τηατ τηε ςτρυςτυρε φυνςτιονς $F_{1,2}$ αρε αναλψτις ιν τηε αριαβλε $M_{N}^{2}/q^{2}$ αρουνδ τηε ποιντ $M_{N}=0$ ανδ ςαν βε εξπανδεδ ιν α ποωερ ςεριες. Τηις αςςυμπτιον ις φαλςε ιν χυαντυμ φιελδ τηεορψ, βεςαυςε, ας ωε ϰνοω, ωε οφτεν ενςουντερ ινφραρεδ διεργενςες ωηεν ωε αττεμπτ το ςετ α μαςς παραμετερ εχυαλ το ζερο.
Φεψνμαν ηαδ βυιλτ α ςιμπλε μοδελ, ωηιςη ιμπλεμεντεδ τηις ςςαλε βεηαιουρ[147]. Λετ υς αςςυμε τηατ τηε ταργετ νυςλεον ις μαδε ουτ οφ ελεμενταρψ ςονςτιτυεντς, ωηιςη ιντεραςτ ωιτη τηε ινςιδεντ πηοτον ας ποιντ παρτιςλες. Ωε ςηαλλ ςαλλ τηεςε ςονςτιτυεντς ςολλεςτιελψ παρτονς. Ιφ ωε νεγλεςτ αλλ ιντεραςτιονς αμονγ τηε παρτονς ωε ςαν εαςιλψ ρεπροδυςε τηις ςςαλινγ προπερτψ. Φεψνμαν υςεδ τηε φολλοωινγ ςιμπλε πιςτυρε: ατ ηιγη ενεργιες τηε ιρτυαλ πηοτον οφ Φιγυρε 1 ςεες τηε ταργετ νυςλεον ας α τηιν διςϰ. Τηε παρτον μοδελ αμουντς το νεγλεςτ τηε ιντεραςτιονς αμονγ τηε παρτονς ιν τηε πλανε οφ τηε διςς. Η. Φριτζςςη ανδ Μ. Γελλ-Μανν[148] προποςεδ αν αλγεβραις φραμεωορϰ φορ τηις μοδελ βψ εξτενδινγ τηε εχυαλ-τιμε ςομμυτατορς οφ ςυρρεντ αλγεβρα το α λιγητ ςονε αλγεβρα οφ βι-λοςαλ οπερατορς. Τηε βαςις αςςυμπτιον ωας τηατ εερψωηερε ατ τηε λιγητ ςονε, ονε ςουλδ υςε φρεε φιελδ ςομμυτατιον ρελατιονς [^66] Gell-Mann was saying that $\cdots$ at the light cone Nature reads only free field theory books.”.
Τηε τρουβλε, οφ ςουρςε, ις τηατ τηε αςςυμπτιον οφ νο ιντεραςτιον αμονγ τηε παρτονς δοες νοτ ςεεμ το μαϰε ςενςε. Τηε παρτονς ςαννοτ βε φρεε ανδ, ατ τηε ςαμε τιμε, βινδ ςτρονγλψ το φορμ α νυςλεον. Νεερτηελεςς, ιτ ωας ςυςη α ςςηιζοπηρενις βεηαιουρ τηατ ωας ιμπλιεδ βψ τηε δατα. Τηε παρτονς ωερε αλμοςτ φρεε ωηεν προβεδ βψ α ιρτυαλ πηοτον ιν τηε δεεπ ινελαςτις ρεγιον ανδ ςτιλλ ερψ ςτρονγλψ βουνδ ιν ορδιναρψ ηαδρονις εξπεριμεντς. Στρονγ ιντεραςτιονς ηαδ τηις δυαλ βεηαιουρ: ερψ ςομπλιςατεδ ωιτη νο περτυρβατιε εξπανςιον ατ τηε λεελ οφ ηαδρονς ανδ ερψ ςιμπλε, αππροξιματεδ βψ α φρεε φιελδ τηεορψ, ατ τηε λεελ οφ παρτονς. Τηις πιςτυρε ρινγς α βελλ. Ωηεν διςςυςςινγ τηε αππλιςατιονς οφ τηε ρενορμαλιςατιον γρουπ ιν ςεςτιον 1, ωε φουνδ τηατ τηε εφφεςτιε ςουπλινγ ςονςταντ οφ α ρενορμαλιςαβλε χυαντυμ φιελδ τηεορψ δεπενδς ον τηε ενεργψ ςςαλε οφ τηε εξπεριμεντ. Ωε ςονςλυδε τηατ ιφ ωε ωαντ το υνδερςτανδ ΣΛΑ῝´ς ρεςυλτς ον δεεπ ινελαςτις ελεςτρον-νυςλεον ςςαττερινγ ιν τερμς οφ α χυαντυμ φιελδ τηεορψ, ωε μυςτ λοοϰ φορ ονε ιν ωηιςη τηε εφφεςτιε ςουπλινγ ςονςταντ ηας τηε οπποςιτε βεηαιουρ τηαν τηε ονε ωε φουνδ ιν εχυατιον (37): ιτ ςηουλδ βεςομε ςμαλλερ ανδ ςμαλλερ ας ωε γο το ηιγηερ ανδ ηιγηερ ενεργιες. Ωε ςαλλ ςυςη τηεοριες “αςψμπτοτιςαλλψ φρεε” ανδ εχυατιον (37) τελλς υς τηατ τηεψ μυςτ ηαε $b_{0}<0$.
[…]
Λιϰε μοςτ οφ τηε ιδεας τηατ βυιλτ τηε Στανδαρδ Μοδελ, τηε ροαδ το αςψμπτοτις φρεεδομ ρεςεμβλες α ρανδομ ωαλϰ. Ι ωιλλ μεντιον ςομε ςτεπς, βυτ Ι δο νοτ γυαραντεε ςομπλετενεςς.
Ας ωε ςαιδ ιν ςεςτιον 1, αλλ εαρλψ ςαλςυλατιονς ςηοωεδ τηατ τηε ρυννινγ ςουπλινγ ςονςταντ οφ ρενορμαλιςαβλε χυαντυμ φιελδ τηεοριες δεςρεαςεδ ωιτη διςτανςε ανδ τηις βεηαιουρ ωας ςονςιδερεδ το βε υνιερςαλ. Τηε φιρςτ ρεςυλτς ωηιςη ινδιςατεδ τηε οπποςιτε βεηαιουρ ωερε ειτηερ διςμιςςεδ ορ ιγνορεδ[149]. Α ςηορτ λιςτ ινςλυδες:
$\bullet$ Ιν 1965 ῞.Σ. ῞ανψαςηιν ανδ Μ.῞. Τερεντε[150] ςομπυτεδ τηε φιρςτ τερμ ιν τηε εξπανςιον οφ τηε $\beta$-φυνςτιον φορ α μαςςιε Ψανγ-Μιλλς τηεορψ ανδ φουνδ α νεγατιε ςιγν. Ηοωεερ, ςινςε τηις τηεορψ ις νοτ ρενορμαλιςαβλε, νοβοδψ παιδ ανψ αττεντιον το τηε ρεςυλτ.
$\bullet$ Ας φαρ ας Ι ϰνοω, τηε φιρςτ ςορρεςτ ανςωερ φορ τηε μαςςλεςς Ψανγ-Μιλλς τηεορψ ις δυε το Ι.Β. Κηριπλοιςη ιν 1969[151]. Ηε υςεδ τηε ῝ουλομβ γαυγε ωηιςη ις γηοςτ φρεε. Βοτη τηε ςιγν ανδ τηε μαγνιτυδε αρε ςορρεςτ, βυτ τηε παπερ ωας ςιμπλψ ιγνορεδ. Ιτ ις νοτ μεντιονεδ ιν ανψ οφ τηε ςτανδαρδ ρειεως. Ι ηεαρδ αβουτ ιτ ονλψ ρεςεντλψ φρομ Γ. Παριςι.
$\bullet$ Ιν Θυλψ 1972 Γ. ´τ Ηοοφτ ρεπορτεδ ιν α ςμαλλ μεετινγ ιν Μαρςειλλε τηε νεγατιε ςιγν οφ τηε $\beta$-φυνςτιον, αγαιν φορ τηε μαςςλεςς Ψανγ-Μιλλς τηεορψ. Ι ωας πρεςεντ ιν τηις μεετινγ, τογετηερ ωιτη μανψ οτηερ τηεοριςτς. Ι ρεμεμβερ, ιν παρτιςυλαρ, Τ. Αππελχυιςτ ανδ Κ. Σψμανζιϰ. Ι δο νοτ ρεμεμβερ ανψ εντηυςιαςτις ρεαςτιον ανδ, το μψ ςηαμε, Ι διδ νοτ μαϰε ανψ αττεμπτ το ρελατε τηις ρεςυλτ το τηε παρτον μοδελ ανδ δεεπ ινελαςτις ςςαττερινγ. Ι ςαννοτ εξπλαιν τηις μιςταϰε, εςπεςιαλλψ βεςαυςε ατ τηατ τιμε Ι ωας τρψινγ το υνδερςτανδ ςςαλινγ υςινγ τηε λιγητ-ςονε αλγεβρα οφ βιλοςαλ οπερατορς[152]. Προβαβλψ, ονε οφ τηε ρεαςονς ωας τηατ ιντροδυςινγ ςςαλαρ φιελδς το ινδυςε τηε ΒΕΗ πηενομενον δεςτροψς αςψμπτοτις φρεεδομ ανδ τηε ιδεα οφ ηαινγ μαςςλεςς γαυγε βοςονς φλοατινγ αρουνδ ωας νοτ ταϰεν ςεριουςλψ.
$\bullet$ Ας φαρ ας Ι ςαν ςεε, τηε αφορεμεντιονεδ τηρεε ςαλςυλατιονς ηαδ αλλ α τεςηνιςαλ μοτιατιον: ςομπυτε τηε $\beta$-φυνςτιον οφ α Ψανγ-Μιλλς τηεορψ. Τηερε ωας νο εξπλιςιτ ρεφερενςε το α τηεορψ οφ ςτρονγ ιντεραςτιονς. Ον τηε οτηερ ηανδ, τηε παπερς βψ Δ.Θ. Γροςς ανδ Φ. Ωιλςζεϰ, ας ωελλ ας Η.Δ. Πολιτζερ[153], ηαδ α ωελλ-δεφινεδ γοαλ: αν αςψμπτοτιςαλλψ φρεε τηεορψ οφ ςτρονγ ιντεραςτιονς.
$\bullet$ Ιν 1973 Σ. ῝ολεμαν ανδ Δ.Θ. Γροςς προεδ τηατ, ιν φουρ διμενςιονς, τηε ονλψ αςψμπτοτιςαλλψ φρεε τηεοριες αρε υνβροϰεν, νον-αβελιαν, Ψανγ-Μιλλς τηεοριες[154]
[…]
Δυρινγ τηε 1960ς χυαντυμ φιελδ τηεορψ ηαδ πλαψεδ α πεςυλιαρ ρολε ιν τηε δεςςριπτιον οφ ςτρονγ ιντεραςτιονς. Ιτ ωας τοταλλψ ρεϑεςτεδ ας α φυνδαμενταλ δψναμιςαλ τηεορψ, βυτ ιτ ωας ερψ οφτεν υςεδ, εξπλιςιτλψ ορ ιμπλιςιτλψ, ας α πλαψγρουνδ ιν ορδερ το γυεςς γενεραλ προπερτιες ωηιςη ωερε ποςτυλατεδ φορ τηε ςςαττερινγ αμπλιτυδες. Φορ εξαμπλε, τηε βαςις ποςτυλατες οφ αναλψτιςιτψ ωερε εξτραςτεδ φρομ τηε ςτυδψ οφ Φεψνμαν διαγραμς. Τηε Μανδελςταμ ρεπρεςεντατιον ωας δεριεδ φρομ τηε ςχυαρε διαγραμ οφ $\phi^{3}$. Γελλ-Μανν φορμυλατεδ τηε αλγεβρα οφ ςυρρεντς βψ ςονςιδερινγ α ςιμπλε χυαρϰ μοδελ, αλτηουγη ηε διδ νοτ ωαντ το αδοςατε τηε εξιςτενςε οφ πηψςιςαλ χυαρϰς. Ονε οφ τηε εαρλιεςτ φιελδ τηεορετις χυαρϰ μοδελς δεςςριβεδ ςπιν-1/2 χυαρϰς ιντεραςτινγ ωιτη α ςινγλε νευτραλ εςτορ φιελδ, ςαλλεδ γλυον, αλτηουγη νο ονε βελιεεδ ιν τηε εξιςτενςε οφ ςυςη α ςπιν-1 παρτιςλε ανδ νο ονε υςεδ τηις μοδελ το ςομπυτε ανψ πηψςιςαλ χυαντιτψ. Τηε χυαρϰς ανδ τηε γλυον ωερε νοτ ςονςιδερεδ ας πηψςιςαλ δψναμιςαλ αριαβλες. Ιν 1972 Η. Φριτζςςη ανδ Μ. Γελλ-Μανν προποςεδ το ενλαργε τηε ςινγλετ γλυον μοδελ το α ςολορ οςτετ οφ γλυονς ενδοωεδ ωιτη Ψανγ-Μιλλς ιντεραςτιονς[155], [156]. Ιτ ωας τηε τηεορψ οφ Χ῝Δ βυτ ςτιλλ, ιτ ωας νοτ μεαντ το βε α δψναμιςαλ χυαντυμ φιελδ τηεορψ. Ιν τηε αβςτραςτ οφ ρεφερενςε [156] ωε ρεαδ: “Ιτ ις ποιντεδ ουτ τηατ τηερε αρε ςεεραλ αδανταγες ιν αβςτραςτινγ προπερτιες οφ ηαδρονς ανδ τηειρ ςυρρεντς φρομ α Ψανγ-Μιλλς γαυγε μοδελ βαςεδ ον ςολορεδ χυαρϰς ανδ ςολορ οςτετ γλυονς.” Τηις ινςιςτενςε το ρεφυςε τηε εξιςτενςε οφ χυαρϰς ας “ηαρδ γραινς” ινςιδε τηε ηαδρονς ηαδ αλςο ινφλυενςεδ τηε εξπεριμενταλ προγραμ, ιν παρτιςυλαρ τηατ οφ ΙΣΡ[157]. Ας λατε ας 1973, προποςαλς το βυιλδ τοταλ αβςορπτιον δετεςτορς ςομπλετελψ ςυρρουνδινγ τηε ιντεραςτιον ρεγιον ωερε ρεϑεςτεδ. Τηε πρεαιλινγ πηιλοςοπηψ αμονγ λεαδινγ τηεοριςτς ανδ εξπεριμενταλιςτς ωας το δενψ ανψ ιντερεςτ το τηε ςτυδψ οφ ςςαττερινγ ατ λαργε ανγλες. “Λετ υς φοργετ 90°. Νοτηινγ ηαππενς ατ 90°”. Προποςαλς βψ ψουνγ εξπεριμενταλιςτς το λοοϰ φορ εεντς ατ λαργε ανγλες ωερε ςςορνφυλλψ διςμιςςεδ.
Ιτ ωας ονλψ δυρινγ τηε λατε 1960ς τηατ χυαντιτατιε χυαντυμ φιελδ τηεορψ ωας ρε-ιντροδυςεδ ιν ηιγη ενεργψ πηψςιςς βψ α νεω γενερατιον οφ τηεοριςτς. Ι ηαε μεντιονεδ αλρεαδψ ςεεραλ οφ τηεμ ανδ Ι ωαντ το αδδ ηερε Κ. Ωιλςον ωηο πλαψεδ α προμινεντ ρολε ιν αδοςατινγ τηε υςε οφ χυαντυμ φιελδ τηεορψ το υνδερςτανδ τηε ςςαλινγ βεηαιορ φουνδ ατ ΣΛΑ῝. Ηις ιδεας ηαε ινφλυενςεδ προφουνδλψ ουρ πρεςεντ ιεως ανδ ηαε βριδγεδ τηε γαπ βετωεεν παρτιςλε πηψςιςς ανδ ςτατιςτιςαλ μεςηανιςς. Ηε δεελοπεδ α νοελ αππροαςη το τηε ρενορμαλιςατιον γρουπ[158] ανδ, ιν 1968, ηε προποςεδ α γενεραλ μετηοδ το ςτυδψ τηε προδυςτ οφ ςυρρεντ οπερατορς ατ ςηορτ διςτανςες, ςαλλεδ τηε Οπερατορ Προδυςτ Εξπανςιον” (ΟΠΕ). Τηις ωορϰ προεδ το βε εξτρεμελψ ποωερφυλ ανδ ιτς ιμπορτανςε γοες βεψονδ τηε αππλιςατιον το δεεπ ινελαςτις ςςαττερινγ[159]. Ωε νοτιςεδ αλρεαδψ τηατ ιν τηε προδυςτ οφ τηε τωο ςυρρεντς οφ εχ. (58) $y^{2}\approx 0$. Ιφ ωε ςοντινυε το Ευςλιδεαν ςπαςε, $y^{2}=0$ ιμπλιες $y_{\mu}=0$, ιν οτηερ ωορδς τηε προδυςτ ις ταϰεν ατ ςηορτ διςτανςε. Λετ $A(x)$ ανδ $B(y)$ βε τωο λοςαλ οπερατορς. Ωιλςον ποςτυλατεδ τηατ ατ ςηορτ διςτανςες, τηε προδυςτ $A\cdot B$ ςαν βε εξπανδεδ ας:
$A(x)B(0)|_{x_{\mu}\rightarrow 0}\sim\sum_{i}C_{i}(x)O^{i}(0)$ (59)
ωηερε τηε ςυμ εξτενδς οερ τηε ινφινιτε ςετ οφ αλλ λοςαλ οπερατορς $O^{i}$ ωηιςη ηαε τηε ςαμε χυαντυμ νυμβερς ας τηε προδυςτ ιν τηε λεφτ. $C_{i}(x)$ αρε $c$-νυμβερ ςοεφφιςιεντ φυνςτιονς ωηιςη μαψ βε ςινγυλαρ ωηεν $x$ γοες το ζερο. Τωο ποιντς αρε ιμπορταντ ιν τηις εξπανςιον: Φιρςτ, ιτ ις αςςυμεδ το βε αν οπερατορ εχυατιον. Τηε ςοεφφιςιεντ φυνςτιονς $C_{i}$ δο νοτ δεπενδ ον τηε παρτιςυλαρ ματριξ ελεμεντ ονε μαψ ςονςιδερ. Σεςονδ, τηε βεηαιουρ οφ $C_{i}(x)$ ατ τηε οριγιν ςαν βε δετερμινεδ βψ διμενςιοναλ αναλψςις. Ιφ $d_{A}$, $d_{B}$ ανδ $d_{i}$ αρε τηε διμενςιονς οφ τηε οπερατορς $A$, $B$ ανδ $O^{i}$ ρεςπεςτιελψ, τηε ςοεφφιςιεντ φυνςτιονς $C_{i}$ αρε αςςυμεδ το βεηαε ατ ςηορτ διςτανςες, υπ το λογαριτημις ςορρεςτιονς, ας:
$C_{i}(x)|_{x_{\mu}\rightarrow 0}\sim|x|^{d_{i}-d_{A}-d_{B}}$ (60)
ωηερε $|x|$ δενοτες τηε μοδυλυς οφ $x_{\mu}$. Ιτ φολλοως τηατ τηε δομιναντ ςοντριβυτιον ςομες φρομ τηε οπερατορς $O^{i}$ ωιτη τηε λοωεςτ διμενςιον $d_{i}$. Ιν α φουρ διμενςιοναλ φιελδ τηεορψ, ιφ ωε ςονςιδερ ονλψ οπερατορς ωηιςη αρε μονομιαλς ιν τηε φιελδς ανδ τηειρ δεριατιες, τηερε ις αλωαψς α φινιτε νυμβερ οφ οπερατορς ωιτη α γιεν $d_{i}$ ανδ τηις ιμπλιες τηατ ανψ προδυςτ (59) ωιλλ βε αππροξιματεδ, ωιτη ανψ δεςιρεδ αςςυραςψ, βψ α φινιτε νυμβερ οφ τερμς. Τηις ςηορτ διςτανςε ΟΠΕ ωας λατερ εξτενδεδ το α λιγητ ςονε ΟΠΕ ιν αςςορδανςε ωιτη εχυατιον (58) ιν Μινϰοωςϰι ςπαςε[160].
Χυαντυμ ςηρομοδψναμιςς, ας α δψναμιςαλ χυαντυμ φιελδ τηεορψ οφ ςτρονγ ιντεραςτιονς, ωας προποςεδ, φολλοωινγ τηε διςςοερψ οφ αςψμπτοτις φρεεδομ, βψ Η.Δ. Πολιτζερ ας ωελλ ας Δ.Θ. Γροςς ανδ Φ. Ωιλςςζεϰ[153]. Ιτ ις αν υνβροϰεν γαυγε τηεορψ βαςεδ ον ςολορ $SU(3)$. Ιτ δεςςριβες τηε δψναμιςς αμονγ $N$ ςολορ τριπλετς οφ ςπιν-1/2 χυαρϰς ιντεραςτινγ ωιτη α ςολορ οςτετ οφ μαςςλεςς ςπιν-1 γλυονς. Ατ πρεςεντ ωε ηαε $N=6$. Τηε τηεορψ ις νοτ ρεςτριςτεδ το ςηορτ διςτανςες. Ιτ ςοερς αλλ ςςαλες, ινςλυδινγ τηε φορματιον οφ ηαδρονς. Ωηατ δεπενδς ον τηε ςςαλε αρε τηε ςομπυτατιοναλ τοολς ωε υςε το εξτραςτ πηψςιςαλ ινφορματιον: περτυρβατιον εξπανςιον τεςηνιχυες ατ ςηορτ διςτανςες ανδ νον-περτυρβατιε μετηοδς, ςυςη ας λαττιςε ςιμυλατιονς, ιν τηε ςτρονγ ςουπλινγ ρεγιμε. Χ῝Δ ηας βεεν ερψ ςυςςεςςφυλ ιν βοτη ρεγιμες.
$\bullet$ $e^{+}e^{-}$ αννιηιλατιον το ηαδρονς. Περτυρβατιε μετηοδς αρε διρεςτλψ αππλιςαβλε ωηεν ςομπυτινγ ςορρελατιον φυνςτιονς $G(p_{1},\cdots p_{n})$ ιν τηε ϰινεματιςαλ ρεγιον ιν ωηιςη αλλ ςςαλαρ προδυςτς $p_{i}\cdot p_{j}$ αρε λαργε ςομπαρεδ το μαςς παραμετερς. Τηε φιρςτ οβιους αππλιςατιον ις τηε τοταλ ηαδρονις προδυςτιον ςροςς ςεςτιον ιν $e^{+}e^{-}$ αννιηιλατιον, ατ τηε ονε-πηοτον εξςηανγε αππροξιματιον. Βψ τηε οπτιςαλ τηεορεμ, ιτ ις γιεν βψ τηε ιμαγιναρψ παρτ οφ τηε πηοτον προπαγατορ, ας ςηοων ιν Φιγ. 1. Τηε ρεςυλτ, νορμαλιςεδ βψ τηε $\mu^{+}\mu^{-}$ προδυςτιον ςροςς ςεςτιον, ις[161]:
Φιγυρε 0: Τηε ηαδρονις ςοντριβυτιονς το τηε πηοτον προπαγατορ ατ ζερο ορδερ ιν $\alpha_{s}$ (α), ανδ τηε ονε-γλυον ςορρεςτιονς (β), (ς) ανδ (δ). Ιν ουρ ςαςε τηε λαςτ τωο διαγραμς γιε εχυαλ ςοντριβυτιονς.
$R(Q^{2})=\frac{\sigma(e^{+}+e^{-}\rightarrow{\rm hadrons})}{\sigma(e^{+}+e^{-}\rightarrow\mu^{+}+\mu^{-})}$ (61)
$R(Q^{2})=\sum_{i}e_{i}^{2}\left(1+\frac{\alpha_{s}(Q^{2})}{\pi}+{\cal O}(\alpha_{s}^{2})\right)\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \alpha_{s}(Q^{2})=\frac{1}{4\pi b_{0}\ln(Q^{2}/\Lambda^{2})}$ (62)
ωιτη $b_{0}=(11-2N_{f}/3)/4\pi^{2}$, $N_{f}$ ις τηε νυμβερ οφ χυαρϰ φλαορς, $e_{i}$ ις τηε ελεςτρις ςηαργε οφ τηε $i$τη χυαρϰ τηατ ςαν βε προδυςεδ ωιτη α πηοτον ενεργψ $Q^{2}$, ανδ $\Lambda\sim{\cal O}$(200 Με῞).
$\bullet$ Δεεπ ινελαςτις ςςαττερινγ. Τηε $e^{+}+e^{-}$ τοταλ ηαδρονις ςροςς ςεςτιον ωας εαςψ το αναλψςε βεςαυςε, βψ αππλψινγ τηε οπτιςαλ τηεορεμ, ωε εξπρεςςεδ ιτ ιν τερμς οφ α χυαντιτψ, τηε πηοτον 2-ποιντ φυνςτιον, ωηιςη δεπενδς ονλψ ον $Q^{2}$. Τηε προπερτψ οφ αςψμπτοτις φρεεδομ ιμπλιες τηατ $\alpha_{s}$ βεςομες ςμαλλ ατ λαργε $Q^{2}$ ανδ ωε ςαν υςε περτυρβατιον τηεορψ. Ουρ νεξτ εξαμπλε ωιλλ βε μορε ςομπλιςατεδ. Ιτ ις τηε ςαλςυλατιον οφ τηε Χ῝Δ ςοντριβυτιονς το τηε δεεπ ινελαςτις ελεςτρον-νυςλεον ςςαττερινγ ςηοων ιν Φιγυρε 1. Ιν εχυατιον (58) ωε εξπρεςςεδ τηε ςροςς ςεςτιον ιν τερμς οφ τηε ματριξ ελεμεντ οφ α προδυςτ οφ τωο ςυρρεντς ταϰεν βετωεεν ονε-προτον ςτατες ωηοςε μομεντυμ ις αλωαψς $P^{2}=M_{N}^{2}$. Τηε δεπενδενςε ον $M_{N}$ βρινγς α ςεςονδ ςςαλε ωηιςη ινολες λαργε διςτανςες, ουτςιδε τηε Χ῝Δ ωεαϰ ςουπλινγ ρεγιμε. Τηερεφορε τηε δεεπ ινελαςτις ςτρυςτυρε φυνςτιονς αρε νοτ διρεςτλψ ςαλςυλαβλε ιν περτυρβατιον. Τηε φιρςτ Χ῝Δ αππλιςατιονς υςεδ τηε Ωιλςον ΟΠΕ ιν ορδερ το φαςτορ τηε λαργε διςτανςε παρτ, γιεν βψ τηε ματριξ ελεμεντ οφ τηε οπερατορς $O_{i}$ βετωεεν προτον ςτατες, φρομ τηε ςηορτ διςτανςε παρτ γιεν βψ τηε φυνςτιονς $C_{i}$[162]. Τηε φορμερ ις $Q^{2}$ ινδεπενδεντ ανδ ις δετερμινεδ βψ εξπεριμεντ ατ ςομε φιξεδ $Q_{0}^{2}$. Τηε λαττερ ις ςαλςυλαβλε ιν α Χ῝Δ περτυρβατιον εξπανςιον ανδ πρεδιςτς ιολατιονς οφ εξαςτ ςςαλινγ. Τηις ρεςυλτ ωας νοτ ιμμεδιατελψ αςςεπτεδ φρομ πεοπλε ωηο ςονςιδερεδ Χ῝Δ ας α μερε μοδελ το εξτραςτ τηε λιγητ ςονε ςυρρεντ αλγεβρα ανδ νοτ ας α ρεαλ δψναμιςαλ τηεορψ [^67] It seems that when K. Wilson attempted to present to his former PhD supervisor M. Gell-Mann the QCD prediction of scaling violations, the latter did not want to listen. In my report at the London Conference[163] I made an analogy with politics and remarked that most people misunderstood the meaning of asymptotic freedom by dropping “asymptotic” and keeping only “freedom”: “As it is often the case, whenever someone talks about freedom, it invariably turns out that he really means something else.”. Ιν λατερ ψεαρς τηε εφφεςτς οφ ςςαλινγ ιολατιονς ωερε ερψ ωελλ μεαςυρεδ, ιν παρτιςυλαρ βψ τηε ΗΕΡΑ ςολλαβορατιονς, ιν φυλλ αγρεεμεντ ωιτη Χ῝Δ πρεδιςτιονς.
$\bullet$ Ηαδρονις προςεςςες – Τηε ΑΠΔΓΛ εχυατιονς. Αππλψινγ Χ῝Δ τεςηνολογψ το ηαδρονις ςολλιςιονς ωας εεν μορε ςομπλιςατεδ βεςαυςε ανψ παρτιςυλαρ ςςαττερινγ αμπλιτυδε ινολες ον-μαςς-ςηελλ ηαδρονς ανδ τηε ΟΠΕ μετηοδ ις νοτ ιμμεδιατελψ αππλιςαβλε. Ονε τριεδ ινςτεαδ το ιςολατε ινφραρεδ-φρεε ινςλυςιε οβςεραβλες, ςυςη ας τηε αεραγε ενεργψ $E(\Omega)$ ςαρριεδ ιντο α γιεν ςολιδ ανγυλαρ ρεγιον $\Omega$[164]. Αν ιμπορταντ ςτεπ, βοτη ςονςεπτυαλ ανδ πραςτιςαλ, ωας τηε ςομβινατιον οφ Χ῝Δ ωιτη τηε ςιμπλε πιςτυρε οφ τηε παρτον μοδελ. Τηεψ αρε τηε Αλταρελλι, Παριςι, Δοϰςηιτζερ, Γριβο, Λιπατο εχυατιονς[165] ωηιςη μαδε ποςςιβλε τηε υςε οφ τηε αςψμπτοτις φρεεδομ τεςηνιχυες το α λαργε ςλαςς οφ ηαδρονις προςεςςες[145]. Τηε εχυιαλενςε οφ τηε τωο αππροαςηες ωας ςηοων ιν 1978[166].
$\bullet$ Χ῝Δ ιν ςτρονγ ςουπλινγ. Ατ ενεργψ ςςαλες $E\leq$ 1Γε῞ ωε εντερ τηε ςτρονγ ςουπλινγ ρεγιμε ανδ περτυρβατιον τηεορψ δοες νοτ αππλψ. Οερ τηε ψεαρς ςεεραλ αππροξιματιον ςςηεμες ηαδ βεεν προποςεδ ωιτη νο ςιγνιφιςαντ ςυςςεςς. Ι δο νοτ ϰνοω ωηο ωας τηε φιρςτ το ςονςιδερ τηε διρεςτ νυμεριςαλ εςτιματιον οφ τηε πατη ιντεγραλ ωηιςη δεφινες α χυαντυμ φιελδ τηεορψ βψ τρυνςατιον ιν α φινιτε ςπαςε-τιμε λαττιςε [^68] Feynman, in his original 1948 paper which introduced the path integral quantisation method, used a time-lattice but as a means to define the integral, not to compute it numerically. βυτ, νοτ ςυρπριςινγλψ, τηις μετηοδ ωας δεελοπεδ φολλοωινγ τηε ινςρεαςινγ ποωερ οφ ςομπυτινγ ςψςτεμς. Αλτηουγη τηε ιδεα ςουνδς ςιμπλε, ιτς πραςτιςαλ αππλιςατιον φορ Χ῝Δ ις νοτ. Μανψ ιμπορταντ προβλεμς οφ πρινςιπλε ηαδ το βε ςολεδ, βεφορε φαςινγ τηε φορμιδαβλε νυμεριςαλ προβλεμς οφ τηε ςιμυλατιονς. Ι ωιλλ μεντιον ονλψ τωο: τηατ οφ γαυγε ιναριανςε ανδ τηατ οφ τηε πρεςενςε οφ ςηιραλ φερμιονς. Κ. Ωιλςον ωας ινςτρυμενταλ ιν τηε ςολυτιον οφ βοτη προβλεμς. Ιν πηψςιςς ωε υςε το τρεατ ματτερ φιελδς ανδ γαυγε φιελδς ιν α ςιμιλαρ ωαψ – ωε ςηοοςε α γαυγε, χυαντιςε ςανονιςαλλψ, ςομπυτε τηε ςορρεςπονδινγ προπαγατορς ανδ οβταιν τηε Φεψνμαν ρυλες – ανδ τηις οβςςυρες τηε φαςτ τηατ, ματηεματιςαλλψ, τηεψ αρε διφφερεντ οβϑεςτς. Ιν τηε Ωιλςονιαν λαττιςε φορμυλατιον[167] τηις διφφερενςε αππεαρς ςλεαρλψ. Λετ υς τρυνςατε τηε φουρ-διμενςιοναλ Ευςλιδεαν ςπαςε βψ α φινιτε λαττιςε οφ $N$ ποιντς αρρανγεδ ιν α ηψπερςυβε ωιτη λαττιςε ςπαςινγ $a$. Τηε ποιντ $x_{\mu}$ ις ρεπλαςεδ βψ αν ιντεγερ $n$, $1\leq n\leq N$ ωηιςη λαβελς τηε λαττιςε ποιντ. Φορ α φιελδ ωε γετ $\phi(x)\rightarrow\phi_{n}$, ι.ε. α φιελδ λιες ον α λαττιςε ποιντ. Ιν ςοντραςτ, Ωιλςον ςηοωεδ τηατ γαυγε φιελδς λιε ον οριεντεδ λαττιςε λινϰς, ωηιςη εξπλαινς ωηψ ματηεματιςιανς ςαλλ γαυγε φιελδς ςοννεςτιονς. Ιν φαςτ, ςτυδψινγ λαττιςε γαυγε τηεοριες ις α ποορ μαν´ς ωαψ το υνδερςτανδ διφφερεντιαλ γεομετρψ. Τηε ςεςονδ προβλεμ ις τηατ οφ ςηιραλ φερμιονς. Ιτ ις εαςψ το υνδερςτανδ ιτς οριγιν. Τηε ιντροδυςτιον οφ α φινιτε λαττιςε ςπαςινγ $a$ προιδες αν υλτραιολετ ςυτ-οφφ ανδ Ωιλςον ςηοωεδ τηατ ιτ ρεςπεςτς γαυγε ιναριανςε. Ωηεν ωε διςςυςςεδ τηε ςηιραλ ανομαλψ ωε ςαιδ τηατ τηερε ις νο ςυτ-οφφ ρεςπεςτινγ βοτη γαυγε ανδ ςηιραλ ςψμμετρψ. Ιτ φολλοως τηατ ςηιραλ φερμιονς ςαννοτ βε διρεςτλψ ιντροδυςεδ ον α λαττιςε[168]. Ωιλςον ςυγγεςτεδ τηε μετηοδ οφ φερμιον δουβλινγ ωηιςη ις ωιδελψ υςεδ. Λαττιςε ςαλςυλατιονς οφ ςτρονγ ιντεραςτιον εφφεςτς ηαε βεςομε α μαϑορ ρεςεαρςη φιελδ ιν ρεςεντ ψεαρς ωιτη μανψ ιμπορταντ, ανδ ερψ πρεςιςε, ρεςυλτς[169]. Τηεψ ηαε ςοντριβυτεδ ιν εςταβλιςηινγ Χ῝Δ ας τηε τηεορψ οφ ςτρονγ ιντεραςτιονς.
[…]
Τηε τηεορετιςαλ ςςηεμε, ωηιςη βεςαμε τηε Στανδαρδ Τηεορψ οφ παρτιςλε πηψςιςς, ωας φυλλψ ωριττεν ιν 1973. Ψετ, ιτ ωας νοτ γενεραλλψ αςςεπτεδ. Φορ μοςτ πηψςιςιςτς ιτ ωας α ωιλδ τηεορετιςαλ ςπεςυλατιον ωιτη νο ςοννεςτιον το τηε ρεαλ ωορλδ. Τηε μαιν ρεαςον φορ τηις νεγατιε αττιτυδε ωας αγαιν τηε μιςτρυςτ τοωαρδς χυαντυμ φιελδ τηεορψ, βυτ ιτ ις αλςο τρυε τηατ τηε μοδελ ςεεμεδ το μαϰε μανψ ςτρανγε πρεδιςτιονς ωιτη λιττλε, ιφ ανψ, εξπεριμενταλ ςυππορτ. Λετ με μεντιον ςομε οφ τηεμ.
$\bullet$ Τηε μοδελ πρεδιςτεδ τηε εξιςτενςε οφ 12 εςτορ βοςονς. Βυτ ονλψ ονε, τηε πηοτον, ωας ϰνοων! Τηρεε, ($W^{\pm},\leavevmode\nobreak\ Z^{0}$) ωερε πρεδιςτεδ το βε ερψ ηεαψ [^69] For the 1973 physicists a mass of 100 GeV was essentially infinite! ανδ τηε οτηερ ειγητ, τηε γλυονς, ωερε δεςλαρεδ υνοβςεραβλε βψ α ςτρανγε προπερτψ οφ ςονφινεμεντ.
$\bullet$ Τηε μοδελ πρεδιςτεδ τηε εξιςτενςε οφ α ςςαλαρ βοςον, τηε ΒΕΗ, ωιτη υνϰνοων μαςς. Φορ μανψ πηψςιςιςτς ιτ ωας α ηερεςψ ςομινγ αφτερ τηε αςςυμεδ τριυμπη οφ τηε $V-A$ τηεορψ οφ ωεαϰ ιντεραςτιονς.
$\bullet$ Νευτραλ ςυρρεντς ωερε πρεδιςτεδ βυτ τηε οβιους ονες, τηε $K^{0}\rightarrow\mu^{+}+\mu^{-}$ δεςαψ, ωερε εξςλυδεδ. Γαργαμελλε ηαδ εςταβλιςηεδ τηε εξιςτενςε οφ ςτρανγενεςς ςονςερινγ νευτραλ ςυρρεντς βυτ, ας Ι ρεμαρϰεδ ιν ςεςτιον 1, νοτ εερψβοδψ ωας ςονινςεδ. Φυρτηερμορε, τηε ποςςιβλε εξιςτενςε οφ ωεαϰ νευτραλ ςυρρεντς ηαδ βεεν ενιςαγεδ βεφορε τηε φορμυλατιον οφ γαυγε τηεοριες, ςο νοτ εερψβοδψ ςονςιδερεδ τηεμ ας α δεςιςιε προοφ οφ τηε Στανδαρδ Μοδελ.
$\bullet$ Προβαβλψ τηε μοςτ “εξτρααγαντ” πρεδιςτιον ωας τηατ οφ τηε ςηαρμεδ χυαρϰ, ιμπλψινγ τηε εξιςτενςε οφ αν εντιρε νεω ηαδρονις ωορλδ οφ ςηαρμεδ παρτιςλες. Φορ μοςτ πεοπλε τηε αργυμεντς ωερε νοτ ςονςιδερεδ ςεριους. Ι ςτιλλ ρεμεμβερ τηε οβϑεςτιονς: ςομε οβςςυρε ηιγηερ ορδερ εφφεςτς – τριανγλε διαγραμς φορ τηε ανομαλιες, ορ ςχυαρε διαγραμς φορ τηε αβςενςε οφ φλαορ ιολατινγ νευτραλ ςυρρεντς – ωουλδ διςτατε τηε ςτρυςτυρε οφ τηε ωορλδ; Τοταλλψ αβςυρδ! Ιν ρετροςπεςτ, Ι τηινϰ τηατ τηε λαργε μαϑοριτψ οφ πηψςιςιςτς ρεϑεςτεδ τηις παρτιςυλαρ πρεδιςτιον βεςαυςε ιτ ωεντ αγαινςτ τηε πρεαιλινγ πηιλοςοπηψ οφ ςομπαρτμενταλιςατιον οφ ηιγη ενεργψ πηψςιςς. Τηεορετιςαλ αργυμεντς μοτιατεδ βψ προπερτιες οφ ωεαϰ ιντεραςτιονς ωερε νοτ αδμιςςιβλε το μαϰε πρεδιςτιονς ον ηαδρονις πηψςιςς, α δομαιν ρεςερεδ εξςλυςιελψ το ςτρονγ ιντεραςτιονς [^70] Sutor, ne supra crepidam. or, Let the cobbler stick to his last..
$\bullet$ Τηε Χ῝Δ πρεδιςτιον φορ τηε ρατιο $R$ οφ εχ. (62) ςεεμεδ το βε ιν ιολεντ ςοντραδιςτιον ωιτη εξπεριμεντ.
Τηις βρινγς με το τηε 17τη Ιντερνατιοναλ ῝ονφερενςε ον Ηιγη Ενεργψ Πηψςιςς ηελδ ιν Λονδον ιν Θυλψ 1974. Ιτ ςαν βε ιεωεδ ας τηε λαςτ ῝ονφερενςε οφ τηε Δαρϰ Αγες, βυτ αλςο τηε ονε ωηιςη αννουνςεδ τηε Νεω Ερα. Ιφ ωε λοοϰ ατ τηε προγραμ οφ τηε πρειους ςονφερενςες ωε φινδ τηε τραδιτιοναλ ςςηεμε οφ ςεςςιονς ον ωεαϰ, ελεςτρομαγνετις ανδ ςτρονγ ιντεραςτιονς. Τηε μαν βεηινδ τηε οργανιςατιον οφ τηε Λονδον ῝ονφερενςε ωας Αβδυς Σαλαμ ωηο ςυςςεεδεδ το ινϑεςτ α ςμαλλ παρτ οφ α νεω ιςιον: τηερε ωερε ςτιλλ ςεςςιονς ον ςτρονγ ιντεραςτιονς, ρεςοναντ πηψςιςς ετς, βυτ ωε φινδ αλςο α ςεςςιον ωιτη α ρεπορτ ον γαυγε τηεοριες ωηιςη ινςλυδεδ τηε αδανςες ιν αλλ τηρεε ιντεραςτιονς, ιν οτηερ ωορδς α ρεπορτ ον ωηατ ωας γοινγ το βεςομε τηε Στανδαρδ Μοδελ. Σεεραλ ολδ ανδ νεω ρεςυλτς ωερε αννουνςεδ ιν τηις ῝ονφερενςε. Ι ωιλλ μεντιον ονλψ τηοςε ωηιςη αρε ιμπορταντ φορ ουρ ςτορψ.
$\bullet$ Δ.῝. ῝υνδψ γαε τηε πλεναρψ ρεπορτ ον νευτρινο πηψςιςς. Οβιουςλψ, τηε Γαργαμελλε ρεςυλτς ωερε τηε ςεντραλ ποιντ. Ι ρεμεμβερ ηιμ ςαψινγ: “Τηοςε ωηο ηαε βετ ον νευτραλ ςυρρεντς, νοω ις τηε τιμε το παψ ανδ ςολλεςτ.”
$\bullet$ Β. Ριςητερ γαε τηε πλεναρψ ρεπορτ ον $e^{+}e^{-}\rightarrow$ ηαδρονς. Ιν Φιγυρε 1 Ι ςηοω τηε γραπη ηε πρεςεντεδ.
Φιγυρε 0: Α ςομπιλατιον οφ αλλ εαρλψ μεαςυρεμεντς οφ τηε ρατιο $R$, ας πρεςεντεδ ιν τηε 1974 Λονδον Ιντερνατιοναλ ῝ονφερενςε ον Ηιγη Ενεργψ Πηψςιςς βψ Βυρτον Ριςητερ.
Ι ρεμινδ ψου τηατ τηε Χ῝Δ πρεδιςτιον οφ εχ. (62) ωας τηατ $R$ ςηουλδ αππροαςη τηε αλυε $R=2$ (τηε ςυμ οφ τηε ελεςτρις ςηαργες οφ $u$, $d$ ανδ $s$ χυαρϰς) φρομ αβοε. Νο ςυςη βεηαιορ ις ιςιβλε ιν Φιγυρε 1.
$\bullet$ Ιν μψ ρεπορτ ον γαυγε τηεοριες[163] Ι ςαιδ: “…τηε ηαδρον προδυςτιον ςροςς ςεςτιον, ωηιςη αβςολυτελψ ρεφυςες το φαλλ, ςρεατες α ςεριους προβλεμ. Τηε βεςτ εξπλανατιον μαψ βε τηατ ωε αρε οβςερινγ τηε οπενινγ οφ τηε ςηαρμεδ τηρεςηολδς, ιν ωηιςη ςαςε εερψτηινγ φιτς τογετηερ ερψ νιςελψ.” Τηε αδδιτιον οφ α ςηαρμεδ χυαρϰ ωουλδ αδδ αν εξτρα 4/3 το $R$ [^71] By a numerical accident, the data of Figure 1 contain also the production of the $\tau$ lepton which was not known at the time.. Σαλαμ ηαδ ρεαδψ α βοττλε οφ α φινε Βορδεαυξ ρεδ ωινε το ρεωαρδ ςπεαϰερς ωηο φινιςη ον τιμε. Νατυραλλψ Ι ωας λατε ανδ Ι ςαιδ ςομετηινγ λιϰε: “Ι ϰνοω Ι αμ αβουτ το λοοςε μψ βοττλε, βυτ Ι αμ ρεαδψ το βετ νοω α ωηολε ςαςε τηατ, ιφ τηε ωεαϰ ιντεραςτιον ςεςςιονς οφ τηις ῝ονφερενςε ωερε δομινατεδ βψ τηε διςςοερψ οφ τηε νευτραλ ςυρρεντς, τηε εντιρε νεξτ ῝ονφερενςε ωιλλ βε δομινατεδ βψ τηε διςςοερψ οφ τηε ςηαρμεδ παρτιςλες.” Τηις ςονινςεδ Σαλαμ το γιε με τηε βοττλε ωηιςη Ι οπενεδ ιμμεδιατελψ, πουρεδ μψςελφ α γενερους λιβατιον, οφφερεδ τηε ρεςτ το τηοςε ςιττινγ ιν τηε φιρςτ ροω, ανδ δρανϰ “το ςηαρμ!”.
[…]
Ιν Νοεμβερ 1974 βοτη Βροοϰηαεν ανδ Στανφορδ πυβλιςηεδ τηειρ ρεςυλτς[2]. ΣΠΕΑΡ δεςιδεδ το ςωεεπ τηε ρεγιον αβοε 3 Γε῞ ιν φινε ςτεπς οφ 1 Με῞. Το τηειρ γρεατ ςυρπριςε τηεψ οβταινεδ α τοταλλψ διφφερεντ πιςτυρε, Φιγυρε 1.
Φιγυρε 0: Τηε διςςοερψ οφ τηε $J/\Psi$ μεςον ιν Νο. 1974 ινδεπενδεντλψ βψ ΣΠΕΑΡ (λεφτ) ανδ ΑΓΣ (ριγητ). Βοτη εξηιβιτ πεαϰς ιν τηε οπποςιτελψ ςηαργεδ διελεςτρον μαςς ςπεςτρυμ ςονςιςτεντ ωιτη τηε $J/\Psi$ μαςς ατ 3.1 Γε῞. Τηις ρεςυλτ ωας αλςο ςονφιρμεδ βψ τηε Φραςςατι γρουπ.
Ι ωας ιν Παρις ατ τηε ενδ οφ 1974 ωηεν Ι ρεςειεδ α τελεπηονε ςαλλ φρομ Α. Λαγαρριγυε ινιτινγ με το αν ινφορμαλ μεετινγ το διςςυςς ιμπορταντ ρεςυλτς φρομ ΣΠΕΑΡ. Γ. ´τ Ηοοφτ ωας ιςιτινγ Εςολε Νορμαλε ανδ Ι τοοϰ ηιμ αλονγ. Ιν τηοςε δαψς νεως διδ νοτ τραελ ωιτη τηε ςπεεδ οφ ιντερνετ. Β. Θεαν-Μαριε ηαδ ςομε φρομ Στανφορδ ωιτη τηε ρεςυλτς. Τηεψ ωερε ινδεεδ ερψ ιμπρεςςιε. Α 3 Γε῞ ηαδρον, δεςαψινγ ιντο πιονς, ωιτη α ωιδτη οφ λεςς τηαν 100 ϰε῞; Ινςρεδιβλε! Λαγαρριγυε αςϰεδ με ωηατ Ι τηουγητ οφ ιτ ανδ Ι ςονφεςς Ι ωας βεωιλδερεδ. Ιτ ωας ´τ Ηοοφτ ωηο φιρςτ γαε με τηε εξπλανατιον. Ιτ ις ςιμπλε το υνδερςτανδ ιν Χ῝Δ. Λετ υς ςονςιδερ τηε ςεριες οφ $1^{-}$ μεςονς: $\rho$ ις α βουνδ ςτατε οφ α χυαρϰ-αντιχυαρϰ παιρ οφ τηε φιρςτ φαμιλψ. Ιτς μαςς λαψς ωελλ αβοε τηε $2\pi$ μαςς ανδ ιτς ωιδτη ις ερψ λαργε $\sim$ 147 Με῞. Τηε $\phi$ ις αν $s\bar{s}$ βουνδ ςτατε ωιτη α μαςς οφ 1020 Με῞. Ιτ λαψς βαρελψ αβοε τηε τηρεςηολδ οφ α $K\bar{K}$ παιρ, ψετ ιτς βρανςηινγ ρατιο το $K\bar{K}$ ις 83% δεςπιτε τηε φαςτ τηατ τηε πηαςε ςπαςε ις τινψ. Τηε πυρε πιονις παρτιαλ ωιδτη ις ονλψ 650 ϰε῞. Ιν τηε ολδ δαψς ωε ηαδ α ρυλε, ςαλλεδ τηε ΟΖΙ (Οϰυβο-Ζωειγ-Ιιζυϰα) ρυλε, ονε οφ τηοςε εμπιριςαλ ρυλες οφ τηε δαρϰ αγες ωιτη νο ρεαλ τηεορετιςαλ ϑυςτιφιςατιον. Ιτ ςτατεδ τηατ ιν α χυαρϰ-αντιχυαρϰ βουνδ ςτατε, τηε δεςαψ μοδες ρεχυιρινγ τηε αννιηιλατιον οφ τηε ινιτιαλ $q\bar{q}$ παιρ, ωερε ηιγηλψ ςυππρεςςεδ [^72] It was called “the re-arrangement model” by H. Rubinstein.. Λετ υς ςομε νοω το $J/\Psi$ ανδ αςςυμε ιτ ις α $c\bar{c}$ βουνδ ςτατε. Τηε $0^{-}$ μεςονς αρε ςυπποςεδ το βε τηε πςευδο-Γολδςτονε βοςονς οφ ςποντανεουςλψ βροϰεν ςηιραλ ςψμμετρψ. Τηε λαττερ ις ερψ γοοδ φορ τηε φιρςτ φαμιλψ, χυεςτιοναβλε φορ τηε $s$ χυαρϰ, ανδ ερψ ποορ φορ ςηαρμ. Τηερεφορε ωε εξπεςτ τηε ςηαρμεδ $0^{-}$ μεςονς $D$ το βε χυιτε ηεαψ ανδ τηε μαςς οφ $J/\Psi$ το λαψ βελοω τηε $D\bar{D}$ τηρεςηολδ. Ας α ρεςυλτ $J/\Psi$ δεςαψς μαινλψ ιντο πιονς. Τηε δεςαψ αμπλιτυδε φορ α $1^{-}$ μεςον γοες τηρουγη τηρεε γλυονς, ςο τηε ωιδτη ις προπορτιοναλ το $\alpha_{s}^{3}$. Βετωεεν 1 ανδ 3 Γε῞ $\alpha_{s}$ ηας δροππεδ βψ α φαςτορ οφ τωο, ςο ωε εξπεςτ τηε $J/\Psi$ ωιδτη το βε 8 τιμες ςμαλλερ τηαν τηε $\phi$ πιονις ωιδτη. Ιτ ις πρεςιςελψ ωηατ ις φουνδ εξπεριμενταλλψ. Ας Ι ςαιδ, Ι φιρςτ ηεαρδ τηις αργυμεντ φρομ ´τ Ηοοφτ, βυτ λατερ Ι φουνδ ιτ ιν παπερς βψ Αππελχυιςτ ανδ Πολιτζερ ας ωελλ ας Δε Ρυϑυλα ανδ Γλαςηοω [170]. Τηε φιρςτ ηας α ςυβμιςςιον δατε οφ Νοεμβερ 19. Τηις μαϰες με βελιεε ιτ ωας φουνδ βεφορε τηε εξπεριμενταλ διςςοερψ.
Ωιτηιν α ψεαρ τηε εντιρε ρεγιον βετωεεν 3 ανδ 5 Γε῞ ωας ςτυδιεδ ιν δεταιλ, ςεε Φιγ. 1.
Φιγυρε 0: Τηε αλυε οφ $R$ φορ ενεργιες βετωεεν 3 ανδ 5 Γε῞.
Ας εξπεςτεδ, τηερε αρε βροαδερ ρεςονανςες ωιτη μαςςες $\geq$ 4Γε῞ ωηιςη αρε τηοςε λψινγ αβοε τηε $D\bar{D}$ τηρεςηολδ. Λο ανδ βεηολδ, τηε παρτιςλες ωιτη ναϰεδ ςηαρμ ωερε φουνδ αμονγ τηε δεςαψ προδυςτς οφ τηεςε ρεςονανςες. Ιτ ωας ιν 1976. Ιν τηε μεαντιμε α ριςη ςηαρμονιυμ ςπεςτροςςοπψ [^73] I believe that the term “charmonium” was coined by A. De Rujula. ωας διςςοερεδ ιν φυλλ αγρεεμεντ ωιτη τηε τηεορετιςαλ πρεδιςτιονς[171]. Αλτηουγη νοβοδψ παιδ τηε βετ Ι οφφερεδ ιν τηε 1974 ῝ονφερενςε, τηε εντιρε 1976 ονε ωας ινδεεδ δομινατεδ βψ ςηαρμεδ παρτιςλες ανδ γαυγε τηεοριες. Τηε ρεπορτ ιν τηις ῝ονφερενςε βψ Α. Δε Ρυϑυλα ηαδ τηε τιτλε: “Τηεορετιςαλ βαςις οφ τηε νεω παρτιςλες.” Τηε φιρςτ ςεντενςε ις “Ι ρειεω τηε φουρ-χυαρϰ ςτανδαρδ γαυγε φιελδ τηεορψ οφ ωεαϰ, ελεςτρομαγνετις ανδ ςτρονγ ιντεραςτιονς.” Νοω ωε ταλϰ αβουτ “τηε φουρ-χυαρϰ ςτανδαρδ γαυγε φιελδ τηεορψ”. Τηε πηαςε τρανςιτιον φρομ Μανψ Μοδελς το Ονε Τηεορψ ωας ςομπλετε. Τηε ορδερ παραμετερ ηας βεεν τηε φραςτιον οφ πηψςιςιςτς ωηο ςηανγεδ τηειρ ιεως: α ςμαλλ μινοριτψ βεφορε 1974 το τηε λαργε μαϑοριτψ αφτερ 1976. Τηε ςομπλετε εριφιςατιον οφ τηε τηεορψ τοοϰ μανψ μορε ψεαρς ανδ μανψ γρεατ διςςοεριες, βυτ τηε μοοδ οφ τηε ςομμυνιτψ ηαδ ςηανγεδ. Τηε φολλοωινγ διςςοεριες οφ τηε εςτορ βοςονς, τηε $b$ ανδ $t$ χυαρϰς ωηιςη ςομπλετε τηε τηιρδ φαμιλψ οφ τηε $\tau$ λεπτον, τηε γλυον ϑετς ανδ τηε ΒΕΗ ςςαλαρ ας ωελλ ας τηε ερψ γοοδ γενεραλ φιτ υςινγ αλλ ααιλαβλε δατα, ωερε νο μορε γρεατ ςυρπριςες, τηεψ ωερε εξπεςτεδ. ΤΗΕ ΣΤΑΝΔΑΡΔ ΜΟΔΕΛ ηαδ βεςομε ΤΗΕ ΣΤΑΝΔΑΡΔ ΤΗΕΟΡΨ.
[…]
Φεψνμαν ηας ςαιδ τηατ προγρεςς ιν πηψςιςς ις το προε ψουρςελφ ωρονγ ας ςοον ας ποςςιβλε. Φορ ηαλφ α ςεντυρψ νοω ωε ηαε νοτ βεεν αβλε το προε τηε Στανδαρδ Τηεορψ ις ωρονγ. Ιτ ηας παςςεδ ςυςςεςςφυλλψ αλλ τεςτς ανδ αλλ ιτς πρεδιςτιονς ηαε βεεν βριλλιαντλψ εριφιεδ. Ωηατ ςομες νεξτ;
Ιν 2011 τηε Ευροπεαν Πηψςιςαλ Σοςιετψ αωαρδεδ το Γλαςηοω, Μαιανι ανδ μψςελφ τηε Ηιγη Ενεργψ Πηψςιςς Πριζε. Ατ τηις οςςαςιον ωε ωερε ινιτεδ το ςπεαϰ ατ τηε Ευροπεαν ῝ονφερενςε ιν Γρενοβλε. Τηε τιτλε οφ μψ ταλϰ ωας “Φολλοωινγ τηε Πατη οφ ῝ηαρμ” ανδ Ι τριεδ το αργυε τηατ πρεςιςιον μεαςυρεμεντς ατ α ςερταιν ενεργψ ςςαλε αλλοω υς το μαϰε πρεδιςτιονς ατ ςομε ηιγηερ ςςαλε. Λετ με ςταρτ φρομ αν εξπανδεδ ερςιον οφ α πλοτ ςηοωινγ τηε αλυε οφ $R$ φρομ λοω ενεργιες υπ το ανδ αβοε τηε $Z^{0}$ μαςς, Φιγυρε 1.
Φιγυρε 0: Τηε ρατιο $R$ φρομ λοω ενεργιες, υπ το ανδ αβοε τηε $Z$ μαςς. Τηε γρεεν ςυρε ις τηε παρτον μοδελ πρεδιςτιον ανδ τηε ρεδ ονε ινςλυδες Χ῝Δ ςορρεςτιονς. Ρεμαρϰαβλε αγρεεμεντ.
Ωηατ ις μοςτ ρεμαρϰαβλε ις τηε πρεςιςιον ωιτη ωηιςη τηεορετιςαλ πρεδιςτιονς φιτ τηε εξπεριμενταλ δατα ιν μοςτ οφ τηε ενεργψ ιντεραλ. Τηε ονλψ ρεγιονς ιν ωηιςη τηερε ις νο αγρεεμεντ αρε τηε ονε ατ λοω ενεργψ βελοω 1 Γε῞, ιν ωηιςη τηε Χ῝Δ ις ιν τηε ςτρονγ ςουπλινγ ρεγιμε, ανδ ςμαλλ, ερψ λοςαλιςεδ ρεγιονς ςιγναλινγ τηε τηρεςηολδς φορ τηε προδυςτιον οφ νεω ςπεςιες οφ ηαδρονς, ςηαρμ ορ $b$. Ουτςιδε τηοςε ρεγιονς φορμυλα (62) γιες αν εξςελλεντ φιτ, αλτηουγη ιτ ις ϑυςτ τηε ρεςυλτ οφ ονε λοοπ περτυρβατιον τηεορψ. Ι ςονςλυδεδ τηατ, φορ ρεαςονς τηατ αρε νοτ φυλλψ υνδερςτοοδ, περτυρβατιον τηεορψ ις ρελιαβλε ουτςιδε τηε ρεγιονς οφ ςτρονγ ιντεραςτιονς.
Ι τριεδ το υςε τηις φαςτ ιν ορδερ το μαϰε πρεδιςτιονς ρεγαρδινγ τηε μυλτι-ηυνδρεδ Γε῞ ςςαλε ωηιςη ωας εξπεςτεδ το βε εξπλορεδ βψ ΛΗ῝. Τηε αργυμεντ ωας βαςεδ ον τηε λοω ενεργψ δατα ωηιςη φαορεδ α Ηιγγς παρτιςλε οφ ρελατιελψ λοω μαςς, $\leq$ 200 Γε῞. Ιτ ωεντ ρουγηλψ ας φολλοως: Τηε εξιςτινγ δατα αρε ςομπατιβλε ωιτη τηε Στανδαρδ Τηεορψ ονλψ ιφ τηε Ηιγγς ις λιγητ. Τηερεφορε, ιφ ωε φινδ α ερψ ηεαψ Ηιγγς ωε μυςτ αλςο φινδ νεω ιντεραςτιονς ωηιςη ιναλιδατε τηε Στανδαρδ Τηεορψ ςαλςυλατιονς. Ιφ ον τηε οτηερ ηανδ ωε φινδ α λιγητ Ηιγγς, ωε μυςτ φινδ νεω ιντεραςτιονς ωηιςη ςταβιλιςε ιτς μαςς ατ τηις λοω αλυε. Ι τηουγητ τηατ βοτη ποςςιβιλιτιες ωερε γοοδ νεως φορ ΛΗ῝. Ιν μψ ταλϰ Ι ςαιδ:
“Ι ωαντ το εξπλοιτ τηις εξπεριμενταλ φαςτ [τηε αλιδιτψ οφ περτυρβατιον τηεορψ] ανδ αργυε τηατ τηε ααιλαβλε πρεςιςιον τεςτς οφ τηε Στανδαρδ Μοδελ αλλοω υς το ςλαιμ ωιτη ςονφιδενςε τηατ νεω πηψςιςς ις πρεςεντ ατ τηε Τε῞ ςςαλε ανδ τηε ΛΗ῝ ςαν, προβαβλψ, διςςοερ ιτ. Τηε αργυμεντ αςςυμες τηε αλιδιτψ οφ περτυρβατιον τηεορψ ανδ ιτ ωιλλ φαιλ ιφ τηε λαττερ φαιλς. Βυτ, ας ωε ϑυςτ ςαω, περτυρβατιον τηεορψ βρεαϰς δοων ονλψ ωηεν ςτρονγ ιντεραςτιονς βεςομε ιμπορταντ. Βυτ νεω ςτρονγ ιντεραςτιονς ιμπλψ νεω πηψςιςς.”
Μψ ςονςλυςιον ωας τηατ, φορ ΛΗ῝, ωηιςη ωας αβουτ το ςταρτ οπερατινγ, νεω πηψςιςς ωας αρουνδ τηε ςορνερ!
Τοδαψ ωε ϰνοω τηατ ΛΗ῝ φουνδ νο ςορνερ!
Βυτ Ι ςεςρετλψ βελιεε τηε αργυμεντ ις ςορρεςτ, ονλψ τηε ςορνερ ις α βιτ φυρτηερ δοων.
Αλτηουγη Ι ωιλλ νοτ ςεε ιτ, Ι αμ ςονφιδεντ ςομε αμονγ ουρ ψουνγ ςολλεαγυες ωιλλ φινδ ιτ.
Τηεςε νοτες ηαε τουςηεδ ςο μανψ ςυβϑεςτς τηατ α ςομπλετε λιςτ οφ ρεφερενςες ις ιμποςςιβλε. Τηε ςελεςτιον ις νεςεςςαριλψ αρβιτραρψ ανδ Ι απολογιςε φορ τηε νυμερους ομιςςιονς.
Ρεφερενςες
  • [1] Φορ α ηιςτοριςαλ αςςουντ οφ ῝ΕΡΝ´ς εςταβλιςημεντ, ας ωελλ ας ιτς φιρςτ δεςαδες οφ οπερατιον, ςεε “Στυδιες ιν ῝ΕΡΝ ηιςτορψ” (1993), ῝ΕΡΝ.
  • [2] Θ.Θ. Αυβερτ ετ αλ. Πηψς. Ρε. Λεττ. 33, 1404 (1974)· Θ.-Ε. Αυγυςτιν ετ αλ. Πηψς. Ρε. Λεττ. 33, 1406 (1974)· ῝. Βαςςι ετ αλ. Πηψς. Ρε. Λεττ. 33, 1408 (1974).
  • [3] Τηερε εξιςτ ςεεραλ βοοϰς ναρρατινγ τηε εαρλψ ηιςτορψ οφ μιςροςςοπις πηψςιςς. Ονε Ι οφτεν υςε ις Ινωαρδ Βουνδ βψ Α. Παις, Οξφορδ ϒνιερςιτψ Πρεςς, 1986.
  • [4] Θ. Στεινβεργερ, Ω.Κ.Η. Πανοφςϰψ, ανδ Θ. Στελλερ, Πηψς. Ρε. 78, 802 (1950).
  • [5] Ν. Κεμμερ, Προς. Ροψ. Σος. Α166, 127 (1938)· Προς. ῝αμβρ. Πηιλ. Σος. 34, 354 (1938).
  • [6] Σεε, φορ εξαμπλε, τηε βοοϰ Ενγινες οφ Διςςοερψ – Α ῝εντυρψ οφ Παρτιςλε Αςςελερατορς βψ Α. Σεςςλερ ανδ Ε. Ωιλςον, Ωορλδ Σςιεντιφις Πυβλιςηινγ ῝ο. Πτε. Λτδ., 2007.
  • [7] Σ. Ωεινβεργ, “Τηε Σεαρςη φορ ϒνιτψ: Νοτες φορ α Ηιςτορψ οφ Χυαντυμ Φιελδ Τηεορψ.” Δαεδαλυς, ῞ολ. ΙΙ. Ιςςυεδ ας ῞ολ. 106 Νο 4 οφ τηε Προςεεδινγς οφ τηε Αμεριςαν Αςαδεμψ οφ Αρτς ανδ Σςιενςες π. 17 (1977).
  • [8] Φ.Θ. Δψςον, Προς. Ροψ. Σος. Α207, 395 (1951· Πηψς. Ρε. 83, 608 (1951)· Πηψς. Ρε. 85, 631 (1952).
  • [9] “Σελεςτεδ παπερς οφ Φρεεμαν Δψςον ωιτη ῝ομμενταρψ”, Αμεριςαν Ματηεματιςαλ Σοςιετψ (1996) π. 15.
  • [10] Φορ α γοοδ εξποςιτιον οφ τηε Ωιγητμαν αξιομς ςεε Ρ. Φ. Στρεατερ ανδ Α.Σ. Ωιγητμαν, “Π῝Τ, Σπιν ανδ Στατιςτιςς ανδ Αλλ Τηατ”, Νεω Ψορϰ, Βενϑαμιν (1964).
  • [11] Τηε Ριςε οφ τηε Στανδαρδ Μοδελ, Εδ. βψ Λ. Ηοδδεςον ετ αλ, ῝αμβριδγε ϒνιερςιτψ Πρεςς, 1997.
  • [12] Τηε Στανδαρδ Τηεορψ οφ Παρτιςλε Πηψςιςς, Εδ. βψ Λ. Μαιανι ανδ Λ. Ρολανδι, Ωορλδ Σςιεντιφις, 2016.
  • [13] Φορ τηε ηιςτορψ οφ ςλαςςιςαλ ελεςτροδψναμιςς ςεε Ο. Δαρριγολ, Ελεςτροδψναμιςς φρομ Αμπὲρε το Εινςτειν, Οξφορδ ϒνιερςιτψ Πρεςς, 2000. Ι ηαε αλςο υςεδ τηε φολλοωινγ ρειεω αρτιςλες φορ τηε ηιςτορψ οφ γαυγε τηεοριες: Λ. Ο´Ραιφεαρταιγη, Τηε Δαωνινγ οφ γαυγε τηεορψ, Πρινςετον ϒνιερςιτψ Πρεςς, 1997· Λ. Ο´Ραιφεαρταιγη ανδ Ν. Στραυμανν, Ρε. Μοδ. Πηψς. 72, 1 (2000)· Θ.Δ. Θαςϰςον ανδ Λ.Β. Οϰυν, Ρε. Μοδ. Πηψς. 73, 663 (2001).
  • [14] Δ. Ηιλβερτ, Γο̈ττ. Ναςηρ. (1915) π. 395.
  • [15] Τη. Καλυζα, Κ. Πρευςς. Αϰαδ. Ωιςς. (1921), π. 966· Ο. Κλειν, Ζ. Πηψς. 37, 895 (1926).
  • [16] Γ. Νορδςτρο̈μ, Πηψς. Ζ. 15, 504 (1914).
  • [17] Η.Κ.Η. Ωεψλ, Ανν. Πηψς. (Λειπζιγ) 59, 101 (1919). Φορ α γοοδ πρεςεντατιον οφ ηις ιδεας ονε μαψ λοοϰ ατ τηε ρειεω αρτιςλες [13] ανδ ρεφερενςες τηερειν. Ηε ηας αλςο ωριττεν τωο βοοϰς ωηιςη ηαε βεςομε ςλαςςις: “Ραυμ, Ζειτ, Ματεριε” ιν 1922, τρανςλατεδ ιν Ενγλιςη ας “Σπαςε, Τιμε, Ματτερ” (Μετηυεν, Λονδον, ανδ Δοερ, Νεω Ψορϰ), ανδ “Γρυππεντηεοριε υνδ Χυαντενμεςηανιϰ” ιν 1928, τρανςλατεδ ιν Ενγλιςη ας “Τηεορψ οφ Γρουπς ανδ Χυαντυμ Μεςηανιςς” (Δοερ, Νεω Ψορϰ, 1950). Βοτη ηαδ ςεεραλ ςυβςεχυεντ εδιτιονς ανδ, ιν εαςη ονε, Ωεψλ ωας βρινγινγ ςιγνιφιςαντ ςηανγες, ςο τηε ςυςςεςςιε εδιτιονς οφφερ α γοοδ πιςτυρε οφ τηε εολυτιον οφ ηις ιδεας.
  • [18] ῞.Α. Φοςϰ, Ζ. Πηψς 39, 226 (1926). Τηις παπερ ωας φιρςτ πυβλιςηεδ ιν Γερμαν. Ιτ ωας ρεωριττεν λατερ ιν Ρυςςιαν ανδ τηεν τρανςλατεδ ιν Ενγλιςη Πηψςιςς-ϒςπεϰηι 53, 839 (2010).
  • [19] Η.Κ.Η. Ωεψλ, Ζ. Πηψς. 56, 330 (1929).
  • [20] Φ. Λονδον, Νατυρωιςςενςςηαφτεν 15, 187 (1927) · Ζ. Πηψς. 42, 375 (1927).
  • [21] Ω. Ηειςενβεργ, Ζ. Πηψς. 77, 1 (1932) · Ζ. Πηψς. 78, 156 (1932) · Ζ. Πηψς. 80, 587 (1933).
  • [22] Ο. Κλειν, ιν Λες Νουελλες Τηέοριες δε λα Πηψςιχυε, Παρις 1939, π. 81. Ρεπορτ ιν α ῝ονφερενςε οργανιςεδ βψ τηε ``Ινςτιτυτ Ιντερνατιοναλ δε ῝οοπέρατιον Ιντελλεςτυελλε´´, Ωαρςαω 1938
  • [23] Ο. Κλειν, Ηελ. Πηψς. Αςτα Συππλ. Ι῞, 58 (1956).
  • [24] Ω. Παυλι, ϒνπυβλιςηεδ. Ιτ ις ςυμμαριςεδ ιν α λεττερ το Α. Παις δατεδ Θυλψ 22-25 1953. Α. Παις, Ρεφερενςε [3], π. 584
  • [25] ῝.Ν. Ψανγ, Σελεςτεδ Παπερς 1945-1980 ωιτη ῝ομμενταρψ Πυβλιςηεδ βψ Φρεεμαν, Σαν Φρανςιςςο, π. 525
  • [26] ῝.Ν. Ψανγ ανδ Ρ.Λ. Μιλλς, Πηψς. Ρε. 96, 191 (1954)· Ιτ ςεεμς τηατ ςιμιλαρ ρεςυλτς ωερε αλςο οβταινεδ βψ Ρ. Σηαω ιν ηις τηεςις ατ τηε ϒνιερςιτψ οφ ῝αμβριδγε.
  • [27] Ρ. ϒτιψαμα, Πηψς. Ρε. 101, 1597 (1956).
  • [28] Σ.Λ. Γλαςηοω ανδ Μ. Γελλ-Μανν, Ανν. οφ Πηψς. 15, 437 (1961).
  • [29] Θ.Σ. Σςηωινγερ, Ανν. οφ Πηψς. 2, 407 (1957).
  • [30] Σ.Α. Βλυδμαν, Νυοο ῝ιμ. 9, 433 (1958).
  • [31] Σ.Λ. Γλαςηοω, Νυςλ. Πηψς. 22, 579 (1961).
  • [32] Ε. Σςηρο̈δινγερ, ῝ολλεςτεδ Παπερς ον Ωαε Μεςηανιςς, τρανςλατεδ φρομ τηε 2νδ Γερμαν εδ., 1928, (῝ηελςεα Πυβλιςηινγ, Νεω Ψορϰ) (1978).
  • [33] Ρ. Σερβερ, Πηψς. Ρε. 49, 545 (1936).
  • [34] Η.Α. Κραμερς, Νυο. ῝ιμ. 15, 108 (1938).
  • [35] Ρ.Π. Φεψνμαν, Ρεπορτ ιν: Ινςτιτυτ ιντερνατιοναλ δε πηψςιχυε Σολαψ. Λα τηéοριε χυαντιχυε δες ςηαμπς: δουζιèμε ῝ονςειλ δε πηψςιχυε, Βρυξελλες (1961). Νεω Ψορϰ Ιντερνατιοναλ Πυβλιςηερς (1962).
  • [36] Τ.Τ. Ωυ, Πηψς. Ρε. 123, 678 (1961).
  • [37] Αλλ εαρλψ παπερς ον χυαντυμ ελεςτροδψναμιςς αρε ινςλυδεδ ιν Θ.Σ. Σςηωινγερ, εδ. Σελεςτεδ Παπερς ον Χυαντυμ Ελεςτροδψναμιςς, Δοερ Πυβλιςατιονς (1958). Φορ α μορε ποπυλαρ αςςουντ, ςεε Σ.Σ. Σςηωεβερ, “ΧΕΔ ανδ τηε Μεν Ωηο Μαδε Ιτ: Δψςον, Φεψνμαν, Σςηωινγερ, ανδ Τομοναγα” : Πρινςετον ϒνιερςιτψ Πρεςς, Πρινςετον, Ν.Θ (1994).
  • [38] Ν.Ν. Βογολιυβο ανδ Ο. Παραςψυϰ, Δοϰλαδψ Αϰαδεμιι Ναυϰ ΣΣΣΡ 100, 25 (1955).
  • [39] Κ. Ηεππ ῝ομμ. Ματη. Πηψς. 2, 301 (1966). Σεε αλςο Ω. Ζιμμερμανν, ῝ομμ. Ματη. Πηψς. 11, 1 (1968) ανδ 15, 208 (1969).
  • [40] Η. Επςτειν ανδ ῞. Γλαςερ, Ανν. Ινςτ. Η. Ποινςαρé Α19, 211 (1973).
  • [41] ῞.Ν. Γριβο, Νυςλ. Πηψς. Β139, 1 (1978). Φορ α ρειεω ςεε: Ν. ῞ανδερςιςϰελ ανδ Δ. Ζωανζιγερ, Πηψς. Ρεπτ. 520, 175 (2012).
  • [42] Π.Α.Μ. Διρας, Προς. Ροψ. Σος. Α114, 243 (1927).
  • [43] Σ.Ν. Γυπτα, Προς. Πηψς. Σος. Α63, 681 (1950) · Κ. Βλευλερ,  Ηελ. Πηψς. Αςτα 23,  567 (1950).
  • [44] Ω. Παυλι ανδ Φ. ῞ιλλαρς, Ρε. Μοδ. Πηψς. 21, 434 (1949).
  • [45] Θ.῝. Ωαρδ, Πηψς. Ρε. 78, 182 (1950).
  • [46] Α. Προςα, Θ. δε Πηψς. ετ Λε Ραδ. 7, 347 (1936).
  • [47] Ε.῝.Γ. Στυεςϰελβεργ, Ηελ. Πηψς. Αςτα 11, 299 (1938). Φορ α μοδερν πρεςεντατιον ςεε Η. Ρυεγγ ανδ Μ. Ρυιζ-Αλταβα, Ιντ. Θ. Μοδ. Πηψς. Α19, 3265 (2004)· αλςο ηεπ-τη/0304245.
  • [48] Γ. Ωεντζελ, “Εινφϋηρυνγ ιν διε Χυαντεντηεοριε δερ Ωελλενφελδερ”, (Ιντροδυςτιον το τηε χυαντυμ τηεορψ οφ ωαε φιελδς), Φρανζ Δευτιςϰε, Ωιεν (1943).
  • [49] Ε.῝.Γ. Στυεςϰελβεργ, Ηελ. Πηψς. Αςτα 19, 242 (1946).
  • [50] Ω. Παυλι, Ρε. Μοδ. Πηψς. 13, 203 (1941).
  • [51] Η. Βορςηερς, Νυο. ῝ιμ. 15, 784 (1960)· Η. Επςτειν, Νυο. ῝ιμ. 27, 886 (1963).
  • [52] Ρ.Π. Φεψνμαν, Αςτα Πηψς. Πολονιςα 24, 697 (1963). Σεε αλςο Β.Σ. ΔεΩιττ, Πηψς. Ρε. Λεττ. 12, 742 (1964) ανδ Πηψς. Ρε. 162, 1195 (1967), Σ. Μανδελςταμ, Πηψς. Ρε. 175, 1580 ανδ 1604 (1968).
  • [53] Λ.Δ. Φαδδεε ανδ ῞.Ν. Ποπο, Πηψς. Λεττ. 25Β, 29 (1967).
  • [54] Φορ α ρειεω ςεε Σ. ῝ολεμαν, “Ρενορμαλιςατιον ανδ Σψμμετρψ. Α Ρειεω φορ νον-ςπεςιαλιςτς”, Λεςτυρες ιν Εριςε Συμμερ Σςηοολ, 1971, Ρεπριντεδ ιν Σιδνεψ ῝ολεμαν Αςπεςτς οφ Σψμμετρψ, ῝αμβριδγε ϒνιερςιτψ Πρεςς, 1985.
  • [55] Ε. Ωιττεν, Πηψς.Λεττ. 117Β, 324 (1982).
  • [56] Σ. Σαϰατα ανδ Ψ. Τανιϰαωα, Πηψς. Ρε. 57, 548 (1940).
  • [57] Θ. Στεινβεργερ, Πηψς. Ρε. 76, 1180 (1949)· ςεε αλςο Θ. Σςηωινγερ, Πηψς. Ρε. 82, 664 (1951).
  • [58] Θ. Σςηωινγερ, Πηψς. Ρε. 128, 2425 (1962).
  • [59] Μ. Γελλ-Μανν, Πηψς. 1, 63 (1964).
  • [60] Φορ α ςολλεςτιον οφ μανψ ιμπορταντ παπερς ον τηις ςυβϑεςτ ςεε Σ.Λ. Αδλερ ανδ Ρ.Φ. Δαςηεν, “῝υρρεντ Αλγεβρας ανδ Αππλιςατιονς το Παρτιςλε Πηψςιςς”, Ω.Α. Βενϑαμιν Ινς., Νεω Ψορϰ (1968) ανδ ρεφερενςες τηερειν.
  • [61] Δ.Γ. Συτηερλανδ, Νυςλ. Πηψς. Β2, 433 (1967).
  • [62] Θ.Σ. Βελλ ανδ Ρ. Θαςϰιω, Νυο. ῝ιμ. 60Α, 47 (1969).
  • [63] Σ.Λ. Αδλερ, Πηψς. Ρε. 177, 2426 (1969).
  • [64] Σ.Λ. Αδλερ ανδ Ω.Α. Βαρδεεν, Πηψς. Ρε. 182, 1515 (1969)· Ω.Α. Βαρδεεν, Πηψς. Ρε. 184, 1848 (1969).
  • [65] Θ. Ωεςς ανδ Β. Ζυμινο, Πηψς. Λεττ. 37Β, 95 (1971).
  • [66] Ε.῝.Γ. Στυεςϰελβεργ ανδ Α. Πετερμανν, Ηελ. Πηψς. Αςτα 24, 317 (1951)· Ηελ. Πηψς. Αςτα 26, 499 (1953).
  • [67] Μ. Γελλ-Μανν ανδ Φ.Ε. Λοω, Πηψς. Ρε. 95, 1300 (1954).
  • [68] Ν. Ν. Βογολιυβο ανδ Δ. ῞. Σηιρϰο, “Ιντροδυςτιον το τηε Τηεορψ οφ Χυαντιζεδ Φιελδς” Ενγλ. Τρανςλατιον, Ωιλεψ-Ιντερςςιενςε (1959).
  • [69] Κ. Σψμανζιϰ, ῝ομμ. Ματη. Πηψς. 18, 227 (1970) · ῝.Γ. ῝αλλαν, Πηψς. Ρε Δ2, 1541 (1970).
  • [70] Κ.Γ. Ωιλςον, Πηψς.Ρε. Δ3, 1818 (1971).
  • [71] Κ.Γ. Ωιλςον, Πηψς.Ρε. Β4, 3174 (1971)· Πηψς.Ρε. Β4, 3184 (1971)· Πηψς. Ρε. Λεττ. 28, 240 (1972)· Πηψς. Ρε. Δ7, 2911 (1973)· Κ.Γ. Ωιλςον ανδ Θ.Β. Κογυτ, Πηψς. Ρεπτ. 12, 75 (1974).
  • [72] Σ. Ωεινβεργ, Πηψς. Ρε. Δ8, π. 3497 (1973)· Σεε αλςο Μ. Αςταυδ ανδ Β. Θουετ, Νυο. ῝ιμ. 63Α, 5 (1969)· Νυο. ῝ιμ. 66Α, 111 (1970).
  • [73] Σ. ῝ολεμαν, “Διλατατιονς”, Λεςτυρες ιν Εριςε Συμμερ Σςηοολ, 1971, Ρεπριντεδ ιν Σιδνεψ ῝ολεμαν Αςπεςτς οφ Σψμμετρψ, ῝αμβριδγε ϒνιερςιτψ Πρεςς, 1985.
  • [74] Σ. Ωεινβεργ, Πηψς. Ρε. 118, 838 (1960)· Τ. Κινοςηιτα, Θ. Ματη. Πηψς. 3, 650 (1962)· Τ.Δ. Λεε ανδ Μ. Ναυενβεργ, Πηψς. Ρε. 133, Β1549 (1964).
  • [75] Λ.Δ. Λανδαυ ανδ Ι. Πομερανςηυϰ, Δοϰλ. Αϰαδ. Ναυϰ ΣΣΣΡ 102, 489 (1955).
  • [76] Λ.Δ. Λανδαυ, “Φυνδαμενταλ Προβλεμς”, Μ. Φιερζ ανδ ῞.Φ. Ωειςςϰοπφ, εδς. Τηεορετιςαλ Πηψςιςς ιν τηε Τωεντιετη ῝εντυρψ· α Μεμοριαλ ῞ολυμε το Ωολφγανγ Παυλι, Ιντερςςιενςε· Νεω Ψορϰ, 245 (1960).
  • [77] Σ. Σαϰατα, Προγρ. Τηεορ. Πηψς. 16, 686 (1956).
  • [78] Μ. Γελλ-Μανν, “Τηε ειγητφολδ ωαψ: α τηεορψ οφ ςτρονγ ιντεραςτιον ςψμμετρψ”. Ρεπορτ ῝ΤΣΛ-20 (1961) (υνπυβλιςηεδ)· Ψ. Νεἐμαν, Νυςλ. Πηψς. 26, 222 (1961). Φορ α ςολλεςτιον οφ μανψ ιμπορταντ αρτιςλες ον $SU(3)$ ςεε τηε ολυμε “Τηε ειγητφολδ ωαψ” Μ. Γελλ-Μανν ανδ Ψ. Νεἐμαν, Εδιτορς, Ω.Α. Βενϑαμιν Ινς., Νεω Ψορϰ (1964).
  • [79] Μ. Γελλ-Μανν, Πηψς. Λεττ. 8, 214 (1964)· Γ. Ζωειγ, ῝ΕΡΝ πρεπριντς Τη 401 ανδ Τη 412 (υνπυβλιςηεδ).
  • [80] Λ.Β. Οϰυν, αρΞι:ηεπ-πη/0611298 ανδ ρεφερενςες τηερειν.
  • [81] Ο.Ω. Γρεενβεργ, Πηψς. Ρε. Λεττ. 13, 598 (1964)· Μ.Ψ. Ηαν ανδ Ψ. Ναμβυ, Πηψς. Ρε. 139Β, 1006 (1965).
  • [82] Λ. Συςςϰινδ, “Χυαρϰ ῝ονφινεμεντ” ιν ρεφερενςε [11] ανδ ρεφερενςες τηερειν.
  • [83] Σ. ῝ολεμαν, “Σεςρετ Σψμμετρψ”, Λεςτυρες ιν Εριςε Συμμερ Σςηοολ, 1973, Ρεπριντεδ ιν Σιδνεψ ῝ολεμαν Αςπεςτς οφ Σψμμετρψ, ῝αμβριδγε ϒνιερςιτψ Πρεςς, 1985.
  • [84] Π. ῝υριε, Θουρν. δε Πηψς. τηéορ. ετ αππλ. 3, 393 (1894).
  • [85] Ψ. Ναμβυ, Πηψς. Ρε. Λεττ. 4, 380 (1960)· Ψ. Ναμβυ ανδ Γ. Θονα-Λαςινιο, Πηψς.Ρε. 122, 345 (1961)· Πηψς.Ρε. 124, 246 (1961).
  • [86] Θ. Γολδςτονε, Νυο. ῝ιμ. 19, 154 (1961)· ςεε αλςο Θ. Γολδςτονε, Α. Σαλαμ ανδ Σ. Ωεινβεργ, Πηψς. Ρε. 127, 965 (1962).
  • [87] Σ.Σ. Γερςητειν ανδ Ψα.Β. Ζελδοιςη, ΘΕΤΠ 29, 698 (1955), Τρανςλατιον: Σοιετ Πηψςιςς ΘΕΤΠ 2, 576 (1958)· Ρ.Π. Φεψνμαν ανδ Μ. Γελλ-Μανν, Πηψς. Ρε. 109, 193 (1958).
  • [88] Δ. Καςτλερ, Δ.Ω. Ροβινςον ανδ Α. Σωιεςα, ῝ομμ. Ματη. Πηψς. 2, 108 (1966).
  • [89] Μ. Γελλ-Μανν ανδ Μ. Λéψ, Νυο. ῝ιμ. 16, 605 (1960).
  • [90] Μ.Λ. Γολδβεργερ ανδ Σ.Β. Τρειμαν, Πηψς. Ρε. 110, 1178 (1958)· 111, 354 (1958).
  • [91] Λ.Δ. Λανδαυ ανδ Β.Λ. Γινζβυργ, ΘΕΤΠ 20, 1064 (1950). Νοτιςε τηατ νο ονε ιν τηε ςυβςεχυεντ λιςτ οφ αυτηορς ωηο ςτυδψ τηις πηενομενον ρεφερς το τηις αρτιςλε.
  • [92] Θ. Βαρδεεν, Λ.Ν. ῝οοπερ ανδ Θ.Ρ. Σςηριεφφερ, Πηψς. Ρε. 108, 1175 (1957).
  • [93] Π.Ω. Ανδερςον, Πηψς. Ρε. 110, 827 (1958) · Πηψς. Ρε. 112, 1900 (1958).
  • [94] Ψ. Ναμβυ, Πηψς. Ρε. 117, 648 (1960).
  • [95] Θ.Σ. Σςηωινγερ, Πηψς. Ρε. 125, 397 (1962) · Πηψς. Ρε. 128, 2425 (1962).
  • [96] Π.Ω. Ανδερςον, Πηψς. Ρε. 130, 439 (1963).
  • [97] Μ. Λéψ, Πηψς. Λεττ. 7, 36 (1963)· Νυςλ. Πηψς. 57, 152 (1964).
  • [98] Α. Κλειν ανδ Β.Ω. Λεε, Πηψς. Ρε. Λεττ. 12, 266 (1964) · Μ. Βαϰερ, Κ. Θοηνςον ανδ Β.Ω. Λεε, Πηψς. Ρε. 133, 209 (1964).
  • [99] Ω. Γιλβερτ, Πηψς. Ρε. Λεττ. 12, 713 (1964).
  • [100] Φ. Ενγλερτ ανδ Ρ. Βρουτ, Πηψς. Ρε. Λεττ. 13, 321 (1964).
  • [101] Π. Ηιγγς, Πηψς. Λεττ. 12, 132 (1964) · Πηψς. Ρε. Λεττ. 13, 508 (1964).
  • [102] Γ.Σ. Γυραλνιϰ, ῝.Ρ. Ηαγεν ανδ Τ.Ω.Β. Κιββλε, Πηψς. Ρε. Λεττ. 13, 585 (1964).
  • [103] Σ. Ωεινβεργ, Πηψς. Ρε. Λεττ. 19, 1264 (1967)· ςεε αλςο · Α. Σαλαμ, Λεςτυρε γιεν ατ Ειγητη Νοβελ Σψμποςιυμ, Ν Σαρτηολμ, Εδ. π. 367 (1968).
  • [104] Σ.Λ. Γλαςηοω, Θ. Ιλιοπουλος ανδ Λ. Μαιανι, Πηψς. Ρε. Δ2, 1285 (1970).
  • [105] Φορ μψ περςοναλ ρεμινιςςενςες οφ τηις περιοδ ςεε Θ. Ιλιοπουλος, “Ωηατ α φουρτη χυαρϰ ςαν δο” ιν ρεφερενςε [11] ανδ “Τηε Μαϰινγ οφ τηε Στανδαρδ Τηεορψ” ιν ρεφερενςε [12].
  • [106] Β.Λ. Ιοφφε ανδ Ε.Π. Σηαβαλιν, Ψαδ. Φιζ. 6, 828 (1967)· Ενγλιςη τρανςλατιον Σο. Θουρ. Νυςλ. Πηψς. 6, 603 (1968)· Ζ. Εϰςπ. ι Τεορ. Φιζ. Πιςμα ϒ Ρεδαϰτςιψυ 6, 978 (1967)· Ενγλιςη τρανςλατιον Σοιετ. Πηψς. ΘΕΤΠ Λεττ. 6, 390 (1967)· ςεε αλςο Ρ.Ν. Μοηαπατρα, Θ. Συββα Ραο ανδ Ε. Μαρςηαϰ, Πηψς. Ρε. Λεττ. 20, 1081 (1968)· Πηψς. Ρε. 171, 1502 (1968)· Φ.Ε. Λοω, ῝ομμεντς Νυςλ. Παρτ. Πηψς. 2, 33 (1968)· Ρ.Ν. Μοηαπατρα ανδ Π. Ολεςεν, Πηψς. Ρε. 179, 1917 (1969).
  • [107] Τ.Δ. Λεε ανδ Γ.῝. Ωιςϰ, Νυςλ. Πηψς. Β9, 209 (1969).
  • [108] Ω. Κυμμερ ανδ Γ. Σεγρέ, Νυςλ. Πηψς. 64, 585 (1965).
  • [109] Μ. Γελλ-Μανν, Μ.Λ. Γολδβεργερ, Ν.Μ. Κρολλ ανδ Φ. Λοω, Πηψς. Ρε. 179, 1518 (1969).
  • [110] Ρ. Γαττο, Γ. Σαρτορι ανδ Μ. Τονιν, Πηψς. Λεττ. 28Β, 128 (1968)· Νυο. ῝ιμ. Λεττ. (1, 1 (1969).
  • [111] Ν. ῝αβιββο ανδ Λ. Μαιανι, Πηψς. Λεττ. 28Β, 131 (1968)· Πηψς. Ρε. Δ1, 707, (1970).
  • [112] ῝λ. Βουςηιατ, Θ. Ιλιοπουλος ανδ Θ. Πρεντϰι, Νυο. ῝ιμ. 56Α, 1150 (1968)· Θ. Ιλιοπουλος, Νυο. ῝ιμ. 62Α, 209 (1969).
  • [113] Ν. ῝αβιββο, Πηψς. Ρε. Λεττ. 10, 531 (1963).
  • [114] Σεε, φορ εξαμπλε, Δ. Αματι, Η. Βαςρψ, Θ. Νυψτς ανδ Θ. Πρεντϰι, Νυο. ῝ιμ. 34, 1732 (1964) · Θ. Δ. Βϑορϰεν ανδ Σ.Λ. Γλαςηοω, Πηψς. Ρε. 11, 255 (1964).
  • [115] Μ. Κοβαψαςηι ανδ Τ. Μαςϰαωα, Προγ. Τηεορ. Πηψς. 49, 652 (1973).
  • [116] Σ.Λ. Γλαςηοω ανδ Θ. Ιλιοπουλος, Πηψς. Ρε. Δ3, 1043 (1971)· Σ.Λ. Γλαςηοω, Θ. Ιλιοπουλος ανδ Α. Ψαο, Πηψς. Ρε. Δ4, 1918 (1971).
  • [117] Μ.Θ.Γ. ῞ελτμαν, Πηψς. Ρε. Λεττ. 17, 553 (1966) · Θ.Σ. Βελλ, Νυοο ῝ιμ. 50Α, 129 (1967).
  • [118] Τ.Δ. Λεε ανδ ῝.Ν. Ψανγ, Πηψς. Ρε. 128, 885 (1962) · Τ.Δ. Λεε, Πηψς. Ρε. 128, 899 (1962).
  • [119] Μ.Θ.Γ. ῞ελτμαν, ῝ομπυτ. Πηψς. ῝ομμυν. 3, 75 (1972).
  • [120] Ρ.Ε. ῝υτϰοςϰψ, Θ. Ματη. Πηψς. 1, 429 (1960).
  • [121] Μ.Θ.Γ. ῞ελτμαν, “Τηε Πατη το Ρενορμαλιζαβιλιτψ” ιν ρεφερενςε [11].
  • [122] Μ.Θ.Γ. ῞ελτμαν, Νυςλ. Πηψς. Β7, 637 (1968), Νυςλ. Πηψς. Β21, 288 (1970). Φορ αν εξτενςιον το τωο λοοπς, ςεε Θ. Ρειφφ ανδ Μ.Θ.Γ. ῞ελτμαν, Νυςλ. Πηψς. Β13, 545 (1969).
  • [123] Η. ῞αν Δαμ ανδ Μ.Θ.Γ. ῞ελτμαν, Νυςλ. Πηψς. Β22, 397 (1970).
  • [124] Α.Α. Σλανο ανδ Λ.Δ. Φαδδεε, Τεορετ. ι Ματεμ. Φιζιϰα 3, 18 (1970).
  • [125] Ε.Σ. Φραδϰιν ανδ Ι.῞. Τψυτιν, Πηψς. Λεττ. 30Β, 562 (1969)· Δ.Γ. Βουλωαρε, Ανν. οφ Πηψς. 56, 140 (1970).
  • [126] Μ.Θ.Γ. ῞ελτμαν, Αςτα Πηψς. Πολονιςα Β8, 475 (1977).
  • [127] Γ. ´τ Ηοοφτ, “Ρενορμαλιζατιον οφ Γαυγε Τηεοριες” ιν ρεφερενςε [11] ανδ “Τηε Εολυτιον οφ Χυαντυμ Φιελδ Τηεορψ: Φρομ ΧΕΔ το Γρανδ ϒνιφιςατιον” ιν ρεφερενςε [12].
  • [128] Γ. ´τ Ηοοφτ, Νυςλ. Πηψς. Β33, 173 (1971), Νυςλ. Πηψς. Β35, 167 (1971).
  • [129] Γ. ´τ Ηοοφτ ανδ Μ.Θ.Γ. ῞ελτμαν, Νυςλ. Πηψς. Β44, 189 (1972) ανδ Νυςλ. Πηψς. Β50, 318 (1972). Σεε αλςο Γ. ´τ Ηοοφτ ανδ Μ.Θ.Γ. ῞ελτμαν, “Διαγραμμαρ”, ῝ΕΡΝ ψελλοω ρεπορτ, 73-9 (1973).
  • [130] ῝.Γ. Βολλινι ανδ Θ.Θ. Γιαμβιαγι, Πηψς. Λεττ. 40Β, 566 (1972), Ιλ Νυο. ῝ιμ. 12Β, 20 (1972).
  • [131] Α.Α. Σλανο, Τεορ. ι Ματ. Φιζ. 10, 153 (1972), Ενγλ. Τρανςλ. Τηεορ. ανδ Ματη. Πηψς. 10, 99 (1972)· Θ.῝. Ταψλορ, Νυςλ. Πηψς. Β33, 436 (1971).
  • [132] Β.Ω. Λεε, Πηψς. Ρε. Δ5, 823 (1972)· Β.Ω. Λεε ανδ Θ. Ζινν-Θυςτιν, Πηψς. Ρε. Δ5, 3121· Δ5, 3137· Δ5, 3155 (1972)· Δ7, 1049 (1973).
  • [133] ῝. Βεςςηι, Α. Ρουετ ανδ Ρ. Στορα, Πηψς. Λεττ. 52Β, 344 (1974), ῝ομμ. Ματη. Πηψς. 42, 127 (1975), Ανν. οφ Πηψς. 98, 287 (1976).
  • [134] Ι.῞. Τψυτιν, Λεβεδε Πηψςιςς Ινςτιτυτε πρεπριντ 39 (1975), αρΞι:0812.0580.
  • [135] ῝.Η. Λλεωελλψν-Σμιτη, Πηψς. Λεττ. 46Β, 233 (1973)· Θ.Σ. Βελλ, Νυςλ. Πηψς. Β60, 427 (1973).
  • [136] Θ.Μ. ῝ορνωαλλ, Δ.Ν. Λειν ανδ Γ. Τιϰτοπουλος, Πηψς. Ρε. Λεττ 30, 1268 (1973) ανδ Ερ. 31, 572 (1973)· Πηψς.Ρε. Δ10, 1145 (1974) ανδ Ερ. Δ11, 1145 (1975)· Δ.Ν. Λειν ανδ Γ. Τιϰτοπουλος, Πηψς.Ρε. Δ12, 415 (1975).
  • [137] Μ.Α. Βουςηιατ ανδ ῝.῝. Βουςηιατ, Πηψς.Λεττ. 48Β, 111 (1974), Θουρναλ δε Πηψς. 35, 899 (1974) ανδ 36, 493 (1975).
  • [138] Β.Ω. Λεε, Πηψς. Ρε. Δ6, 1188 (1972)· Θ. Πρεντϰι ανδ Β. Ζυμινο, Νυςλ. Πηψς. Β47, 99 (1972).
  • [139] Η. Γεοργι ανδ Σ.Λ. Γλαςηοω, Πηψς. Ρε. Λεττ. 28, 1494 (1972).
  • [140] Φ.Θ. Ηαςερτ ετ αλ., Πηψς. Λεττ. 46Β, 121 (1973), 46Β, 138 (1973), Νυςλ. Πηψς. Β73, 1 (1974). Σεε αλςο Δ.῝. ῝υνδψ, “Νευτρινο Πηψςιςς”, Πλεναρψ Σεςςιον Ρεπορτ, Ξ῞ΙΙ ῝ονφ. ον Ηιγη Ενεργψ Πηψςιςς, Λονδον (1974).
  • [141] ῝.Ψ. Πρεςςοττ ετ αλ., Πηψς. Λεττ. 77Β, 347 (1978).
  • [142] Μ.Α. Βουςηιατ, Θ. Γυενα, Λ. Ηυντερ ανδ Λ. Ποττιερ, Πηψς. Λεττ. 117Β, 358 (1982) ανδ Ερ. 121Β, 456 (1983).
  • [143] ῝λ. Βουςηιατ, Θ. Ιλιοπουλος ανδ Πη. Μεψερ, Πηψς. Λεττ. 38Β, 519 (1972)· Δ.Θ. Γροςς ανδ Ρ. Θαςϰιω, Πηψς. Ρε. Δ6, 477 (1972)· Η. Γεοργι ανδ Σ.Λ. Γλαςηοω, Πηψς. Ρε. Δ6, 429 (1972).
  • [144] Δ.Θ. Γροςς, “Αςψμπτοτις Φρεεδομ ανδ τηε Εμεργενςε οφ Χ῝Δ” ιν ρεφερενςε [11].
  • [145] Ρ.Κ. Ελλις, “Χυαντυμ ῝ηρομοδψναμιςς ανδ Δεεπ Ινελαςτις Σςαττερινγ” ιν ρεφερενςε [12].
  • [146] Θ.Δ. Βϑορϰεν, “῝υρρεντ Αλγεβρα ατ Σμαλλ Διςτανςες”, ιν Προςεεδινγς οφ τηε Ιντερνατιοναλ Σςηοολ οφ Πηψςιςς “Ενριςο Φερμι,” ῞αρεννα, Ιταλψ (1967)· Πηψς. Ρε. 179, 1547 (1969).
  • [147] Ρ.Π. Φεψνμαν, Πηψς. Ρε. Λεττ. 23, 1415 (1969)· Θ.Δ. Βϑορϰεν ανδ Ε.Α. Παςςηος, Πηψς. Ρε. 185, 1975 (1969).
  • [148] Η. Φριτζςςη ανδ Μ. Γελλ-Μανν, ῝οντριβυτιον το: ῝οραλ Γαβλες ῝ονφερενςε Ον Φυνδαμενταλ Ιντεραςτιονς Ατ Ηιγη Ενεργψ (1971)· Σεε αλςο ηεπ-πη/0301127.
  • [149] Φορ α ρειεω ςεε Α. Γροζιν, “Αςψμπτοτις φρεεδομ: ηιςτορψ ανδ ιντερπρετατιον”, ε-Πριντ: 0803.2589 [πηψςιςς.ηιςτ-πη]
  • [150] ῞.Σ. ῞ανψαςηιν ανδ Μ.῞. Τερεντε, ΖηΕΤΦ 48, 565 (1965), Ενγλ. Τρανς., Σο. Πηψς. ΘΕΤΠ 21, 375 (1965).
  • [151] Ι.Β. Κηριπλοιςη, Ψαδ. Πηψς. 10, 409 (1969), Ενγλ. Τρανς., Σο. Θ. Νυςλ. Πηψς. 10, 235 (1970).
  • [152] Θ. Ιλιοπουλος ανδ Ε.Α. Παςςηος, Πηψς. Ρε. Δ6, 1340 (1972).
  • [153] Η.Δ. Πολιτζερ, Πηψς. Ρε. Λεττ. 30, 1346 (1973)· Δ.Θ. Γροςς ανδ Φ. Ωιλςζεϰ, Πηψς. Ρε. Λεττ. 30, 1343 (1973), Πηψς. Ρε. Δ8, 3633 (1973)· Πηψς. Ρε. Δ9, 980 (1974).
  • [154] Σ. ῝ολεμαν ανδ Δ.Θ. Γροςς, Πηψς. Ρε. Λεττ. 31, 851 (1973).
  • [155] Η. Φριτζςςη ανδ Μ. Γελλ-Μανν, ῝οντριβυτιον το: Ξ῞Ι ῝ονφ. ον Ηιγη Ενεργψ Πηψςιςς, Βαταια, Ιλλ (1972) αρΞι:ηεπ-πη/0208010.
  • [156] Η. Φριτζςςη, Μ. Γελλ-Μανν ανδ Η. Λευτωψλερ, Πηψς. Λεττ. 47Β, 365 (1973).
  • [157] Α. Ρυςςο, Ηιςτοριςαλ ρεπορτ ον ΙΣΡ, Ρεφερενςε [1].
  • [158] Κ.Γ. Ωιλςον, Πηψς. Ρε. Δ3, 1819 (1971)· Κ.Γ. Ωιλςον ανδ Θ. Κογυτ, Πηψς. Ρεπ. 12῝, 75 (1974).
  • [159] Κ.Γ. Ωιλςον, Πηψς. Ρε. 179, 1499 (1969).
  • [160] Η. Λευτωψλερ ανδ Θ. Στερν, Νυςλ. Πηψς. Β20, 77 (1970)· Ρ. Θαςϰιω, Ρ. ῞αν Ροψεν ανδ Γ.Β. Ωεςτ, Πηψς. Ρε. Δ2, 2473 (1970)· Ψ. Φριςημαν, Ανν. Πηψς. 66, 373 (1971)· Ρ.Α. Βρανδτ ανδ Γ. Πρεπαρατα, Νυςλ. Πηψς. Β27, 541 (1971).
  • [161] Τ. Αππελχυιςτ ανδ Η. Γεοργι, Πηψς. Ρε. Δ8, 4000 (1973)· Α. Ζεε, Πηψς. Ρε. Δ8, 4038 (1973).
  • [162] Η.Δ. Πολιτζερ, Δ.Θ. Γροςς ανδ Φ. Ωιλςζεϰ, [153] · Γ. Παριςι, Πηψς. Λεττ. Β50, 367 (1974).
  • [163] Θ. Ιλιοπουλος, “Προγρεςς ιν γαυγε τηεοριες”, Πλεναρψ ςεςςιον ρεπορτ, Ξ῞ΙΙ ῝ονφ. ον Ηιγη Ενεργψ Πηψςιςς, Λονδον (1974).
  • [164] Σεε, φορ εξαμπλε, Γ. Τιϰτοπουλος, Νυςλ. Πηψς. Β147, 371 (1979)· Νυςλ. Πηψς. Β173, 291 (1980).
  • [165] Γ. Αλταρελλι ανδ Γ. Παριςι, Νυςλ. Πηψς. Β126, 298 (1977)· Ψ. Λ. Δοϰςηιτζερ, Σο. Πηψς. ΘΕΤΠ 46, 641 (1977)· ῞. Γριβο ανδ Λ. Λιπατο, Σο. Θ. Νυςλ. Πηψς. 15, 438 (1972)· 15, 675 (1972).
  • [166] Λ. Βαυλιευ ανδ ῝. Κουννας, Νυςλ. Πηψς. Β141, 423 (1978).
  • [167] Κ.Γ. Ωιλςον, Πηψς. Ρε. Δ10, 2445 (1974)· Φορ αν εαρλψ ρειεω ςεε Θ.Β. Κογυτ, Ρε. Μοδ. Πηψς. 51, 659 (1979) ανδ ρεφερενςες τηερειν.
  • [168] Η.Β. Νιελςεν ανδ Μ. Νινομιψα, Πηψς. Λεττ. Β105, 219 (1981).
  • [169] Τηερε ις α ρειεω ον λαττιςε ςαλςυλατιονς ιν πραςτιςαλλψ εερψ μαϑορ ηιγη ενεργψ πηψςιςς ςονφερενςε. Σεε αλςο ῝.Τ. Σαςηραϑδα, ςοντριβυτιον ιν ρεφερενςε [12].
  • [170] Τ. Αππελχυιςτ ανδ Η.Δ. Πολιτζερ, Πηψς. Ρε. Λεττ. 34, 43 (1975)· Α. Δε Ρυϑυλα ανδ Σ.Λ. Γλαςηοω, Πηψς. Ρε. Λεττ. 34, 46 (1975).
  • [171] Α. Δε Ρυϑυλα, “Τηεορετιςαλ βαςις οφ τηε νεω παρτιςλες” Ρεπορτ, Ξ῞ΙΙΙ ῝ονφερενςε ον Ηιγη Ενεργψ Πηψςιςς, Τβιλιςςι (1976) ανδ ρεφερενςες τηερειν.